### v1

# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

Abstract
Seymour’s Second Neighborhood Conjecture asserts that in the square of any oriented graph, there exists a node whose out-degree at least doubles. This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph-Level Oriented Vertex Expansion and Reduction) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions. By leveraging this structure, we construct a decreasing sequence of subsets that establish a well-ordering of nodes, ensuring that no counterexample can exist. This approach not only confirms the conjecture for all oriented graphs but also provides a novel framework for analyzing degrees and arcs in complex networks. The findings have implications for theoretical graph studies and practical applications in network optimization and algorithm design.
1 Introduction
Seymour’s Second Neighborhood Conjecture, proposed by Paul Seymour in 1990, seeks to identify a node in an oriented graph whose out-degree at least doubles in the graph’s square. The conjecture was first published by Nathaniel Dean and Brenda J. Latka in 1995 [4], where they posed a similar conjecture specific to tournaments. This version of the conjecture was first proven by Fischer [7], with a second proof provided by Havet and Thomasse [10].
Conjecture 1.1.
(Seymour’s Second Neighborhood Conjecture). For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}$ represent the first and second neighborhoods of $v$ respectively.
This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [8]) to encompass a broader class of oriented graphs, thereby providing a framework applicable across various minimum degree scenarios.
The Second Neighborhood Conjecture presents a compelling challenge within the context of graph theory, focusing on oriented graphs and the behavior of second neighbors in the graph’s square. Efforts to resolve this conjecture reflect the deep intrigue and mathematical beauty inherent in such problems. Beyond its theoretical importance, the conjecture has potential applications in network theory, algorithm design, and combinatorial optimization—fields where optimizing connections between nodes is crucial.
While Seymour’s Conjecture remains an open problem, various efforts have tackled special cases. For example, Kaneko and Locke [11] demonstrated its truth for graphs with a minimum degree of at most six. Brantner et al. [1] showed that, if the conjecture is false, infinitely many counterexamples must exist. Additionally, Chen et al. [3] provided a lower bound for how much a node’s out-degree increases in the graph’s square, concluding that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1. In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture. This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
In this paper, we will prove Seymour’s Conjecture by assuming it is false and then attempting to construct a counterexample. Specifically, we explore the conditions under which no node’s out-degree doubles in the square of an oriented graph. Our approach focuses initially on minimum out-degree nodes, analyzing how their out-degree could fail to double, and gradually extending this analysis to other nodes based on their distance from the minimum out-degree node.
The Seymour Second Neighborhood Conjecture has long been considered an intractable problem in graph theory, with some experts even doubting its validity. This paper presents a rigorous proof of the conjecture, accompanied by extensive supporting materials, including appendices, detailing definitions, lemmas, examples and an algorithm, as well as an interactive website for visualizing key ideas. These resources aim to make the proof as transparent and accessible as possible, acknowledging the high standards of scrutiny such results rightly deserve.
The paper is organized as follows: Section 2 begins with some common terminology used in graph theory. Section 3 introduces our approach to the problem. In section 4 we look at some small examples of this style of thinking. Section 5 introduces the terminology and identifying the conditions under which their out-degree might fail to double. Section 6 examines the back arcs and their place in oriented graphs. Section 7 looks at the conditions which would cause an interior degree to double, and its importance to the Seymour Second Neighborhood Conjecture. Section 8 further looks at containers under the context of density constraints. In Section 9 considers transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm. Section 10 presents this process the full Decreasing Sequence Property algorithm as well as the theorem stating it’s solution to the Conjecture.
2 Graph Theory Terminology
A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(x,x)$ where $x$ is a node in $G$) and no symmetric arcs, that is no arcs of the form $(x,y)$ and $(y,x)$ where $x$ and $y$ are nodes in $G$. The square of a graph $G$, denoted $G^{2}$, consists of the same node set as $G$ and an arcs exists from node $x$ to node $z$ if there is a node $y$ such that both arcs $(x,y)$ and $(y,z)$ exist in $G$ as well as the arcs already in $G$. Essentially, $G^{2}$ captures two-step connections from $x$ to $z$.
Definition 2.1.
Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of edges defined as:
$E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$, which are nodes that $v$ has arcs pointing to. Similarly, $N^{++}(v)$ refers to the second neighbors of $v$–that is, nodes that are two steps away from $v$ in $G$.
Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices. The induced subgraph $G[S]$ is the graph whose vertices are the set $S$, and whose edge set consists of all edges in $E$ that have both endpoints in $S$. For any node $v$, we refer to the neighbor-induced subgraph on $N^{+}(v)$, consisting of $v$’s out-neighbors. Let $u,v$ be nodes in $G$. The distance between $dist(u,v)$, is $k$ if the shortest path from $u$ to $v$ has length $k$.
Example 2.1.
[…]
Figure 1: A five-vertex graph illustrating node out-degrees and distances. Out-degrees are labeled near each node, and the dashed rectangle indicates an induced subgraph on vertices A, B, and C. Previously used in [8]
3 Approach and Motivation
We begin our analysis by considering general oriented graphs without making any initial assumptions about node degrees or graph properties. There are, however, fundamental characteristics common to all graphs, such as the existence of an out-degree distribution. Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes. By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
From this foundation, we approach the problem from two perspectives. The first is the graph $k$-partition problem [9], where the original oriented graph $G$ is partitioned into $d_{v_{0}}+1$ subgraphs, corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$. This partitioning enables us to leverage the divide-and-conquer paradigm, which is widely recognized in computer science for simplifying complex problems by tackling smaller, independent subproblems. These partitions enable us to decouple the analysis of interior relationships within each partition from the exterior interactions between partitions, allowing us to address each separately before examining how they integrate within the overall structure. Although the partitions remain connected through certain edges, each can be analyzed independently, allowing us to address subproblems in isolation before combining results to solve the overall problem.
By analyzing each partition individually, we aim to gain localized insights into the graph’s structure and degree dynamics, which we can then synthesize to form a cohesive solution. This divide-and-conquer approach has demonstrated considerable success in algorithms such as mergesort, quicksort, and matrix multiplication. While divide-and-conquer is not commonly applied in graph theory—largely because graphs are typically viewed as unordered collections of vertices and edges—we seek to view this differently by introducing a total ordering on the nodes of $G$. This will allow us to treat the graph in a way that is amenable to divide-and-conquer strategies.
While previous approaches to Seymour’s Second Neighborhood Conjecture often assume the location of a degree-doubling node and focus their analysis around this fixed point, our approach adopts a search-based strategy. Rather than assuming where the doubling will occur, this paper introduces an algorithm that systematically searches the graph, identifying degree-doubling nodes based on structural properties. This method offers a flexible and robust framework that does not depend on preexisting assumptions, allowing for a comprehensive analysis applicable to a broader range of oriented graphs.
This fundamentally shifts the way we approach graph theory problems by introducing a hierarchical structure that is conducive to divide-and-conquer techniques. This reimagining of graph structure allows us to efficiently isolate subproblems, analyze them independently, and then integrate results—offering both computational efficiency and scalability.
One of the key advantages of the structure is its capacity to handle multiple metrics: an outer metric (e.g., distance) for partitioning the graph, and an inner metric (e.g., degree) for analyzing each partition’s internal structure. This dual-metric approach is powerful because it allows for a more nuanced exploration of graph properties, particularly in cases where traditional methods would be insufficient.
The algorithm introduced in this paper systematically searches the graph, leveraging the graph’s structure, total ordering and degree metrics. By not presupposing the location of degree-doubling nodes, the algorithm dynamically identifies these nodes, ensuring a comprehensive exploration of the graph’s structure.
Previous work on Seymour’s Second Neighborhood Conjecture often centers around assuming the position of a degree-doubling node and anchoring the analysis around this fixed point. In contrast, our approach eliminates this assumption, offering a more flexible framework that dynamically adapts to the structure of the graph. This ability to apply a divide-and-conquer strategy to a broader set of graphs marks a significant step forward in graph theory.
4 Initial Lemmas
These first lemmas will seek to place a lower bound on the out-degree of a minimum out-degree node in a counterexample. For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$. We will also introduce some standard notation from graph theory. The work in this section agrees with what was already done by [11] and is not meant to serve as original content. It does however, provide insight into our reasoning.
We begin our investigation of oriented graphs by searching for a node whose degree could potentially double. While many candidates exist, our goal is to find a node such that, if its degree does not double, we still gain useful information about the structure of the graph. This approach allows us to proceed iteratively, analyzing the graph node by node until we either exhaust all possibilities or discover a counterexample.
Example 4.1.
$v_{0}$$v_{2}$$v_{3}$$v_{1}$$v_{6}$$v_{5}$$v_{4}$
$\Rightarrow H$
Figure 2: Example illustrating how a minimum out-degree node in $G$ has its degree doubled in $G^{2}$ under certain conditions.
Lemma 4.1.
Minimum Out-Degree ¡ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
Proof.
We will consider the cases where $d^{+}(v_{0})=0$, $1$, and $2$.
Case 1: Assume that $d^{+}(v_{0})=0$. Then $v_{0}$ has no outgoing edges in $G$. In $G^{2}$, it will still have out-degree 0, and hence its out-degree trivially satisfies the condition of at least doubling.
Case 2: Assume that $d^{+}(v_{0})=1$. Then $v_{0}$ is connected to some node $v_{1}\in G$. Since $v_{0}$ is a minimum out-degree node, we know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$. Because $G$ is an oriented graph and $(v_{0},v_{1})\in G$, it follows that $(v_{1},v_{0})\notin G$. Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$, meaning $(v_{1},v_{2})\in G$. This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
Case 3: Assume that $d^{+}(v_{0})=2$. Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$. Since $v_{0}$ has the minimum out-degree, we know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$. As $G$ is an oriented graph, $(v_{1},v_{0})$ and $(v_{2},v_{0})$ are not in $G$. Consider the subgraph $G[v_{1},v_{2}]$, induced on the nodes ${v_{1},v_{2}}$. There is at most one directed edge between $v_{1}$ and $v_{2}$ in $G$, meaning one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes. Let these nodes be $v_{3}$ and $v_{4}$. Since $G$ is oriented, $v_{0}$ cannot be one of these nodes. Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4, thus doubling its out-degree.
Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
Lemma 4.1 Minimum Out-Degree ¡ 3 gives us a glimpse on the picture of an oriented graphs. It tells us that, even with just the concept of a minimum degree node we are able to impose restrictions on not only that node $v_{0}$, but also on every other node $v\in G$. This later part is true because $v_{0}$ is a minimum degree node and so every other node must have at least the out-degree of $v_{0}$. What this lemma is saying is that cases where $v_{0}$ has very few out-neighbors, its second neighborhood will expand significantly. This aligns with the idea that low out-degree nodes should see a substantial increase in neighbors when considering the second neighborhood.
Lemma 4.2.
Minimum Out-Degree 3 with Neighbor Out-Degree 0 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3, and at least one of its neighbors has out-degree 0 in the induced subgraph of its neighbors, then the out-degree of $v_{0}$ will at least double in $G^{2}$.
Proof.
Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$. Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$ Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to nodes outside of $\{v_{1},v_{2},v_{3}\}$, as $v_{1}$ has no out-neighbors within the induced subgraph.
Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in G.
Thus, $v_{0}$’s out-degree will at least double in $G^{2}$.
Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors is somewhat ”isolated” within the neighborhood, leading to an increased second neighborhood for $v_{0}$.
Example 4.2.
0
$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$ H
Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$. Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$. The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$. While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$. Thus, in $G^{2}$, the degree of $v_{0}$ will double.
Lemma 4.3.
Minimum Out-Degree 3 with Neighbors 1 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3 and all of its neighbors have an out-degree of 1 in the induced subgraph of its neighbors, then $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}.$
Proof.
Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ Since $v_{0}$ is a minimum out-degree node, we know $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$. Let $G[v_{1},v_{2},v_{3}]$ represent the subgraph induced by $v_{1},v_{2},v_{3}$.
We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$, and by Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0) , none of $v_{1}$,$v_{2},$ or $v_{3}$ have out-degree 0 in $G[v_{1},v_{2},v_{3}]$.
Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$. Therefore, assume they connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree ¡ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case. There is at most one edge between $v_{4}$ and $v_{5}$. One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$. As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
Thus, either $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}$. ∎
This lemma underscores a crucial insight: while a node’s out-degree doubles, it is not necessarily the initial minimum out-degree node $v_{0}$ whose degree doubles. Instead, the doubling occurs in one of its neighbors. This distinction is significant because it challenges prior approaches that focus on selecting the ”right” starting node for analysis.
Our approach does not depend on starting with a specific node but instead leverages a search process that inherently reveals the doubling behavior in the graph structure. This framework allows for a more general and robust analysis.
For the remainder of this paper, we will refer to graphs like those in Lemma 4.2 using double arrows to indicate that a node or set of nodes connects to every node in the subsequent subgraph. The graph can be illustrated as follows:
The initial lemmas serve as essential building blocks in our exploration. By establishing these results, we create a solid foundation for the more profound theoretical analysis that follows. Each lemma reinforces the central theme of understanding the interplay between degrees and connectivity in oriented graphs.
5 Theoretical Foundation
Building on the initial lemmas, this section develops the theoretical foundation necessary for addressing the conjecture. We explore key properties such as containers, the Decreasing Sequence Property and the concepts of interior and exterior degrees, which are vital to our argument.
Example 5.1.
[…]
Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from the minimum degree node, with no assumptions on container relationships or sizes.
Definition 5.1.
Given a minimum degree node $v_{0}$, a container $C_{i}$ is the set of nodes that are at distance $i$ from $v_{0}$, along with the arcs between those nodes. More formally, $C_{i}=\{u\in V(G)\text{ .s.t. }dist(v_{0},u)=i\}\text{ and A}(C_{i})=\{(u,v)|u,v\in C_{i}\}$ ,where $V(G)$ represents the vertex set of the graph $G$, and $dist(v_{0},u)$ is the shortest path distance between $v_{0}$ and $u$.
As the previous definition aligns with the concept of double arrows, the term ”containers” allows us to explore the structure of oriented graphs and their interactions with minimum degree nodes and distances from these nodes. By introducing containers, we can focus not only on the properties of the nodes relative to a minimum degree node, but also on their interactions within and across containers.
This framework enables us to examine node behavior in three key aspects. First, we can study node behavior relative to a minimum degree node. Containers group nodes based on their distance from a minimum degree node $v_{0}$, making it easier to analyze how these nodes behave and how their properties evolve as we move further from $v_{0}$.
Second, we can investigate internal interactions within containers. Once nodes are grouped into containers, we can study how nodes within the same container influence one another. These internal interactions may reveal patterns related to degree growth, connectivity, and second-degree effects.
Containers are an attempt to bring order to an oriented graph. We mean that in the literal sense of the word, order. For nodes belonging to two different containers, they can now be compared by their distance to a minimum degree node $v_{0}$ which we call in container 0. Since distance is a total ordering, this allows to compare over every other node in $G$
The next few chapters will speak about how these containers communicate with each other and within themselves. No nodes or arcs are removed from the graph to create these containers. This is just a way of thinking about the graph so as to perform our analysis in search of a counterexample.
Finally, we can explore inter-container interactions. By analyzing interactions between nodes from different containers, we gain insight into how graph components expand or contract, how nodes influence one another across distances, and how arc directions impact the convergence of the algorithm or counterexample we aim to develop.
The case where $d^{+}(v_{0})=3$, and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) in particular, is an important one because it provides a glimpse into what a counterexample to the conjecture may look like. We were able to construct a graph $G$ where $v_{0}$’s out-degree does not double in $G^{2}$. Simple graphs of out-degree 3 are the smallest possibilities of cycles in undirected graphs. So if we orient the edges in the same direction we can get a cycle of order 3 among the nodes adjacent to $v_{0}$. If we look at this situation further we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double. In this section we will begin exploring the sufficient conditions for other nodes in the graph to not have their degrees double.
To continue along this line of reasoning though we will need to define new terms. First we will need a term for a node’s out-degree not doubling. A node in $G$ whose out-degree does not double in $G^{2}$ is said to have the decreasing sequence property. We have seen that examples exist where $v_{0}$ does not have its out-degree double in $G^{2}$. We will look into what other nodes can have the decreasing sequence property, and what properties must exist in the graph in order for this to take place.
Example 5.2.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Figure 5: This is an illustration where node $v_{0}$ has the Decreasing Sequence Property (DSP). There are more first neighbors $d$ of $v_{0}$ than second neighbors.
Definition 5.2.
For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$,where $N^{+}(x)$ denotes the set of out-neighbors of $x$, and $N^{++}(x)$ denotes the set of out-neighbors of the nodes in $N^{+}(x)$.
In previous sections, we discussed two fundamental concepts: the minimum degree node and the decreasing sequence property (DSP). A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point for connectivity and flow within the graph. As the graph’s structural anchor, the minimum degree node drives the development of relationships and properties that will emerge in the counterexample construction. It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
In contrast, the DSP introduces constraints on connectivity by systematically limiting potential connections beyond nodes that exhibit this property. This concept is akin to a bottleneck: even nodes with multiple connections may still encounter restricted reachability in the graph. A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP.
Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively. They jointly reveal scenarios where expected connectivity may falter, highlighting vulnerabilities that could challenge established conjectures. The DSP, in particular, illustrates how exterior nodes in an oriented graph may systematically decrease in number, creating a structure where connectivity flow is constrained despite high out-degrees of certain nodes.
By examining the DSP, we underscore a critical aspect of connectivity, resembling bottlenecks in network theory. Diestel’s Graph Theory [6] addresses the relationship between node degree and connectivity, demonstrating that a node $x$ can maintain many outgoing neighbors while still limiting reachability. This discrepancy signals inefficiencies and sets the stage for potential counterexamples to existing conjectures, thus motivating our investigation of graph properties and their broader implications.
Lemma 5.1.
Total Order Induced by Containers Relative to the Minimum Degree Node
Let $G$ be a graph with a minimum degree node $v_{0}$, where all nodes satisfy the Decreasing Sequence Property (DSP). Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ relative to $v_{0}$, where each container $C_{i}$ consists of nodes ordered by their adjacency and connectivity properties with respect to $v_{0}$ and preceding containers. This structure imposes a total order on the nodes in $G$, where every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$, establishing a ranking aligned with the DSP.
Proof.
To prove a total order, we need to show that every node in $G$ can be compared to every other node in $G$. Let $u,v$ be two nodes in $G$ such that $u\neq v$. It is assumed that $dist(v_{0},u)$ and $dist(v_{0},v)$ are finite for vertices $u$ and $v$ in the containers $C_{i}$. This assumption implies that both vertices are reachable from $v_{0}$, although it does not necessitate that all vertices are connected. We know that distance is a total ordering, so we know that one of three things are true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$. The situation where $dist(v_{0},u)<dist(v_{0},v)$, implies that the node $u$ is in an closer container to $v_{0}$ than the node $v$. The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container. And the situation $dist(v_{0},u)>dist(v_{0},v)$, implies that the node $u$ is in a further container to $v_{0}$. This shows that any two nodes can be compared and shows that it is a total order. ∎
This is a good start as it imposes order on our graph, but the containers do more than just impose a total order. They essentially allow for grouped comparisons among nodes at the same distance from a reference node, $v_{0}$. This is something that a standard total order does not account for. This ”containerized ordering” enables us to distinguish nodes not only by their distance from $v_{0}$ but also by specific structural or adjacency properties within each distance group.
Unlike a simple total order, which could be any sequence with a single, consistent ordering (like lexicographic ordering), the container structure allows nodes at the same distance to be further organized. This organization is significant because it facilitates comparisons within each distance layer (container) by additional criteria, such as adjacency or neighborhood properties. This layered organization is especially useful in proofs, where we might need to argue about the relative positions of nodes within the same distance from $v_{0}$.
This structure allows us to establish a refined form of comparability among nodes at equal distance that is crucial for analyzing specific properties. For example, in certain proofs, we might need to show that nodes within the same container have certain relationships that would not necessarily follow from a mere total order.
Definition 5.3.
A Level Total Order (LTO) on a set $S$ can be defined as follows:
  • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
    $u$ is less than $v$ (denoted as $u<v$), $u$ is greater than $v$ (denoted as $u>v$), or $u$ is equal to $v$ (denoted as $u=v$).
  • Leveled Structure: The elements of $S$ are partitioned into a leveling of ”containers” or ”levels,” $C_{1},C_{2},,,,,C_{n}$, with each container itself forming a subset of $S$ that maintains a total order. The containers are sequentially ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
  • Comparability within Containers: For any two elements $u,v\in Ci$, their order can be determined based on a specific criterion, such as a designated metric or property relevant to the application.
  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
A leveled total order not only provides an ordering across distance groups but also enables detailed, structure-based comparisons within each group.
By defining both inter- and intra-container comparisons, this ordering allows arguments to leverage container-level properties (like distance) and within-container properties (like adjacency) simultaneously. This flexibility is particularly useful in graph theoretic proofs like the SSNC, where it may be necessary to combine distance-based and structural properties.
Defining a leveled total order gives rigor to proofs, as it offers a structured, multi-layered way to approach ordered sets that are partitioned by one property but require further comparability within each partition.
This formalization could also have broader applications beyond the SSNC. It might apply to any problem where elements need to be organized hierarchically but also require nuanced relationships within each level of the hierarchy—such as layered networks, hierarchical clustering, and complex sorting problems.
Lemma 5.2.
Decreasing Sequence Property Lemma If a node in an oriented graph $G$ does not have the decreasing sequence property, then its out-degree at least doubles in $G^{2}$.
If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$, and those that are not neighbors of $x$. In the proofs of Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0), and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor induced subgraph (of $v_{0}$) and in the graph with the neighbor induced subgraph removed. We will now give terms to these concepts.
Definition 5.4.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
$int(x,y)=N^{+}(x)\cap N^{+}(y)$
,where $N^{+}(z)$ is the set of out-neighbors of $z$. The number of interior neighbors is the interior degree of $y$ with respect to $x$.
We previously introduced the concept of containers, which are crucial for understanding the structure of the graph. Interior neighbors are specifically those nodes for which both endpoints are located within the same container. By examining these interior neighbors, we gain insights into how containers relate to one another and how they function internally.
Definition 5.5.
Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
$ext(x,y)=N^{+}(y)\cap N^{++}(x)$
, where $N^{+}(z)$ is the set of out-neighbors of $z$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers. This interaction is critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container. However, it is essential to consider other possibilities that may also influence this relationship.
To better understand the distinction between interior and exterior neighbors, consider Example 5.2. In this example, arcs that connect nodes within the same container (or box) represent interior neighbors. In contrast, arcs that extend to nodes in the next container illustrate exterior neighbors.
Definition 5.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
Containers are ordered based on their distance from the chosen minimum degree node. Up to this point, all arcs have either been interior or forward-facing. However, it is important to consider the possibility that some arcs may also direct back toward the chosen minimum degree node. Addressing this potential will help us understand how these back arcs could impact the Seymour Second Neighborhood Conjecture.
Now that we have established the definitions of interior and exterior nodes, we can delve into the decreasing sequence property. This property refers specifically to the sizes of the exterior sets within the oriented graph. As the DSP path algorithm progresses, the exterior nodes — which are not part of the neighborhood of the minimum degree node — will systematically decrease in number. This phenomenon reflects a yin and yang relationship between the interior and exterior nodes: while the interior nodes maintain and potentially increase their interconnections, the exterior nodes steadily diminish, ensuring a balanced growth within the graph.
In this dynamic, the minimum degree node serves as a catalyst for this process, driving the algorithm forward by imposing constraints and guiding the restructuring of connections. In contrast, the decreasing size of the exterior nodes acts as a bottleneck, effectively limiting the expansion of the graph’s out-degrees. This interplay is crucial for the success of the DSP path algorithm, as it maintains the structural integrity of the graph while facilitating necessary transformations.
This relationship between the interior and exterior nodes will be the focal point of our proof throughout the rest of this paper, providing a framework for understanding how the DSP path algorithm operates effectively.
With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the
Definition 5.7.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Leveled Structure: The elements of $S$ are partitioned into levels of ”containers” or ”rankings,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Theorem 5.1.
Given an oriented graph $G$ and a minimum degree node $v_{0}$, where the nodes of $G$ are partitioned into containers $C_{1},C_{2},\dots,C_{k}$ based on their distance from $v_{0}$ and assuming no back arcs, and the interior arcs in each container and exterior arcs between each container forms a Graph Level Order (GLOVER) under the distance total order.
Proof.
To establish that the structure $\{C_{1},C_{2},\dots,C_{k}\}$ satisfies the GLOVER properties, we verify each condition in the GLOVER definition:
  • Total Order on Elements within $G$: - By partitioning nodes based on their distance from $v_{0}$, we establish a total order on the elements of $G$: for any two nodes $u$ and $v$, their order is determined by their respective containers. Nodes in closer containers (e.g., $C_{i}$) are considered ”less than” nodes in further containers (e.g., $C_{j}$ with $i<j$).
    - Within each container, nodes may be ordered based on additional criteria (such as their individual distances to $v_{0}$ if there are multiple paths). Thus, we achieve a total order among all nodes in $G$.
  • Leveled Structure with Containers: - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$. The elements within each container are in a total order relative to each other, and the containers themselves are ordered by distance from $v_{0}$, establishing a multi-level ranked structure.
  • Non-triviality Requirement: - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node. This fulfills the non-triviality requirement, as we avoid having all singleton containers.
  • Comparability within Containers: - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$ (for example, their shortest-path distance if there are multiple paths). Thus, comparability is ensured within containers.
  • Consistency of Comparisons Across Containers: - Since the graph is oriented and there are no back arcs (i.e., no edges from nodes in $C_{j}$ to nodes in $C_{i}$ for $j>i$), the ordering between containers is respected. This guarantees that if $C_{i}$ is ”less than” $C_{j}$, then every element in $C_{i}$ is ”less than” every element in $C_{j}$, ensuring consistency across containers.
  • Interior and Exterior Neighbors: Nodes $u$ and $v$ of the same distance from the chosen minimum degree node $v_{0}$ would be in the same container and thus be interior neighbors. If these nodes were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance different by 1.
    Each criterion for an GLOVER is satisfied by the structure defined by partitioning nodes based on distance from $v_{0}$. Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a Ranked Total Order under the distance total order, completing the proof.
This lemma aligns with the ranked structure we’re building, where each container $C_{i}$ is distinct in terms of reachability and influence. By enforcing this strict order under the absence of back arcs, we’re creating a clear layering within the graph, which further justifies the ”multi-level total order” approach.
Back arcs, if present, could disrupt the unidirectional progression implied by containers, potentially creating feedback loops. By assuming no back arcs, we ensure that each container maintains its place in the total ordering, which adds robustness to the ranked structure we’re developing.
Containers, as defined in the GLOVER framework, serve not only as organizational units but also as bottlenecks that regulate the flow of connectivity within the graph. By grouping nodes based on their shortest-path distance from the minimum degree node $v_{0}$, containers segment the graph into hierarchical levels. The decreasing sequence property ensures that the size of exterior neighborhoods diminishes as we move outward from $v_{0}$, creating a natural bottleneck effect. This reduction in connectivity simplifies the analysis of degree growth and neighbor interactions while preserving the structural integrity of the graph. Such bottlenecks are critical for maintaining the layered structure of the GLOVER framework, particularly in problems like the Seymour Second Neighborhood Conjecture, where second-level connectivity is a key focus.
Next, we will show the power of this data structure and why is introduced in this paper.
To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
Lemma 5.3.
GLOVER Supports Induction: In any oriented graph, the minimum degree node $v_{0}$ within a Glover structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
Proof.
Let $v_{0}$ be the node with the minimum out-degree in the oriented graph $G$. - By definition, for any node $v\in G$, $d^{+}(v_{0})\leq d^{+}(v)$, where $d^{+}(v)$ is the out-degree of node $v$. - This ordering provides a baseline for comparison because all other nodes are either equal to or have greater out-degree than $v_{0}$.
The analysis begins at $v_{0}$, which serves as the foundation of the GLOVER structure. - All outgoing arcs from $v_{0}$ lead to nodes at the next level (or ”container”) $C_{1}$. - By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
Assume that for any node $v$ in a container $C_{i}$, the following properties hold: - $d^{+}(v)\geq d^{+}(v_{0})$. - All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$. - By the inductive hypothesis, $v$ satisfies all conditions derived from $v_{0}$. - Since $u$ receives its inbound arcs from $v$, and $d^{+}(v)\geq d^{+}(v_{0})$, the relationships in $C_{i+1}$ are consistent with the properties of the Glover structure. - Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that: - The out-degree of any node is anchored to $v_{0}$. - There is no upward deviation in degree that contradicts the structure’s constraints.
Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph.
Example 5.3.
[…]
Figure 6: This example illustrates back arcs in a graph. The two blue arcs from $x3$ to $y3$ and $y3$ to $z3$ represent exterior neighbors indicate forward directing arcs. Additionally, a red arc points backwards from $z5$ to $x1$. Although the containers are not outlined, they are implied by the difference in variables for each group.
Lemma 5.4.
Interior Degree Greater than 1 In any oriented graph $G$, represented by a GLOVER with $v_{0}$ as a minimum out-degree node satisfying the decreasing sequence property, In an attempt to prevent a degree doubling node, every node $x$ in the neighboring container $N^{+}(v_{0})$ has an interior degree of at least 1. That is, for $x\in N^{+}(v_{0})$,
$|\text{int}(v_{0},x)|\geq 1.$
Proof.
By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$
where $N^{+}(x)$ denotes the set of first neighbors (out-neighbors) of $x$, and $N^{++}(x)$ denotes the set of second neighbors (out-neighbors of $N^{+}(x)$).
For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $y$, and then from $y$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $y\in N^{+}(v_{0})$.
If $y$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $y$’s neighbors, which are $v_{0}$’s second neighbors. If $y$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $y$’s first neighbors.
To avoid this doubling, $y$ must connect to at least one interior arc. This means that the interior degree of $y$ with respect to $v_{0}$ must satisfy
$|int(v_{0},y)|\geq 1$
This argument holds for any neighbor $y\in N^{+}(v_{0})$, ensuring that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
Thus, for all $x\in N^{+}(v_{0})$,$|int(v_{0},x)|\geq 1$, completing the proof. ∎
Lemma 5.4 (Interior Degree Greater than 1) is important because it constrains the out-neighbor structure of the graph when the minimum out-degree $v_{0}$ exhibits the decreasing sequence property. It guarantees that each out-neighbor of $v_{0}$ must have at least one other out-neighbor within $N^{+}(v_{0})$ , which helps to prevent the out-degree of nodes from being too heavily concentrated in the second neighborhood.
The GLOVER structure partitions the neighbors of nodes into interior and exterior categories in a systematic way. This partitioning provides a powerful framework for analyzing oriented graphs, particularly in controlling and understanding degree properties.
A notable consequence of this structured representation is the ability to guarantee that for any oriented graph $G$ represented as a GLOVER, the degree of a chosen minimum out-degree node $v_{0}$ will not double due to cycles in container $C_{1}$. More formally:
Proposition 5.1.
Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node, and if there exists a cycle in the first container $C_{1}$, then the degree of $v_{0}$ will not double.
Example 5.4.
0
$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
Lemma 5.5.
Exterior Less than Degree In any oriented graph $G$, represented by a GLOVER In an attempt to prevent a degree doubling node, if a node $x_{i}$ is in container $C_{i}$ has the decreasing sequence property, In an attempt to prevent a degree doubling node, then for all $y\in N^{+}(x)\subseteq C_{i+1}$, $|ext(x,y)|<d^{+}(x)$.
Proof.
Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|ext(x,y)|\geq d^{+}(x)$. By definition, $ext(x,y)$ consists of the set of out-neighbors of $y$, specifically:
$ext(x,y)=\{z\in N^{++}(x)\mid z\text{ is an out-neighbor of }y\}.$
Thus, $|ext(x,y)|\subseteq N^{++}(x)$, where $N^{++}(x)$ denotes the set of second out-neighbors of $x$. If $|ext(x,y)|\geq d^{+}(x)$, then:
$|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x),$
which suggests that the out-degree of $x$ would at least double in the graph $G^{2}$ (the second power of the graph, where edges represent paths of length 2). This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property, where the degree of nodes cannot increase as we move through the containers.
Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|ext(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
This proof directly shows that a contradiction arises if the exterior degree exceeds the out-degree, which would violate the decreasing sequence property. The argument about the doubling of the out-degree in $G^{2}$ is key in demonstrating that the structure defined by the GLOVER representation prevents such a situation, thereby ensuring that $|ext(x,y)|<d^{+}(x)$.
Lemma 5.6.
[…]
Then, for all $z\in ext(x,y)$, we have $|ext(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
Proof.
We begin by recalling the conditions for a node to have the decreasing sequence property (DSP). A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
Now, assume the conditions of the lemma hold: 1. The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increases significantly in $G[N^{+}(x)]$. 2. For all $z\in ext(x,y)$, $ext(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
Now, suppose for contradiction that there exists $z\in ext(x,y)$ such that $|ext(y,z)|\geq|ext(x,y)|$. This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$. However, this contradicts the assumptions about the DSP of $y$, which requires $y$ to have fewer second neighbors than first neighbors.
From $|int(x,y)|$ doubling, $y$’s influence in $G[N^{+}(x)]$ grows, yet its exterior degree should decrease due to the DSP. The absence of back arcs prevents any increase in $y$’s exterior degree from its second neighbors.
Thus, the assumption $|ext(y,z)|\ \geq|\text{ext}(x,y)|$ leads to a contradiction,
confirming that $|\text{ext}(y,z)|<|\text{ext}(x,y)|$ for all $z\in\text{ext}(x,y)$.
Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
The central concept of this proof revolves around the relationship between decreasing exteriors and back arcs. This relationship is not merely general; it specifically addresses the concern of potential double counting involving second neighbors that arise from back arcs through $y$.
The additional condition - ensuring that a neighbor of $y$ does not also serve as a first neighbor of $x$ is vital. This condition helps to prevent the overlap in counting, particularly in situations where $z$, a second neighbor of $y$, could inadvertently be counted as a first neighbor of $x$.
Given that containers are structured by their distance from the chosen minimum degree node, it follows that if a node $z$ is a second neighbor of $y$ and also a first neighbor of $x$, this necessitates the presence of a back arc from $z$ to a neighbor of $x$. Therefore, the lemma’s validity heavily relies on the configurations of back arcs in our construction, particularly in the context of a potential counterexample.
In summary, the interplay between back arcs and the classification of neighbors is crucial in Lemma 5.6 (Weak Decreasing Exteriors Lemma). By maintaining the distinction between first and second neighbors through careful conditions, we can mitigate the risk of double counting and strengthen the argument presented in the proof.
Example 5.5.
0
$\Rightarrow$ 111122223333444455556666 $\Rightarrow$ 7777888899991010101011111111 $\Rightarrow$ H
Figure 7: Nodes in each container connect to all nodes in the subsequent container. Here, the minimum out-degree node has out-degree 6, with connectivity patterns shown to demonstrate degree doubling in $G^{2}$.
Table 1: Intersection Results for Triplets of Nodes in 5.5
Triplet(A, B, C) int(A, B) ext(B, C) Intersection Result
[…]
The theoretical foundation established in this section underpins the subsequent arguments and proofs. By thoroughly examining the properties of oriented graphs, we prepare to tackle the complexities of back arcs and their implications for the overall graph structure.
6 Back Arcs
In this section, we analyze the role of back arcs in oriented graphs and their effects on graph properties. Understanding back arcs is crucial for assessing how they influence the interior and exterior degrees, as well as the overall structure of the graph.
It is important to restate that back arcs and exterior neighbors are distinct concepts. Back arcs are not a subset of exterior neighbors. While back arcs connect to any previous container, exterior neighbors are limited to connecting with the next container. Additionally, exterior neighbors do not represent all possible external neighbors. Rather, an exterior neighbor ext(x, y) specifically refers to nodes at a distance of two from $x$ that are also neighbors of $y$.
If we look at Example 4.2 again we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$. Similarly we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$. In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double. This next lemma formalizes that concept.
Example 6.1.
[…]
Figure 8: Graph representation of back arcs (shown in red) and forward-directed arcs (in blue) between containers, illustrating how these arcs impact connectivity patterns in an oriented graph.
Lemma 6.1.
(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with minimum degree node $v_{0}$ partitioned by containers $C_{i}$. Suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles. If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
Proof.
Assume $u$ and $v$ are two nodes in container $C_{i}$, both with a common parent $v_{p}$ in container $C_{i-1}$, and let $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$, where $\text{int}(v_{p},x)$ denotes the interior degree of $x$ relative to $v_{p}$’s container $C_{i-1}$.
By the Decreasing Sequence Property (DSP), if the interior degree of $v$ is higher than $u$, then $v$ will require fewer exterior connections to maintain the required degree restrictions in the graph. Meanwhile, $u$, with fewer interior connections, will require more exterior connections to compensate.
Now, the imbalance in interior degrees means that $u$’s exterior connections are effectively increasing the degree of $v_{p}$. This is because $u$ must connect to more external nodes to maintain its degree, leading to an increase in the number of exterior neighbors of $v_{p}$.
Consequently, the exterior degree of $v_{p}$ must double to accommodate the increased number of connections needed by $u$ and $v$ in the graph, violating the degree constraints intended by the partitioning into containers.
Thus, the differing interior degrees of $u$ and $v$ in container $C_{i}$, combined with the effects of the DSP, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double. This leads to a contradiction, so the claim holds. ∎
Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples. Specifically, it shows that all the out-neighbors of interior arcs within a container must have the same degree, which is known as out-degree regularity. If, additionally, the nodes in this container have equal in-degrees, then the container achieves total regularity. It is important to note that this regularity emerges as a counterexample strictly due to the interaction between the Decreasing Sequence Property (DSP) and the minimum degree node $v_{0}$.
In Example 5.5 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph. The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth. This outward growth reflects the structure of the graph, where nodes are grouped into containers based on their distance from $v_{0}$.
A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$, effectively pointing against the distance-based order. If such a back arc were introduced, it would imply that the source node $y$ gains new neighbors at a lower level, potentially doubling its degree. This contradicts the hierarchical organization of containers, where each level is strictly defined by shortest-path distances from $v_{0}$.
Challenge:
To better understand why back arcs are not possible in this graph, consider the following questions:
  • Can you construct a back arc in this graph that respects the container construction rules and distance-based ordering from $v_{0}$?
  • If a back arc is introduced, does it lead to contradictions or disrupt the hierarchical structure of containers?
Explanation:
By the construction of the GLOVER framework, all nodes in $C_{i}$ are farther from $v_{0}$ than those in $C_{j}$ for $j<i$. A back arc $(y,z)$ would imply that $z$ is closer to $v_{0}$ than $y$, violating this ordering.
Additionally, introducing a back arc disrupts the unidirectional flow of the graph by creating a feedback loop. This contradicts the fundamental property of containers: that each level represents a discrete, non-overlapping set of nodes ordered by distance from $v_{0}$. Without back arcs, the GLOVER structure preserves a clear layering, enabling rigorous analysis of degree growth and neighbor interactions.
Lemma 6.2.
No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $ext(u_{i},v_{k})$ is at least the degree of $v_{k}$. Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
Proof.
Assume that $u_{i}\in C_{i}$ has its first back arc to $v_{k}\in C_{k}$ where $k<i-1$. By definition of exterior neighbors:
$|ext(u_{i},v_{k})|=|N^{+}(v_{k})\cap N^{++}(u_{i})|$
where $N^{+}(v_{k})$ is the set of immediate neighbors of $v_{k}$, and $N^{++}(u_{i})$ is the set of nodes at distance two from $u_{i}$.
Since $N^{+}(v_{k})$ includes all direct neighbors of $v_{k}$, it follows that:
$|N^{+}(v_{k})|=deg(v_{k}).$
By the structure of GLOVER and the absence of back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, the size of $ext(u_{i},v_{k})$ satisfies:
$|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
Now, consider the relationship between interior and total degree. If the interior degree of $u_{i}$ doubles, the total degree of $u_{i}$ must also double. This is because the exterior neighbors are tightly constrained by the back arc and are not impacted by the increase in interior connections. Thus, any increase in the interior degree directly translates into a proportional increase in the total degree.
Finally, by the Weak Decreasing Exteriors Lemma, the size of $ext(u_{i},v_{k})$ decreases as we move further in the sequence of containers. This ensures that the total degree of $u_{i}$ remains bounded and behaves as expected under the given conditions, completing the proof. ∎
Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container structured graph. It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container.
Lemma 6.3.
First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph represented by a GLOVER with minimum degree $d$ in an attempt to prevent a degree doubling node, and let $v_{0}$ be a node of minimum degree. Suppose $G$ is an oriented graph, where arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$. If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will violate the structure of the graph.
Proof.
Assume that the oriented graph $G$ is partitioned into containers $C_{1},C_{2},\dots,C_{i}$ such that nodes in each container $C_{j}$ (for $j<i$) only have either arcs to fellow nodes in $C_{j}$, or they map forward to nodes in $C_{j+1}$, with no back arcs present between containers up to $C_{i}$.
Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$. In this oriented graph, by definition, all arcs are supposed to point in the forward direction—from container $C_{j}$ to container $C_{j+1}$. By Lemma 6.1 (Regular Interior Degrees), we know that all nodes inside container $C_{i-1}$ have the same interior degree. Because of their matching interior degrees, they will also share the same exterior degrees. Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes. If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist{v_{0},u_{i}}-i$ so $u_{i}$ cannot belong to container $C_{i}$.
Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ leads to a fundamental contradiction in the oriented nature of the graph, and no back arcs can exist prior to $C_{i}$. ∎
The concept of a first back arc inherently assumes that no back arcs exist up to a certain point. Lemma 6.2 (No Back Arc to Early Container) supports this assumption, relying on the doubling of nodes’ interior degrees, a condition not necessary for Lemma 6.3. The strength of this lemma lies in its capacity to identify the first back arc; if such an arc cannot be established, it reinforces the graph’s structural integrity. This lemma is founded on our ability to well-order the graph using containers and the Decreasing Sequence Property (DSP). With this ordering, we highlighted regularity within these containers through Lemma 6.1 (Regular Interior Degrees), which proved essential in showing that these back arcs would lead to a node’s degree doubling if a node’s interior degree was already doubling.
Lemma 6.4.
No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$,$C_{2}$, $\dots$,$C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
Proof.
This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.3 (First Back Arc Between Consecutive Containers) . ∎
[…]
Figure 9: Illustration of the proof of 7.2 (Interior Degrees Double). Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
[…]
The exploration of back arcs reveals their significant impact on the dynamics of oriented graphs. By removing them, we can better navigate the structures of oriented graphs and in particular it paves the way to investigate interior degree dynamics.
7 Interior Degrees Doubling
This section focuses on the important aspect of interior degree doubling within oriented graphs. We investigate the conditions under which this occurs and the implications for the overall degree distribution of nodes.
Lemma 7.1.
Indistinguishable Regular Interior Nodes In a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER in an attempt to prevent a degree doubling node. If a node $v_{p}$’s interior degree doubles then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
Proof.
Let $C_{i}$ be a regular container of size $n$, and let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$. The interior degree of a node $v_{i}$ in container is defined as the number of arcs shared $v_{i}$ between $v_{i}$ and a neighbor or parent node. In a regular container, all nodes are indistinguishable in terms of their degree. They also have the same parents because of the DSP. Any automorhism we construct including a random automorphism $\phi$ of the container $C_{i}$ will preserve the degree structures. This means any node will have its interior degree double. ∎
The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others. Initially, we accounted for nodes where interior degrees had doubled and ensured consistency when exterior degrees doubled as well. This distinction was critical for our algorithm up to this point. However, with regular containers, interior nodes are now indistinguishable from one another. This means we can swap a node $v_{i}$ with an exterior node $v_{e}$ if their degrees have doubled, ensuring that we have a node whose overall degree has doubled. This flexibility simplifies the process and ensures that the algorithm maintains consistency.
Lemma 7.2.
Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP). Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
Proof.
We prove this by strong induction on the size of the graph.
Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3 there exists at least one node in each container with a doubled interior degree. These cases form our base, supporting the induction.
Inductive Hypothesis: Assume that for any graph of size $n^{\prime}<n$, each container $C_{i}$ contains at least one node whose interior degree doubles, thus satisfying the DSP by proportionally reducing exterior degrees.
Inductive Step: Consider a graph of size $n$ where we need to show that at least one node in each container $C_{i}$ has its interior degree doubled.
Suppose we construct $G$ such that no node at distance $i$ from $v_{0}$ initially has a doubled interior degree in $C_{i}$. Define a sequence of subgraphs $H_{0}$,$H_{1}$,…,$H_{k}$ for $C_{i}$ as follows:
$H_{0}=\{u_{i}\}$
$H_{1}=N^{+}(u_{i})$
with nodes in $H_{1}$ having doubled interior degrees by the I.H., as applied to each subgraph. By the Induction Hypothesis (I.H.),there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$. We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}inH_{2}$ whose interior degree doubles in $H_{2}$. We can define $H_{3}=ext(y_{1},y_{2})$. Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
By the Weak Exterior Size Decreasing lemma, we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers. By the DSP, an increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$, preventing exterior degrees from exceeding interior degrees. This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$, as required by DSP.
The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$. At this point, the remaining nodes in $H_{i}$ either have no further exterior neighbors or their degrees lie entirely within the interior set.
Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$. Thus, the parent of $y^{*}$, denoted $y^{*}-1$, in $G$ achieves a doubled degree within $C_{i}$.
This doubling process terminates as DSP constraints prevent exterior degrees from surpassing interior degrees without contradicting the DSP.
Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
Considering Lemma 7.2 (Interior Degree Double), we observe that it represents a specific instance of the Seymour Second Neighborhood Conjecture. Here, each container functions as its own self-contained oriented subgraph, where we pose the same core question: is there a node whose degree doubles? The affirmative answer we obtain, even when constructing a counterexample, suggests promising progress toward resolving the overall conjecture.
Furthermore, we can expand upon the process illustrated in Lemma 7.2 (Interior Degree Double)). Within the proof, we constructed a monotonically decreasing sequence of subsets, which cannot continue indefinitely. This sequence reflects the essence of the Decreasing Sequence Property (DSP), where we observe that as interior nodes increase, the exterior nodes necessarily decrease. This requirement is what we meant by “decreasing” in DSP: the exterior set gradually diminishes as nodes transition to the interior.
This dynamic can be interpreted as a form of equilibrium, reminiscent of the principle of yin and yang. Each node integrated into the interior necessitates a corresponding reduction in the exterior set, thereby preserving a balance. This interplay ensures that interior growth does not occur in isolation; instead, it aligns directly with constraints on the exterior, maintaining the graph’s structural integrity.
Such interaction is essential for upholding the properties defined by DSP. It underscores the inherent relationship between interior and exterior neighbors, reinforcing the notion that the expansion of one aspect imposes natural limitations on the other.
In summary, examining the doubling of interior degrees within this context provides critical insights into the degree distribution in oriented graphs. This understanding informs the proofs that will be discussed in subsequent sections, enhancing our grasp of both local and global properties of the graph.
8 Container Density
Here, we delve into the concept of container density and its relevance to the properties of oriented graphs. We explore how regular exterior degrees interact with the overall density of containers and the implications for graph structure.
Having proved that the interior degrees double within a container and that is linked to container size decreasing, we are ready to tackle the overall Seymour Second Neighborhood Conjecture. First we need to establish some fundamental truths about containers and oriented graphs in these counterexamples to set up the rest of the paper and things that must be true.
Lemma 8.1.
Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node. Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP). Then the regular interior degrees will result in regular exterior degrees in $G$.
Proof.
Let $x_{i}$ and $x_{j}$ be two nodes at the same distance $i$ from the minimum degree node $v_{0}$ in the oriented graph $G$, both contained in the same container $C_{i}$.
Assume, for contradiction, that the sizes of their exterior sets differ by at least one. Without loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|=|ext(x_{j},y)|+1$. This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
Since the size of the next container $C_{i+1}$ is determined by the exterior degrees of nodes in $C_{i}$, this irregularity in the exterior degrees would force the size of $C_{i+1}$ to differ, contradicting the assumption that the containers are of regular size.
By Lemmas 6.1 (Regular Interior Degrees) and 7.2 (Interior Degree Double), we know that if the interior degrees within a container are regular, the exterior degrees must also follow this regularity. Specifically, if there is an irregularity, any parent node of $x_{i}$ and $x_{j}$ will have its exterior degree double, which would in turn double their total degree.
This doubling of the parents’ total degree would violate the Decreasing Sequence Property (DSP).
Therefore, the assumption that $|ext(x_{i},y)|$ and $|ext(x_{j},y)|$ differ is false. Hence, the exterior degrees in $G$ must be regular if the interior degrees are regular. ∎
What we saw in Lemma 6.1 (Regular Interior Degrees) that the regularity of interior degrees was something we were able to prove while we were trying to prove that back arcs did not exist. This both made sense and was very helpful in that environment. We see here that the regularity of interior degrees is not restricted to the container itself, but also impacts the exterior arcs to the next container - forcing them to be regular as well.
Lemma 8.2.
Regular Graph All Nodes Have Equal Our-Degree Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with $v_{0}$ as a minimum degree node. Further, let all nodes satisfy the Decreasing Sequence Property (DSP). Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $d_{v_{0}}=d^{+}(v_{0})$.
Proof.
We will prove this lemma by induction on the number of nodes in the graph $G$.
For the base case, consider the node $v_{0}$, which, by definition, has the minimum degree $d$, so $d^{+}(v_{0})=d_{v_{0}}$.
Assume that for some node $v_{k}$ in $G$, the degree of $v_{k}$ satisfies $d^{+}(v_{k})=d_{v_{0}}$.
Now consider an arbitrary node $v_{k+1}$ in $G$. Suppose, for contradiction, that $d^{+}(v_{k+1})>d_{v_{0}}$.
Consider a neighbor $v_{n}$ of $v_{k+1}$. By the inductive hypothesis, the degree of $v_{n}$ is exactly $d_{v_{0}}$. Since $v_{k+1}$ has more than $d_{v_{0}}$ neighbors, the neighbors of $v_{k+1}$ are second neighbors of $v_{n}$ in the graph $G^{2}$. Therefore, $v_{n}$ would have more than $d_{v_{0}}$ second neighbors in $G^{2}$.
This implies that $v_{n}$’s degree has effectively doubled, which contradicts the DSP, which requires that $v_{n}$ must have degree $d_{v_{0}}$. Hence, the assumption that $d^{+}(v_{k+1})>d_{v_{0}}$ leads to a contradiction.
Thus, we must conclude that:
$d^{+}(v_{k+1})=d_{v_{0}}.$
By the principle of mathematical induction, it follows that for every node $v$ in $G$, the degree of $v$ is equal to $d_{v_{0}}$. ∎
We are incrementally refining the properties that must hold for our counterexamples. The two most crucial aspects are the Decreasing Sequence Property and the identification of the minimum degree node. However, additional constraints, such as the overall graph regularity established by Lemma 8.2 (Regular Graph), help simplify future arguments. This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
Lemma 8.3.
Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER In an attempt to prevent a degree doubling node, with no back arcs, and let $v_{0}$ be the chosen minimum degree node in $G$. For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
for some $z\in N^{+}(x_{i})$ in order to keep any node’s degree from doubling in the oriented graph.
Proof.
We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
$\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
is bounded by $d^{+}(v_{0})-1$, since each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2, as there are no back arcs to $v_{0}$.
Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
Inductive Step: We need to show that this bound holds for nodes at distance $i+1$ from $v_{0}$. Let $x_{i+1}$ be a node in container $C_{i+1}$ relative to $v_{0}$. Suppose, for contradiction, that
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|>d^{+}(v_{0})-i.$
This would imply that the exterior neighbors of $x_{i+1}$ exceed the allowed bound, violating the decreasing sequence property (DSP) conditions of $x_{i}$. By Lemmas 7.2 (Interior Degrees Double), 6.1(Regular Interior Degree), and 7.1 (Indistinguishable Regular Interior Nodes), this would cause the interior degree of $x_{i}$ to double, thereby forcing an overall doubling of the degree for $x_{i}$.
This contradiction implies that our assumption must be false. Therefore, we conclude:
|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
This lemma is a more extended form of Lemma 5.6 (Weak Decreasing Exteriors Lemma), but without the added dependencies. It gives us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named. Several examples and observations brought this to a conjecture, which brought it here. The question now becomes is there any way we can strengthen this result?
We demonstrated that the Decreasing Exteriors Lemma relies on two key factors: the elimination of back arcs and the doubling of interior degrees within each container. Furthermore, we established that the elimination of back arcs is directly dependent on the interior degrees doubling within those containers.
By proving the previous lemma, we not only confirm the doubling of interior degrees but also conclude two additional results: (1) back arcs are entirely eliminated, and (2) the exterior arcs exhibit the desired decreasing pattern.
Lemma 8.4.
Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER, In an attempt to prevent a degree doubling node, and let $v_{0}$ be a minimum degree node in $G$, where every node in $G$ has the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$. Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
Proof.
Let $x$ be a node at distance $i$ from $v_{0}$. By Lemma 8.3 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$, $|ext(v_{i-1},x)|$, is bounded by $d^{+}(v_{0})-i$. This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
Now, since $x$ has at least the same out-degree as $v_{0}$ (by the assumption of DSP), $x$ must map to at least $i$ other nodes. Given that back arcs are not allowed (as shown previously), $x$ cannot map to nodes at earlier distances without violating the structure of the graph.
Therefore, the only remaining possibility is that $x$ maps to interior nodes. Since $x$’s degree must account for at least $i$ mappings, and these must be to interior nodes, we conclude that $x$ must map to at least $i$ interior nodes.
Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
Proof.
Let $G$ be an oriented graph represented by a GLOVER. Let $v_{0}$ be a node with the minimum degree in $G$, and assume that every node in $G$ satisfies the Decreasing Sequence Property (DSP). Let $x$ be a node at distance $i$ from $v_{0}$, and let $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$.
Our goal is to show that the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided that $(v_{0},x)\in G$ (i.e., $x$ is reachable from $v_{0}$).
By Lemma 8.3 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$ with respect to $x$, $|ext(v_{i-1},x)|$, is bounded above by:
$|ext(v_{i-1},x)|\leq d^{+}(v_{0})-i$
where $d^{+}(v_{0})$ is the out-degree of $v_{0}$. This gives a limit on the number of nodes that $x$ can connect to at distance $i+1$ from $v_{0}$.
Since the graph satisfies the Decreasing Sequence Property (DSP), we know that all nodes in the graph have a degree at least as large as $v_{0}$’s degree. Therefore, the degree of node $x$ satisfies:
$d(x)\geq d(v_{0})$
This means that node $x$ must have at least $i$ neighbors at distance $i+1$ from $v_{0}$.
The structure of the graph explicitly disallows back arcs, i.e., edges pointing to nodes at earlier distances from $v_{0}$. Thus, $x$ cannot connect to nodes at distances less than $i$ without violating the structure of the graph. This constraint eliminates the possibility of $x$ mapping to earlier-distance nodes.
Given that $x$ must connect to at least $i$ nodes, and since it cannot connect to earlier-distance nodes (due to the disallowance of back arcs), the only remaining option for $x$ is to map to interior nodes—nodes that are at distance $i$ or higher.
Thus, since $x$ must map to at least $i$ distinct nodes, and all these must be interior nodes (by the elimination of back arcs and exterior connections), we conclude that:
$|int(v_{i-1},x)|\geq i$
This shows that the interior degree of $x$ with respect to $v_{i-1}$ is at least $i$.
Therefore, we have shown that for a node $x$ at distance $i$ from $v_{0}$, the interior degree of $x$ with respect to $v_{i-1}$, $|int(v_{i-1},x)|$, is at least $i$, as required.
$|int(v_{i-1},x)|\geq i$
Now that the exterior has an upper bound on it’s size we can see how that impacts the interior degree. Because the degree of each node is still impacted by the minimum degree node, they still need to have their out-degrees go somewhere. Lemma 8.4 (Interior Degrees Lower Bound) compliments 8.3 (Strong Decreasing Exteriors Lemma) by saying thaw each decrease in the exterior degree will be matched by an increase in the interior degree.
9 Transitive Triangles and Seymour Diamonds
Before we continue with the development of our DSP algorithm and show how it solves the Seymour Second Neighborhood Conjecture, we need to address some potential problems and show that they will not hinder our algorithm.
Definition 9.1.
Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Here, $x$ points to both $y$ and $u$, and $y$ also points to $u$. The ”transitivity” comes from the fact that the direction of the edges follows a logical progression: $x\rightarrow y$ and $y\rightarrow u$ implies a pathway $x\rightarrow u$ via $y$. The concept of a transitive triangle is significant in understanding how the structure of the graph influences properties like out-degrees in $G^{2}$.
In a transitive triangle, $x$ has a direct path to both $y$ and $u$, and the existence of the edge $(y,u)$ implies that the out-degree of $x$ in $G^{2}$ could potentially increase because the paths from $x$ to nodes reachable from $y$ and $u$ are being counted. Understanding how transitive triangles are distributed within the graph can help in analyzing the neighborhood structures and their effects on the square graph, is critical for analyzing the Seymour’s Second Neighborhood Conjecture.
Definition 9.2.
Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $u_{1}$ and $u_{2}$, and these two out-neighbors both share the same out-neighbor, $y$. This is basically a four vertex oriented graph. What this situation presents is redundant second neighbors. There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond. As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and more importantly the Seymour’s Second Neighborhood Conjecture.
The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles. They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e. one where every node has the decreasing sequence property. We provided a similar result with Lemma 5.4 (Interior Degree Greater than 1). When a minimum degree node $v_{0}$ has the decreasing sequence property, all of its neighbors have interior degree of at least 1. That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$ will form a transitive triangle. We can extend that further for any two constructive containers.
xyz
Lemma 9.1.
Transitive Triangles from Containers In a graph $G$ represented by a GLOVER, for any two consecutive containers $C_{i}$ and $C_{i+1}$ in a graph satisfying the Decreasing Sequence Property (DSP), there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle. That is, there exist directed edges from $u$ to $w$, $v$ to $w$, and from $u$ to $v$, or from $v$ to $u$.
This lemma highlights the relationship between a minimum degree node and the formation of transitive triangles, providing a foundation for discussing the implications related to the Seymour conjecture.
Lemma 9.2.
Seymour Diamonds from Containers
In a graph $G$ represented by a GLOVER, for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
$u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
The following arcs hold:
1. There are forward arcs $u\to v$ and $u\to w$, 2. There are forward arcs $v\to y$ and $w\to y$, 3. No back arcs exist between containers.
This lemma effectively describes the conditions for forming a Seymour diamond and links it to the implications regarding the Seymour conjecture. Because we can extend the previous lemmas to encompass any nodes that have the DSP. This inclusion emphasizes that transitive triangles and Seymour diamonds have been integral to our analysis rather than overlooked.
xywu
Lemma 9.3.
Seymour Diamonds and Decreasing Path Algorithm Let $G$ be an oriented graph represented by the GLOVER, in an attempt to prevent a degree during node, where we are applying the DSP path algorithm. Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond. Then, the presence of the Seymour diamond $D$ does not disrupt the DSP path algorithm.
Proof.
  • A Seymour diamond consists of two vertices, $x$ and $u$, connected by two parallel paths via $y$ and $w$.
  • The DSP path algorithm can utilize one of these paths while disregarding the other, as both connect $x$ and $u$, thus maintaining degree constraints.
  • The presence of redundant paths does not diminish the outgoing neighbors of any node, ensuring that the minimum degree condition holds throughout the process.
  • Consequently, these additional connections offer alternative routes without introducing cycles or otherwise disrupting the algorithm’s operation.
None of these case are a deterrent to the execution of the DSP Path algorithm. ∎
Lemma 9.4.
Interaction of Minimum Degree Nodes and Transitive Triangles Let $G$ be an oriented graph represented by a GLOVER, containing a minimum degree node $v_{0}$ with degree $d$. Consider a transitive triangle formed by nodes $z$, $y$, and another vertex $x$, where $z$ and $y$ are part of this triangle. The following cases are possible:
  • If $d^{+}(y)=d$ and $d^{+}g(z)=d$
  • If $d^{+}(y)=d$ and $d^{+}(z)>d$
  • If $d^{+}(y)>d$
  • If $d^{+}(y)<d$
Proof.
  • Case 1: If $d^{+}(y)=d$ and $d^{+}g(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
  • Case 2: If $d^{+}(y)=d$ and $d^{+}(z)>d$: $z$ covers some neighbors of $y$, but since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
  • Case 3: If $d^{+}(y)>d$: $y$ has more outgoing connections than required. It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
  • Case 4: If $d^{+}(y)<d$: This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
Understanding transitive triangles and Seymour diamonds enriches our analysis of oriented graphs, demonstrating the interplay between different graph structures. This exploration contributes to the overall argument and leads us to our main theorem.
10 Main Theorem: Seymour’s Conjecture is True
10.1 Introduction
In this pivotal section, we present the main theorem derived from our analysis of oriented graphs. We articulate the theorem clearly, outlining its implications and the conditions under which it holds.
This paper has been a battle between the chosen minimum degree node and the Decreasing Sequence Property (DSP), with the former acting as a catalyst pushing things forward and the latter acting as a bottleneck, trying to hold things back. Along the way, we have proved several key results, most notably that:
  • Back arcs cannot exist,
  • Interior degrees double inside each container,
  • The interior degree doubling causes a decrease in exterior size.
This interaction between interior degree increases and exterior size decreases drives the structure of the graph. However, this degree increase is bounded by the bound $\frac{n-1}{2}$ because of the size of the containers, which provides a limit on the maximum average degree inside a container. Once the system reaches this maximum bound, the exterior degree of a parent container will double, causing a chain reaction where both interior and exterior degrees double, ultimately leading to the node’s overall degree doubling as well.
We first restate the Lemma 6.4 (Non-Existence of Back-Arcs), the No Back Arcs Lemma from section 6, but slightly rephrased.
Lemma 10.1.
Let $G$ be an oriented graph represented by a GLOVER, in an attempt to prevent a degree doubling node, with all nodes satisfying the Decreasing Sequence Property (DSP), where $v_{0}$ is a minimum degree node in $G$, and all nodes are placed into containers. If there are an back arcs in $G$, then a node’s degree will double.
10.2 Theorem Statement
Degree Doubling by Exterior Decreasing
Theorem 10.1.
Let $G$ be an oriented graph represented by a GLOVER with all nodes the Decreasing Sequence Property (DSP). If all nodes in $G$ are placed into containers, where interior degrees are maximized, then Seymour’s conjecture holds true. Specifically, a node in $G$ will have its exterior size decrease to less than 2.
Algorithm 1 Algorithm for Analyzing Degree Doubling in Containers
1:Initialization:
2:Choose the minimum degree node $v_{0}$.
3:Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$.
4:for each container $C_{i}$ in $G$ do:
5:     Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles.
6:     Move to a neighboring container $C_{i+1}$.
7:     if there is a back arc from $C_{i+1}$ to a prior container then
8:         Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double.
9:         Halt.
10:     else
11:         Since there is no back arc, the container $C_{i+1}$ will be bound in size to be 1 less than the size of container $C_{i}$.
12:         Repeat from Step 4.
13:     end if
14:end for
Start with graph size $n$Select minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots$Identify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to prior containerLocate child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
Figure 10: Flowchart depicting the degree-doubling algorithm applied within containers, demonstrating the process of identifying nodes with the Decreasing Sequence Property.
Corollary 10.1.
Degree Doubling by Container Density In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},C_{2},...,C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable. This collision can manifest in two forms: either the size of the container reaches a point where it can no longer support the increasing interior degrees (resulting in nodes that cannot maintain their minimum degree requirements), or smaller containers force a doubling of degrees that contradicts the capacity for interior arcs. The structure of the graph, particularly in small sizes, inherently limits these growth dynamics, leading to contradictions that must be addressed.
Corollary 10.1 (Degree Doubling by Container Density) captures the insight exploration our interior degree growth dynamics within shrinking containers in an oriented graph. The language highlights the inevitable ”collision” and ties it back to fundamental graph constraints. Hopefully, this framing gives a precise summary of why these graphs have a limited capacity to sustain both shrinking containers and rising interior degrees without contradictions.
The main theorem, algorithm, and corollary encapsulates the findings of our research, providing a crucial result that advances the understanding of oriented graphs. The proofs and implications of this theorem will be further explored in the following sections.
Corollary 10.2.
[…]
$i=\frac{k-1}{2}$
where $k$ is the minimum degree of $G$. For even values of $k$, the complete graph $K_{n}$ will not occur on any actual even number of nodes but may appear in the interval between two odd numbers.
Proof.
We begin by restating some previously shown examples:
  • For $k=3$: $K_{3}$ (triangle) occurs in $C_{1}$.
  • For $k=5$: $K_{5}$ occurs in $C_{2}$.
  • For $k=7$: $K_{7}$ occurs in $C_{3}$.
These observations hold because the GLOVER structure ensures that each container $C_{i}$ increases in regularity and density as the degrees of nodes double, creating the conditions necessary for complete graphs to emerge.
For even values of $k$, such as $k=4$ or $k=6$, the symmetry and density required for a complete graph $K_{n}$ cannot form directly at any container level. This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the structural requirements for even $n$ are never satisfied in full. Instead, the structure transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
The relationship $i=\frac{k-1}{2}$ reflects the interplay between the minimum degree $k$ and the growth of density in containers. This relationship confirms that the GLOVER structure consistently generates complete graphs $K_{2i+1}$ in odd-numbered increments, with no direct occurrence for even $n$.
Corollary 10.3.
Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the complete graph occur in container $C_{i}$. Then every node in container $C_{i}$ has its degree doubled in the representation.
The identification of regularity was key to understanding why multiple nodes’ degrees can double in certain graphs. This concept offers a powerful explanation for the phenomenon. The regularity is the key factor that helps explain why multiple nodes can have their degrees double, rather than the search for multiple nodes being inherently complex. The regularity is the central feature that ties everything together.
Degree doubling is essentially a consequence of the underlying regularity. When a graph’s structure is highly regular, particularly in terms of interior and exterior degrees, it becomes easier to predict and verify that certain structural properties (like degree doubling) will occur at specific stages.
Regularity implies that as nodes increase in distance from the starting node, their connectivity and interior degrees will follow a pattern. This pattern is key to understanding why degree doubling happens more naturally and why multiple nodes can experience it simultaneously, if the conditions allow.
10.3 Decreasing Sequence Property Path Algorithm
We now present a precise algorithm to identify nodes whose out-degree doubles in the square of an oriented graph, as per Seymour’s Second Neighborhood Conjecture (SSNC). This algorithm synthesizes results from the lemmas, theorem, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
  • The containers reach a size two causing a node in the previous container to have its degree double, as stated in Theorem 10.2 (Degree Doubling by Exterior Decreasing).
  • The interior degree required for nodes inside containers will be greater than the possible for a container of that size causing a node in the previous container to have its degree double, as stated in Corollary 10.1
In both these cases we are able to find a node whose degree doubles.
Here is the algorithm in pseudocode. Input: An oriented graph $G$ with minimum out-degree $d$.
Output: A node $v$ in $G^{2}$ whose out-degree at least doubles.
Algorithm:
1. Initialize the Graph Structure: - Identify the minimum degree node $v_{0}$ in $G$. - Partition $G$ into containers $C_{1},C_{2},\ldots,C_{d}$, where each container $C_{i}$ groups nodes by their distance $i$ from $v_{0}$.
2. Check for Back Arcs: - For each container $C_{i}$, check for arcs leading to any container $C_{j}$ where $j<i$. - If a back arc exists from a node $u_{i}\in C_{i}$ to a node $v_{j}\in C_{j}$, record $u_{i}$ as a degree-doubling node by Lemma 7.2 (Interior Degree Doubling) and Lemma 6.4 (Non-Existence of Back Arcs). - If no back arcs exist, proceed with Step 3.
3. Establish Connectivity Properties: - For each node $v$ in container $C_{i}$, define: - Interior neighbors $\text{int}(v)$: Nodes within the same container. - Exterior neighbors $\text{ext}(v)$: Nodes in the next container.
4. Iterative Degree-Doubling Check: - For each container $C_{i}$: - Calculate the out-degree of each node in $C_{i}$. - If a node in $C_{i}$ has fewer second neighbors than first neighbors (DSP violation), record it as a potential degree-doubling node.
5. Decreasing Sequence Property Verification: - Verify that the size of each exterior set $\text{ext}(v)$ is strictly less than the size of the corresponding interior set $\text{int}(v)$ in $G^{2}$. - If not, apply the conditions set in Lemma 7.2 to ensure any doubling in interior degrees results in a matching exterior degree.
6. Termination and Output: - Return the identified node(s) with doubled degrees as proof of degree-doubling in $G^{2}$.
11 Conclusion and Future Work
Solving a conjecture with such a long and debated history requires thorough validation. This paper represents a comprehensive effort to meet the rigorous standards of mathematical proof, supported by tools designed to aid understanding. Acknowledging that such results are subject to detailed scrutiny, this process is viewed as an essential step in establishing confidence and recognition within the mathematical community.
The GLOVER structure—short for Graph Level Order—represents a significant advancement in addressing long-standing open problems in graph theory, such as the Seymour conjecture, which remained unresolved for 34 years. While not all oriented graphs are inherently well-ordered, the GLOVER structure reveals that, within the context of the Seymour conjecture, oriented graphs become effectively well-ordered. This property arises because any deviation from such an ordering would lead to a degree-doubling effect, rendering the conjecture invalid.
By leveraging this induced well-ordering, the GLOVER structure enables tools like mathematical induction to be applied effectively, facilitating a solution to the Seymour conjecture. Beyond this specific case, the framework provides a foundation for exploring dual metrics in graph theory, such as prioritizing distance and degree, and opens the door to applications in clustering, ranking, and multi-metric optimization.
The GLOVER structure emerges as a groundbreaking framework for addressing problems that rely on dual metrics, offering a versatile tool for organizing and analyzing data across theoretical and practical contexts. While its utility was demonstrated here in the context of the Seymour conjecture—where the dual metrics of distance and degree played a pivotal role—it is important to recognize that the structure’s applications extend far beyond regularity or degree-based problems.
At its core, the GLOVER structure’s ability to prioritize and organize relationships based on dual metrics allows it to address challenges that require both exhaustive and hierarchical searches. Unlike local search methods traditionally constrained by single metrics, the GLOVER approach enables partitioning and clustering in ways that preserve relationships and provide deeper insights. For instance, applications in clustering, ranking, and network analysis—whether in social networks, employee rankings, or other multivariate datasets—can benefit from this structure.
Looking ahead, the GLOVER structure challenges us to investigate its applicability to other open problems in graph theory and computer science. The approach may uncover new insights in contexts where dual or multi-metric prioritization is key, highlighting the broader potential of this versatile framework.
Future research and development should explore these broader applications. Questions of complexity, scalability, and adaptability to non-graph-based datasets remain promising areas for investigation. Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
12 Acknowledgments
I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:6 has been a guiding light throughout this journey. I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
Special thanks to the late Dr. Nate Dean for his contributions to this work and to Dr. Michael Ball for his support. I want to sincerely thank everyone who has taken the time to pray for me, share encouraging words, and offer constructive feedback. Your support has meant so much to me, and I’m truly grateful for the kindness and thoughtfulness you’ve shown.
Finally, I extend heartfelt gratitude to my family and friends for their patience and belief in me throughout this journey.
The GLOVER data structure is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Ranked Structure: The elements of $S$ are partitioned into a ranking of ”containers” or ”levels,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order. The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$. This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
Interior and Exterior Neighbors:
  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
Definition 13.4.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the interior neighbors of $y$ with respect to $x$, denoted as $int(x,y)$, are the out-neighbors of $y$ that are also out-neighbors of $x$. Formally,
$int(x,y)=N^{+}(x)\cap N^{+}(y),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$ in the graph. The number of interior neighbors of $y$ with respect to $x$ is called the interior degree of $y$.
Definition 13.5.
If the node $x$ has the decreasing sequence property and $y\in N^{+}(x)$, then the exterior neighbors of $y$ with respect to $x$, denoted as $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$. Formally,
$ext(x,y)=N^{+}(y)\cap N^{++}(x),$
where $N^{+}(z)$ denotes the set of out-neighbors of $z$, and $N^{++}(x)$ denotes the set of out-neighbors of the out-neighbors of $x$ (i.e., the second neighbors of $x$). The number of exterior neighbors of $y$ with respect to $x$ is called the exterior degree of $y$.
Definition 13.6.
Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$. A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
Definition 13.7.
Definition: Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
14 Appendix 2: Applicability of the Graph Level Order
Traditionally, a total order on a set is a linear, one-dimensional comparison between elements. The key properties of total orders—reflexivity, transitivity, and antisymmetry—give us a clear, unambiguous ranking where every pair of elements is comparable. Graph Level Orders (GLOVER) extend this concept to n-dimensions within a leveled framework.
GLOVER accomplishes this through the use of containers that not only arrange elements in sequence but also enable multiple levels of organization. Elements within each level maintain an internal order, while the containers themselves are comparable across ranked layers. This structure allows for sophisticated comparisons both within a single level and between levels, making GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
The examples provided in this paper illustrate concrete cases of GLOVER. However, it’s the abstract nature of GLOVER that sets it apart as a transformative mathematical concept. This structure allows us to conceptualize any set of ranked organized objects—not just graphs with regularly structured nodes, but also diverse entities, such as sets with multiple attributes or time-sequenced data. Within discrete mathematics, the ability to impose a multi-dimensional total order provides a richer understanding of relationships among elements. The order goes beyond pairwise comparisons to establish a multi-layered framework capable of supporting complex operations.
The true strength of GLOVER lies in its flexibility. It accommodates inherent ordering within single-level sets while also adapting to more intricate structures. By generalizing beyond domain-specific contexts like graph theory, GLOVER becomes a universal framework for imposing total order on any ranked structure, from simple ordered pairs to multi-dimensional data sets in fields like machine learning or clustering rankings.
The GLOVER structure also uniquely supports induction in graph-based proofs due to its total order, providing a new way to approach graph theory problems. Unlike traditional graph structures, which typically do not allow for induction, the total ordering of nodes and containers in GLOVER makes it possible to apply inductive reasoning effectively.
This expanded understanding of total orders paves the way for developing algorithms, proofs, and applications capable of handling complex, scalable data structures. GLOVER is not merely a form of ordering; it is a system of relationships that facilitates the manipulation, comparison, and analysis of multi-dimensional data through a clear and consistent layering.
The GLOVER framework, as initially defined, is tailored to address the specific requirements of the Seymour Second Neighborhood Conjecture. Its rigid structure-classifying arcs either as interior or exterior, and strictly prohibiting skipping levels-ensures clarity and sufficiency for analyzing second neighborhoods, However the framework can be extended to accomodate more general problems, including those with nested containers or more flexible neighbor relationships .
Nested containers introduce a hierarchy within individual containers, grouping nodes based on finer structural or relational properties. This extension enables the framework to address problems where additional levels of granularity within the same partition group are required.
Definition 14.1.
A nested container $C_{i,j}$ is a subset of a container $C_{i}$ defined by additional criteria beyond the qualifying metric that determined it into $C_{i}$. For example, nodes in $C_{i,j}$ might share specific adjacency patterns or connectivity property properties. Use Cases:
  • Clustering Problems: Nested Containers can represent hierarchal clustering within each distance group, allowing comparisons between subgroups.
  • Weighted Graphs: In graphs with edge weights, nested containers could group nodes based on weight thresholds within the same distance level.
Example:
Suppose $C_{2}$ contains nodes $x,y,z$ all at distance 2 from $v_{0}$. If $x$ and $y$ are connected but not $z$, $x$ and $y$ might form a nested container $C_{2,1}$, while $z$ forms $C_{2,2}$
In certain problems, arcs may not connect nodes in non-adjacent containers, effectively skipping levels in the distance hierarchy. While the SSNC prohibits such arcs to maintain a second-neighbor focus, other graph probems may allow or even require these connections.
  • Extension - Relax the exterior neighbors to include $C_{j})$ for $j>i+1$, introducing a new category of arcs, long-distance exterior arcs.
  • Use Cases -
    • Network Design - In communication networks, skipping levels might model direct long-range connections (for example satellite links).
    • Flow Problems - In transportation or flow networks, skipping levels could represent shortcuts or direct transfers.
  • Impact on GLOVER Properties - Allowing for skipping levels introducing new challenges
    • Order Maintenance - Total order within containers must be preserved despite the additional complexity of the long-distance connections.
    • New Arc Classifications - Introduce distinctions like short range exterior arcs (j = i + 1) and long range exterior arcs (j ¿ i + 1).
Exterior neighbors, as defined in the SSNC, connect nodes in adjacent containers. Expanding this definition allows for more flexible interactions between containers, broadening the applicability of the GLOVER framework.
  • Generalized Exterior Neighbors - for $x\in C_{i}$, an exterior neighbor $y$ can belong to any $j>i$. The specific range of $j$ will depend on the problem context.
  • Example Problem: Consider a graph where connections beyond second neighborhoods are of interest. Exterior neighbors could include nodes up to $k$-steps away, accomodating problems requiring $k$-neighborhood analysis.
While these extensions offer exciting generalizations, the SSNC’s specific focus on second neighbors necessittates a stricter interpretation of the GLOVER framework.
  • No skipping levels - Ensures that second neighbors are strictly defined nodes in $C_{2}$.
  • Rigid container definitions - prevents ambiguity in arc classifications or neighbor interactions.
  • Limited use of nested containers - The SSNC does not require additional granularity within distance groups, making nested containers optional for this specific problem.
Example 14.1.
Graph Theory - Seymour Second Neighborhood Conjecture: Graphs, which are normally seen as an unordered arrangement of nodes and edges, were able to be ordered with the GLOVER. We initially chose a minimum degree node and partitioned the nodes of the graph into containers based on their proximity to this node.
Partitioning nodes by distance allowed for definition of the interior degree and exterior degree which were very important to showing that a node’s overall degree would double.
Example 14.2.
Engineering - Dynamic Scheduling: In fields such as manufacturing, task scheduling often involves prioritizing independent tasks. With GLOVER, tasks are grouped into containers based on dependency levels and ordered according to priority. For example, preparatory tasks are prioritized in early containers, while finishing tasks are sequenced last, minimizing delays and optimizing task flow.
Partitioning groups tasks by dependency, while ordering between containers ensures essential tasks are completed first. Within each container, tasks can be further organized by resource requirements, promoting efficiency at each priority level.
Edges introduced at the container level to represent this dependency. Interior edges would show common dependencies between tasks in a common container, whereas exterior edges represent tasks of different priorities with common dependencies.
From this we can use graph theory concepts like centrality to better understand things like important tasks, important dependencies, how the containers (task tiers relate to one another) etc.
Example 14.3.
Engineering - Resource Management: In network resource allocation or energy grid management, nodes can be clustered by proximity to demand or criticality and ordered to prioritize high-need resources. This ordering ensures that essential nodes remain accessible, while resource usage remains balanced across the network.
Partitioning groups resources based on proximity to critical nodes, while ordering between containers directs allocations toward high-demand areas. Within each container, resources are accessed in order of current availability, optimizing responsiveness to fluctuating demands.
Example 14.4.
Engineering - Project Management: In complex projects involving multiple teams and dependencies, tasks can be clustered by project phase and ordered sequentially to prevent bottlenecks. Foundational tasks, such as design and planning, are prioritized early, while execution tasks follow, ensuring smooth progression through each phase.
Partitioning groups tasks by project phase, while ordering between containers facilitates the transition of tasks from one stage to the next. This ranked order remains stable despite potential project changes, enabling efficient coordination across large teams.
Example 14.5.
Cybersecurity - Threat Detection and Prioritization: In cybersecurity, ranked structures are often essential for organizing and prioritizing threats. Using an GLOVER framework, threats can be grouped into containers based on their urgency or potential impact. For example, critical threats with known vulnerabilities might be prioritized in early containers, while less immediate risks or less probable attack vectors are sequenced into subsequent layers. This ranked ordering helps ensure that the most pressing issues are handled first, thereby minimizing potential breaches or disruptions.
Partitioning into containers groups threats by their severity and probability, while ordering between containers facilitates a systematic approach to addressing risks from the highest priority down. Within each container, threats are further organized based on attributes like type, location in the network, or historical occurrence, allowing security teams to deploy resources effectively. This layered approach enhances the overall defense strategy by ensuring structured response and continuous monitoring across multiple risk levels.
Example 14.6.
Natural Language Processing - Ranked Text Classification: In NLP, ranked total order graphs can improve text classification by organizing data through multi-level categories. For instance, documents can be grouped into containers based on broad topics (e.g., news, entertainment, scientific research) with each container further ordered by subtopics like sports or politics under news, and biology or physics under science. This layered structure allows models to systematically refine classifications, progressing from general to specific categories, improving the accuracy and interpretability of the classification model.
Partitioning into containers groups text by broad categories, while ordering between containers prioritizes classification tasks from general to specific. Within each container, documents can be ordered by relevance, keywords, or thematic clustering, which supports tasks like topic modeling or intent recognition. This approach provides a clear and adaptable framework for complex, multi-layered categorization, enhancing the model’s ability to handle diverse and evolving datasets.
The versatility of the Graphs Level Order (GLOVER) structure hints at its applicability beyond this proof of the Seymour Second Neighborhood Conjecture. GLOVER has the potential to address several open problems and conjectures within discrete mathematics and graph theory, where ranked or layered ordering of elements can provide new insights or simplify complex relationships. Future work may reveal more connections between GLOVER and these longstanding challenges, offering new paths for research and exploration.
15 Appendix 3: Visualization and Accessibility of Oriented Graphs
In exploring oriented graphs, it became evident that fully grasping these structures often requires more than static illustrations in a paper can offer. While oriented graphs can be defined and studied abstractly, their complexity and behavior are best understood through dynamic, interactive representations. The Seymour Second Neighborhood Conjecture, in particular, necessitated defining paths, locating squares, and identifying degree-doubling nodes—tasks that involved intricate arc counting and relationships that benefit from visualization.
This realization motivated the creation of a complementary website. The site serves as a tool to enhance the accessibility of key definitions, lemmas, and proofs by providing visualizations and interactive examples. Our goal is to bridge the gap between abstract theory and intuitive understanding, making the intricate aspects of the conjecture approachable for a wider audience.
For example, consider the concept of a degree-doubling node or a lemma describing staircase-like structures in oriented graphs. While the paper describes these ideas in formal mathematical terms, the website illustrates them dynamically. This approach allows readers to ”see” the behavior and implications of these structures, offering a visual intuition that complements the formal reasoning presented in the paper.
This methodology draws inspiration from James Robert Brown’s insights in Philosophy of Mathematics, where he emphasizes the dual role of visualization in mathematics—as both a powerful tool for intuition and a potential pitfall if over-relied upon. By carefully balancing these aspects, we aim to enhance comprehension without compromising the rigor essential to mathematical proofs.
The website also provides an opportunity to include numerous examples and extended visual explanations that are not feasible within the space constraints of a traditional journal format. Through this medium, we invite readers to engage more interactively with the conjecture and its underlying theories, fostering a deeper understanding and encouraging constructive feedback.
Built using technologies like D3.js, JavaScript, and PHP, the website features dynamic visualizations of proof steps, interactive diagrams, and supplementary explanations. By demonstrating the synergy between mathematical results and computational tools, we aim to make these concepts not only accessible but also engaging for a broader audience.
Readers are encouraged to explore the website for additional insights and examples at:
https://glovermethod.com
While the website is already live and functional, it is an evolving project. Additional examples, applications, and refinements are being added continuously.
Readers are invited to explore the website and are encouraged to provide feedback on any aspects of the site or the content itself, including suggestions for improvements, corrections, or additional examples.
The goal is to create a resource that not only supports this paper but also serves as a tool for further exploration and discussion in the field of oriented graph theory, the Seymour Second Neighborhood Conjecture, and the Graph Level Order (GLOVER) data structure. Thank you for your engagement and contributions to this ongoing effort.

### v2

@@
 Seymour’s Second Neighborhood Conjecture asserts that in the square of any oriented graph, there exists a node whose out-degree at least doubles.
-This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph-Level Oriented Vertex Expansion and Reduction) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions.
+This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph Level Order) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions.
 By leveraging this structure, we construct a decreasing sequence of subsets that establish a well-ordering of nodes, ensuring that no counterexample can exist.
@@
 This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
+The problem of identifying structural properties in oriented graphs has been studied extensively, with particular focus on $m$-free graphs and Seymour vertices.
+Daamouch’s [7] work has been instrumental in this area, introducing significant results regarding induced subgraphs and their role in determining key vertex properties.
+This paper builds on these foundational ideas while presenting a novel approach that extends these concepts to interconnected container structures.
 In this paper, we will prove Seymour’s Conjecture by assuming it is false and then attempting to construct a counterexample.
@@
 }dist(v_{0},u)=i\}\text{ and A}(C_{i})=\{(u,v)|u,v\in C_{i}\}$ ,where $V(G)$ represents the vertex set of the graph $G$, and $dist(v_{0},u)$ is the shortest path distance between $v_{0}$ and $u$.
+This definition of a container is conceptually related to Daamouch’s use of induced subgraphs in his study of $m$-free graphs [7].
+However, the key difference lies in the formation of connected containers (see image), or neighboring containers, which allows for a more dynamic and scalable analysis of graph properties.
+This approach facilitates the identification of structural patterns across broader graph classes, a task that would otherwise be computationally prohibitive.
 As the previous definition aligns with the concept of double arrows, the term ”containers” allows us to explore the structure of oriented graphs and their interactions with minimum degree nodes and distances from these nodes.
@@
 Proposition 5.1.
-Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node, and if there exists a cycle in the first container $C_{1}$, then the degree of $v_{0}$ will not double.
+Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node of degree $\delta$, and if there exists a cycle in the first container $C_{1}$ of length $\delta$, then the degree of $v_{0}$ will not double.
 Example 5.4.
@@
 The argument about the doubling of the out-degree in $G^{2}$ is key in demonstrating that the structure defined by the GLOVER representation prevents such a situation, thereby ensuring that $|ext(x,y)|<d^{+}(x)$.
+The results presented in 5.1(Cycle in Container 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1, which demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
+While our work is similar to this result, we will extend the framework by leveraging container-based structures to provide a polynomial-time solution, highlighting the interplay between neighboring containers and their shared arcs.
 Lemma 5.6.
@@
 By maintaining the distinction between first and second neighbors through careful conditions, we can mitigate the risk of double counting and strengthen the argument presented in the proof.
+This next lemma looks into the Weak Decreasing Exteriors Lemma and considers the assumption of back arcs.
+There is a particular scenario that needs to be evaluated where a second neighbor of $y$ and first neighbor of $z$ can also be a first neighbor of x.
+The above lemma, called 5.6(Weak Decreasing Exteriors Lemma), has an explicit assumption that gets rid of this.
+In the next chapter we will do a lot of work to make sure that these back arcs are no longer a problem.
+However we will consider a second version of this lemma that gets rid of the back arcs assumption.
 Example 5.5.
@@
 It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container.
+This is important because the previous section introduced the GLOVER framework as a total order of containers.
+Back arcs can disrupt this order.
+What we see next is a special case of how back arcs could disrput this order.
+Lemma 6.3.
+Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
+1.
+Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP).
+2.
+The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$.
+3.
+There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
+Then, for all $z\in\text{ext}(x,y)$, we have:
+$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
+First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
+This gives the only possibility for $z$ as being an interior neighbor of $x$.
+Consequently, the arc $w,z$ is an early back arc as we saw in 6.2(No Back Arc to Early Container).
+These arcs can lead to double counting.
+An out-neighbor of $w$ might be an out-neighbor of $x$, and a second out-neighbor of $y$.
+This implies that under the assumption that $w$’s interior degree double’s, $w$’s overall degree will double through that early back arc.
+The framework under 6.2 (No Back Arc to Early Container) generalizes the solution to handle all potential back arc scenarios in the graph.
+Thus, we have strengthened the overall structure, making the algorithm robust against these issues and ensuring that the exterior degree conditions are preserved across all containers.
 Lemma 6.4.
@@
 Proof.
-This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.3 (First Back Arc Between Consecutive Containers) . ∎
+This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.4 (First Back Arc Between Consecutive Containers) . ∎
 […]
@@
 We can extend that further for any two constructive containers.
-xyz
+xyu
 Lemma 9.1.
@@
 Proof.
-  • Case 1: If $d^{+}(y)=d$ and $d^{+}g(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections.
+  • Case 1: If $d^{+}(y)=d$ and $d^{+}(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections.
 For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$.
@@
 This relationship confirms that the GLOVER structure consistently generates complete graphs $K_{2i+1}$ in odd-numbered increments, with no direct occurrence for even $n$.
+This emergence of orientations of complete graphs as containers decrease in size adds a layer of symmetry and beauty to the Seymour Second Neighborhood Conjecture.
+There was never a question of these arising.
+Some questioned multiple nodes.
+Some questioned zero nodes.
+A complete graph occurrence though is a thing of beauty and it helps give rise to another result.
 Corollary 10.3.
@@
 Then every node in container $C_{i}$ has its degree doubled in the representation.
+The discovery of multiple degree-doubling nodes transforms the Seymour Second Neighborhood Conjecture from a question of existence to one of abundance.
+This result not only confirms the conjecture but reveals an inherent richness in the structure of oriented graphs.
+The presence of multiple nodes is tied to the emergence of complete graph structures as containers decrease in size, showcasing an element of mathematical beauty.
+Moreover, this has real-world applications: in network analysis, it suggests that influence or importance may not rest with a single node but instead within a clique or tightly connected group, offering new perspectives for studying collaborative dynamics and group influence.
 The identification of regularity was key to understanding why multiple nodes’ degrees can double in certain graphs.

### v3

@@
 Abstract
-Seymour’s Second Neighborhood Conjecture asserts that in the square of any oriented graph, there exists a node whose out-degree at least doubles.
-This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph Level Order) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions.
-By leveraging this structure, we construct a decreasing sequence of subsets that establish a well-ordering of nodes, ensuring that no counterexample can exist.
-This approach not only confirms the conjecture for all oriented graphs but also provides a novel framework for analyzing degrees and arcs in complex networks.
-The findings have implications for theoretical graph studies and practical applications in network optimization and algorithm design.
+Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph.
+This paper first presents a data structure GLOVER (Graph Level Order) which is a partitioning of the nodes into a total order of neighborhoods.
+This data structure allows us to construct a decreasing sequence of subsets, ensuring that no counterexample can exist.
+Further, this algorithmic approach allows us to know where the node is.
+Later findings extend the results to multiple nodes.
+This is true for all oriented graphs.
+These findings also have practical applications in network optimization and algorithm design.
 1 Introduction
-Seymour’s Second Neighborhood Conjecture, proposed by Paul Seymour in 1990, seeks to identify a node in an oriented graph whose out-degree at least doubles in the graph’s square.
+Seymour’s Second Neighborhood Conjecture (SSNC) was originally proposed by Paul Seymour in 1990.
+It asks the question of the existence of a node in an oriented graph whose out-degree at least doubles in the graph’s square.
 The conjecture was first published by Nathaniel Dean and Brenda J.
@@
 For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}$ represent the first and second neighborhoods of $v$ respectively.
-This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]) to encompass a broader class of oriented graphs, thereby providing a framework applicable across various minimum degree scenarios.
-The Second Neighborhood Conjecture presents a compelling challenge within the context of graph theory, focusing on oriented graphs and the behavior of second neighbors in the graph’s square.
-Efforts to resolve this conjecture reflect the deep intrigue and mathematical beauty inherent in such problems.
-Beyond its theoretical importance, the conjecture has potential applications in network theory, algorithm design, and combinatorial optimization—fields where optimizing connections between nodes is crucial.
-While Seymour’s Conjecture remains an open problem, various efforts have tackled special cases.
-For example, Kaneko and Locke [12] demonstrated its truth for graphs with a minimum degree of at most six.
+This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]).
+Here, we hope to tackle the entire class of oriented graphs based on the structure laid out in that paper.
+The Seymour Second Neighborhood Challenge is a difficult graph theory problem.
+The problem itself asks us to focus on the question of oriented graphs and second neighbors.
+There have been several papers that have looked into special cases.
+Kaneko and Locke [11] showed that the conjecture was true for graphs with a minimum degree of at most six.
 Brantner et al.
@@
 Additionally, Chen et al.
-[3] provided a lower bound for how much a node’s out-degree increases in the graph’s square, concluding that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
+[3] provided a lower bound for how much a node’s out-degree increases in the graph’s square.
+They concluded that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
 In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture.
 This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
-The problem of identifying structural properties in oriented graphs has been studied extensively, with particular focus on $m$-free graphs and Seymour vertices.
-Daamouch’s [7] work has been instrumental in this area, introducing significant results regarding induced subgraphs and their role in determining key vertex properties.
-This paper builds on these foundational ideas while presenting a novel approach that extends these concepts to interconnected container structures.
-In this paper, we will prove Seymour’s Conjecture by assuming it is false and then attempting to construct a counterexample.
-Specifically, we explore the conditions under which no node’s out-degree doubles in the square of an oriented graph.
-Our approach focuses initially on minimum out-degree nodes, analyzing how their out-degree could fail to double, and gradually extending this analysis to other nodes based on their distance from the minimum out-degree node.
-The Seymour Second Neighborhood Conjecture has long been considered an intractable problem in graph theory, with some experts even doubting its validity.
-This paper presents a rigorous proof of the conjecture, accompanied by extensive supporting materials, including appendices, detailing definitions, lemmas, examples and an algorithm, as well as an interactive website for visualizing key ideas.
-These resources aim to make the proof as transparent and accessible as possible, acknowledging the high standards of scrutiny such results rightly deserve.
+The topic of $m$-free graphs is of particular interest to this paper because Daamouch’s [7] work is very similar to what we see here.
+In his paper, he introduces induced subgraphs and speaks of their roles in degree properties on vertices, which will also be a key thing in this paper.
+The Seymour Second Neighborhood Conjecture is not strictly a graph theory problem.
+Although it has historically been studied within the field of mathematics, it has practical applications across several domains, including social media, social networking, epidemiology, disease control, and many other fields.
+In the critical context of disease control, the nodes in this conjecture can represent individuals, and the orientation of the arcs symbolizes one person infecting another.
+Here, the application of SSNC becomes relevant by asking: Is there a person whose influence in their second radius of interaction (i.e., people they indirectly connect with) exceeds the number of people in their first radius (i.e., direct contacts)?
+In this paper, we will prove the Seymour Second Neighborhood Conjecture is true.
+We will do this by first assuming that it is false and attempting to construct a counterexample.
+We will explore the conditions under which no node’s out-degree doubles will double in the square of the oriented graph.
+This approach will focus initially on minimum out-degree nodes.
+We will analyze how their out-degree could fail to double.
+We will then extend out algorithm to other nodes based on their distance from the minimum out-degree node.
 The paper is organized as follows: Section 2 begins with some common terminology used in graph theory.
@@
 $E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
-We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$, which are nodes that $v$ has arcs pointing to.
-Similarly, $N^{++}(v)$ refers to the second neighbors of $v$–that is, nodes that are two steps away from $v$ in $G$.
+We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$.
+These are the nodes that $v$ has arcs pointing to.
+Similarly, $N^{++}(v)$ refers to the second neighbors of $v$ that is, nodes that are two steps away from $v$ in $G$.
 Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices.
@@
 3 Approach and Motivation
-We begin our analysis by considering general oriented graphs without making any initial assumptions about node degrees or graph properties.
-There are, however, fundamental characteristics common to all graphs, such as the existence of an out-degree distribution.
+We begin our analysis by considering general oriented graphs.
+We make no initial assumptions about node degrees or graph properties.
+There are, however, fundamental characteristics common to all graphs.
+We know that there is an out-degree distribution.
 Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes.
 By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
-From this foundation, we approach the problem from two perspectives.
-The first is the graph $k$-partition problem [10], where the original oriented graph $G$ is partitioned into $d_{v_{0}}+1$ subgraphs, corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$.
-This partitioning enables us to leverage the divide-and-conquer paradigm, which is widely recognized in computer science for simplifying complex problems by tackling smaller, independent subproblems.
-These partitions enable us to decouple the analysis of interior relationships within each partition from the exterior interactions between partitions, allowing us to address each separately before examining how they integrate within the overall structure.
-Although the partitions remain connected through certain edges, each can be analyzed independently, allowing us to address subproblems in isolation before combining results to solve the overall problem.
-By analyzing each partition individually, we aim to gain localized insights into the graph’s structure and degree dynamics, which we can then synthesize to form a cohesive solution.
-This divide-and-conquer approach has demonstrated considerable success in algorithms such as mergesort, quicksort, and matrix multiplication.
-While divide-and-conquer is not commonly applied in graph theory—largely because graphs are typically viewed as unordered collections of vertices and edges—we seek to view this differently by introducing a total ordering on the nodes of $G$.
-This will allow us to treat the graph in a way that is amenable to divide-and-conquer strategies.
-While previous approaches to Seymour’s Second Neighborhood Conjecture often assume the location of a degree-doubling node and focus their analysis around this fixed point, our approach adopts a search-based strategy.
-Rather than assuming where the doubling will occur, this paper introduces an algorithm that systematically searches the graph, identifying degree-doubling nodes based on structural properties.
-This method offers a flexible and robust framework that does not depend on preexisting assumptions, allowing for a comprehensive analysis applicable to a broader range of oriented graphs.
-This fundamentally shifts the way we approach graph theory problems by introducing a hierarchical structure that is conducive to divide-and-conquer techniques.
-This reimagining of graph structure allows us to efficiently isolate subproblems, analyze them independently, and then integrate results—offering both computational efficiency and scalability.
-One of the key advantages of the structure is its capacity to handle multiple metrics: an outer metric (e.g., distance) for partitioning the graph, and an inner metric (e.g., degree) for analyzing each partition’s internal structure.
-This dual-metric approach is powerful because it allows for a more nuanced exploration of graph properties, particularly in cases where traditional methods would be insufficient.
-The algorithm introduced in this paper systematically searches the graph, leveraging the graph’s structure, total ordering and degree metrics.
-By not presupposing the location of degree-doubling nodes, the algorithm dynamically identifies these nodes, ensuring a comprehensive exploration of the graph’s structure.
-Previous work on Seymour’s Second Neighborhood Conjecture often centers around assuming the position of a degree-doubling node and anchoring the analysis around this fixed point.
-In contrast, our approach eliminates this assumption, offering a more flexible framework that dynamically adapts to the structure of the graph.
-This ability to apply a divide-and-conquer strategy to a broader set of graphs marks a significant step forward in graph theory.
+From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$.
+This partitioning enables us to leverage the divide-and-conquer paradigm.
+No nodes are being removed to form these partitions.
+This is a conceptual idea.
+Instead, Think of it as putting a box around the partition.
+Now the nodes inside the box are can work with the other nodes inside the box and solve their problem independent of the other partitions.
+These partitions are still connected to one another, but thinking of them in this way, allows us to do more.
+This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecure.
+Many of these approaxhes were conceptual or theoretical, saying only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper.
+Secondly, most of these paper focused on local search techniques that statrted at or near the desired degree doubling node.
+That presents a problem because oriented graphs can be varying in nature, This is the first exhaustive search technique on the SSNC problem.
+We have laid the groundwork for our algorithm to tackle the oriented graphs.
+With a minimum degree node, we can partition this graph into independent subproblems.
+This is now ready for an exhaustive search algorithm.
 4 Initial Lemmas
@@
 It does however, provide insight into our reasoning.
-We begin our investigation of oriented graphs by searching for a node whose degree could potentially double.
-While many candidates exist, our goal is to find a node such that, if its degree does not double, we still gain useful information about the structure of the graph.
-This approach allows us to proceed iteratively, analyzing the graph node by node until we either exhaust all possibilities or discover a counterexample.
 Example 4.1.
@@
 Lemma 4.1.
-Minimum Out-Degree ¡ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
+Minimum Out-Degree $<$ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
 Proof.
@@
 Then $v_{0}$ has no outgoing edges in $G$.
-In $G^{2}$, it will still have out-degree 0, and hence its out-degree trivially satisfies the condition of at least doubling.
+In $G^{2}$, it will still have out-degree 0.
+Hence its out-degree trivially satisfies the condition of at least doubling.
 Case 2: Assume that $d^{+}(v_{0})=1$.
@@
 Because $G$ is an oriented graph and $(v_{0},v_{1})\in G$, it follows that $(v_{1},v_{0})\notin G$.
-Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$, meaning $(v_{1},v_{2})\in G$.
+Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$.
+This means that $(v_{1},v_{2})\in G$.
 This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
@@
 Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
-Lemma 4.1 Minimum Out-Degree ¡ 3 gives us a glimpse on the picture of an oriented graphs.
-It tells us that, even with just the concept of a minimum degree node we are able to impose restrictions on not only that node $v_{0}$, but also on every other node $v\in G$.
-This later part is true because $v_{0}$ is a minimum degree node and so every other node must have at least the out-degree of $v_{0}$.
-What this lemma is saying is that cases where $v_{0}$ has very few out-neighbors, its second neighborhood will expand significantly.
-This aligns with the idea that low out-degree nodes should see a substantial increase in neighbors when considering the second neighborhood.
+Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs.
+We are learning that, in order to construct our counter-example, we need out minimum degree to be at least 3.
+This will also place that same minimum value restriction on every other node in the graph.
+In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
 Lemma 4.2.
@@
 Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$.
-Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$ Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$.
+Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$.
+Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$.
 This implies that $v_{1}$ must connect to nodes outside of $\{v_{1},v_{2},v_{3}\}$, as $v_{1}$ has no out-neighbors within the induced subgraph.
 Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in G.
-Thus, $v_{0}$’s out-degree will at least double in $G^{2}$.
-Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors is somewhat ”isolated” within the neighborhood, leading to an increased second neighborhood for $v_{0}$.
+Thus, $v_{0}$’s out-degree will at least double in $G^{2}$. ∎
+Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$ shares no neighbors with $v_{0}$.
+This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
 Example 4.2.
@@
 Therefore, assume they connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
-This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree ¡ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case.
+This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree $<$ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case.
 There is at most one edge between $v_{4}$ and $v_{5}$.
@@
 Thus, either $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}$. ∎
-This lemma underscores a crucial insight: while a node’s out-degree doubles, it is not necessarily the initial minimum out-degree node $v_{0}$ whose degree doubles.
-Instead, the doubling occurs in one of its neighbors.
-This distinction is significant because it challenges prior approaches that focus on selecting the ”right” starting node for analysis.
-Our approach does not depend on starting with a specific node but instead leverages a search process that inherently reveals the doubling behavior in the graph structure.
-This framework allows for a more general and robust analysis.
+Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) answers a question that we will come back to throughout this paper - this is an exhaustive search algorithm.
+In order for that to work, it has to allow some node’s out-degree to not double.
+In this case, what we see is that the initial node in the example $v_{0}$ does not have its degree double.
+Instead, we are forced to search for the next node.
+That comes at a neighbor of $v_{0}$, $v_{1}$ whose degree does double in this example.
+This is an important distinction because we did not need the ”right” starting node.
+We chose a minimum degree node because of its ability to partition the nodes in the graph, not for its potential closeness to the double degree node.
 For the remainder of this paper, we will refer to graphs like those in Lemma 4.2 using double arrows to indicate that a node or set of nodes connects to every node in the subsequent subgraph.
-The graph can be illustrated as follows:
-The initial lemmas serve as essential building blocks in our exploration.
-By establishing these results, we create a solid foundation for the more profound theoretical analysis that follows.
-Each lemma reinforces the central theme of understanding the interplay between degrees and connectivity in oriented graphs.
+This can be done because the nodes within a common induced subgraph are all relating to the same set of nodes.
+These lemmas provide insight for our motivation, and a framework for theoretical foundation. What is coming in the next sections will build on these results and will refer back to them.
 5 Theoretical Foundation
-Building on the initial lemmas, this section develops the theoretical foundation necessary for addressing the conjecture.
-We explore key properties such as containers, the Decreasing Sequence Property and the concepts of interior and exterior degrees, which are vital to our argument.
+In order to understand this next section, we will introduce some new terms.
+We will give a formal name to the partitions we spoke of about previously - containers.
+We will also speak about the decreasing sequence property and talk about its importance.
+Then we will talk about interior, exterior and back arcs.
+Then we will close with the GLOVER data structure.
 Example 5.1.
@@
 }dist(v_{0},u)=i\}\text{ and A}(C_{i})=\{(u,v)|u,v\in C_{i}\}$ ,where $V(G)$ represents the vertex set of the graph $G$, and $dist(v_{0},u)$ is the shortest path distance between $v_{0}$ and $u$.
-This definition of a container is conceptually related to Daamouch’s use of induced subgraphs in his study of $m$-free graphs [7].
-However, the key difference lies in the formation of connected containers (see image), or neighboring containers, which allows for a more dynamic and scalable analysis of graph properties.
-This approach facilitates the identification of structural patterns across broader graph classes, a task that would otherwise be computationally prohibitive.
-As the previous definition aligns with the concept of double arrows, the term ”containers” allows us to explore the structure of oriented graphs and their interactions with minimum degree nodes and distances from these nodes.
-By introducing containers, we can focus not only on the properties of the nodes relative to a minimum degree node, but also on their interactions within and across containers.
-This framework enables us to examine node behavior in three key aspects.
-First, we can study node behavior relative to a minimum degree node.
-Containers group nodes based on their distance from a minimum degree node $v_{0}$, making it easier to analyze how these nodes behave and how their properties evolve as we move further from $v_{0}$.
-Second, we can investigate internal interactions within containers.
-Once nodes are grouped into containers, we can study how nodes within the same container influence one another.
-These internal interactions may reveal patterns related to degree growth, connectivity, and second-degree effects.
-Containers are an attempt to bring order to an oriented graph.
-We mean that in the literal sense of the word, order.
-For nodes belonging to two different containers, they can now be compared by their distance to a minimum degree node $v_{0}$ which we call in container 0.
-Since distance is a total ordering, this allows to compare over every other node in $G$
-The next few chapters will speak about how these containers communicate with each other and within themselves.
-No nodes or arcs are removed from the graph to create these containers.
-This is just a way of thinking about the graph so as to perform our analysis in search of a counterexample.
-Finally, we can explore inter-container interactions.
-By analyzing interactions between nodes from different containers, we gain insight into how graph components expand or contract, how nodes influence one another across distances, and how arc directions impact the convergence of the algorithm or counterexample we aim to develop.
+Daamouch used induced subgraphs similar in his study of $m$-free graphs [7]. The primary differece here is the neighboring containers (see image) which will be important in the proofs.
+With the concept of containers we can explore the structure of oriented graphs from two perspectives.
+We have the option of exploring it from the perspective of the minimum degree node, or from the perspective of containers.
+  • Containers are seen based on their distance from the
+  • Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
+minimum degree node.
 The case where $d^{+}(v_{0})=3$, and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) in particular, is an important one because it provides a glimpse into what a counterexample to the conjecture may look like.
@@
 For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$,where $N^{+}(x)$ denotes the set of out-neighbors of $x$, and $N^{++}(x)$ denotes the set of out-neighbors of the nodes in $N^{+}(x)$.
-In previous sections, we discussed two fundamental concepts: the minimum degree node and the decreasing sequence property (DSP).
-A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point for connectivity and flow within the graph.
-As the graph’s structural anchor, the minimum degree node drives the development of relationships and properties that will emerge in the counterexample construction.
-It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
-In contrast, the DSP introduces constraints on connectivity by systematically limiting potential connections beyond nodes that exhibit this property.
-This concept is akin to a bottleneck: even nodes with multiple connections may still encounter restricted reachability in the graph.
-A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP.
+We have now introduced two concepts: the minimum degree node and the decreasing sequence property (DSP).
+A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point both for our algorithm and degree values.
+By selecting it as the starting point it, it serves as the anchor.
+Now, the minimum degree node is forcing the remaining nodes into the properties that will emerge in the counterexample construction.
+It is acting as a catalyst.
+In contrast, the DSP introduces constraints on connectivity.
+This concept is akin to a bottleneck.
 Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
-They jointly reveal scenarios where expected connectivity may falter, highlighting vulnerabilities that could challenge established conjectures.
-The DSP, in particular, illustrates how exterior nodes in an oriented graph may systematically decrease in number, creating a structure where connectivity flow is constrained despite high out-degrees of certain nodes.
-By examining the DSP, we underscore a critical aspect of connectivity, resembling bottlenecks in network theory.
-Diestel’s Graph Theory [6] addresses the relationship between node degree and connectivity, demonstrating that a node $x$ can maintain many outgoing neighbors while still limiting reachability.
-This discrepancy signals inefficiencies and sets the stage for potential counterexamples to existing conjectures, thus motivating our investigation of graph properties and their broader implications.
 Lemma 5.1.
@@
 This organization is significant because it facilitates comparisons within each distance layer (container) by additional criteria, such as adjacency or neighborhood properties.
-This layered organization is especially useful in proofs, where we might need to argue about the relative positions of nodes within the same distance from $v_{0}$.
-This structure allows us to establish a refined form of comparability among nodes at equal distance that is crucial for analyzing specific properties.
-For example, in certain proofs, we might need to show that nodes within the same container have certain relationships that would not necessarily follow from a mere total order.
 Definition 5.3.
@@
 This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
-A leveled total order not only provides an ordering across distance groups but also enables detailed, structure-based comparisons within each group.
-By defining both inter- and intra-container comparisons, this ordering allows arguments to leverage container-level properties (like distance) and within-container properties (like adjacency) simultaneously.
-This flexibility is particularly useful in graph theoretic proofs like the SSNC, where it may be necessary to combine distance-based and structural properties.
-Defining a leveled total order gives rigor to proofs, as it offers a structured, multi-layered way to approach ordered sets that are partitioned by one property but require further comparability within each partition.
-This formalization could also have broader applications beyond the SSNC.
-It might apply to any problem where elements need to be organized hierarchically but also require nuanced relationships within each level of the hierarchy—such as layered networks, hierarchical clustering, and complex sorting problems.
+This order means that we have order in two ways.
+First, the containers themselves are ordered.
+Second, the nodes within the containers are ordered.
+This ordering must also be consistent within the containers.
+The notion of a non-triviality requirement is to distinguish this ordering from other traditional total orderings on graphs.
 Lemma 5.2.
@@
 ,where $N^{+}(z)$ is the set of out-neighbors of $z$. The number of interior neighbors is the interior degree of $y$ with respect to $x$.
-We previously introduced the concept of containers, which are crucial for understanding the structure of the graph.
-Interior neighbors are specifically those nodes for which both endpoints are located within the same container.
-By examining these interior neighbors, we gain insights into how containers relate to one another and how they function internally.
+Now that we have introduced containers, we can talk arcs where both nodes are inside the containers, these are interior neighbors.
+Interior neighbors will be important to our study on Seymour’s Second Neighborhood Conjecture and our search for a counterexample.
+In particular we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1).
+The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1 and we see that when all the neighbors in that subgraph had outdegree of 1, $v_{0}$’s degree did not double.
 Definition 5.5.
@@
 $ext(x,y)=N^{+}(y)\cap N^{++}(x)$
-, where $N^{+}(z)$ is the set of out-neighbors of $z$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
+, where $N^{+}(y)$ is the set of out-neighbors of $y$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
 We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers.
-This interaction is critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container.
-However, it is essential to consider other possibilities that may also influence this relationship.
-To better understand the distinction between interior and exterior neighbors, consider Example 5.2.
-In this example, arcs that connect nodes within the same container (or box) represent interior neighbors.
-In contrast, arcs that extend to nodes in the next container illustrate exterior neighbors.
+This interaction is again critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container.
+As an exercise to visualize interior and exterior neighbors, consider example 5.2.
+Exterior arcs are the double arrows between the boxes.
+Interior arcs are not drawn but can be mentally placed either between any two nodes in the red or blue box.
 Definition 5.6.
@@
 A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
-Containers are ordered based on their distance from the chosen minimum degree node.
-Up to this point, all arcs have either been interior or forward-facing.
-However, it is important to consider the possibility that some arcs may also direct back toward the chosen minimum degree node.
-Addressing this potential will help us understand how these back arcs could impact the Seymour Second Neighborhood Conjecture.
-Now that we have established the definitions of interior and exterior nodes, we can delve into the decreasing sequence property.
-This property refers specifically to the sizes of the exterior sets within the oriented graph.
-As the DSP path algorithm progresses, the exterior nodes — which are not part of the neighborhood of the minimum degree node — will systematically decrease in number.
-This phenomenon reflects a yin and yang relationship between the interior and exterior nodes: while the interior nodes maintain and potentially increase their interconnections, the exterior nodes steadily diminish, ensuring a balanced growth within the graph.
-In this dynamic, the minimum degree node serves as a catalyst for this process, driving the algorithm forward by imposing constraints and guiding the restructuring of connections.
-In contrast, the decreasing size of the exterior nodes acts as a bottleneck, effectively limiting the expansion of the graph’s out-degrees.
-This interplay is crucial for the success of the DSP path algorithm, as it maintains the structural integrity of the graph while facilitating necessary transformations.
-This relationship between the interior and exterior nodes will be the focal point of our proof throughout the rest of this paper, providing a framework for understanding how the DSP path algorithm operates effectively.
-With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the
+We have spoken of so much of our algorithm depending on the distance property of this oriented graph.
+We would like to continue this.
+In order to do so though, we need to consider the option of back-arcs and see how they might hurt our algorithm.
+These have the possibility of altering a container’s distance to the minimum degree node.
+Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture where they may force us to double count nodes.
+Having defined our arcs, we can talk about the decreasing sequence property.
+This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decreases as we get further from the minimum degree node.
+Exterior degrees, though, are a part of a node’s degree equation.
+So there’s another part of it.
+As we move further from the minimum degree node, the size of the interior degrees increases.
+This phenomenon reflects a yin and yang relationship between the interior and exterior degrees.
+In all of this, the minimum degree node is serving as a catalyst and driving force for the process.
+At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s outdegrees.
+With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the Graph Level Order (GLOVER) data structure.
 Definition 5.7.
@@
   • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
+To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
 Theorem 5.1.
@@
 Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a Ranked Total Order under the distance total order, completing the proof.
-This lemma aligns with the ranked structure we’re building, where each container $C_{i}$ is distinct in terms of reachability and influence.
-By enforcing this strict order under the absence of back arcs, we’re creating a clear layering within the graph, which further justifies the ”multi-level total order” approach.
-Back arcs, if present, could disrupt the unidirectional progression implied by containers, potentially creating feedback loops.
-By assuming no back arcs, we ensure that each container maintains its place in the total ordering, which adds robustness to the ranked structure we’re developing.
-Containers, as defined in the GLOVER framework, serve not only as organizational units but also as bottlenecks that regulate the flow of connectivity within the graph.
-By grouping nodes based on their shortest-path distance from the minimum degree node $v_{0}$, containers segment the graph into hierarchical levels.
-The decreasing sequence property ensures that the size of exterior neighborhoods diminishes as we move outward from $v_{0}$, creating a natural bottleneck effect.
-This reduction in connectivity simplifies the analysis of degree growth and neighbor interactions while preserving the structural integrity of the graph.
-Such bottlenecks are critical for maintaining the layered structure of the GLOVER framework, particularly in problems like the Seymour Second Neighborhood Conjecture, where second-level connectivity is a key focus.
+What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers.
+This ordering is also ordering the nodes within the containers, so the GLOVER data structure is a multi-level total order.
+The one thing that could disrupt this total ordering is a back arc.
+This could potentially lead to a feedback loop.
+By assuming no back arcs, we ensure that each container maintains its place in the total ordering.
 Next, we will show the power of this data structure and why is introduced in this paper.
-To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
 Lemma 5.3.
@@
 By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
-Assume that for any node $v$ in a container $C_{i}$, the following properties hold: - $d^{+}(v)\geq d^{+}(v_{0})$. - All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
+Assume that for any node $v$ in a container $C_{i}$, the following properties hold:
+  • $d^{+}(v)\geq d^{+}(v_{0})$.
+  • All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
 Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$.
@@
 Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
-Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that: - The out-degree of any node is anchored to $v_{0}$.
+Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that:
+  • The out-degree of any node is anchored to $v_{0}$.
   • There is no upward deviation in degree that contradicts the structure’s constraints.
-Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph.
+Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph. ∎
 Example 5.3.
@@
 Thus, for all $x\in N^{+}(v_{0})$,$|int(v_{0},x)|\geq 1$, completing the proof. ∎
-Lemma 5.4 (Interior Degree Greater than 1) is important because it constrains the out-neighbor structure of the graph when the minimum out-degree $v_{0}$ exhibits the decreasing sequence property.
-It guarantees that each out-neighbor of $v_{0}$ must have at least one other out-neighbor within $N^{+}(v_{0})$ , which helps to prevent the out-degree of nodes from being too heavily concentrated in the second neighborhood.
-The GLOVER structure partitions the neighbors of nodes into interior and exterior categories in a systematic way.
-This partitioning provides a powerful framework for analyzing oriented graphs, particularly in controlling and understanding degree properties.
-A notable consequence of this structured representation is the ability to guarantee that for any oriented graph $G$ represented as a GLOVER, the degree of a chosen minimum out-degree node $v_{0}$ will not double due to cycles in container $C_{1}$.
-More formally:
-Proposition 5.1.
-Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node of degree $\delta$, and if there exists a cycle in the first container $C_{1}$ of length $\delta$, then the degree of $v_{0}$ will not double.
+Lemma 5.4 (Interior Degree Greater than 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling.
+In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
 Example 5.4.
@@
 Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|ext(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
-This proof directly shows that a contradiction arises if the exterior degree exceeds the out-degree, which would violate the decreasing sequence property.
-The argument about the doubling of the out-degree in $G^{2}$ is key in demonstrating that the structure defined by the GLOVER representation prevents such a situation, thereby ensuring that $|ext(x,y)|<d^{+}(x)$.
-The results presented in 5.1(Cycle in Container 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1, which demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
-While our work is similar to this result, we will extend the framework by leveraging container-based structures to provide a polynomial-time solution, highlighting the interplay between neighboring containers and their shared arcs.
+Lemma 5.5 is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree Greater than 1), with the notable exception that Lemma LABEL:extbd1 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
+The results presented in 5.4(Interior Degree Greater then 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1.
+There, he demonstrated that a minimum degree vertex with no cycles in its induced graph has a degree that effectively doubles.
+Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we’re developing.
 Lemma 5.6.
@@
 Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
-The central concept of this proof revolves around the relationship between decreasing exteriors and back arcs.
-This relationship is not merely general; it specifically addresses the concern of potential double counting involving second neighbors that arise from back arcs through $y$.
-The additional condition - ensuring that a neighbor of $y$ does not also serve as a first neighbor of $x$ is vital.
-This condition helps to prevent the overlap in counting, particularly in situations where $z$, a second neighbor of $y$, could inadvertently be counted as a first neighbor of $x$.
-Given that containers are structured by their distance from the chosen minimum degree node, it follows that if a node $z$ is a second neighbor of $y$ and also a first neighbor of $x$, this necessitates the presence of a back arc from $z$ to a neighbor of $x$.
-Therefore, the lemma’s validity heavily relies on the configurations of back arcs in our construction, particularly in the context of a potential counterexample.
-In summary, the interplay between back arcs and the classification of neighbors is crucial in Lemma 5.6 (Weak Decreasing Exteriors Lemma).
-By maintaining the distinction between first and second neighbors through careful conditions, we can mitigate the risk of double counting and strengthen the argument presented in the proof.
-This next lemma looks into the Weak Decreasing Exteriors Lemma and considers the assumption of back arcs.
-There is a particular scenario that needs to be evaluated where a second neighbor of $y$ and first neighbor of $z$ can also be a first neighbor of x.
-The above lemma, called 5.6(Weak Decreasing Exteriors Lemma), has an explicit assumption that gets rid of this.
-In the next chapter we will do a lot of work to make sure that these back arcs are no longer a problem.
-However we will consider a second version of this lemma that gets rid of the back arcs assumption.
+The first thing to notice in this proof are the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions has an important purpose.
+First we will talk about the lack of back arcs.
+We have already talked about how back arcs can be a hindrance to our algorithm overall.
+This is one of the main points where they can be a hindrance.
+In particular there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist.
+We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
+Secondly, we assumed that a node’s interior degree doubled.
+This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem.
+We are assuming that this not $y$ was in another container and its interior degree doubled in that container.
+In doing so, $y$ is now in a better position to have its overall degree double, needing only to have its exterior degree double.
 Example 5.5.
@@
 […]
-The theoretical foundation established in this section underpins the subsequent arguments and proofs.
-By thoroughly examining the properties of oriented graphs, we prepare to tackle the complexities of back arcs and their implications for the overall graph structure.
+We have now defined the data structure that we will use throughout the rest of this paper.
+In doing so, we have a better understanding of interior and exterior arcs which will help guide us in our search for a counterexample.
 6 Back Arcs
-In this section, we analyze the role of back arcs in oriented graphs and their effects on graph properties.
-Understanding back arcs is crucial for assessing how they influence the interior and exterior degrees, as well as the overall structure of the graph.
+With a theoretical foundation underway, we are ready to launch an investigation into back arcs.
+We will look into whether these arcs can exist in a counterexample, and what other properties of oriented graphs must be true.
 It is important to restate that back arcs and exterior neighbors are distinct concepts.
@@
 Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples.
-Specifically, it shows that all the out-neighbors of interior arcs within a container must have the same degree, which is known as out-degree regularity.
-If, additionally, the nodes in this container have equal in-degrees, then the container achieves total regularity.
-It is important to note that this regularity emerges as a counterexample strictly due to the interaction between the Decreasing Sequence Property (DSP) and the minimum degree node $v_{0}$.
+In particular, it shows that all the out-neighbors of interior arcs within a container must have the same degree.
 In Example 5.5 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph.
 The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth.
-This outward growth reflects the structure of the graph, where nodes are grouped into containers based on their distance from $v_{0}$.
-A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$, effectively pointing against the distance-based order.
+A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$.
 If such a back arc were introduced, it would imply that the source node $y$ gains new neighbors at a lower level, potentially doubling its degree.
 This contradicts the hierarchical organization of containers, where each level is strictly defined by shortest-path distances from $v_{0}$.
-Challenge:
-To better understand why back arcs are not possible in this graph, consider the following questions:
-  • Can you construct a back arc in this graph that respects the container construction rules and distance-based ordering from $v_{0}$?
-  • If a back arc is introduced, does it lead to contradictions or disrupt the hierarchical structure of containers?
-Explanation:
-By the construction of the GLOVER framework, all nodes in $C_{i}$ are farther from $v_{0}$ than those in $C_{j}$ for $j<i$.
-A back arc $(y,z)$ would imply that $z$ is closer to $v_{0}$ than $y$, violating this ordering.
-Additionally, introducing a back arc disrupts the unidirectional flow of the graph by creating a feedback loop.
-This contradicts the fundamental property of containers: that each level represents a discrete, non-overlapping set of nodes ordered by distance from $v_{0}$.
-Without back arcs, the GLOVER structure preserves a clear layering, enabling rigorous analysis of degree growth and neighbor interactions.
 Lemma 6.2.
@@
 This implies that under the assumption that $w$’s interior degree double’s, $w$’s overall degree will double through that early back arc.
-The framework under 6.2 (No Back Arc to Early Container) generalizes the solution to handle all potential back arc scenarios in the graph.
-Thus, we have strengthened the overall structure, making the algorithm robust against these issues and ensuring that the exterior degree conditions are preserved across all containers.
 Lemma 6.4.
@@
 Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ leads to a fundamental contradiction in the oriented nature of the graph, and no back arcs can exist prior to $C_{i}$. ∎
-The concept of a first back arc inherently assumes that no back arcs exist up to a certain point.
-Lemma 6.2 (No Back Arc to Early Container) supports this assumption, relying on the doubling of nodes’ interior degrees, a condition not necessary for Lemma 6.4.
-The strength of this lemma lies in its capacity to identify the first back arc; if such an arc cannot be established, it reinforces the graph’s structural integrity.
-This lemma is founded on our ability to well-order the graph using containers and the Decreasing Sequence Property (DSP).
-With this ordering, we highlighted regularity within these containers through Lemma 6.1 (Regular Interior Degrees), which proved essential in showing that these back arcs would lead to a node’s degree doubling if a node’s interior degree was already doubling.
+Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees).
+From this, we can see that a back arc to a parent container violates the orientated nature of the graph.
 Lemma 6.5.
@@
 […]
-The exploration of back arcs reveals their significant impact on the dynamics of oriented graphs.
-By removing them, we can better navigate the structures of oriented graphs and in particular it paves the way to investigate interior degree dynamics.
+We stated that back arcs could be detrimental to our algorithm, so the removal of them in this section was important. Now we can better navigate the other properties of oriented graphs.
 7 Interior Degrees Doubling
-This section focuses on the important aspect of interior degree doubling within oriented graphs.
-We investigate the conditions under which this occurs and the implications for the overall degree distribution of nodes.
+This section introduces the concept of interior degree doubling.
+This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s the interior degree inside of a container.
+This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
 Lemma 7.1.
@@
 The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others.
-Initially, we accounted for nodes where interior degrees had doubled and ensured consistency when exterior degrees doubled as well.
-This distinction was critical for our algorithm up to this point.
-However, with regular containers, interior nodes are now indistinguishable from one another.
-This means we can swap a node $v_{i}$ with an exterior node $v_{e}$ if their degrees have doubled, ensuring that we have a node whose overall degree has doubled.
-This flexibility simplifies the process and ensures that the algorithm maintains consistency.
+Previously we were searching for a node in a container whose interior degree had doubled.
+That would be the key.
+What indistinguishabable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double.
+That means that we can select any of those nodes.
 Lemma 7.2.
@@
 Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
-Considering Lemma 7.2 (Interior Degree Double), we observe that it represents a specific instance of the Seymour Second Neighborhood Conjecture.
-Here, each container functions as its own self-contained oriented subgraph, where we pose the same core question: is there a node whose degree doubles?
-The affirmative answer we obtain, even when constructing a counterexample, suggests promising progress toward resolving the overall conjecture.
-Furthermore, we can expand upon the process illustrated in Lemma 7.2 (Interior Degree Double)).
-Within the proof, we constructed a monotonically decreasing sequence of subsets, which cannot continue indefinitely.
-This sequence reflects the essence of the Decreasing Sequence Property (DSP), where we observe that as interior nodes increase, the exterior nodes necessarily decrease.
-This requirement is what we meant by “decreasing” in DSP: the exterior set gradually diminishes as nodes transition to the interior.
-This dynamic can be interpreted as a form of equilibrium, reminiscent of the principle of yin and yang.
-Each node integrated into the interior necessitates a corresponding reduction in the exterior set, thereby preserving a balance.
-This interplay ensures that interior growth does not occur in isolation; instead, it aligns directly with constraints on the exterior, maintaining the graph’s structural integrity.
-Such interaction is essential for upholding the properties defined by DSP.
-It underscores the inherent relationship between interior and exterior neighbors, reinforcing the notion that the expansion of one aspect imposes natural limitations on the other.
-In summary, examining the doubling of interior degrees within this context provides critical insights into the degree distribution in oriented graphs.
-This understanding informs the proofs that will be discussed in subsequent sections, enhancing our grasp of both local and global properties of the graph.
+Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture.
+We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles.
+By finding a positive answer to this, and not just a positive answer, but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
+The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets.
+This cannot continue indefinitely.
+Once again, we see why the Decreasing Sequence Property (DSP) received its name.
+This doubling of interior degrees will help us with the construction of our algorithm in search for a counterexample.
+It does, however, give us a bit of light into the side of the Conjecture being true.
 8 Container Density
-Here, we delve into the concept of container density and its relevance to the properties of oriented graphs.
-We explore how regular exterior degrees interact with the overall density of containers and the implications for graph structure.
-Having proved that the interior degrees double within a container and that is linked to container size decreasing, we are ready to tackle the overall Seymour Second Neighborhood Conjecture.
-First we need to establish some fundamental truths about containers and oriented graphs in these counterexamples to set up the rest of the paper and things that must be true.
+Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean.
+The Weak Exteriors Decreasing Lemma implies that containers will decrease in size.
+We need to see how that decrease in size will impact the degrees of the nodes.
 Lemma 8.1.
@@
 Therefore, the assumption that $|ext(x_{i},y)|$ and $|ext(x_{j},y)|$ differ is false. Hence, the exterior degrees in $G$ must be regular if the interior degrees are regular. ∎
-What we saw in Lemma 6.1 (Regular Interior Degrees) that the regularity of interior degrees was something we were able to prove while we were trying to prove that back arcs did not exist.
-This both made sense and was very helpful in that environment.
-We see here that the regularity of interior degrees is not restricted to the container itself, but also impacts the exterior arcs to the next container - forcing them to be regular as well.
+We proved Lemma 6.1 (Regular Interior Degrees) on our way to proving that back arcs did not exist.
+LABEL:refintdeg is an important property of the graph though that we are going to use to get other properties of the graph, like 8.1.
 Lemma 8.2.
@@
 |$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
-This lemma is a more extended form of Lemma 5.6 (Weak Decreasing Exteriors Lemma), but without the added dependencies.
-It gives us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
-The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove.
-It is for this sequence that the term ”decreasing sequence property” is named.
-Several examples and observations brought this to a conjecture, which brought it here.
-The question now becomes is there any way we can strengthen this result?
-We demonstrated that the Decreasing Exteriors Lemma relies on two key factors: the elimination of back arcs and the doubling of interior degrees within each container.
-Furthermore, we established that the elimination of back arcs is directly dependent on the interior degrees doubling within those containers.
-By proving the previous lemma, we not only confirm the doubling of interior degrees but also conclude two additional results: (1) back arcs are entirely eliminated, and (2) the exterior arcs exhibit the desired decreasing pattern.
+Lemma (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma), to give us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
+The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named.
 Lemma 8.4.
@@
 Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
-Proof.
-Let $G$ be an oriented graph represented by a GLOVER.
-Let $v_{0}$ be a node with the minimum degree in $G$, and assume that every node in $G$ satisfies the Decreasing Sequence Property (DSP).
-Let $x$ be a node at distance $i$ from $v_{0}$, and let $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$.
-Our goal is to show that the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided that $(v_{0},x)\in G$ (i.e., $x$ is reachable from $v_{0}$).
-By Lemma 8.3 (Strong Decreasing Exteriors Lemma), we know that the exterior degree of $v_{i-1}$ with respect to $x$, $|ext(v_{i-1},x)|$, is bounded above by:
-$|ext(v_{i-1},x)|\leq d^{+}(v_{0})-i$
-where $d^{+}(v_{0})$ is the out-degree of $v_{0}$. This gives a limit on the number of nodes that $x$ can connect to at distance $i+1$ from $v_{0}$.
-Since the graph satisfies the Decreasing Sequence Property (DSP), we know that all nodes in the graph have a degree at least as large as $v_{0}$’s degree. Therefore, the degree of node $x$ satisfies:
-$d(x)\geq d(v_{0})$
-This means that node $x$ must have at least $i$ neighbors at distance $i+1$ from $v_{0}$.
-The structure of the graph explicitly disallows back arcs, i.e., edges pointing to nodes at earlier distances from $v_{0}$.
-Thus, $x$ cannot connect to nodes at distances less than $i$ without violating the structure of the graph.
-This constraint eliminates the possibility of $x$ mapping to earlier-distance nodes.
-Given that $x$ must connect to at least $i$ nodes, and since it cannot connect to earlier-distance nodes (due to the disallowance of back arcs), the only remaining option for $x$ is to map to interior nodes—nodes that are at distance $i$ or higher.
-Thus, since $x$ must map to at least $i$ distinct nodes, and all these must be interior nodes (by the elimination of back arcs and exterior connections), we conclude that:
-$|int(v_{i-1},x)|\geq i$
-This shows that the interior degree of $x$ with respect to $v_{i-1}$ is at least $i$.
-Therefore, we have shown that for a node $x$ at distance $i$ from $v_{0}$, the interior degree of $x$ with respect to $v_{i-1}$, $|int(v_{i-1},x)|$, is at least $i$, as required.
-$|int(v_{i-1},x)|\geq i$
-Now that the exterior has an upper bound on it’s size we can see how that impacts the interior degree.
-Because the degree of each node is still impacted by the minimum degree node, they still need to have their out-degrees go somewhere.
-Lemma 8.4 (Interior Degrees Lower Bound) compliments 8.3 (Strong Decreasing Exteriors Lemma) by saying thaw each decrease in the exterior degree will be matched by an increase in the interior degree.
+Having already placed an upper bound on the exterior sizes and containers, it is only natural that the there is an opposite bound for the interiors.
+Once again the minimum degree node is the driving force behind this.
+The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
 9 Transitive Triangles and Seymour Diamonds
@@
 Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
-Here, $x$ points to both $y$ and $u$, and $y$ also points to $u$.
-The ”transitivity” comes from the fact that the direction of the edges follows a logical progression: $x\rightarrow y$ and $y\rightarrow u$ implies a pathway $x\rightarrow u$ via $y$.
-The concept of a transitive triangle is significant in understanding how the structure of the graph influences properties like out-degrees in $G^{2}$.
-In a transitive triangle, $x$ has a direct path to both $y$ and $u$, and the existence of the edge $(y,u)$ implies that the out-degree of $x$ in $G^{2}$ could potentially increase because the paths from $x$ to nodes reachable from $y$ and $u$ are being counted.
-Understanding how transitive triangles are distributed within the graph can help in analyzing the neighborhood structures and their effects on the square graph, is critical for analyzing the Seymour’s Second Neighborhood Conjecture.
+Here, we have $x$ pointing to $y$ and $u$; and $y$ pointing to $u$.
+The concept of a transitive triangle is an important one because it involves the risk of double counting.
+The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles we might improperly count out second neighbors and thus not have a correct algorithm.
 Definition 9.2.
 Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
-The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $u_{1}$ and $u_{2}$, and these two out-neighbors both share the same out-neighbor, $y$.
+The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$.
 This is basically a four vertex oriented graph.
@@
 That is, there exist directed edges from $u$ to $w$, $v$ to $w$, and from $u$ to $v$, or from $v$ to $u$.
-This lemma highlights the relationship between a minimum degree node and the formation of transitive triangles, providing a foundation for discussing the implications related to the Seymour conjecture.
+Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm.
 Lemma 9.2.
@@
 The following arcs hold:
-1. There are forward arcs $u\to v$ and $u\to w$, 2. There are forward arcs $v\to y$ and $w\to y$, 3. No back arcs exist between containers.
-This lemma effectively describes the conditions for forming a Seymour diamond and links it to the implications regarding the Seymour conjecture.
-Because we can extend the previous lemmas to encompass any nodes that have the DSP.
-This inclusion emphasizes that transitive triangles and Seymour diamonds have been integral to our analysis rather than overlooked.
+1. There are forward arcs $x\to y$ and $y\to u$, 2. There are forward arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
+Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it.
+We didn’t know it because they they did not disrupt our algorithm and they are not a hindrance to our algorithm.
 xywu
@@
 Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
-Understanding transitive triangles and Seymour diamonds enriches our analysis of oriented graphs, demonstrating the interplay between different graph structures.
-This exploration contributes to the overall argument and leads us to our main theorem.
+Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs.
+Next we will blend everything together into our main results.
 10 Main Theorem: Seymour’s Conjecture is True
 10.1 Introduction
-In this pivotal section, we present the main theorem derived from our analysis of oriented graphs.
-We articulate the theorem clearly, outlining its implications and the conditions under which it holds.
-This paper has been a battle between the chosen minimum degree node and the Decreasing Sequence Property (DSP), with the former acting as a catalyst pushing things forward and the latter acting as a bottleneck, trying to hold things back.
-Along the way, we have proved several key results, most notably that:
-  • Back arcs cannot exist,
-  • Interior degrees double inside each container,
-  • The interior degree doubling causes a decrease in exterior size.
-This interaction between interior degree increases and exterior size decreases drives the structure of the graph.
-However, this degree increase is bounded by the bound $\frac{n-1}{2}$ because of the size of the containers, which provides a limit on the maximum average degree inside a container.
-Once the system reaches this maximum bound, the exterior degree of a parent container will double, causing a chain reaction where both interior and exterior degrees double, ultimately leading to the node’s overall degree doubling as well.
-We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), the No Back Arcs Lemma from section 6, but slightly rephrased.
+Here is where we sum up all the lemmas we have derived from our analysis on oriented graphs. We will state the final algorithm and some final results.
+So far, this battle has been a battle between a catalyst - the chosen minimum degree node, and the bottleneck - the Decreasing Sequence Property (DSP).
+We have proved some important things - Back arcs cannot exist; interior degrees must double inside each container, the exterior degrees decrease in size; interior degrees proportionally increase in size.
+There is a limit though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
+We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
 Lemma 10.1.
@@
 The structure of the graph, particularly in small sizes, inherently limits these growth dynamics, leading to contradictions that must be addressed.
-Corollary 10.1 (Degree Doubling by Container Density) captures the insight exploration our interior degree growth dynamics within shrinking containers in an oriented graph.
-The language highlights the inevitable ”collision” and ties it back to fundamental graph constraints.
-Hopefully, this framing gives a precise summary of why these graphs have a limited capacity to sustain both shrinking containers and rising interior degrees without contradictions.
-The main theorem, algorithm, and corollary encapsulates the findings of our research, providing a crucial result that advances the understanding of oriented graphs.
-The proofs and implications of this theorem will be further explored in the following sections.
+Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes.
+This was represented by the yin and yang effects early on.
+We see it full circle here.
+Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbor neighbors represented the size of the containers.
 Corollary 10.2.
@@
 Then every node in container $C_{i}$ has its degree doubled in the representation.
-The discovery of multiple degree-doubling nodes transforms the Seymour Second Neighborhood Conjecture from a question of existence to one of abundance.
-This result not only confirms the conjecture but reveals an inherent richness in the structure of oriented graphs.
-The presence of multiple nodes is tied to the emergence of complete graph structures as containers decrease in size, showcasing an element of mathematical beauty.
-Moreover, this has real-world applications: in network analysis, it suggests that influence or importance may not rest with a single node but instead within a clique or tightly connected group, offering new perspectives for studying collaborative dynamics and group influence.
-The identification of regularity was key to understanding why multiple nodes’ degrees can double in certain graphs.
-This concept offers a powerful explanation for the phenomenon.
-The regularity is the key factor that helps explain why multiple nodes can have their degrees double, rather than the search for multiple nodes being inherently complex.
+The discovery of multiple degree-doubling nodes lifts the Seymour Second Neighborhood Conjecture from a mere question of existence to one in abundance.
+There were many who doubted that one even existed.
+Now there are many.
+Second is the way we find it. It’s the presence of an oriented complete graph as these containers decrease in size. There’s an element of mathematical beauty there.
+Then third we can look at the applications.
+We may be looking for the person who has the most influence, but this is saying that there may not be just one person there may be a whole set of people, a tightly connected group.
+Identifying regularity was key to understanding why multiple nodes’ degrees can double in certain graphs.
+The regularity is the key factor that helps explain why multiple nodes can have their degrees double.
 The regularity is the central feature that ties everything together.
@@
 When a graph’s structure is highly regular, particularly in terms of interior and exterior degrees, it becomes easier to predict and verify that certain structural properties (like degree doubling) will occur at specific stages.
-Regularity implies that as nodes increase in distance from the starting node, their connectivity and interior degrees will follow a pattern.
-This pattern is key to understanding why degree doubling happens more naturally and why multiple nodes can experience it simultaneously, if the conditions allow.
 10.3 Decreasing Sequence Property Path Algorithm
-We now present a precise algorithm to identify nodes whose out-degree doubles in the square of an oriented graph, as per Seymour’s Second Neighborhood Conjecture (SSNC).
-This algorithm synthesizes results from the lemmas, theorem, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
+We will now give a pseudocode algorithm to identify the node whose out-degree doubles in the square of an oriented graph.
+This is the culmination of the lemmas, theorems, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
   • The containers reach a size two causing a node in the previous container to have its degree double, as stated in Theorem 10.2 (Degree Doubling by Exterior Decreasing).
@@
 11 Conclusion and Future Work
-Solving a conjecture with such a long and debated history requires thorough validation.
-This paper represents a comprehensive effort to meet the rigorous standards of mathematical proof, supported by tools designed to aid understanding.
-Acknowledging that such results are subject to detailed scrutiny, this process is viewed as an essential step in establishing confidence and recognition within the mathematical community.
-The GLOVER structure—short for Graph Level Order—represents a significant advancement in addressing long-standing open problems in graph theory, such as the Seymour conjecture, which remained unresolved for 34 years.
-While not all oriented graphs are inherently well-ordered, the GLOVER structure reveals that, within the context of the Seymour conjecture, oriented graphs become effectively well-ordered.
-This property arises because any deviation from such an ordering would lead to a degree-doubling effect, rendering the conjecture invalid.
-By leveraging this induced well-ordering, the GLOVER structure enables tools like mathematical induction to be applied effectively, facilitating a solution to the Seymour conjecture.
-Beyond this specific case, the framework provides a foundation for exploring dual metrics in graph theory, such as prioritizing distance and degree, and opens the door to applications in clustering, ranking, and multi-metric optimization.
-The GLOVER structure emerges as a groundbreaking framework for addressing problems that rely on dual metrics, offering a versatile tool for organizing and analyzing data across theoretical and practical contexts.
-While its utility was demonstrated here in the context of the Seymour conjecture—where the dual metrics of distance and degree played a pivotal role—it is important to recognize that the structure’s applications extend far beyond regularity or degree-based problems.
-At its core, the GLOVER structure’s ability to prioritize and organize relationships based on dual metrics allows it to address challenges that require both exhaustive and hierarchical searches.
-Unlike local search methods traditionally constrained by single metrics, the GLOVER approach enables partitioning and clustering in ways that preserve relationships and provide deeper insights.
-For instance, applications in clustering, ranking, and network analysis—whether in social networks, employee rankings, or other multivariate datasets—can benefit from this structure.
-Looking ahead, the GLOVER structure challenges us to investigate its applicability to other open problems in graph theory and computer science.
-The approach may uncover new insights in contexts where dual or multi-metric prioritization is key, highlighting the broader potential of this versatile framework.
-Future research and development should explore these broader applications.
-Questions of complexity, scalability, and adaptability to non-graph-based datasets remain promising areas for investigation.
+The Seymour Second Neighborhood Conjecture started as a question in 1990.
+Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers.
+It has been talked about in books.
+on message boards, in classrooms, and at conferences.
+This problem has a long history.
+With a solution now in hand, its time to begin looking forward.
+The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph.
+Here, the nodes are people, where noone follows back.
+The same question is of interest, is there anyone whose degree doubles, except we know the answer is yes.
+We just need to implement the algorithm to find the person
+We could have the graph represent epidemiology.
+Here the nodes be infected people, and the arcs would be one way infections.
+The question of interest would be can we find the person who is a super-spreader.
+Secondly, we introduced a data structure, Graph Level Order-GLOVER.
+This data structure allowed us to get a lot done on the conjecture.
+Many of the things inherent to SSNC made it easier for GLOVER - the well ordering and the dual metrics of distance and degrees.
+Future research and development should explore questions of complexity, scalability.
 Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
@@
 14 Appendix 2: Applicability of the Graph Level Order
-Traditionally, a total order on a set is a linear, one-dimensional comparison between elements.
-The key properties of total orders—reflexivity, transitivity, and antisymmetry—give us a clear, unambiguous ranking where every pair of elements is comparable.
-Graph Level Orders (GLOVER) extend this concept to n-dimensions within a leveled framework.
-GLOVER accomplishes this through the use of containers that not only arrange elements in sequence but also enable multiple levels of organization.
-Elements within each level maintain an internal order, while the containers themselves are comparable across ranked layers.
-This structure allows for sophisticated comparisons both within a single level and between levels, making GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
-The examples provided in this paper illustrate concrete cases of GLOVER.
-However, it’s the abstract nature of GLOVER that sets it apart as a transformative mathematical concept.
-This structure allows us to conceptualize any set of ranked organized objects—not just graphs with regularly structured nodes, but also diverse entities, such as sets with multiple attributes or time-sequenced data.
-Within discrete mathematics, the ability to impose a multi-dimensional total order provides a richer understanding of relationships among elements.
-The order goes beyond pairwise comparisons to establish a multi-layered framework capable of supporting complex operations.
-The true strength of GLOVER lies in its flexibility.
-It accommodates inherent ordering within single-level sets while also adapting to more intricate structures.
-By generalizing beyond domain-specific contexts like graph theory, GLOVER becomes a universal framework for imposing total order on any ranked structure, from simple ordered pairs to multi-dimensional data sets in fields like machine learning or clustering rankings.
-The GLOVER structure also uniquely supports induction in graph-based proofs due to its total order, providing a new way to approach graph theory problems.
-Unlike traditional graph structures, which typically do not allow for induction, the total ordering of nodes and containers in GLOVER makes it possible to apply inductive reasoning effectively.
-This expanded understanding of total orders paves the way for developing algorithms, proofs, and applications capable of handling complex, scalable data structures.
-GLOVER is not merely a form of ordering; it is a system of relationships that facilitates the manipulation, comparison, and analysis of multi-dimensional data through a clear and consistent layering.
-The GLOVER framework, as initially defined, is tailored to address the specific requirements of the Seymour Second Neighborhood Conjecture.
-Its rigid structure-classifying arcs either as interior or exterior, and strictly prohibiting skipping levels-ensures clarity and sufficiency for analyzing second neighborhoods, However the framework can be extended to accomodate more general problems, including those with nested containers or more flexible neighbor relationships .
-Nested containers introduce a hierarchy within individual containers, grouping nodes based on finer structural or relational properties.
-This extension enables the framework to address problems where additional levels of granularity within the same partition group are required.
-Definition 14.1.
-A nested container $C_{i,j}$ is a subset of a container $C_{i}$ defined by additional criteria beyond the qualifying metric that determined it into $C_{i}$.
-For example, nodes in $C_{i,j}$ might share specific adjacency patterns or connectivity property properties.
-Use Cases:
-  • Clustering Problems: Nested Containers can represent hierarchal clustering within each distance group, allowing comparisons between subgroups.
-  • Weighted Graphs: In graphs with edge weights, nested containers could group nodes based on weight thresholds within the same distance level.
-Example:
-Suppose $C_{2}$ contains nodes $x,y,z$ all at distance 2 from $v_{0}$. If $x$ and $y$ are connected but not $z$, $x$ and $y$ might form a nested container $C_{2,1}$, while $z$ forms $C_{2,2}$
-In certain problems, arcs may not connect nodes in non-adjacent containers, effectively skipping levels in the distance hierarchy.
-While the SSNC prohibits such arcs to maintain a second-neighbor focus, other graph probems may allow or even require these connections.
-  • Extension - Relax the exterior neighbors to include $C_{j})$ for $j>i+1$, introducing a new category of arcs, long-distance exterior arcs.
-  • Use Cases -
-    • Network Design - In communication networks, skipping levels might model direct long-range connections (for example satellite links).
-    • Flow Problems - In transportation or flow networks, skipping levels could represent shortcuts or direct transfers.
-  • Impact on GLOVER Properties - Allowing for skipping levels introducing new challenges
-    • Order Maintenance - Total order within containers must be preserved despite the additional complexity of the long-distance connections.
-    • New Arc Classifications - Introduce distinctions like short range exterior arcs (j = i + 1) and long range exterior arcs (j ¿ i + 1).
-Exterior neighbors, as defined in the SSNC, connect nodes in adjacent containers.
-Expanding this definition allows for more flexible interactions between containers, broadening the applicability of the GLOVER framework.
-  • Generalized Exterior Neighbors - for $x\in C_{i}$, an exterior neighbor $y$ can belong to any $j>i$. The specific range of $j$ will depend on the problem context.
-  • Example Problem: Consider a graph where connections beyond second neighborhoods are of interest.
-Exterior neighbors could include nodes up to $k$-steps away, accomodating problems requiring $k$-neighborhood analysis.
-While these extensions offer exciting generalizations, the SSNC’s specific focus on second neighbors necessittates a stricter interpretation of the GLOVER framework.
-  • No skipping levels - Ensures that second neighbors are strictly defined nodes in $C_{2}$.
-  • Rigid container definitions - prevents ambiguity in arc classifications or neighbor interactions.
-  • Limited use of nested containers - The SSNC does not require additional granularity within distance groups, making nested containers optional for this specific problem.
-Example 14.1.
-Graph Theory - Seymour Second Neighborhood Conjecture: Graphs, which are normally seen as an unordered arrangement of nodes and edges, were able to be ordered with the GLOVER.
-We initially chose a minimum degree node and partitioned the nodes of the graph into containers based on their proximity to this node.
-Partitioning nodes by distance allowed for definition of the interior degree and exterior degree which were very important to showing that a node’s overall degree would double.
-Example 14.2.
-Engineering - Dynamic Scheduling: In fields such as manufacturing, task scheduling often involves prioritizing independent tasks.
-With GLOVER, tasks are grouped into containers based on dependency levels and ordered according to priority.
-For example, preparatory tasks are prioritized in early containers, while finishing tasks are sequenced last, minimizing delays and optimizing task flow.
-Partitioning groups tasks by dependency, while ordering between containers ensures essential tasks are completed first.
-Within each container, tasks can be further organized by resource requirements, promoting efficiency at each priority level.
-Edges introduced at the container level to represent this dependency.
-Interior edges would show common dependencies between tasks in a common container, whereas exterior edges represent tasks of different priorities with common dependencies.
-From this we can use graph theory concepts like centrality to better understand things like important tasks, important dependencies, how the containers (task tiers relate to one another) etc.
-Example 14.3.
-Engineering - Resource Management: In network resource allocation or energy grid management, nodes can be clustered by proximity to demand or criticality and ordered to prioritize high-need resources.
-This ordering ensures that essential nodes remain accessible, while resource usage remains balanced across the network.
-Partitioning groups resources based on proximity to critical nodes, while ordering between containers directs allocations toward high-demand areas.
-Within each container, resources are accessed in order of current availability, optimizing responsiveness to fluctuating demands.
-Example 14.4.
-Engineering - Project Management: In complex projects involving multiple teams and dependencies, tasks can be clustered by project phase and ordered sequentially to prevent bottlenecks.
-Foundational tasks, such as design and planning, are prioritized early, while execution tasks follow, ensuring smooth progression through each phase.
-Partitioning groups tasks by project phase, while ordering between containers facilitates the transition of tasks from one stage to the next.
-This ranked order remains stable despite potential project changes, enabling efficient coordination across large teams.
-Example 14.5.
-Cybersecurity - Threat Detection and Prioritization: In cybersecurity, ranked structures are often essential for organizing and prioritizing threats.
-Using an GLOVER framework, threats can be grouped into containers based on their urgency or potential impact.
-For example, critical threats with known vulnerabilities might be prioritized in early containers, while less immediate risks or less probable attack vectors are sequenced into subsequent layers.
-This ranked ordering helps ensure that the most pressing issues are handled first, thereby minimizing potential breaches or disruptions.
-Partitioning into containers groups threats by their severity and probability, while ordering between containers facilitates a systematic approach to addressing risks from the highest priority down.
-Within each container, threats are further organized based on attributes like type, location in the network, or historical occurrence, allowing security teams to deploy resources effectively.
-This layered approach enhances the overall defense strategy by ensuring structured response and continuous monitoring across multiple risk levels.
-Example 14.6.
-Natural Language Processing - Ranked Text Classification: In NLP, ranked total order graphs can improve text classification by organizing data through multi-level categories.
-For instance, documents can be grouped into containers based on broad topics (e.g., news, entertainment, scientific research) with each container further ordered by subtopics like sports or politics under news, and biology or physics under science.
-This layered structure allows models to systematically refine classifications, progressing from general to specific categories, improving the accuracy and interpretability of the classification model.
-Partitioning into containers groups text by broad categories, while ordering between containers prioritizes classification tasks from general to specific.
-Within each container, documents can be ordered by relevance, keywords, or thematic clustering, which supports tasks like topic modeling or intent recognition.
-This approach provides a clear and adaptable framework for complex, multi-layered categorization, enhancing the model’s ability to handle diverse and evolving datasets.
-The versatility of the Graphs Level Order (GLOVER) structure hints at its applicability beyond this proof of the Seymour Second Neighborhood Conjecture.
-GLOVER has the potential to address several open problems and conjectures within discrete mathematics and graph theory, where ranked or layered ordering of elements can provide new insights or simplify complex relationships.
-Future work may reveal more connections between GLOVER and these longstanding challenges, offering new paths for research and exploration.
+A total order is a linear ranking.
+There are no containers or buckets to make extra comparisons.
+Everything is sorted.
+That may not be what we desite though.
+Suppose we have two metrics, like age and income.
+Those two metrics cannot be compared on the same scale.
+They’d need to be grouped by one metric and compared by the other.
+The Graph Level Order (GLOVER) is able to do these multi level comparisons.
+GLOVER does this by the usage of containers.
+These containers arrange elements in sequence.
+Elements within each level maintain an internal order.
+The containers themselves are comparable across ranked layers.
+This structure allows for multi-level comparisons.
+These can happen both within a single level and between levels.
+This makes GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
+Just as the SSNC was proved with the GLOVER data structure, there are more proofs and algorithms that this method is capable of handling.
 15 Appendix 3: Visualization and Accessibility of Oriented Graphs
-In exploring oriented graphs, it became evident that fully grasping these structures often requires more than static illustrations in a paper can offer.
-While oriented graphs can be defined and studied abstractly, their complexity and behavior are best understood through dynamic, interactive representations.
-The Seymour Second Neighborhood Conjecture, in particular, necessitated defining paths, locating squares, and identifying degree-doubling nodes—tasks that involved intricate arc counting and relationships that benefit from visualization.
-This realization motivated the creation of a complementary website.
-The site serves as a tool to enhance the accessibility of key definitions, lemmas, and proofs by providing visualizations and interactive examples.
-Our goal is to bridge the gap between abstract theory and intuitive understanding, making the intricate aspects of the conjecture approachable for a wider audience.
-For example, consider the concept of a degree-doubling node or a lemma describing staircase-like structures in oriented graphs.
-While the paper describes these ideas in formal mathematical terms, the website illustrates them dynamically.
-This approach allows readers to ”see” the behavior and implications of these structures, offering a visual intuition that complements the formal reasoning presented in the paper.
-This methodology draws inspiration from James Robert Brown’s insights in Philosophy of Mathematics, where he emphasizes the dual role of visualization in mathematics—as both a powerful tool for intuition and a potential pitfall if over-relied upon.
-By carefully balancing these aspects, we aim to enhance comprehension without compromising the rigor essential to mathematical proofs.
-The website also provides an opportunity to include numerous examples and extended visual explanations that are not feasible within the space constraints of a traditional journal format.
-Through this medium, we invite readers to engage more interactively with the conjecture and its underlying theories, fostering a deeper understanding and encouraging constructive feedback.
-Built using technologies like D3.js, JavaScript, and PHP, the website features dynamic visualizations of proof steps, interactive diagrams, and supplementary explanations.
-By demonstrating the synergy between mathematical results and computational tools, we aim to make these concepts not only accessible but also engaging for a broader audience.
+When we take a course on graph theory, oriented graphs are probably not on the course load.
+Second neighbors or squares of graphs are probably not either.
+To fully grasp this problem though we need to be able to delve into these definitions through visualizations.
+For example, research on the Seymour Second Neighborhood Conjecture led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs.
+The paper addressed how these terms were intertwined with one another.
+One example is where the paper says that there is a monotonically decreasing sequence.
+James Robert Brown’s book Philosophy of Mathematics speaks of the importance of visualizations in mathematics.
+Proofs are still the foundation, but sometimes visualizations can serve as an inspiration or a spark for those proofs.
+This realization motivated the creation of a complimentary website.
+The site serves as an interactive tool to understand research.
+It also has things like mathjax installed to read lemmas, definitions and proofs from this paper.
+Where possible, things like HTML canvas and JavaScript D3.js told are used to visualize these concepts and help improve understanding.
 Readers are encouraged to explore the website for additional insights and examples at:
 https://glovermethod.com
-While the website is already live and functional, it is an evolving project. Additional examples, applications, and refinements are being added continuously.
-Readers are invited to explore the website and are encouraged to provide feedback on any aspects of the site or the content itself, including suggestions for improvements, corrections, or additional examples.
-The goal is to create a resource that not only supports this paper but also serves as a tool for further exploration and discussion in the field of oriented graph theory, the Seymour Second Neighborhood Conjecture, and the Graph Level Order (GLOVER) data structure.
-Thank you for your engagement and contributions to this ongoing effort.
+This is an evolving project. Additional examples and applications will continuously be added
+Readers are invited to explore the website. Please provide feedback on any aspects of the site or the content for improvements.
+Our goal here was to help users connect with the Seymour Second Neighborhood Conjecture.
+Hopefully, this tool not only serves as support for this paper, but as an educational resource in the fields of graph theory and computer science.
+Thank you for your engagement

### v4

@@
 This paper first presents a data structure GLOVER (Graph Level Order) which is a partitioning of the nodes into a total order of neighborhoods.
-This data structure allows us to construct a decreasing sequence of subsets, ensuring that no counterexample can exist.
+This data structure allows us to introduce a total order on oriented graphs and construct a decreasing sequence of subsets.
+This decreasing sequence of subsets is what ensures that no counterexample can exist.
 Further, this algorithmic approach allows us to know where the node is.
@@
 Then $v_{0}$ is connected to some node $v_{1}\in G$.
-Since $v_{0}$ is a minimum out-degree node, we know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$.
-Because $G$ is an oriented graph and $(v_{0},v_{1})\in G$, it follows that $(v_{1},v_{0})\notin G$.
-Therefore, $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$.
+We know that $d^{+}(v_{1})\geq d^{+}(v_{0})=1$.
+This node $v_{1}$ must be connected to another node $v_{2}\neq v_{0}$.
 This means that $(v_{1},v_{2})\in G$.
@@
 Then $v_{0}$ is connected to two nodes, $v_{1}$ and $v_{2}\in G$.
-Since $v_{0}$ has the minimum out-degree, we know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$.
-As $G$ is an oriented graph, $(v_{1},v_{0})$ and $(v_{2},v_{0})$ are not in $G$.
-Consider the subgraph $G[v_{1},v_{2}]$, induced on the nodes ${v_{1},v_{2}}$.
-There is at most one directed edge between $v_{1}$ and $v_{2}$ in $G$, meaning one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes.
+We know $d^{+}(v_{1})\geq 2$ and $d^{+}(v_{2})\geq 2$.
+Consider the induced subgraph $G[v_{1},v_{2}]$.
+There is at most one directed edge between $v_{1}$ and $v_{2}$ in this induced subgraph.
+This means that one of these nodes has an out-degree of at least 2 outside $G[v_{1},v_{2}]$, connecting to other nodes.
 Let these nodes be $v_{3}$ and $v_{4}$.
-Since $G$ is oriented, $v_{0}$ cannot be one of these nodes.
-Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4, thus doubling its out-degree.
+The node $v_{0}$ cannot be one of these nodes.
+Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4.
+Thus, its out-degree will double.
 Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
@@
 Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$.
-Consider the subgraph $G[v_{1},v_{2},v_{3}]$, which is the subgraph induced by the set $\{v_{1},v_{2},v_{3}\}$.
-Now, assume that one of $v_{1},v_{2},v_{3}$, say $v_{1}$, has an out-degree of 0 within the subgraph $G[v_{1},v_{2},v_{3}]$.
-This implies that $v_{1}$ must connect to nodes outside of $\{v_{1},v_{2},v_{3}\}$, as $v_{1}$ has no out-neighbors within the induced subgraph.
-Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in G.
-Thus, $v_{0}$’s out-degree will at least double in $G^{2}$. ∎
+Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$.
+Now, assume that $v_{1}$, has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$.
+This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
+Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
 Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$ shares no neighbors with $v_{0}$.
@@
 Proof.
-Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ Since $v_{0}$ is a minimum out-degree node, we know $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$.
-Let $G[v_{1},v_{2},v_{3}]$ represent the subgraph induced by $v_{1},v_{2},v_{3}$.
+Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$.
+Let $G[v_{1},v_{2},v_{3}]$ represent the induced subgraph by $v_{1},v_{2},v_{3}$.
 We can reason that if any of $v_{1},$ $v_{2},$ or $v_{3}$ has an out-degree greater than 3, the out-degree of $v_{0}$ will at least double in $G^{2}$.
-Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$, and by Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0) , none of $v_{1}$,$v_{2},$ or $v_{3}$ have out-degree 0 in $G[v_{1},v_{2},v_{3}]$.
+Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$.
 Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$.
-Therefore, assume they connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
+Therefore, assume they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
 This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree $<$ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case.
@@
 As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
-Thus, either $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}$. ∎
 Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) answers a question that we will come back to throughout this paper - this is an exhaustive search algorithm.
@@
 Definition 5.2.
-For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$,where $N^{+}(x)$ denotes the set of out-neighbors of $x$, and $N^{++}(x)$ denotes the set of out-neighbors of the nodes in $N^{+}(x)$.
+For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$.
 We have now introduced two concepts: the minimum degree node and the decreasing sequence property (DSP).
@@
 Total Order Induced by Containers Relative to the Minimum Degree Node
-Let $G$ be a graph with a minimum degree node $v_{0}$, where all nodes satisfy the Decreasing Sequence Property (DSP).
-Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ relative to $v_{0}$, where each container $C_{i}$ consists of nodes ordered by their adjacency and connectivity properties with respect to $v_{0}$ and preceding containers.
-This structure imposes a total order on the nodes in $G$, where every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$, establishing a ranking aligned with the DSP.
+Let $G$ be a graph with a minimum degree node $v_{0}$.
+Further, let all nodes in $G$ satisfy the Decreasing Sequence Property (DSP).
+Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ ordered by distance relative to $v_{0}$.
+Each container $C_{i}$ will consist of nodes at that respective distance from $v_{0}$.
+Then this structure imposes a total order on the nodes in $G$ i.e.
+every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
 Proof.
@@
 Degree Doubling by Container Density In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},C_{2},...,C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
-This collision can manifest in two forms: either the size of the container reaches a point where it can no longer support the increasing interior degrees (resulting in nodes that cannot maintain their minimum degree requirements), or smaller containers force a doubling of degrees that contradicts the capacity for interior arcs.
-The structure of the graph, particularly in small sizes, inherently limits these growth dynamics, leading to contradictions that must be addressed.
+There are two possibilities:
+  1. 1.
+    As the containers shrink in size and the interior degrees grow in size, we reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
+  2. 2.
+    For small minimum degree nodes $v_{0}$, The container’s shrinkage is not fast enough to cause this collision.
 Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes.
@@
 Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbor neighbors represented the size of the containers.
+In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur.
+To understand the reasoning behind this, understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1.
+The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$.
+So for small graphs, we are just not going out far enough to catch up.
+For larger graphs though, we do start to see this.
 Corollary 10.2.
@@
 For even values of $k$, such as $k=4$ or $k=6$, the symmetry and density required for a complete graph $K_{n}$ cannot form directly at any container level.
-This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the structural requirements for even $n$ are never satisfied in full.
-Instead, the structure transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
+This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the requirements for even $n$ are never satisfied in full.
+Instead, it transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
 The relationship $i=\frac{k-1}{2}$ reflects the interplay between the minimum degree $k$ and the growth of density in containers.
 This relationship confirms that the GLOVER structure consistently generates complete graphs $K_{2i+1}$ in odd-numbered increments, with no direct occurrence for even $n$.
-This emergence of orientations of complete graphs as containers decrease in size adds a layer of symmetry and beauty to the Seymour Second Neighborhood Conjecture.
+This emergence of orientations of complete graphs as containers decrease in size adds a layer of beauty to the Seymour Second Neighborhood Conjecture.
 There was never a question of these arising.
@@
 We may be looking for the person who has the most influence, but this is saying that there may not be just one person there may be a whole set of people, a tightly connected group.
-Identifying regularity was key to understanding why multiple nodes’ degrees can double in certain graphs.
-The regularity is the key factor that helps explain why multiple nodes can have their degrees double.
-The regularity is the central feature that ties everything together.
-Degree doubling is essentially a consequence of the underlying regularity.
-When a graph’s structure is highly regular, particularly in terms of interior and exterior degrees, it becomes easier to predict and verify that certain structural properties (like degree doubling) will occur at specific stages.
 10.3 Decreasing Sequence Property Path Algorithm
@@
 11 Conclusion and Future Work
+In this work, we explored the Seymour Second Neighborhood Conjecture in oriented graphs.
+We did so through the lenses of graph theory, combinatorics, algorithms and data structures.
+This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide and conquer paradigm.
+One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph.
+This effectively solves the Seymour Second Neighborhood Conjecture with an answer in the affirmative.
+The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree doubling node was in the graph.
+Instead the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
+This approach introduced a new perspective on the Seymour Second Neighborhood Conjecture.
+It was no longer framed as merely a graph theory problem.
+By utilizing the divide and conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effective be treated with computer science techniques.
+This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists.
+Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
 The Seymour Second Neighborhood Conjecture started as a question in 1990.
@@
 Here, the nodes are people, where no one follows back.
-The same question is of interest, is there anyone whose degree doubles, except we know the answer is yes.
-We just need to implement the algorithm to find the person
+Now instead of wondering if there is a person whose sphere of influence doubles, we would simply be searching for that person.
+We just need to implement the algorithm to find the person.
 We could have the graph represent epidemiology.
 Here the nodes be infected people, and the arcs would be one way infections.
-The question of interest would be can we find the person who is a super-spreader.
+The question of interest would be can we implement the algorithm to find the person who is a super-spreader, and do so quickly?
+There are many more real life examples where we can apply this problem and this algorithm.
+There are fields like security, telecommunications, urban planning, neural networks.
+There are too many to name in this short paper.
 Secondly, we introduced a data structure, Graph Level Order-GLOVER.
@@
 14 Appendix 2: Applicability of the Graph Level Order
-A total order is a linear ranking.
-There are no containers or buckets to make extra comparisons.
-Everything is sorted.
-That may not be what we desite though.
-Suppose we have two metrics, like age and income.
-Those two metrics cannot be compared on the same scale.
-They’d need to be grouped by one metric and compared by the other.
-The Graph Level Order (GLOVER) is able to do these multi level comparisons.
-GLOVER does this by the usage of containers.
-These containers arrange elements in sequence.
-Elements within each level maintain an internal order.
-The containers themselves are comparable across ranked layers.
-This structure allows for multi-level comparisons.
-These can happen both within a single level and between levels.
-This makes GLOVER a higher-level mathematical framework suitable for multi-layered, multi-dimensional data.
-Just as the SSNC was proved with the GLOVER data structure, there are more proofs and algorithms that this method is capable of handling.
+The deciding factor that held the algorithm together and helped solve the Seymour Second Neighborhood Conjecture was the Graph Level Order (GLOVER) data structure.
+This research did not begin as an investigation into data structures.
+Instead, a series of operations were conducted on the oriented graphs - partitioning, ordering those partitions, adding interior and exterior arcs, until what was left was an ordering of both the partition and nodes within the partition.
+Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric.
+Unlike the Seymour Second Neighborhood Conjecture, exterior arcs do not always have to only be defined to the next container.
+That was a problem specific definition.
+Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond.
+By introducing containers, another way of doing induction on graphs has been introduced.
+The concept of divide and conquer, which has dominated much of computer science, shows up in this paper as well.
+We are able to treat these containers independent of one another and solve their problems locally before bringing them back to the global problem.
+This shows the how versatile the GLOVER data structure is moving forward.
+What makes GLOVER special are things like total order on containers.
+Now there is a two way ranking of nodes, as opposed to traditional one way rankings like lexicographical sorting.
+This agrees with many real world systems where we have dual metrics, that are often competing, like price and performance.
+This is hard to measure on a single scale.
+Next these nodes are not just treated like clusters inside these containers, but like nodes in graph theory.
+The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth first search, or centrality to be called.
+What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
+This is still a new data structure, so we are learning a lot about it.
+There are many applications that have not been tried yet.
+However, since it brings the strengths of a total order and a graph into one unified structure, it should be able to model complex relationships.
+This should include domains of machine learning like clustering, dynamic scheduling, resource management, project management, and natural language processing.
 15 Appendix 3: Visualization and Accessibility of Oriented Graphs

### v6

@@
 Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph.
-This paper first presents a data structure GLOVER (Graph Level Order) which is a partitioning of the nodes into a total order of neighborhoods.
+This paper first presents a data structure, GLOVER (Graph Level Order), which is a partitioning of the nodes into a total order of neighborhoods.
 This data structure allows us to introduce a total order on oriented graphs and construct a decreasing sequence of subsets.
 This decreasing sequence of subsets is what ensures that no counterexample can exist.
-Further, this algorithmic approach allows us to know where the node is.
+Further, this algorithmic approach lets us know where the node is.
 Later findings extend the results to multiple nodes.
@@
 In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture.
-This was an extension of the work done by Bolter et al [2] where their work confirms that almost all orientations of $G(n,p)$ are satisfy the Seymour Second Neighborhood Conjecture.
+This was an extension of the work done by Bolter et al.
+[2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the Seymour Second Neighborhood Conjecture.
 The topic of $m$-free graphs is of particular interest to this paper because Daamouch’s [7] work is very similar to what we see here.
@@
 We will analyze how their out-degree could fail to double.
-We will then extend out algorithm to other nodes based on their distance from the minimum out-degree node.
+We will then extend our algorithm to other nodes based on their distance from the minimum out-degree node.
 The paper is organized as follows: Section 2 begins with some common terminology used in graph theory.
@@
 In section 4 we look at some small examples of this style of thinking.
-Section 5 introduces the terminology and identifying the conditions under which their out-degree might fail to double.
+Section 5 introduces the terminology and identifies the conditions under which their out-degree might fail to double.
 Section 6 examines the back arcs and their place in oriented graphs.
-Section 7 looks at the conditions which would cause an interior degree to double, and its importance to the Seymour Second Neighborhood Conjecture.
+Section 7 looks at the conditions that would cause an interior degree to double and its importance to the Seymour Second Neighborhood Conjecture.
 Section 8 further looks at containers under the context of density constraints.
-In Section 9 considers transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm.
-Section 10 presents this process the full Decreasing Sequence Property algorithm as well as the theorem stating it’s solution to the Conjecture.
+In Section 9, we consider transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm.
+Section 10 presents this process, the full Decreasing Sequence Property algorithm, as well as the theorem stating its solution to the Conjecture.
 2 Graph Theory Terminology
-A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(x,x)$ where $x$ is a node in $G$) and no symmetric arcs, that is no arcs of the form $(x,y)$ and $(y,x)$ where $x$ and $y$ are nodes in $G$.
-The square of a graph $G$, denoted $G^{2}$, consists of the same node set as $G$ and an arcs exists from node $x$ to node $z$ if there is a node $y$ such that both arcs $(x,y)$ and $(y,z)$ exist in $G$ as well as the arcs already in $G$.
+A directed graph $G$ is called oriented if it has no self-loops (i.e., no arcs of the form $(x,x)$ where $x$ is a node in $G$) and no symmetric arcs, that is, no arcs of the form $(x,y)$ and $(y,x)$ where $x$ and $y$ are nodes in $G$.
+The square of a graph $G$, denoted $G^{2}$, consists of the same node set as $G$, and an arc exists from node $x$ to node $z$ if there is a node $y$ such that both arcs $(x,y)$ and $(y,z)$ exist in $G$ as well as the arcs already in $G$.
 Essentially, $G^{2}$ captures two-step connections from $x$ to $z$.
@@
 These are the nodes that $v$ has arcs pointing to.
-Similarly, $N^{++}(v)$ refers to the second neighbors of $v$ that is, nodes that are two steps away from $v$ in $G$.
+Similarly, $N^{++}(v)$ refers to the second neighbors of $v$, that is, nodes that are two steps away from $v$ in $G$.
 Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices.
-The induced subgraph $G[S]$ is the graph whose vertices are the set $S$, and whose edge set consists of all edges in $E$ that have both endpoints in $S$.
+The induced subgraph $G[S]$ is the graph whose vertices are the set $S$ and whose edge set consists of all edges in $E$ that have both endpoints in $S$.
 For any node $v$, we refer to the neighbor-induced subgraph on $N^{+}(v)$, consisting of $v$’s out-neighbors.
 Let $u,v$ be nodes in $G$.
-The distance between $dist(u,v)$, is $k$ if the shortest path from $u$ to $v$ has length $k$.
+The distance between $dist(u,v)$ is $k$ if the shortest path from $u$ to $v$ has length $k$.
 Example 2.1.
@@
 Previously used in [9]
+Conjecture 2.1.
+(Square version of Seymour’s Second Neighborhood Conjecture). Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
+Now that we have two different conjectures, we can investigate how they relate to one another.
+Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods, while conjecture 2.1 relates this same concept to the square graph.
+The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first.
+By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph.
+This is why these two conjectures are seen as identical.
+We will focus our attention on Conjecture 1.1 for the reason that the conjecture 2.1 requires two different graphs - the original graph and the square graph.
+While the construction of the square graph is possible, we have found that it is not as helpful to our research at this time.
+Instead, focusing on the neighborhood analysis brought many insights and characteristics that helped to develop our algorithm.
+Next, we turn our attention to the term degree-doubling nodes.
+By this term, we aim to get to the heart of conjecture 1.1.
+That conjecture is looking for a node whose degree at least doubles, and we are doing so by looking at the first and second neighborhoods of nodes and finding one whose second neighborhood is at least as large as the first, as the conjecture asks for.
+A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree.
+We could equivalently search for a Seymour vertex or a degree-doubling node.
 3 Approach and Motivation
@@
 Instead, think of it as putting a box around the partition.
-Now the nodes inside the box are can work with the other nodes inside the box and solve their problem independent of the other partitions.
-These partitions are still connected to one another, but thinking of them in this way, allows us to do more.
-This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecure.
-Many of these approaxhes were conceptual or theoretical, saying only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper.
-Secondly, most of these paper focused on local search techniques that statrted at or near the desired degree doubling node.
-That presents a problem because oriented graphs can be varying in nature, This is the first exhaustive search technique on the SSNC problem.
+Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions.
+These partitions are still connected to one another, but thinking of them in this way allows us to do more.
+This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecture.
+Many of these previous approaches were conceptual or theoretical.
+They would say only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper.
+Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node.
+That presents a problem because oriented graphs can be varying in nature.
+This is the first exhaustive search technique on the SSNC problem.
+When speaking of an exhaustive search algorithm, complexity becomes a factor and questions of how the algorithms will handle large graphs compared to smaller graphs.
+This should not be a problem as we will see in the following sections; the slowest part of this algorithm is building the data structure.
+That will require partitioning the nodes and edges of the graph, in a worst case complexity of $O(V+E)$.
+By partitioning the graph, even larger graphs, often with higher minimum degree nodes and thus more partitions, are able to be broken into sub-problems and solved independent of one another.
+These subproblems are each their own instances of the Seymour Second Neighborhood Conjecture, just a localized version.
+What the partition is doing is breaking a node’s individual degree into two parts-the degree within the container and the degree to the next container.
+The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
+We still need to show that this is a formidable foundation.
 We have laid the groundwork for our algorithm to tackle the oriented graphs.
@@
 The work in this section agrees with what was already done by [11] and is not meant to serve as original content.
-It does however, provide insight into our reasoning.
+It does, however, provide insight into our reasoning.
 Example 4.1.
@@
 Lemma 4.1.
-Minimum Out-Degree $<$ 3 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3, then $v_{0}$’s out-degree will at least double in $G^{2}$.
+Minimum Out-Degree $<$ 3 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double in $G^{2}$.
 Proof.
@@
 Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4.
-Thus, its out-degree will double.
+Thus, it’s out-degree will double.
 Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
 Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs.
-We are learning that, in order to construct our counter-example, we need out minimum degree to be at least 3.
+We are learning that, in order to construct our counterexample, we need our minimum degree to be at least 3.
 This will also place that same minimum value restriction on every other node in the graph.
@@
 Lemma 4.2.
-Minimum Out-Degree 3 with Neighbor Out-Degree 0 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3, and at least one of its neighbors has out-degree 0 in the induced subgraph of its neighbors, then the out-degree of $v_{0}$ will at least double in $G^{2}$.
+Minimum Out-Degree 3 with Neighbor Out-Degree 0 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3.
+Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0; then the out-degree of $v_{0}$ will at least double in $G^{2}$.
 Proof.
-Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, $v_{3}\in V(G)$.
+Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, and $v_{3}\in V(G)$.
 Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$.
-Now, assume that $v_{1}$, has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$.
+Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$.
 This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
-Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$, As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
-Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$ shares no neighbors with $v_{0}$.
+Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$.
+As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
+Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$.
 This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
@@
 Thus, in $G^{2}$, the degree of $v_{0}$ will double.
+Minimum Out-Degree 3 with Neighbors 1
 Lemma 4.3.
-Minimum Out-Degree 3 with Neighbors 1 If a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3 and all of its neighbors have an out-degree of 1 in the induced subgraph of its neighbors, then $v_{0}$ or one of its neighbors will have its out-degree at least double in $G^{2}.$
+Assume that in an oriented graph, the minimal out-degree node $v_{0}$ has an out-degree of 3.
+Additionally, suppose that in the induced subgraph of its neighbors, every neighbor has an out-degree of 1.
+The out-degree of $v_{0}$ or one of its neighbors will then at least double in $G^{2}.$.
 Proof.
-Assume $d^{+}(v_{0})=3,$ so $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$.
+Assume $d^{+}(v_{0})=3$.
+So $v_{0}$ is connected to nodes $v_{1},v_{2},v_{3}\in V(G).$ We know that $d^{+}(v_{i})\geq 3$ for each $v_{i}\in\{v_{1},v_{2},v_{3}\}$.
 Let $G[v_{1},v_{2},v_{3}]$ represent the induced subgraph by $v_{1},v_{2},v_{3}$.
@@
 As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
-Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) answers a question that we will come back to throughout this paper - this is an exhaustive search algorithm.
-In order for that to work, it has to allow some node’s out-degree to not double.
-In this case, what we see is that the initial node in the example $v_{0}$ does not have its degree double.
-Instead, we are forced to search for the next node.
-That comes at a neighbor of $v_{0}$, $v_{1}$ whose degree does double in this example.
-This is an important distinction because we did not need the ”right” starting node.
-We chose a minimum degree node because of its ability to partition the nodes in the graph, not for its potential closeness to the double degree node.
-For the remainder of this paper, we will refer to graphs like those in Lemma 4.2 using double arrows to indicate that a node or set of nodes connects to every node in the subsequent subgraph.
-This can be done because the nodes within a common induced subgraph are all relating to the same set of nodes.
-These lemmas provide insight for our motivation, and a framework for theoretical foundation. What is coming in the next sections will build on these results and will refer back to them.
+∎
+Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) provides an exhaustive search strategy, which is a question we shall revisit throughout this paper.
+It must prevent a node’s out-degree from doubling in order for that to function.
+In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled.
+We are compelled to look for the next node instead.
+In this case, the degree of the neighbor of $v_{0}$ $v_{1}$ doubles.
+We did not require the ”right” starting node; therefore, this is a crucial distinction.
+Not because of its possible proximity to the double degree node, but rather because of its capacity to divide the nodes in the graph, we selected a minimal degree node.
+We shall use double arrows to show that a node or collection of nodes connects to every node in the next subgraph for graphs similar to those in Lemma 4.2 throughout the rest of this work.
+Because all of the nodes in a shared induced subgraph relate to the same set of nodes, this is possible.
+These lemmas offer a theoretical framework and insight into our rationale. These findings will be expanded upon and referenced in the upcoming sections.
 5 Theoretical Foundation
-In order to understand this next section, we will introduce some new terms.
-We will give a formal name to the partitions we spoke of about previously - containers.
-We will also speak about the decreasing sequence property and talk about its importance.
-Then we will talk about interior, exterior and back arcs.
-Then we will close with the GLOVER data structure.
+We will introduce some new terms to help comprehend this section.
+The partitions we discussed earlier will now be formally referred to as containers.
+Additionally, we shall discuss the decreasing sequence property and its significance.
+After that, we will discuss interior, exterior, and back arcs.
+We will show the importance of these arcs with some foundational lemmas.
+The GLOVER data structure will then be used to wrap up.
 Example 5.1.
@@
 Definition 5.1.
-Given a minimum degree node $v_{0}$, a container $C_{i}$ is the set of nodes that are at distance $i$ from $v_{0}$, along with the arcs between those nodes.
-More formally, $C_{i}=\{u\in V(G)\text{ .s.t.
-}dist(v_{0},u)=i\}\text{ and A}(C_{i})=\{(u,v)|u,v\in C_{i}\}$ ,where $V(G)$ represents the vertex set of the graph $G$, and $dist(v_{0},u)$ is the shortest path distance between $v_{0}$ and $u$.
-Daamouch used induced subgraphs similar in his study of $m$-free graphs [7]. The primary differece here is the neighboring containers (see image) which will be important in the proofs.
-With the concept of containers we can explore the structure of oriented graphs from two perspectives.
-We have the option of exploring it from the perspective of the minimum degree node, or from the perspective of containers.
-  • Containers are seen based on their distance from the
+A container $C_{i}$ is the set of nodes at distance $i$ from $v_{0}$, together with the arcs connecting those nodes, given a minimum degree node $v_{0}$. Formally speaking,
+The formula $C_{i}=\{u\in V(G)\text{.s.t.
+}A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where $V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
+Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7]. The neighboring containers (see image), which will be crucial in the proofs, are the main distinction here.
+We will investigate the structure of oriented graphs from two angles by using the idea of containers. Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node.
+  • Containers are seen based on their distance from the minimum degree node.
   • Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
-minimum degree node.
-The case where $d^{+}(v_{0})=3$, and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) in particular, is an important one because it provides a glimpse into what a counterexample to the conjecture may look like.
-We were able to construct a graph $G$ where $v_{0}$’s out-degree does not double in $G^{2}$.
-Simple graphs of out-degree 3 are the smallest possibilities of cycles in undirected graphs.
-So if we orient the edges in the same direction we can get a cycle of order 3 among the nodes adjacent to $v_{0}$.
-If we look at this situation further we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double.
-In this section we will begin exploring the sufficient conditions for other nodes in the graph to not have their degrees double.
-To continue along this line of reasoning though we will need to define new terms.
-First we will need a term for a node’s out-degree not doubling.
-A node in $G$ whose out-degree does not double in $G^{2}$ is said to have the decreasing sequence property.
-We have seen that examples exist where $v_{0}$ does not have its out-degree double in $G^{2}$.
-We will look into what other nodes can have the decreasing sequence property, and what properties must exist in the graph in order for this to take place.
+Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve.
+A graph $G$ was created in which the out-degree of $v_{0}$ does not double in $G^{2}$.
+A simple graph of degree three corresponds to the lowest cycle that can exist in an undirected graph.
+Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$.
+A closer examination of this case reveals that the out-degree of a node that is an out-neighbor of $v_{0}$ is (at least) twice in this situation.
+This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
+However, we will need to define new concepts in order to proceed with this line of thinking.
+To prevent a node’s out-degree from doubling, we must first find a term.
+The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$.
+There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen.
+We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
 Example 5.2.
@@
 Definition 5.2.
-For a node $x\in V(G)$ in an oriented graph $G$, we say that x has the decreasing sequence property if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e.,$|N^{+}(x)|>|N^{++}(x)|$.
-We have now introduced two concepts: the minimum degree node and the decreasing sequence property (DSP).
-A minimum degree node, defined as the node with the fewest connections, serves as a critical starting point both for our algorithm and degree values.
-By selecting it as the starting point it, it serves as the anchor.
-Now, the minimum degree node is forcing the remaining nodes into the properties that will emerge in the counterexample construction.
-It is acting as a catalyst.
+If the size of a node $x\in V(G)$ in an oriented graph $G$ is strictly bigger than the size of its second out-neighborhood, that is, $|N^{+}(x)|>|N^{++}(x)|$, then x is said to have the decreasing sequence property.
+Two new ideas have been presented thus far: the decreasing sequence property (DSP) and the minimal degree node.
+A crucial starting point for our approach and degree values is a minimal degree node, which is the node with the fewest connections.
+It acts as the anchor by being chosen as the beginning point.
+The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node.
+It is serving as a stimulant.
 In contrast, the DSP introduces constraints on connectivity.
@@
 Total Order Induced by Containers Relative to the Minimum Degree Node
-Let $G$ be a graph with a minimum degree node $v_{0}$.
-Further, let all nodes in $G$ satisfy the Decreasing Sequence Property (DSP).
-Construct a series of containers $C_{1},C_{2},\ldots,C_{k}$ ordered by distance relative to $v_{0}$.
-Each container $C_{i}$ will consist of nodes at that respective distance from $v_{0}$.
-Then this structure imposes a total order on the nodes in $G$ i.e.
-every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
+Let $G$ be a graph with a minimum-degree-node $v_{0}$.
+Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$.
+Each container $C_{i}$ is a set of nodes at distance $i$ from $v_{0}$.
+This imposes a total order on the nodes in $G$.
+This means that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
 Proof.
-To prove a total order, we need to show that every node in $G$ can be compared to every other node in $G$.
+To prove a total order, we need to show that any two nodes in $G$ can be compared.
 Let $u,v$ be two nodes in $G$ such that $u\neq v$.
-It is assumed that $dist(v_{0},u)$ and $dist(v_{0},v)$ are finite for vertices $u$ and $v$ in the containers $C_{i}$.
-This assumption implies that both vertices are reachable from $v_{0}$, although it does not necessitate that all vertices are connected.
-We know that distance is a total ordering, so we know that one of three things are true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$.
-The situation where $dist(v_{0},u)<dist(v_{0},v)$, implies that the node $u$ is in an closer container to $v_{0}$ than the node $v$.
+We assume that both $u$ and $v$ are reachable from $v_{0}$.
+Distance is a total ordering, so one of three things is true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$.
+The situation where $dist(v_{0},u)<dist(v_{0},v)$ implies that $u$ is in a closer container than $v$.
 The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container.
-And the situation $dist(v_{0},u)>dist(v_{0},v)$, implies that the node $u$ is in a further container to $v_{0}$.
-This shows that any two nodes can be compared and shows that it is a total order.
+And the situation $dist(v_{0},u)>dist(v_{0},v)$ implies that $u$ is in a further container.
+This shows that it is a total order.
 ∎
-This is a good start as it imposes order on our graph, but the containers do more than just impose a total order.
-They essentially allow for grouped comparisons among nodes at the same distance from a reference node, $v_{0}$.
-This is something that a standard total order does not account for.
-This ”containerized ordering” enables us to distinguish nodes not only by their distance from $v_{0}$ but also by specific structural or adjacency properties within each distance group.
-Unlike a simple total order, which could be any sequence with a single, consistent ordering (like lexicographic ordering), the container structure allows nodes at the same distance to be further organized.
-This organization is significant because it facilitates comparisons within each distance layer (container) by additional criteria, such as adjacency or neighborhood properties.
+While imposing order on our graph is a solid first step, the containers do more than just enforce overall order.
+In essence, they enable group comparisons between nodes that are equally spaced from a reference node, $v_{0}$.
+A typical total order, such as lexicographical order, does not take this into consideration.
+This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
 Definition 5.3.
@@
   • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
-    $u$ is less than $v$ (denoted as $u<v$), $u$ is greater than $v$ (denoted as $u>v$), or $u$ is equal to $v$ (denoted as $u=v$).
-  • Leveled Structure: The elements of $S$ are partitioned into a leveling of ”containers” or ”levels,” $C_{1},C_{2},,,,,C_{n}$, with each container itself forming a subset of $S$ that maintains a total order.
-The containers are sequentially ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
-  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
-  • Comparability within Containers: For any two elements $u,v\in Ci$, their order can be determined based on a specific criterion, such as a designated metric or property relevant to the application.
+    • $u$ is less than $v$ (denoted as $u<v$),
+    • $u$ is greater than $v$ (denoted as $u>v$), or
+    • $u$ is equal to $v$ (denoted as $u=v$).
+  • Leveled Structure: The elements of $S$ are partitioned into a leveling of containers.
+$C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$ Each container itself forms a subset of $S$.
+This maintains a total order.
+The containers are sequentially ordered relative to each other.
+This establishes a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
+  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$.
+This avoids trivial total orderings where all elements are isolated in singleton containers.
+  • Comparability within Containers: For any two elements $u,v\in C_{i}$, their order can be determined based on a specific criterion.
   • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
-This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
+This enforces that the order of elements between containers aligns with the order of the containers themselves.
 This order means that we have order in two ways.
@@
 Lemma 5.2.
-Decreasing Sequence Property Lemma If a node in an oriented graph $G$ does not have the decreasing sequence property, then its out-degree at least doubles in $G^{2}$.
-If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$, and those that are not neighbors of $x$.
-In the proofs of Lemma 4.2(Minimum Out-Degree 3 with Neighbor Out-Degree 0), and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor induced subgraph (of $v_{0}$) and in the graph with the neighbor induced subgraph removed.
+Decreasing Sequence Property Lemma Suppose a node in an oriented graph $G$ does not have the decreasing sequence property. Then this node will have its out-degree at least double in $G^{2}$.
+If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$.
+In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed.
 We will now give terms to these concepts.
 Definition 5.4.
-Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$.
+Let $x$ be a node with the decreasing sequence property and let $y\in N^{+}(x)$.
 The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
 $int(x,y)=N^{+}(x)\cap N^{+}(y)$
-,where $N^{+}(z)$ is the set of out-neighbors of $z$. The number of interior neighbors is the interior degree of $y$ with respect to $x$.
-Now that we have introduced containers, we can talk arcs where both nodes are inside the containers, these are interior neighbors.
-Interior neighbors will be important to our study on Seymour’s Second Neighborhood Conjecture and our search for a counterexample.
-In particular we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1).
-The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1 and we see that when all the neighbors in that subgraph had outdegree of 1, $v_{0}$’s degree did not double.
+The number of interior neighbors is the interior degree of $y$ with respect to $x$.
+Now that we have introduced containers, we can talk about arcs where both nodes are inside the containers; these are interior neighbors.
+Interior neighbors will be important to our study on SSNC and our search for a counterexample.
+In particular, we saw things about interior neighbors in 4 (Minimum Out-Degree 3 with Neighbors 1).
+The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
 Definition 5.5.
@@
 $ext(x,y)=N^{+}(y)\cap N^{++}(x)$
-, where $N^{+}(y)$ is the set of out-neighbors of $y$, and $N^{++}(x)$ is the set of second neighbors of $x$. The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
+The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
 We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers.
@@
 Exterior arcs are the double arrows between the boxes.
-Interior arcs are not drawn but can be mentally placed either between any two nodes in the red or blue box.
+Interior arcs are not drawn but can be mentally placed either between any two nodes in the container boxes in example 5.2 or example 5.1.
 Definition 5.6.
-Let $v_{0}$ be a minimum degree node, $x$ be a node in $G$, and $y\in N^{+}(x)$ be in the component $C_{i}$.
-A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
+Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the component $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
 We have spoken of so much of our algorithm depending on the distance property of this oriented graph.
 We would like to continue this.
-In order to do so though, we need to consider the option of back-arcs and see how they might hurt our algorithm.
+In order to do so, though, we need to consider the option of back-arcs and see how they might hurt our algorithm.
 These have the possibility of altering a container’s distance to the minimum degree node.
-Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture where they may force us to double count nodes.
+Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
 Having defined our arcs, we can talk about the decreasing sequence property.
-This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decreases as we get further from the minimum degree node.
+This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from the minimum degree node.
 Exterior degrees, though, are a part of a node’s degree equation.
-So there’s another part of it.
+There’s another part of it though.
 As we move further from the minimum degree node, the size of the interior degrees increases.
@@
 In all of this, the minimum degree node is serving as a catalyst and driving force for the process.
-At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s outdegrees.
-With this definition of interior and exterior nodes now introduced, we are now ready to now introduce the Graph Level Order (GLOVER) data structure.
+At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees.
+With this definition of interior and exterior nodes now introduced, we are now ready to introduce the Graph Level Order (GLOVER) data structure.
 Definition 5.7.
-A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
-Leveled Structure: The elements of $S$ are partitioned into levels of ”containers” or ”rankings,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
-The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
-Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
-Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
+A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
+Leveled Structure: The elements of $S$ are partitioned into levels of containers $C_{1},C_{2},\ldots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
+Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$.
+Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion.
 Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
-This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
 Interior and Exterior Neighbors:
@@
 Theorem 5.1.
-Given an oriented graph $G$ and a minimum degree node $v_{0}$, where the nodes of $G$ are partitioned into containers $C_{1},C_{2},\dots,C_{k}$ based on their distance from $v_{0}$ and assuming no back arcs, and the interior arcs in each container and exterior arcs between each container forms a Graph Level Order (GLOVER) under the distance total order.
+Assume we have an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER.
 Proof.
-To establish that the structure $\{C_{1},C_{2},\dots,C_{k}\}$ satisfies the GLOVER properties, we verify each condition in the GLOVER definition:
-  • Total Order on Elements within $G$: - By partitioning nodes based on their distance from $v_{0}$, we establish a total order on the elements of $G$: for any two nodes $u$ and $v$, their order is determined by their respective containers.
-Nodes in closer containers (e.g., $C_{i}$) are considered ”less than” nodes in further containers (e.g., $C_{j}$ with $i<j$).
-    - Within each container, nodes may be ordered based on additional criteria (such as their individual distances to $v_{0}$ if there are multiple paths).
-Thus, we achieve a total order among all nodes in $G$.
-  • Leveled Structure with Containers: - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$.
-The elements within each container are in a total order relative to each other, and the containers themselves are ordered by distance from $v_{0}$, establishing a multi-level ranked structure.
-  • Non-triviality Requirement: - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node.
-This fulfills the non-triviality requirement, as we avoid having all singleton containers.
-  • Comparability within Containers: - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$ (for example, their shortest-path distance if there are multiple paths).
-Thus, comparability is ensured within containers.
-  • Consistency of Comparisons Across Containers: - Since the graph is oriented and there are no back arcs (i.e., no edges from nodes in $C_{j}$ to nodes in $C_{i}$ for $j>i$), the ordering between containers is respected.
-This guarantees that if $C_{i}$ is ”less than” $C_{j}$, then every element in $C_{i}$ is ”less than” every element in $C_{j}$, ensuring consistency across containers.
-  • Interior and Exterior Neighbors: Nodes $u$ and $v$ of the same distance from the chosen minimum degree node $v_{0}$ would be in the same container and thus be interior neighbors.
-If these nodes were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance different by 1.
-    Each criterion for an GLOVER is satisfied by the structure defined by partitioning nodes based on distance from $v_{0}$.
-Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a Ranked Total Order under the distance total order, completing the proof.
+To establish a Graph Level Order, we first need a minimum degree node. We can obtain this by taking the minimum value on the degree vector of nodes. We will call that in the node $v_{0}$.
+Now that we know the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
+Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes whose distance from the node $v_{0}$ is equal.
+These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
+  • Total Order on Elements within $G$ Because nodes are partitioned based on their distance from $v_{0}$, we have a total order on the elements of $G$.
+For any two nodes $u$ and $v$, this total order is determined by their containers.
+  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$.
+  • Non-triviality Requirement - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node.
+  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$.
+  • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
+  • Interior and Exterior Neighbors:
+    • Nodes $u$ and $v$ of the same distance from $v_{0}$ would be in the same container and thus be interior neighbors.
+    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1.
+    Each criterion for a GLOVER is satisfied by partitioning nodes based on distance from $v_{0}$.
+Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a GLOVER under the distance total order, completing the proof.
 What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers.
@@
 By assuming no back arcs, we ensure that each container maintains its place in the total ordering.
-Next, we will show the power of this data structure and why is introduced in this paper.
+Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering.
 Lemma 5.3.
-GLOVER Supports Induction: In any oriented graph, the minimum degree node $v_{0}$ within a Glover structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
+GLOVER Supports Induction
+In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
 Proof.
-Let $v_{0}$ be the node with the minimum out-degree in the oriented graph $G$.
-By definition, for any node $v\in G$, $d^{+}(v_{0})\leq d^{+}(v)$, where $d^{+}(v)$ is the out-degree of node $v$.
-This ordering provides a baseline for comparison because all other nodes are either equal to or have greater out-degree than $v_{0}$.
-The analysis begins at $v_{0}$, which serves as the foundation of the GLOVER structure.
-All outgoing arcs from $v_{0}$ lead to nodes at the next level (or ”container”) $C_{1}$.
-By the properties of the GLOVER structure (e.g., no back arcs), the relationships between nodes are strictly directed, ensuring no circular dependencies.
-Assume that for any node $v$ in a container $C_{i}$, the following properties hold:
-  • $d^{+}(v)\geq d^{+}(v_{0})$.
-  • All properties derived from $v_{0}$ apply recursively to nodes in $C_{i}$.
-Consider any node $u\in C_{i+1}$ that is connected to a node $v\in C_{i}$.
-By the inductive hypothesis, $v$ satisfies all conditions derived from $v_{0}$.
-Since $u$ receives its inbound arcs from $v$, and $d^{+}(v)\geq d^{+}(v_{0})$, the relationships in $C_{i+1}$ are consistent with the properties of the Glover structure.
-Furthermore, no node $u$ in $C_{i+1}$ can violate the inductive assumptions because $v_{0}$ ensures a well-ordering, propagating constraints downward.
-Starting from $v_{0}$, the well-ordering propagates through all containers $C_{i}$, ensuring that:
-  • The out-degree of any node is anchored to $v_{0}$.
-  • There is no upward deviation in degree that contradicts the structure’s constraints.
-Thus, the GLOVER structure supports induction by leveraging the minimum degree node $v_{0}$ as an anchor, ensuring consistency throughout the graph. ∎
+We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
+A set is considered well-ordered if every non-empty subset of that set has a least element.
+A discrete total order is a total order where each element has a distinct ”next” element in the order.
+Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order.
+This fulfills the condition of being well-ordered.
+Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set.
 Example 5.3.
@@
 Figure 6: This example illustrates back arcs in a graph.
-The two blue arcs from $x3$ to $y3$ and $y3$ to $z3$ represent exterior neighbors indicate forward directing arcs.
-Additionally, a red arc points backwards from $z5$ to $x1$.
+The two blue arcs from $x_{3}$ to $y_{3}$ and $y_{3}$ to $z_{3}$ represent exterior neighbors indicating forward-directing arcs.
+Additionally, a red arc points backwards from $z_{5}$ to $x_{1}$.
 Although the containers are not outlined, they are implied by the difference in variables for each group.
 Lemma 5.4.
-Interior Degree Greater than 1 In any oriented graph $G$, represented by a GLOVER with $v_{0}$ as a minimum out-degree node satisfying the decreasing sequence property, In an attempt to prevent a degree doubling node, every node $x$ in the neighboring container $N^{+}(v_{0})$ has an interior degree of at least 1.
+Interior Degree at least 1
+Suppose we have an oriented graph $G$, represented by a GLOVER.
+Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property.
+In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $N^{+}(v_{0})$ must have an interior degree of at least 1.
 That is, for $x\in N^{+}(v_{0})$,
@@
 Proof.
-By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$
-where $N^{+}(x)$ denotes the set of first neighbors (out-neighbors) of $x$, and $N^{++}(x)$ denotes the set of second neighbors (out-neighbors of $N^{+}(x)$).
-For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $y$, and then from $y$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
-Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $y\in N^{+}(v_{0})$.
-If $y$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $y$’s neighbors, which are $v_{0}$’s second neighbors.
-If $y$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $y$’s first neighbors.
-To avoid this doubling, $y$ must connect to at least one interior arc. This means that the interior degree of $y$ with respect to $v_{0}$ must satisfy
-$|int(v_{0},y)|\geq 1$
-This argument holds for any neighbor $y\in N^{+}(v_{0})$, ensuring that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
-Thus, for all $x\in N^{+}(v_{0})$,$|int(v_{0},x)|\geq 1$, completing the proof. ∎
-Lemma 5.4 (Interior Degree Greater than 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling.
+By definition, a node satisfies the decreasing sequence property if
+$|N^{+}(x)|>N^{++}(x)|$
+For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $x$, and then from $x$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
+Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $x\in N^{+}(v_{0})$.
+If $x$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $x$’s neighbors, which are $v_{0}$’s second neighbors.
+If $x$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $x$’s first neighbors.
+To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
+$|int(v_{0},x)|\geq 1$
+This argument holds for any neighbor $x\in N^{+}(v_{0})$. This ensures that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
+Thus, for all $x\in N^{+}(v_{0})$, $|int(v_{0},x)|\geq 1$.
+Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling.
 In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
@@
 Lemma 5.5.
-Exterior Less than Degree In any oriented graph $G$, represented by a GLOVER In an attempt to prevent a degree doubling node, if a node $x_{i}$ is in container $C_{i}$ has the decreasing sequence property, In an attempt to prevent a degree doubling node, then for all $y\in N^{+}(x)\subseteq C_{i+1}$, $|ext(x,y)|<d^{+}(x)$.
+Exterior Less than Degree
+Suppose we have an oriented graph $G$, represented by a GLOVER.
+In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i}),|\text{ext}(x,y)|<d^{+}(x)$.
 Proof.
-Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|ext(x,y)|\geq d^{+}(x)$.
-By definition, $ext(x,y)$ consists of the set of out-neighbors of $y$, specifically:
-$ext(x,y)=\{z\in N^{++}(x)\mid z\text{ is an out-neighbor of }y\}.$
-Thus, $|ext(x,y)|\subseteq N^{++}(x)$, where $N^{++}(x)$ denotes the set of second out-neighbors of $x$. If $|ext(x,y)|\geq d^{+}(x)$, then:
+Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$.
+By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$
+Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$, then:
 $|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x),$
-which suggests that the out-degree of $x$ would at least double in the graph $G^{2}$ (the second power of the graph, where edges represent paths of length 2).
-This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property, where the degree of nodes cannot increase as we move through the containers.
-Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|ext(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
-Lemma 5.5 is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree Greater than 1), with the notable exception that Lemma LABEL:extbd1 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
-The results presented in 5.4(Interior Degree Greater then 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1.
-There, he demonstrated that a minimum degree vertex with no cycles in its induced graph has a degree that effectively doubles.
-Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we’re developing.
+suggesting that the degree of outwardness of $x$ would at least double on graph $G^{2}$. This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property.
+Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$.
+Lemma 5.5 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree at least 1), with the notable exception that Lemma 5.5 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
+The results presented in 5.4 (Interior Degree At Least 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1.
+There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
+Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we are developing.
 Lemma 5.6.
-[…]
-Then, for all $z\in ext(x,y)$, we have $|ext(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
+Weak Decreasing Exteriors Lemma
+Suppose that $G$ is an oriented graph represented by a GLOVER, and that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
+Suppose that:
+  1. 1.
+    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the subgraph $G[N^{+}(x)]^{2}$.
+  2. 2.
+    For all $z\in ext(x,y)$, we have $ext(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
+Then, for all $z\in\text{ext}(x,y)$, we have $|\text{ext}(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
 Proof.
@@
 A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
-Now, assume the conditions of the lemma hold: 1.
-The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increases significantly in $G[N^{+}(x)]$.
-2.
-For all $z\in ext(x,y)$, $ext(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
-Now, suppose for contradiction that there exists $z\in ext(x,y)$ such that $|ext(y,z)|\geq|ext(x,y)|$.
+Now, assume the conditions of the lemma hold:
+  1. 1.
+    The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
+  2. 2.
+    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
+Now, suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$.
+The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees.
+By the assumption that $y$’s interior degree doubles, if $y$’s exterior degree were to double, it would lead to a contradiction.
+If there exists a node $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$, then $y$’s out-degree will be impacted by $\text{ext}(x,y)$ and $\text{ext}(y,z)$.
+$\begin{split}|\text{ext}(y,z)|&\geq|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|ext(x,y)|&\geq|\text{ext}(x,y)|+|\text{ext}(x,y)|\\ |\text{ext}(y,z)|+|\text{ext}(x,y)|&\geq 2*|\text{ext}(x,y)|\end{split}$ (1)
 This would mean that the exterior degree of $y$ is at least as large as the exterior degree of $x$.
@@
 The absence of back arcs prevents any increase in $y$’s exterior degree from its second neighbors.
-Thus, the assumption $|ext(y,z)|\ \geq|\text{ext}(x,y)|$ leads to a contradiction,
-confirming that $|\text{ext}(y,z)|<|\text{ext}(x,y)|$ for all $z\in\text{ext}(x,y)$.
-Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
-The first thing to notice in this proof are the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions has an important purpose.
-First we will talk about the lack of back arcs.
+Thus, the assumption $|\text{ext}(y,z)|\allowbreak\geq\allowbreak|\text{ext}(x,y)|$ leads to a contradiction, confirming that $|\text{ext}(y,z)|\allowbreak<\allowbreak|\text{ext}(x,y)|$ for all $z\allowbreak\in\allowbreak\text{ext}(x,y)$.
+Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$.
+The first thing to notice in this proof is the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions are important.
+First, we will talk about the lack of back arcs.
 We have already talked about how back arcs can be a hindrance to our algorithm overall.
 This is one of the main points where they can be a hindrance.
-In particular there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist.
+In particular, there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist.
 We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
-Secondly, we assumed that a node’s interior degree doubled.
+Secondly, we assumed that the interior degree of a node doubled.
 This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem.
-We are assuming that this not $y$ was in another container and its interior degree doubled in that container.
-In doing so, $y$ is now in a better position to have its overall degree double, needing only to have its exterior degree double.
+We are assuming that this node $y$ was in another container and its interior degree doubled in that container.
+In doing so, $y$ is now in a better position to have its overall degree double, only needing to have its exterior degree double.
+The more popular representations of graphs are adjacency matrices, arraylists and edge lists.
+These suffer from the same limitations of graphs as mentioned above, They scatter the nodes along the two-dimensional plane without attempts to group into an ordering.
+The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way.
+For the SSNC, that reasonable way was the distance metric.
+In order to represent graphs in a GLOVER we would use any markup language that can encode graphs.
+There needs to be three types of data- containers, nodes, and arcs.
+The containers will be given ids, the nodes will then be assigned to their unique containers, and arcs will be determined to be interior, exterior, or back, depending on the containers of the endpoints of the arcs.
+This determination of the number of containers and the assignment of nodes to containers does not need to be declared beforehand.
+Instead, just as we have been talking about the distance of a node from a minimum degree node $v_{0}$ as the determination of container assignment, we would have the same or a similar function to assign nodes to containers.
+With this container representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas.
+For example, the lemma 5.6 (Weak Decreasing Exteriors Lemma) can be visualized by example 5.5.
+This shows a series of ordered containers.
+Each container has fewer nodes.
+The fewer nodes implies that the container circle can be drawn with a smaller radius, giving light to the decrease in size.
 Example 5.5.
-0
-$\Rightarrow$ 111122223333444455556666 $\Rightarrow$ 7777888899991010101011111111 $\Rightarrow$ H
-Figure 7: Nodes in each container connect to all nodes in the subsequent container.
-Here, the minimum out-degree node has out-degree 6, with connectivity patterns shown to demonstrate degree doubling in $G^{2}$.
-Table 1: Intersection Results for Triplets of Nodes in 5.5
+[…]
+To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another but no-one follows back.
+This network is then an oriented graph.
+Running the Seymour Second Neighborhood Conjecture in this context means looking for an account whose indirect sphere of influence is greater than their direct sphere of influence.
+Here, we are defining the indirect sphere of influence as those accounts they can reach through others, and we are defining the direct sphere of influence as those accounts they can directly connect with.
+Essentially, these are the same definitions as the second neighbor and first neighbor respectively that we have been working with.
+With this in mind, after finding the person who connects to the fewest accounts in this social media graph, we could set up containers just as before.
+Now, lemma 5.4 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1.
+What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about, or a number of disjoint cycles where messages are spread among themselves.
+Considering example 5.6, which takes the former of these approaches.
+These cycles in the first container are what stops the user Andi from being an influencer.
+This brings us to the next container.
+Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people - (Hugo, Isla, Jade, Kai, and Lei).
+If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3).
+From this we would conclude that a person in container 1, whose indirect sphere of influence is this persons sphere of influence, would have its first degree double and its decond degree double, implying that its overall degree doubles.
+Example 5.6.
+Andi
+$\Rightarrow$ BellChadDawnElseFayeGus $\Rightarrow$ HugoIslaJadeKaiLei $\Rightarrow$ H
+Figure 7: This is an example of a social media setup of the Seymour Second Neighborhood Conjecture.
+The Second Container has ten arcs and 5 nodes, which represents a perfect graph.
+This implies that in the next container $C_{3}$ (not shown), with 4 people can only have 6 arcs.
+Due to this fact, at least one person in $C_{3}$ will relate to at least five people in $C_{4}$.
+We can conclude then that at least one person in $C_{1}$ will be an influencer and have their degree double.
+Table 1: Intersection Results for Triplets of Nodes in 5.6
 Triplet(A, B, C) int(A, B) ext(B, C) Intersection Result
@@
 We have now defined the data structure that we will use throughout the rest of this paper.
-In doing so, we have a better understanding of interior and exterior arcs which will help guide us in our search for a counterexample.
+In doing so, we have a better understanding of interior and exterior arcs, which will help guide us in our search for a counterexample.
 6 Back Arcs
 With a theoretical foundation underway, we are ready to launch an investigation into back arcs.
-We will look into whether these arcs can exist in a counterexample, and what other properties of oriented graphs must be true.
+We will look into whether these arcs can exist in a counterexample and what other properties of oriented graphs must be true.
 It is important to restate that back arcs and exterior neighbors are distinct concepts.
@@
 Rather, an exterior neighbor ext(x, y) specifically refers to nodes at a distance of two from $x$ that are also neighbors of $y$.
-If we look at Example 4.2 again we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$.
-Similarly we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
-One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$.
+If we look at Example 4.2 again, we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$.
+Similarly, we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
+One of the things we notice about Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4 allowed for a cycle to exist in the out-neighbors of $v_{0}$.
 In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double.
@@
 Lemma 6.1.
-(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with minimum degree node $v_{0}$ partitioned by containers $C_{i}$.
-Suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles.
+(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER.
+Let $v_{0}$ be a minimum degree node of $G$.
+Suppose that the containers of $G$ are $C_{i}$.
+In an attempt to prevent a degree-doubling node, suppose that two nodes $u$ and $v$ in a container $C_{i}$ have a common parent $v_{p}$ in the preceding container $C_{i-1}$, and that $v_{p}$’s interior degree doubles.
 If the interior degrees of $u$ and $v$ differ by at least 1, then the exterior degree of $v_{p}$ in container $C_{i-1}$ will double as well.
 Proof.
-Assume $u$ and $v$ are two nodes in container $C_{i}$, both with a common parent $v_{p}$ in container $C_{i-1}$, and let $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$, where $\text{int}(v_{p},x)$ denotes the interior degree of $x$ relative to $v_{p}$’s container $C_{i-1}$.
-By the Decreasing Sequence Property (DSP), if the interior degree of $v$ is higher than $u$, then $v$ will require fewer exterior connections to maintain the required degree restrictions in the graph.
-Meanwhile, $u$, with fewer interior connections, will require more exterior connections to compensate.
-Now, the imbalance in interior degrees means that $u$’s exterior connections are effectively increasing the degree of $v_{p}$.
-This is because $u$ must connect to more external nodes to maintain its degree, leading to an increase in the number of exterior neighbors of $v_{p}$.
-Consequently, the exterior degree of $v_{p}$ must double to accommodate the increased number of connections needed by $u$ and $v$ in the graph, violating the degree constraints intended by the partitioning into containers.
-Thus, the differing interior degrees of $u$ and $v$ in container $C_{i}$, combined with the effects of the DSP, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double.
-This leads to a contradiction, so the claim holds.
-∎
+Assume that $u$ and $v$ are two nodes in container $C_{i}$, and that they share a parent $v_{p}$ in container $C_{i-1}$. Suppose that $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$
+In order to preserve the necessary degree limits in the graph, $v$ will need fewer external connections if its inner degree is greater than $u$.
+$u$, on the other hand, will need more external connections to make up for its lack of inner connections.
+As a result of the imbalance in internal degrees, the external connections of $u$ are essentially raising the second neighbors of degree of $v_{p}$.
+As a result, the exterior degree of $v_{p}$ will double to make room for the additional connections that $u$ in the graph requires.
+Thus, these different interior degrees of $u$ and $v$ in container $C_{i}$, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double. This leads to a contradiction, so the claim holds. ∎
 Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples.
@@
 The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth.
-A back arc is defined as an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$.
-If such a back arc were introduced, it would imply that the source node $y$ gains new neighbors at a lower level, potentially doubling its degree.
-This contradicts the hierarchical organization of containers, where each level is strictly defined by shortest-path distances from $v_{0}$.
+The definition of a back arc is an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$ respectively.
+It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included.
+The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which is in contrast to this.
 Lemma 6.2.
-No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node.
-If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $ext(u_{i},v_{k})$ is at least the degree of $v_{k}$.
+No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node.
+If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $\text{ext}(u_{i},v_{k})$ is at least the degree of $v_{k}$.
 Furthermore, if the interior degree of $u_{i}$ doubles, this will cause the total degree of $u_{i}$ to double as well.
@@
 $|ext(u_{i},v_{k})|=|N^{+}(v_{k})\cap N^{++}(u_{i})|$
-where $N^{+}(v_{k})$ is the set of immediate neighbors of $v_{k}$, and $N^{++}(u_{i})$ is the set of nodes at distance two from $u_{i}$.
 Since $N^{+}(v_{k})$ includes all direct neighbors of $v_{k}$, it follows that:
@@
 $|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
-Now, consider the relationship between interior and total degree.
-If the interior degree of $u_{i}$ doubles, the total degree of $u_{i}$ must also double.
-This is because the exterior neighbors are tightly constrained by the back arc and are not impacted by the increase in interior connections.
-Thus, any increase in the interior degree directly translates into a proportional increase in the total degree.
-Finally, by the Weak Decreasing Exteriors Lemma, the size of $ext(u_{i},v_{k})$ decreases as we move further in the sequence of containers.
-This ensures that the total degree of $u_{i}$ remains bounded and behaves as expected under the given conditions, completing the proof.
-∎
-Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container structured graph.
+Each node in $u_{i}$’s exterior is matched up with a node in $v_{k}$’s neighborhood.
+This means that the back arc will double $u_{i}$’s exterior degree.
+Then, because we’ve already assumed that $u_{i}$’s interior degree will double, we can say that $u_{i}$’s overall degree will double.
+This is true because its interior degree and its exterior degree will double.
+Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container-structured graph.
 It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container.
@@
 Back arcs can disrupt this order.
-What we see next is a special case of how back arcs could disrput this order.
+What we see next is a special case of how back arcs could disrupt this order.
 Lemma 6.3.
-Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
-1.
+Back Arcs Special Case
+Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
+  1. 1.
     Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP).
-2.
+  2. 2.
     The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$.
-3.
+  3. 3.
     There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
@@
 $|\text{ext}(y,z)|<|\text{ext}(x,y)|$
-First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
+First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e.
+from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e.
+not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
 This gives the only possibility for $z$ as being an interior neighbor of $x$.
@@
 These arcs can lead to double counting.
-An out-neighbor of $w$ might be an out-neighbor of $x$, and a second out-neighbor of $y$.
-This implies that under the assumption that $w$’s interior degree double’s, $w$’s overall degree will double through that early back arc.
+An out-neighbor of $w$ might be an out-neighbor of $x$ and a second out-neighbor of $y$.
+This implies that under the assumption that $w$’s interior degree doubles, $w$’s overall degree will double through that early back arc.
 Lemma 6.4.
-First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph represented by a GLOVER with minimum degree $d$ in an attempt to prevent a degree doubling node, and let $v_{0}$ be a node of minimum degree.
-Suppose $G$ is an oriented graph, where arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$.
-If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will violate the structure of the graph.
+First Back Arc Between Consecutive Containers in an Oriented Graph:
+Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$.
+Suppose that we represent this graph by a GLOVER.
+Suppose we are attempting to prevent a degree-doubling node and that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$.
+If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
 Proof.
-Assume that the oriented graph $G$ is partitioned into containers $C_{1},C_{2},\dots,C_{i}$ such that nodes in each container $C_{j}$ (for $j<i$) only have either arcs to fellow nodes in $C_{j}$, or they map forward to nodes in $C_{j+1}$, with no back arcs present between containers up to $C_{i}$.
+Assume that the oriented graph $G$ is partitioned into containers $C_{1},C_{2},\ldots,C_{i}$ such that nodes in each container $C_{j}$ (for $j<i$) only have interior arcs in $C_{j}$ or exterior arcs to nodes in $C_{j+1}$, with no back arcs present between containers up to $C_{i}$.
 Now, let the first back arc appear from a node $u_{i}\in C_{i}$ to a node $v_{i-1}\in C_{i-1}$.
@@
 If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist{v_{0},u_{i}}-i$ so $u_{i}$ cannot belong to container $C_{i}$.
-Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ leads to a fundamental contradiction in the oriented nature of the graph, and no back arcs can exist prior to $C_{i}$. ∎
+Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$.
 Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees).
@@
 Lemma 6.5.
-No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$,$C_{2}$, $\dots$,$C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
+No Back Arcs Non-Existence of Back Arc
+In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
 Proof.
-This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.4 (First Back Arc Between Consecutive Containers) . ∎
+This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers).
+[…]
+Figure 9: Illustration of the proof of 7.2 (Interior Degrees Double).
+This is another illustration of the Interior Degree Doubling.
+Each node inside a container represents a node whose degree is doubled by the inductive hypothesis.
+This process cannot continue forever so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
 […]
@@
 7 Interior Degrees Doubling
-This section introduces the concept of interior degree doubling.
-This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s the interior degree inside of a container.
+This section introduces the concept of interior degree-doubling.
+This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container.
 This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
 Lemma 7.1.
-Indistinguishable Regular Interior Nodes In a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER in an attempt to prevent a degree doubling node.
-If a node $v_{p}$’s interior degree doubles then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
+Indistinguishable Regular Interior Nodes
+Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER.
+We are attempting to prevent a degree-doubling node.
+If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree doubled.
 Proof.
-Let $C_{i}$ be a regular container of size $n$, and let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$.
-The interior degree of a node $v_{i}$ in container is defined as the number of arcs shared $v_{i}$ between $v_{i}$ and a neighbor or parent node.
+Let $C_{i}$ be a regular container of size $n$.
+Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$.
+The interior degree of a node $v_{i}$ in a container is defined as the number of arcs shared by $v_{i}$ between $v_{i}$ and a neighbor or parent node.
 In a regular container, all nodes are indistinguishable in terms of their degree.
 They also have the same parents because of the DSP.
-Any automorhism we construct including a random automorphism $\phi$ of the container $C_{i}$ will preserve the degree structures.
-This means any node will have its interior degree double.
-∎
+Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures.
+This means any node will have its interior degree doubled.
 The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others.
@@
 That would be the key.
-What indistinguishabable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double.
+What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double.
 That means that we can select any of those nodes.
 Lemma 7.2.
-Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node.
+Interior Degrees Double
+Let $G$ be an oriented graph represented by a GLOVER.
+We are attempting to prevent a degree-doubling node.
 Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP).
@@
 We prove this by strong induction on the size of the graph.
-Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3 there exists at least one node in each container with a doubled interior degree.
+Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree.
 These cases form our base, supporting the induction.
-Inductive Hypothesis: Assume that for any graph of size $n^{\prime}<n$, each container $C_{i}$ contains at least one node whose interior degree doubles, thus satisfying the DSP by proportionally reducing exterior degrees.
-Inductive Step: Consider a graph of size $n$ where we need to show that at least one node in each container $C_{i}$ has its interior degree doubled.
+Inductive Hypothesis: Assume that for any graph of size $n^{\prime}<n$, each container $C_{i}$ contains at least one node whose interior degree doubles.
+Inductive Step: Consider a graph of size $n$. We need to show that at least one node in each container $C_{i}$ has its interior degree doubled.
 Suppose we construct $G$ such that no node at distance $i$ from $v_{0}$ initially has a doubled interior degree in $C_{i}$.
@@
 $H_{0}=\{u_{i}\}$
+, where $u_{i}$ is the minimum degree node in $C_{i}$.
 $H_{1}=N^{+}(u_{i})$
@@
 We can now define $H_{2}=ext(u_{i},y_{1})$.
-Likewise, by the I.H., there will be a node $y_{2}inH_{2}$ whose interior degree doubles in $H_{2}$.
+Likewise, by the I.H., there will be a node $y_{2}\in H_{2}$ whose interior degree doubles in $H_{2}$.
 We can define $H_{3}=ext(y_{1},y_{2})$.
 Continuing this process, we can define $H_{j}=ext(y_{j-2},y_{j-1})$, where $y_{j-1}$ is a node whose interior degree doubles in $H_{j-1}$.
-By the Weak Exterior Size Decreasing lemma, we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers.
-By the DSP, an increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$, preventing exterior degrees from exceeding interior degrees.
-This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$, as required by DSP.
+By Lemma 5.6 (Weak Decreasing Exteriors Lemma), we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers.
+An increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$.
+This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$.
 The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$.
-At this point, the remaining nodes in $H_{i}$ either have no further exterior neighbors or their degrees lie entirely within the interior set.
-Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$.
+At this point, there are not enough arcs (1) in $H_{i}$ for both nodes to have an interior degree of 2.
+So, one node must have all its degree be exterior.
+Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i-1}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$.
 Thus, the parent of $y^{*}$, denoted $y^{*}-1$, in $G$ achieves a doubled degree within $C_{i}$.
-This doubling process terminates as DSP constraints prevent exterior degrees from surpassing interior degrees without contradicting the DSP.
-Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
+Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step.
 Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture.
@@
 Once again, we see why the Decreasing Sequence Property (DSP) received its name.
-This doubling of interior degrees will help us with the construction of our algorithm in search for a counterexample.
-It does, however, give us a bit of light into the side of the Conjecture being true.
+This double of interior degrees will help us to build our algorithm in the search for a counterexample. It does, however, give us a bit of light into the side of the Conjecture being true.
 8 Container Density
 Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean.
-The Weak Exteriors Decreasing Lemma implies that containers will decrease in size.
+The Weak Decreasing Exteriors Lemma implies that containers will decrease in size.
 We need to see how that decrease in size will impact the degrees of the nodes.
 Lemma 8.1.
-Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node.
+Regular Exterior Degrees
+Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node.
 Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP).
-Then the regular interior degrees will result in regular exterior degrees in $G$.
+Then the nodes in a container in $G$ will result in regular exterior degrees.
 Proof.
-Let $x_{i}$ and $x_{j}$ be two nodes at the same distance $i$ from the minimum degree node $v_{0}$ in the oriented graph $G$, both contained in the same container $C_{i}$.
+Let $x_{i}$ and $x_{j}$ be two nodes in the same container $C_{i}$.
 Assume, for contradiction, that the sizes of their exterior sets differ by at least one.
@@
 This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
-Since the size of the next container $C_{i+1}$ is determined by the exterior degrees of nodes in $C_{i}$, this irregularity in the exterior degrees would force the size of $C_{i+1}$ to differ, contradicting the assumption that the containers are of regular size.
-By Lemmas 6.1 (Regular Interior Degrees) and 7.2 (Interior Degree Double), we know that if the interior degrees within a container are regular, the exterior degrees must also follow this regularity.
-Specifically, if there is an irregularity, any parent node of $x_{i}$ and $x_{j}$ will have its exterior degree double, which would in turn double their total degree.
-This doubling of the parents’ total degree would violate the Decreasing Sequence Property (DSP).
-Therefore, the assumption that $|ext(x_{i},y)|$ and $|ext(x_{j},y)|$ differ is false. Hence, the exterior degrees in $G$ must be regular if the interior degrees are regular. ∎
-We proved Lemma 6.1 (Regular Interior Degrees) on our way to proving that back arcs did not exist.
-LABEL:refintdeg is an important property of the graph though that we are going to use to get other properties of the graph, like 8.1.
+By Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs), we can imply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here.
+This says that with each distance from the minimum degree node, our exterior size is decreased by 1.
+This stopped at the container $C_{i}$.
+That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$.
+So, a node in $C-{i-2}$ can have its exterior degree double.
+By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
+So, this assumption led to a contradiction.
+This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the Weak Decreasing Exteriors Lemma in this proof because the two go hand in hand.
+Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can take place in this graph.
 Lemma 8.2.
-Regular Graph All Nodes Have Equal Our-Degree Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree doubling node, with $v_{0}$ as a minimum degree node.
-Further, let all nodes satisfy the Decreasing Sequence Property (DSP).
-Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $d_{v_{0}}=d^{+}(v_{0})$.
-Proof.
-We will prove this lemma by induction on the number of nodes in the graph $G$.
-For the base case, consider the node $v_{0}$, which, by definition, has the minimum degree $d$, so $d^{+}(v_{0})=d_{v_{0}}$.
-Assume that for some node $v_{k}$ in $G$, the degree of $v_{k}$ satisfies $d^{+}(v_{k})=d_{v_{0}}$.
-Now consider an arbitrary node $v_{k+1}$ in $G$. Suppose, for contradiction, that $d^{+}(v_{k+1})>d_{v_{0}}$.
-Consider a neighbor $v_{n}$ of $v_{k+1}$.
-By the inductive hypothesis, the degree of $v_{n}$ is exactly $d_{v_{0}}$.
-Since $v_{k+1}$ has more than $d_{v_{0}}$ neighbors, the neighbors of $v_{k+1}$ are second neighbors of $v_{n}$ in the graph $G^{2}$.
-Therefore, $v_{n}$ would have more than $d_{v_{0}}$ second neighbors in $G^{2}$.
-This implies that $v_{n}$’s degree has effectively doubled, which contradicts the DSP, which requires that $v_{n}$ must have degree $d_{v_{0}}$.
-Hence, the assumption that $d^{+}(v_{k+1})>d_{v_{0}}$ leads to a contradiction.
-Thus, we must conclude that:
-$d^{+}(v_{k+1})=d_{v_{0}}.$
-By the principle of mathematical induction, it follows that for every node $v$ in $G$, the degree of $v$ is equal to $d_{v_{0}}$. ∎
-We are incrementally refining the properties that must hold for our counterexamples.
-The two most crucial aspects are the Decreasing Sequence Property and the identification of the minimum degree node.
-However, additional constraints, such as the overall graph regularity established by Lemma 8.2 (Regular Graph), help simplify future arguments.
-This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
-Lemma 8.3.
-Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER In an attempt to prevent a degree doubling node, with no back arcs, and let $v_{0}$ be the chosen minimum degree node in $G$.
-For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
+Strong Decreasing Exteriors Lemma
+Let $G$ be an oriented graph represented by a GLOVER.
+Let $v_{0}$ be a minimum degree node in $G$.
+We are attempting to prevent a degree-doubling node, so for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
 |$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
-for some $z\in N^{+}(x_{i})$ in order to keep any node’s degree from doubling in the oriented graph.
+for some $z\in N^{+}(x_{i})$.
 Proof.
 We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
-Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
+Base Case ($i=1$):
+Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
 $\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
-is bounded by $d^{+}(v_{0})-1$, since each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2, as there are no back arcs to $v_{0}$.
-Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
+is bounded by $d^{+}(v_{0})-1$. Each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2. We have already proved that there are no back arcs to $v_{0}$.
+Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
 |$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
@@
 |$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
-Lemma (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma), to give us a way of calculating the maximum size of a container based only in its distance from the minimum degree node and the minimum degree node’s degree.
-The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove. It is for this sequence that the term ”decreasing sequence property” is named.
+Lemma 8.2 (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma) to give us a way of calculating the maximum size of a container based only on its distance from the minimum degree node and the minimum degree node’s degree.
+The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove.
+It is for this sequence that the term ”decreasing sequence property” was originally named.
 Lemma 8.3.
-Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER, In an attempt to prevent a degree doubling node, and let $v_{0}$ be a minimum degree node in $G$, where every node in $G$ has the Decreasing Sequence Property (DSP).
+Interior Degrees Lower Bound
+Let $G$ be an oriented graph represented by a GLOVER.
+Let $v_{0}$ be a minimum degree node in $G$.
+Suppose every node in $G$ has the Decreasing Sequence Property (DSP).
 Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$.
@@
 This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
-Now, since $x$ has at least the same out-degree as $v_{0}$ (by the assumption of DSP), $x$ must map to at least $i$ other nodes.
-Given that back arcs are not allowed (as shown previously), $x$ cannot map to nodes at earlier distances without violating the structure of the graph.
+Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not asked, $x$ cannot map to nodes at earlier distances.
 Therefore, the only remaining possibility is that $x$ maps to interior nodes.
-Since $x$’s degree must account for at least $i$ mappings, and these must be to interior nodes, we conclude that $x$ must map to at least $i$ interior nodes.
-Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
-Having already placed an upper bound on the exterior sizes and containers, it is only natural that the there is an opposite bound for the interiors.
-Once again the minimum degree node is the driving force behind this.
+Since $x$’s degree must account for at least $i$ mappings.
+These must be to interior nodes.
+So, we conclude that $x$ must map to at least $i$ interior nodes.
+Thus, $|int(v_{i-1},x)|\geq i$, as required.
+Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is an opposite bound for the interiors.
+Once again, the minimum degree node is the driving force behind this.
 The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
+Lemma 8.4.
+Regular Graph
+All Nodes Have Equal Out-Degree
+Let $G$ be an oriented graph represented by a GLOVER.
+Suppose $v_{0}$ is a minimum degree node of $G$.
+Further, let all nodes satisfy the Decreasing Sequence Property (DSP).
+We are attempting to prevent a degree-doubling node.
+Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
+Proof.
+With the proofs on the lack of back arcs and the regularity of interior arcs and the regularity of exterior arcs, we arrive at the following equation for each container:
+$\delta=|int(v_{p},u_{i})|+|ext(v_{p},u_{i})|$
+where $u_{i}$ is in container $C_{i}$, and $v_{p}$ is a parent of $u_{i}$ in container $C_{i-1}$.
+For a node $u_{i}$ in container $C_{i}$, with degree $\delta+1$, there would be the possible places for this degree to be located—back arcs, interior arcs, or exterior arcs.
+We have already shown that back arcs cannot exist.
+If this extra degree went to the exterior arcs, then it would cause the degree of a parent node to double.
+The only remaining possibility is for this extra degree to go to the interior arcs of $u_{i}$.
+While this is not a problem, it will cause the interior degrees to increase at a faster rate than we just showed in Lemma 8.3 (Interior Degree Lower Bound), as this would increase the number of interior arcs.
+This more rapid reduction in the interior degree leads to container density faster. To avoid this, regularity is key.
+Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph.
+We are incrementally refining the properties that must hold for our counterexamples.
+The two most crucial aspects are the Decreasing Sequence Property and the identification of the minimum degree node.
+However, additional constraints, such as the overall graph regularity established by Lemma 8.4 (Regular Graph), help simplify future arguments.
+This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
 9 Transitive Triangles and Seymour Diamonds
@@
 Definition 9.1.
-Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
-Here, we have $x$ pointing to $y$ and $u$; and $y$ pointing to $u$.
+Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, and $u$ form a transitive triangle.
+Here, we have $x$ pointing to $y$ and $u$ and $y$ pointing to $u$.
 The concept of a transitive triangle is an important one because it involves the risk of double counting.
-The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles we might improperly count out second neighbors and thus not have a correct algorithm.
+The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
 Definition 9.2.
-Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
-The Seymour Diamond is a graph theory concept in graph there where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$.
-This is basically a four vertex oriented graph.
+Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, and $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
+The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$.
+This is a four-vertex-oriented graph.
 What this situation presents is redundant second neighbors.
 There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond.
-As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and more importantly the Seymour’s Second Neighborhood Conjecture.
+As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
 The notion of interior neighbors is another way of describing what [1] referred to as transitive triangles.
-They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e.
-one where every node has the decreasing sequence property.
-We provided a similar result with Lemma 5.4 (Interior Degree Greater than 1).
-When a minimum degree node $v_{0}$ has the decreasing sequence property, all of its neighbors have interior degree of at least 1.
-That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$ will form a transitive triangle.
+They showed that these triangles are always present in a graph that doesn’t satisfy Seymour’s conjecture, i.e., one where every node has the decreasing sequence property.
+We provided a similar result with Lemma 5.4 (Interior Degree at least 1).
+When a minimum degree node $v_{0}$ has the Decreasing Sequence Property, all of its neighbors have an interior degree of at least 1.
+That means that every node $x\in N^{+}(v_{0})$ will relate to some other node $y\in N^{+}(v_{0})$ and these three nodes, $(v_{0},x,y)$, will form a transitive triangle.
 We can extend that further for any two constructive containers.
@@
 Lemma 9.1.
-Transitive Triangles from Containers In a graph $G$ represented by a GLOVER, for any two consecutive containers $C_{i}$ and $C_{i+1}$ in a graph satisfying the Decreasing Sequence Property (DSP), there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
-That is, there exist directed edges from $u$ to $w$, $v$ to $w$, and from $u$ to $v$, or from $v$ to $u$.
+Transitive Triangles from Containers
+Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP).
+For any two consecutive containers $C_{i}$ and $C_{i+1}$, there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
 Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm.
@@
 Seymour Diamonds from Containers
-In a graph $G$ represented by a GLOVER, for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
+Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). for any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
 $u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
-The following arcs hold:
-1. There are forward arcs $x\to y$ and $y\to u$, 2. There are forward arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
+The following arcs exist:
+1. There are forward arcs $x\to y$ and $y\to u$,
+2. There are forward arcs $x\to w$ and $w\to u$,
+3. No back arcs exist between containers.
 Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it.
-We didn’t know it because they they did not disrupt our algorithm and they are not a hindrance to our algorithm.
+We didn’t know it because they did not disrupt our algorithm.
 xywu
 Lemma 9.3.
-Seymour Diamonds and Decreasing Path Algorithm Let $G$ be an oriented graph represented by the GLOVER, in an attempt to prevent a degree during node, where we are applying the DSP path algorithm.
+Seymour Diamonds and Decreasing Path Algorithm
+Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP).
+We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm.
 Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond.
-Then, the presence of the Seymour diamond $D$ does not disrupt the DSP path algorithm.
+Then, the presence of the Seymour diamond $D$ does not disrupt the DSP algorithm.
 Proof.
 A Seymour diamond consists of two vertices, $x$ and $u$, connected by two parallel paths via $y$ and $w$.
-  • The DSP path algorithm can utilize one of these paths while disregarding the other, as both connect $x$ and $u$, thus maintaining degree constraints.
-  • The presence of redundant paths does not diminish the outgoing neighbors of any node, ensuring that the minimum degree condition holds throughout the process.
-  • Consequently, these additional connections offer alternative routes without introducing cycles or otherwise disrupting the algorithm’s operation.
-None of these case are a deterrent to the execution of the DSP Path algorithm. ∎
+The DSP algorithm can utilize one of these paths while disregarding the other. Both paths connect $x$ and $u$, thus maintaining degree constraints.
+The presence of redundant paths does not diminish the outgoing neighbors of any node. This ensures that the minimum degree condition holds throughout the process.
+Consequently, these additional connections offer alternative routes without disrupting the algorithm’s operation.
+None of these cases are a deterrent to the execution of the DSP algorithm.
 Lemma 9.4.
-Interaction of Minimum Degree Nodes and Transitive Triangles Let $G$ be an oriented graph represented by a GLOVER, containing a minimum degree node $v_{0}$ with degree $d$.
-Consider a transitive triangle formed by nodes $z$, $y$, and another vertex $x$, where $z$ and $y$ are part of this triangle.
-The following cases are possible:
-  • If $d^{+}(y)=d$ and $d^{+}g(z)=d$
-  • If $d^{+}(y)=d$ and $d^{+}(z)>d$
-  • If $d^{+}(y)>d$
-  • If $d^{+}(y)<d$
+Interaction of Minimum Degree Nodes and Transitive Triangles
+Suppose $G$ is an oriented graph represented by a GLOVER.
+Let $v_{0}$ be a minimum degree node with degree $d$.
+Consider a transitive triangle formed by nodes $z$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(z)=0$.
+Then the following cases are possible:
+  • $d^{+}(y)=d$ and $d^{+}g(z)=d$
+  • $d^{+}(y)=d$ and $d^{+}(z)>d$
+  • $d^{+}(y)>d$
+  • $d^{+}(y)<d$
 Proof.
-  • Case 1: If $d^{+}(y)=d$ and $d^{+}(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections.
-For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$.
+  • Case 1: $d^{+}(y)=d$ and $d^{+}(z)=d$:
+    Both $y$ and $z$ have exactly $d$ outgoing connections.
+For the triangle to be transitive, they must not share all neighbors.
+This means that there exists at least one neighbor of $y$ that is not a neighbor of $z$.
 Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
-  • Case 2: If $d^{+}(y)=d$ and $d^{+}(z)>d$: $z$ covers some neighbors of $y$, but since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
-  • Case 3: If $d^{+}(y)>d$: $y$ has more outgoing connections than required.
+  • Case 2: $d^{+}(y)=d$ and $d^{+}(z)>d$:
+    Then $z$ may cover some neighbors of $N^{+}(y)$. Since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
+  • Case 3: If $d^{+}(y)>d$:
+    $y$ has more outgoing connections than required.
 It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
-  • Case 4: If $d^{+}(y)<d$: This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections.
-Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
+  • Case 4: If $d^{+}(y)<d$:
+    This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
 Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs.
-Next we will blend everything together into our main results.
+Next, we will blend everything together into our main results.
 10 Main Theorem: Seymour’s Conjecture is True
 10.1 Introduction
-Here is where we sum up all the lemmas we have derived from our analysis on oriented graphs. We will state the final algorithm and some final results.
-So far, this battle has been a battle between a catalyst - the chosen minimum degree node, and the bottleneck - the Decreasing Sequence Property (DSP).
-We have proved some important things - Back arcs cannot exist; interior degrees must double inside each container, the exterior degrees decrease in size; interior degrees proportionally increase in size.
-There is a limit though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
+Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some final results.
+So far, this battle has been a battle between a catalyst, the chosen minimum degree node and the bottleneck, the Decreasing Sequence Property (DSP).
+We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
+There is a limit, though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
 We first restate the Lemma 6.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
 Lemma 10.1.
-Let $G$ be an oriented graph represented by a GLOVER, in an attempt to prevent a degree doubling node, with all nodes satisfying the Decreasing Sequence Property (DSP), where $v_{0}$ is a minimum degree node in $G$, and all nodes are placed into containers.
-If there are an back arcs in $G$, then a node’s degree will double.
+Let $G$ be an oriented graph represented by a GLOVER, with all nodes satisfying the Decreasing Sequence Property (DSP).
+Let $v_{0}$ be a minimum degree node in $G$.
+We are attempting to prevent a degree-doubling node.
+If there are back arcs in $G$, then a node’s degree will double.
 10.2 Theorem Statement
@@
 Theorem 10.1.
-Let $G$ be an oriented graph represented by a GLOVER with all nodes the Decreasing Sequence Property (DSP).
-If all nodes in $G$ are placed into containers, where interior degrees are maximized, then Seymour’s conjecture holds true.
-Specifically, a node in $G$ will have its exterior size decrease to less than 2.
+Let $G$ be an oriented graph represented by a GLOVER with all nodes having the Decreasing Sequence Property (DSP).
+Let $v_{0}$ be a minimum degree node where we set up our partition.
+Then Seymour’s conjecture holds true.
+Proof.
+We have set up the containers by a GLOVER data structure.
+There are no back arcs, else we have a degree-doubling node.
+Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
+These lemmas are what allowed us to set up and use lemma 5.6 (Weak Decreasing Exteriors Lemma).
+That lemma gives us a decreasing set of positive numbers.
+This set is bound to reach the size of two.
+When it does, one of the nodes in that set will have all of its arcs to the next container (thereby being exterior arcs).
+This will cause a parent of that node’s degree to double.
 Algorithm 1 Algorithm for Analyzing Degree Doubling in Containers
@@
 Corollary 10.1.
-Degree Doubling by Container Density In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},C_{2},...,C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
+Degree Doubling by Container Density
+In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},\allowbreak C_{2},\allowbreak\ldots\allowbreak C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
 There are two possibilities:
-  1. 1.
-    As the containers shrink in size and the interior degrees grow in size, we reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
-  2. 2.
-    For small minimum degree nodes $v_{0}$, The container’s shrinkage is not fast enough to cause this collision.
+  • Case 1: The containers continue to shrink in size.
+Simultaneously, the interior degrees continue to grow in size.
+We reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
+  • Case 2: For small minimum degree nodes, the container’s shrinkage is not fast enough to cause this collision.
 Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes.
 This was represented by the yin and yang effects early on.
-We see it full circle here.
-Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbor neighbors represented the size of the containers.
-In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur.
+We see it come full circle here.
+Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbors represented the size of the containers.
+In small graphs like 4 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur.
 To understand the reasoning behind this, understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1.
 The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$.
-So for small graphs, we are just not going out far enough to catch up.
-For larger graphs though, we do start to see this.
+So, for small graphs, we are just not going out far enough to catch up.
+For larger graphs, though, we do start to see this.
 Corollary 10.2.
-[…]
-$i=\frac{k-1}{2}$
-where $k$ is the minimum degree of $G$. For even values of $k$, the complete graph $K_{n}$ will not occur on any actual even number of nodes but may appear in the interval between two odd numbers.
+Occurrence of Last Dense Container
+Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$.
+Then there will be a container that reaches maximal density.
+Further, if that container is $C_{i}$, it will be $i$-out-regular.
 Proof.
-We begin by restating some previously shown examples:
-  • For $k=3$: $K_{3}$ (triangle) occurs in $C_{1}$.
-  • For $k=5$: $K_{5}$ occurs in $C_{2}$.
-  • For $k=7$: $K_{7}$ occurs in $C_{3}$.
-These observations hold because the GLOVER structure ensures that each container $C_{i}$ increases in regularity and density as the degrees of nodes double, creating the conditions necessary for complete graphs to emerge.
-For even values of $k$, such as $k=4$ or $k=6$, the symmetry and density required for a complete graph $K_{n}$ cannot form directly at any container level.
-This is because the degree-doubling effect and the odd-number dependency of $K_{2i+1}$ mean that the requirements for even $n$ are never satisfied in full.
-Instead, it transitions through intermediate states between odd complete graphs, and complete graphs may appear to ”exist” in a fractional or approximate form between $K_{2i-1}$ and $K_{2i+1}$.
-The relationship $i=\frac{k-1}{2}$ reflects the interplay between the minimum degree $k$ and the growth of density in containers.
-This relationship confirms that the GLOVER structure consistently generates complete graphs $K_{2i+1}$ in odd-numbered increments, with no direct occurrence for even $n$.
-This emergence of orientations of complete graphs as containers decrease in size adds a layer of beauty to the Seymour Second Neighborhood Conjecture.
-There was never a question of these arising.
-Some questioned multiple nodes.
-Some questioned zero nodes.
-A complete graph occurrence though is a thing of beauty and it helps give rise to another result.
+We begin by stating lemma 6.1 (Regular Interior Degrees).
+While these are called interior degrees, these are the out-degrees of the nodes in a container.
+This lemma says that every container must have regular out-degrees.
+The second part of this proof is lemma 8.2 (Strong Exterior Decreasing), which says that in container $i$, the size of the exterior is at most $d^{+}(v_{0})-i$.
+Lemma 8.3 (Interior Degree Lower Bound) tells us that each node must have degree $i$ in this container.
+Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree none.
+At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs.
+This is progressively getting to a denser container.
+As Corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever.
+There must be a last, most dense container.
+The Seymour Second Neighborhood Conjecture is made more beautiful by the appearance of these orientations of out-regular graphs as containers get smaller.
+Not only has the conjecture been proven true, but with an inherent symmetry and order.
+There is a visual pattern emerging inside these containers that is not just beautiful to graph theorists, but there is a structure that is easily appealing to non-mathematicians.
 Corollary 10.3.
-Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the complete graph occur in container $C_{i}$.
+Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the most dense graph in container $C_{i}$.
 Then every node in container $C_{i}$ has its degree doubled in the representation.
@@
 Now there are many.
-Second is the way we find it. It’s the presence of an oriented complete graph as these containers decrease in size. There’s an element of mathematical beauty there.
+Second is the way we find it. It’s the presence of an oriented regular graph as these containers decrease in size. There’s an element of mathematical beauty there.
 Then third we can look at the applications.
-We may be looking for the person who has the most influence, but this is saying that there may not be just one person there may be a whole set of people, a tightly connected group.
-10.3 Decreasing Sequence Property Path Algorithm
-We will now give a pseudocode algorithm to identify the node whose out-degree doubles in the square of an oriented graph.
-This is the culmination of the lemmas, theorems, and corollaries discussed throughout the paper, providing a cohesive framework for finding degree-doubling nodes.
-  • The containers reach a size two causing a node in the previous container to have its degree double, as stated in Theorem 10.2 (Degree Doubling by Exterior Decreasing).
-  • The interior degree required for nodes inside containers will be greater than the possible for a container of that size causing a node in the previous container to have its degree double, as stated in Corollary 10.1
-In both these cases we are able to find a node whose degree doubles.
-Here is the algorithm in pseudocode. Input: An oriented graph $G$ with minimum out-degree $d$.
-Output: A node $v$ in $G^{2}$ whose out-degree at least doubles.
-Algorithm:
-1.
-Initialize the Graph Structure: - Identify the minimum degree node $v_{0}$ in $G$.
-- Partition $G$ into containers $C_{1},C_{2},\ldots,C_{d}$, where each container $C_{i}$ groups nodes by their distance $i$ from $v_{0}$.
-2.
-Check for Back Arcs: - For each container $C_{i}$, check for arcs leading to any container $C_{j}$ where $j<i$.
-- If a back arc exists from a node $u_{i}\in C_{i}$ to a node $v_{j}\in C_{j}$, record $u_{i}$ as a degree-doubling node by Lemma 7.2 (Interior Degree Doubling) and Lemma 6.5 (Non-Existence of Back Arcs).
-- If no back arcs exist, proceed with Step 3.
-3.
-Establish Connectivity Properties: - For each node $v$ in container $C_{i}$, define: - Interior neighbors $\text{int}(v)$: Nodes within the same container.
-- Exterior neighbors $\text{ext}(v)$: Nodes in the next container.
-4.
-Iterative Degree-Doubling Check: - For each container $C_{i}$: - Calculate the out-degree of each node in $C_{i}$.
-- If a node in $C_{i}$ has fewer second neighbors than first neighbors (DSP violation), record it as a potential degree-doubling node.
-5.
-Decreasing Sequence Property Verification: - Verify that the size of each exterior set $\text{ext}(v)$ is strictly less than the size of the corresponding interior set $\text{int}(v)$ in $G^{2}$.
-- If not, apply the conditions set in Lemma 7.2 to ensure any doubling in interior degrees results in a matching exterior degree.
-6. Termination and Output: - Return the identified node(s) with doubled degrees as proof of degree-doubling in $G^{2}$.
+We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group.
+Theorem 10.2.
+Suppose we have a graph $G$.
+The Decreasing Sequence Property Algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ represents the number of vertices in $G$ and $|E|$ represents the number of edges in $G$.
+Proof.
+We will first show that the complexity of building the partitions has a worst-case complexity of $O(|V|+|E|)$.
+Given a graph $G$, the search for a minimum degree node that’s $O(|V|)$ nodes.
+If a priority queue is used, this can be optimized further.
+Then, to evaluate the degrees of each of those nodes is a constant factor.
+This makes the search for the minimum degree node $O(|V|)$.
+Once we have this value, we will place every node and every edge of $G$ into a preprocessing array ready to be partitioned.
+Each edge is examined, contributing $O(|E|)$ complexity, while nodes contribute $O(|V|)$.
+Each node will individually be processed based on its distance from the minimum degree node, along with the connected outgoing arcs.
+This will require $O(|V|)$ comparisons and $O(|V|+|E|)$ stores into containers.
+This means that the partitioning has a complexity of $O(|V|+|E|)$.
+Next, we will look into the complexity of the run-time analysis of the Decreasing Sequence Algorithm.
+First, remember that we proved in lemma 7.2 (Interior Degree Doubles) that every container has a node whose interior degree doubles.
+Secondly, we also note that the lemma 6.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container.
+Third, lemma 7.1 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another.
+The combination of these three lemmas let us know that every node inside a container has their interior degree double.
+This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
+What remains is to consider the complexity necessary to combine these interior degree nodes to solve the overall problem.
+Solving the full problem requires determining the number of containers necessary for this partition, which is bounded above by $O(|V|)$.
+The total number of containers cannot exceed the number of nodes, $|V|$, since containers are based on partitions.
+Processing each container’s exterior arcs and remaining possible nodes is bounded by $O(|V|+|E|)$.
+This gives the run time a complexity of $O(|V|+|E|)$.
+This complexity analysis shows that the Decreasing Sequence Property Algorithm can effectively find a degree-doubling node in polynomial time.
+Moreover, this will work for any oriented graph.
+Complexity analysis shows that it is not hindered by larger graphs.
+A further point is to consider the concept in theorem 10.2 of all nodes inside a container having their degrees double.
+It was not spoken about in the paper because the focus was entirely set on proving the SSNC, but this is another example of multiple degree-doubling nodes, just a localized instance.
 11 Conclusion and Future Work
 In this work, we explored the Seymour Second Neighborhood Conjecture in oriented graphs.
-We did so through the lenses of graph theory, combinatorics, algorithms and data structures.
-This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide and conquer paradigm.
+We did so through the lenses of graph theory, combinatorics, algorithms, and data structures.
+This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
 One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph.
 This effectively solves the Seymour Second Neighborhood Conjecture with an answer in the affirmative.
-The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree doubling node was in the graph.
-Instead the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
+The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph.
+Instead, the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
 This approach introduced a new perspective on the Seymour Second Neighborhood Conjecture.
 It was no longer framed as merely a graph theory problem.
-By utilizing the divide and conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effective be treated with computer science techniques.
+By utilizing the divide-and-conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effectively be treated with computer science techniques.
 This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists.
@@
 This problem has a long history.
-With a solution now in hand, its time to begin looking forward.
+With a solution now in hand, it’s time to begin looking forward.
 The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph.
@@
 We could have the graph represent epidemiology.
-Here the nodes be infected people, and the arcs would be one way infections.
-The question of interest would be can we implement the algorithm to find the person who is a super-spreader, and do so quickly?
-There are many more real life examples where we can apply this problem and this algorithm.
-There are fields like security, telecommunications, urban planning, neural networks.
+Here the nodes will be infected people, and the arcs would be one-way infections.
+The question of interest would be, can we implement the algorithm to find the person who is a super-spreader and do so quickly?
+There are many more real-life examples where we can apply this problem and this algorithm.
+There are fields like security, telecommunications, urban planning, and neural networks.
 There are too many to name in this short paper.
@@
 This data structure allowed us to get a lot done on the conjecture.
-Many of the things inherent to SSNC made it easier for GLOVER - the well ordering and the dual metrics of distance and degrees.
-Future research and development should explore questions of complexity, scalability.
+Many of the things inherent to SSNC made it easier for GLOVER—the well ordering and the dual metrics of distance and degrees.
+Future research and development should explore questions of complexity and scalability.
 Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
@@
 By introducing containers, another way of doing induction on graphs has been introduced.
-The concept of divide and conquer, which has dominated much of computer science, shows up in this paper as well.
+The concept of divide-and-conquer, which has dominated much of computer science, shows up in this paper as well.
 We are able to treat these containers independent of one another and solve their problems locally before bringing them back to the global problem.
-This shows the how versatile the GLOVER data structure is moving forward.
+This shows how versatile the GLOVER data structure is moving forward.
 What makes GLOVER special are things like total order on containers.
-Now there is a two way ranking of nodes, as opposed to traditional one way rankings like lexicographical sorting.
-This agrees with many real world systems where we have dual metrics, that are often competing, like price and performance.
+Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting.
+This agrees with many real-world systems where we have dual metrics that are often competing, like price and performance.
 This is hard to measure on a single scale.
@@
 When we take a course on graph theory, oriented graphs are probably not on the course load.
-Second neighbors or squares of graphs are probably not either.
-To fully grasp this problem though we need to be able to delve into these definitions through visualizations.
+Second neighbors or squares of graphs are probably not, either.
+To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
 For example, research on the Seymour Second Neighborhood Conjecture led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs.
-The paper addressed how these terms were intertwined with one another.
-One example is where the paper says that there is a monotonically decreasing sequence.
+The paper addressed how these terms were intertwined with each other.
+An example is where the paper says that there is a monotonically decreasing sequence.
 James Robert Brown’s book Philosophy of Mathematics speaks of the importance of visualizations in mathematics.
-Proofs are still the foundation, but sometimes visualizations can serve as an inspiration or a spark for those proofs.
+Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
 This realization motivated the creation of a complimentary website.
-The site serves as an interactive tool to understand research.
-It also has things like mathjax installed to read lemmas, definitions and proofs from this paper.
-Where possible, things like HTML canvas and JavaScript D3.js told are used to visualize these concepts and help improve understanding.
-Readers are encouraged to explore the website for additional insights and examples at:
-https://glovermethod.com
-This is an evolving project. Additional examples and applications will continuously be added
-Readers are invited to explore the website. Please provide feedback on any aspects of the site or the content for improvements.
+The site serves as an interactive tool to understand the research.
+It also has things like mathjax installed to read lemmas, definitions, and proofs from this paper.
+Where possible, things like HTML canvas and JavaScript D3.js are used to visualize these concepts and help improve understanding.
+Readers are encouraged to explore the website for additional information and examples at: https://glovermethod.com This is an evolving project.
+Additional examples and applications will continue to be added.
+Readers are invited to explore the website.
+Please provide feedback on any aspects of the site or content for improvement.
 Our goal here was to help users connect with the Seymour Second Neighborhood Conjecture.

### v7

@@
 Seymour’s Second Neighborhood Conjecture claims that there will always exist a node whose out-degree doubles in the square of an oriented graph.
-This paper first presents a data structure, GLOVER (Graph Level Order), which is a partitioning of the nodes into a total order of neighborhoods.
-This data structure allows us to introduce a total order on oriented graphs and construct a decreasing sequence of subsets.
-This decreasing sequence of subsets is what ensures that no counterexample can exist.
-Further, this algorithmic approach lets us know where the node is.
-Later findings extend the results to multiple nodes.
-This is true for all oriented graphs.
-These findings also have practical applications in network optimization and algorithm design.
+In this paper, we first present a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers.
+This data structure establishes a well-ordering on oriented graphs and allows for the construction of a decreasing sequence of subsets of nodes.
+This sequence proves the non-existence of counterexamples to the SSNC and precisely identifies the required node.
+Further, our approach finds the occurrence of dense regular graphs inside containers.
+This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property.
+Beyond theoretical implications, the algorithm and data structure have practical applications in network optimization and algorithm design.
 1 Introduction
-Seymour’s Second Neighborhood Conjecture (SSNC) was originally proposed by Paul Seymour in 1990.
-It asks the question of the existence of a node in an oriented graph whose out-degree at least doubles in the graph’s square.
-The conjecture was first published by Nathaniel Dean and Brenda J.
-Latka in 1995 [4], where they posed a similar conjecture specific to tournaments.
-This version of the conjecture was first proven by Fischer [8], with a second proof provided by Havet and Thomasse [10].
+The Seymour Second Neighborhood Conjecture (SSNC), proposed by Paul Seymour in 1990, posits that in every oriented graph, there exists a node whose out-degree in the graph’s square is at least twice its out-degree in the original graph.
+This conjecture was first formally published by Dean and Latka in 1995 [4], where they proposed a related version specific to tournaments.
+The tournament case was proved by Fisher [8], with an alternative proof later provided by Havet and Thomassé [10].
 Conjecture 1.1.
 (Seymour’s Second Neighborhood Conjecture).
-For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}$ represent the first and second neighborhoods of $v$ respectively.
-This work generalizes an approach initially developed for the minimum degree 7 case (outlined in Glover [9]).
-Here, we hope to tackle the entire class of oriented graphs based on the structure laid out in that paper.
-The Seymour Second Neighborhood Challenge is a difficult graph theory problem.
-The problem itself asks us to focus on the question of oriented graphs and second neighbors.
-There have been several papers that have looked into special cases.
-Kaneko and Locke [11] showed that the conjecture was true for graphs with a minimum degree of at most six.
+For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}(v)$ denote the first and second neighborhoods of $v$, respectively.
+Despite decades of research, the SSNC remains one of graph theory’s challenging open problems.
+Special cases of the conjecture have been explored extensively.
+Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six.
 Brantner et al.
-[1] showed that, if the conjecture is false, infinitely many counterexamples must exist.
-Additionally, Chen et al.
-[3] provided a lower bound for how much a node’s out-degree increases in the graph’s square.
-They concluded that every oriented graph has a vertex whose second neighborhood is at least $\lambda$ times larger than the first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
+[1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist.
+Chen et al.
+[3] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
 In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture.
-This was an extension of the work done by Bolter et al.
-[2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the Seymour Second Neighborhood Conjecture.
-The topic of $m$-free graphs is of particular interest to this paper because Daamouch’s [7] work is very similar to what we see here.
-In his paper, he introduces induced subgraphs and speaks of their roles in degree properties on vertices, which will also be a key thing in this paper.
-The Seymour Second Neighborhood Conjecture is not strictly a graph theory problem.
-Although it has historically been studied within the field of mathematics, it has practical applications across several domains, including social media, social networking, epidemiology, disease control, and many other fields.
-In the critical context of disease control, the nodes in this conjecture can represent individuals, and the orientation of the arcs symbolizes one person infecting another.
-Here, the application of SSNC becomes relevant by asking: Is there a person whose influence in their second radius of interaction (i.e., people they indirectly connect with) exceeds the number of people in their first radius (i.e., direct contacts)?
-In this paper, we will prove the Seymour Second Neighborhood Conjecture is true.
-We will do this by first assuming that it is false and attempting to construct a counterexample.
-We will explore the conditions under which no node’s out-degree doubles will double in the square of the oriented graph.
-This approach will focus initially on minimum out-degree nodes.
-We will analyze how their out-degree could fail to double.
-We will then extend our algorithm to other nodes based on their distance from the minimum out-degree node.
-The paper is organized as follows: Section 2 begins with some common terminology used in graph theory.
-Section 3 introduces our approach to the problem.
-In section 4 we look at some small examples of this style of thinking.
-Section 5 introduces the terminology and identifies the conditions under which their out-degree might fail to double.
-Section 6 examines the back arcs and their place in oriented graphs.
-Section 7 looks at the conditions that would cause an interior degree to double and its importance to the Seymour Second Neighborhood Conjecture.
-Section 8 further looks at containers under the context of density constraints.
-In Section 9, we consider transitive triangles and Seymour diamonds and show that they are not a hindrance to our algorithm.
-Section 10 presents this process, the full Decreasing Sequence Property algorithm, as well as the theorem stating its solution to the Conjecture.
+This was an extension of the work done by Botler et al.
+[2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
+The concept of $m$-free graphs is particularly relevant to this paper, as it closely aligns with the work of Daamouch [7].
+A directed graph $D$ is $m$-free if it contains no directed cycles of length at most $m$.
+In his paper, Daamouch introduces induced subgraphs and explores their role in the degree properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
+This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9].
+Our key contribution is the introduction of a novel data structure, GLOVER (Graph Level Order), which partitions nodes into a total ordering of containers.
+This structure enables the construction of a decreasing sequence of subsets of nodes, forming the basis of our algorithmic proof of the SSNC.
+Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
+Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, and algorithm design.
+For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections surpasses their direct contacts, offering insights into the spread of infections.
+Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node.
+We also analyze the complexity of our algorithm, demonstrating its practical feasibility for large-scale problems.
+The paper includes extensive supporting materials, such as detailed proofs, examples, and an interactive website to visualize key ideas, ensuring accessibility and transparency.
+The remainder of this paper is structured as follows: Section 2 introduces common terminology used in graph theory.
+Section 3 outlines our approach to the problem, while Section 4 provides illustrative examples.
+Section 5 focuses on conditions under which a node’s out-degree may fail to double, and Section 6 examines the role of back arcs in oriented graphs.
+Section 7 explores the conditions that lead to doubling of interior degrees, examining their significance in relation to the SSNC.
+It introduces a divide-and-conquer approach, where the graph is partitioned, and each subproblem is solved independently before being integrated into the broader structure.
+Section 8 explores density constraints in containers.
+Section 9 investigates the impact of transitive triangles and Seymour diamonds on our algorithm.
+Finally, Section 10 presents the full Decreasing Sequence Property algorithm and its proof of the conjecture.
 2 Graph Theory Terminology
@@
 Now that we have two different conjectures, we can investigate how they relate to one another.
-Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods, while conjecture 2.1 relates this same concept to the square graph.
+Conjecture 1.1 relates the notion of a node’s degree doubling to its neighborhoods.
+Similarly, conjecture 2.1 relates this same concept to the square graph.
 The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first.
@@
 This is why these two conjectures are seen as identical.
-We will focus our attention on Conjecture 1.1 for the reason that the conjecture 2.1 requires two different graphs - the original graph and the square graph.
-While the construction of the square graph is possible, we have found that it is not as helpful to our research at this time.
-Instead, focusing on the neighborhood analysis brought many insights and characteristics that helped to develop our algorithm.
+We will focus our attention on Conjecture 1.1 and Conjecture 2.1.
+Conjecture 1.1 is easier to work with initially and understand why a node’s degree will double in the square graph.
+Concentrating on neighborhood analysis yielded numerous insights and features that aided in the development of our algorithm.
 Next, we turn our attention to the term degree-doubling nodes.
 By this term, we aim to get to the heart of conjecture 1.1.
-That conjecture is looking for a node whose degree at least doubles, and we are doing so by looking at the first and second neighborhoods of nodes and finding one whose second neighborhood is at least as large as the first, as the conjecture asks for.
+That conjecture is looking for a node whose degree at least doubles.
+In order to do this, we examine the node’s first and second neighborhoods.
+Then identify a node whose second neighborhood is at least as large as the first, as required by the conjecture.
 A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree.
@@
 This partitioning enables us to leverage the divide-and-conquer paradigm.
-No nodes are being removed to form these partitions.
+No nodes or arcs are being removed to form these partitions.
 This is a conceptual idea.
-Instead, think of it as putting a box around the partition.
+Instead, think of it as putting a box or circle around the nodes and arcs.
 Now the nodes inside the box can work with the other nodes inside the box and solve their problem independent of the other partitions.
-These partitions are still connected to one another, but thinking of them in this way allows us to do more.
-This represents a stark contrast to most previous approaches to the Seymour Second Neighborhood Conjecture.
+These partitions are still connected to one another and will have to reconnect to solve the global problem.
+However, thinking of them in this way allows us to use their independent nature.
+This represents a stark contrast to most previous approaches to the SSNC.
 Many of these previous approaches were conceptual or theoretical.
-They would say only that a node whose degree doubled existed, but the notion of a path to that node was not part of the paper.
+Papers would mention the existence of a node whose degree doubled.
+The notion, though, of a path to that node was not part of the paper.
 Secondly, most of these papers focused on local search techniques that started at or near the desired degree-doubling node.
 That presents a problem because oriented graphs can be varying in nature.
-This is the first exhaustive search technique on the SSNC problem.
-When speaking of an exhaustive search algorithm, complexity becomes a factor and questions of how the algorithms will handle large graphs compared to smaller graphs.
-This should not be a problem as we will see in the following sections; the slowest part of this algorithm is building the data structure.
-That will require partitioning the nodes and edges of the graph, in a worst case complexity of $O(V+E)$.
-By partitioning the graph, even larger graphs, often with higher minimum degree nodes and thus more partitions, are able to be broken into sub-problems and solved independent of one another.
-These subproblems are each their own instances of the Seymour Second Neighborhood Conjecture, just a localized version.
-What the partition is doing is breaking a node’s individual degree into two parts-the degree within the container and the degree to the next container.
+This is the first global search technique on the SSNC problem.
+Complexity enters the picture when discussing global search algorithms.
+This raises concerns about how well they will handle larger graphs in comparison to smaller ones.
+This should not be a problem, as we will see in the following sections.
+The slowest part of this algorithm is building the data structure.
+That will require partitioning the nodes and edges of the graph, in a worst-case complexity of $O(V+E)$.
+When the graph is partitioned, even larger graphs, often with higher value minimum degree nodes and thus more partitions, are able to be broken into subproblems.
+These subproblems are then able to be solved independent of one another.
+These subproblems are each their own instances of the SSNC, just a localized version.
+What the partition is doing is breaking a node’s individual degree into two parts—the degree within the container and the degree to the next container.
 The degree within the container is the subproblem that will be solved separately, and the degree to the next container is how they connect back to the global problem.
@@
 With a minimum degree node, we can partition this graph into independent subproblems.
-This is now ready for an exhaustive search algorithm.
+This is now ready for a global search algorithm.
 4 Initial Lemmas
@@
 For this section, unless otherwise noted, assume that $v_{0}$ is a node with minimum out-degree in our oriented graph $G$.
-We will also introduce some standard notation from graph theory.
-The work in this section agrees with what was already done by [11] and is not meant to serve as original content.
-It does, however, provide insight into our reasoning.
+The work in this section agrees with what was already done by [11] and is not intended as original content.
+However, it does provide insight into our reasoning.
 Example 4.1.
@@
 Lemma 4.1.
-Minimum Out-Degree $<$ 3 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double in $G^{2}$.
+Minimum Out-Degree $<$ 3 Suppose that a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
 Proof.
@@
 Then $v_{0}$ has no outgoing edges in $G$.
-In $G^{2}$, it will still have out-degree 0.
-Hence its out-degree trivially satisfies the condition of at least doubling.
+Then $v_{0}$’s first and second neighborhoods are of size 0.
+Hence, its out-degree trivially satisfies the condition of at least doubling.
 Case 2: Assume that $d^{+}(v_{0})=1$.
@@
 This means that $(v_{1},v_{2})\in G$.
-This implies that in $G^{2}$, both $(v_{0},v_{1})$ and $(v_{0},v_{2})$ exist, giving $v_{0}$ an out-degree of at least 2, thus doubling its out-degree.
+Hence $v_{2}$ is a second neighbor of $v_{0}$, thus making it at least as large as $v_{0}$’s first neighborhood.
 Case 3: Assume that $d^{+}(v_{0})=2$.
@@
 The node $v_{0}$ cannot be one of these nodes.
-Hence, in $G^{2}$, $v_{0}$ will connect to $v_{1}$, $v_{2}$, $v_{3}$, and $v_{4}$, giving it an out-degree of at least 4.
-Thus, it’s out-degree will double.
-Therefore, in all cases, $v_{0}$’s out-degree at least doubles in $G^{2}$.
+Hence $v_{0}$ will have at least two second neighbors - $v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first.
+Thus, its out-degree will double in $G^{2}$.
+Therefore, in all cases, $v_{0}$’s out-degree at least doubles.
 Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs.
-We are learning that, in order to construct our counterexample, we need our minimum degree to be at least 3.
-This will also place that same minimum value restriction on every other node in the graph.
+We are learning that, to construct our counterexample, we need the minimum degree of our graph to be at least 3.
 In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
@@
 Minimum Out-Degree 3 with Neighbor Out-Degree 0 Suppose a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree of 3.
-Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0; then the out-degree of $v_{0}$ will at least double in $G^{2}$.
+Suppose also that at least one of $v_{0}$’s neighbors has out-degree 0.
+Then the out-degree of $v_{0}$ will at least double in $G^{2}$.
 Proof.
-Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, and $v_{3}\in V(G)$.
-Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$.
-Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$.
-This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
-Since $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$, the distance-2 neighbors of $v_{0}$ in $G^{2}$ will include these additional neighbors of $v_{1}$.
-As a result, $v_{0}$ will have at least twice as many out-neighbors in $G^{2}$ compared to its out-degree in $G$.
+Assume $d^{+}(v_{0})=3$, and let $v_{0}$ be connected to nodes $v_{1}$, $v_{2}$, and $v_{3}\in V(G)$. Consider the induced subgraph $G[v_{1},v_{2},v_{3}]$.
+Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
+We know that $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$.
+This means that $v_{0}$ has at least three second neighbors.
+As a result, $v_{0}$ has at least as many second neighbors as first neighbors, and equivalently, $v_{0}$’s out-degree doubles in $G^{2}$.
+∎
 Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$.
@@
 While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$.
-Thus, in $G^{2}$, the degree of $v_{0}$ will double.
-Minimum Out-Degree 3 with Neighbors 1
+Thus, in $G^{2}$, the degree of $v_{1}$ will double.
+See lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1)
 Lemma 4.3.
-Assume that in an oriented graph, the minimal out-degree node $v_{0}$ has an out-degree of 3.
+Minimum Out-Degree 3 with Neighbors 1 Assume that in an oriented graph, the minimal out-degree node $v_{0}$ has an out-degree of 3.
 Additionally, suppose that in the induced subgraph of its neighbors, every neighbor has an out-degree of 1.
-The out-degree of $v_{0}$ or one of its neighbors will then at least double in $G^{2}.$.
+Then the out-degree of $v_{0}$ or one of its neighbors will then at least double in $G^{2}$.
 Proof.
@@
 Therefore, assume that all three nodes have an out-degree of exactly 3 in $G$.
-Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}].$ If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$.
-Therefore, assume they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
-This reduces the situation to Case 3 in Lemma 4.1 Minimum Out-Degree $<$ 3, where $v_{1}$ acts similarly to how $v_{0}$ did in that case.
+Each node $v_{1},v_{2},v_{3}$ must connect to two nodes in $G-G[v_{1},v_{2},v_{3}]$.
+If the set of nodes to which $v_{1},v_{2},v_{3}$ connect in $G-G[v_{1},v_{2},v_{3}]$ contains more than 2 distinct nodes, the out-degree of $v_{0}$ will at least double in $G^{2}$.
+Therefore, assume that they all connect to exactly 2 nodes outside of $G[v_{1},v_{2},v_{3}]$, say $v_{4}$ and $v_{5}$.
+This reduces the situation to Case 3 in Lemma 4.1 (Minimum Out-Degree $<$ 3), where $v_{1}$ acts similarly to how $v_{0}$ did in that case.
 There is at most one edge between $v_{4}$ and $v_{5}$.
 One of these nodes, say $v_{4}$, must connect to three additional nodes, $v_{6},v_{7}$, and $v_{8}$.
-As a result, in $G^{2}$, we will have the edges $(v_{1},v_{6})$, $(v_{1},v_{7})$, and $(v_{1},v_{8})$, meaning that the out-degree of $v_{1}$ will at least double in $G^{2}$.
+As a result, $v_{4}$ will have three second neighbors, and that is at least as many as its first neighbors.
+This means that degree doubles in $G^{2}$.
 ∎
-Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) provides an exhaustive search strategy, which is a question we shall revisit throughout this paper.
-It must prevent a node’s out-degree from doubling in order for that to function.
+Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) provides a global search strategy, which is a question we shall revisit throughout this paper.
+It must prevent a node’s out-degree from doubling for that to function.
 In this instance, we observe that the degree of the starting node in the example $v_{0}$ is not doubled.
 We are compelled to look for the next node instead.
-In this case, the degree of the neighbor of $v_{0}$ $v_{1}$ doubles.
-We did not require the ”right” starting node; therefore, this is a crucial distinction.
-Not because of its possible proximity to the double degree node, but rather because of its capacity to divide the nodes in the graph, we selected a minimal degree node.
-We shall use double arrows to show that a node or collection of nodes connects to every node in the next subgraph for graphs similar to those in Lemma 4.2 throughout the rest of this work.
-Because all of the nodes in a shared induced subgraph relate to the same set of nodes, this is possible.
+In this case, the degree of the neighbor of $v_{0}$, $v_{1}$, doubles.
+We did not require the ”right” starting node.
+Therefore, this is a crucial distinction.
+Not because of its possible proximity to the degree-doubling node.
+Instead, we chose a minimal degree node due to its ability to split the graph’s nodes.
+We shall use double arrows to show that a collection of nodes connects to every node in the next induced subgraph as shown in example 4.2.
+Because all nodes in a shared induced subgraph relate to the same set of nodes, this is possible.
 These lemmas offer a theoretical framework and insight into our rationale. These findings will be expanded upon and referenced in the upcoming sections.
 5 Theoretical Foundation
-We will introduce some new terms to help comprehend this section.
+We will introduce some new terms to help with the rest of this paper.
 The partitions we discussed earlier will now be formally referred to as containers.
@@
 A container $C_{i}$ is the set of nodes at distance $i$ from $v_{0}$, together with the arcs connecting those nodes, given a minimum degree node $v_{0}$. Formally speaking,
-The formula $C_{i}=\{u\in V(G)\text{.s.t.
-}A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where $V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
-Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7]. The neighboring containers (see image), which will be crucial in the proofs, are the main distinction here.
-We will investigate the structure of oriented graphs from two angles by using the idea of containers. Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node.
+The formula
+$C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where
+$V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
+Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7].
+The neighboring containers (see example 5.1), which will be crucial in the proofs, are the main distinction here.
+We will investigate the structure of oriented graphs from two angles by using the idea of containers.
+Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node.
 Containers are seen based on their distance from the minimum degree node.
-  • Nodes are seen based on their position relative to containers. They are either before a container, in a container, or after a container.
-Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve.
-A graph $G$ was created in which the out-degree of $v_{0}$ does not double in $G^{2}$.
-A simple graph of degree three corresponds to the lowest cycle that can exist in an undirected graph.
+Nodes are seen based on their position relative to containers.
+They are either before a container, in a container, or after a container.
+Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve.
+A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in $G^{2}$.
+A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph.
 Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$.
-A closer examination of this case reveals that the out-degree of a node that is an out-neighbor of $v_{0}$ is (at least) twice in this situation.
+If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double.
 This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
-However, we will need to define new concepts in order to proceed with this line of thinking.
-To prevent a node’s out-degree from doubling, we must first find a term.
+However, we will need to define new concepts to proceed with this line of reasoning.
+To prevent a node’s out-degree from doubling, we must first define a term.
 The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$.
@@
 If the size of a node $x\in V(G)$ in an oriented graph $G$ is strictly bigger than the size of its second out-neighborhood, that is, $|N^{+}(x)|>|N^{++}(x)|$, then $x$ is said to have the decreasing sequence property.
-Two new ideas have been presented thus far: the decreasing sequence property (DSP) and the minimal degree node.
-A crucial starting point for our approach and degree values is a minimal degree node, which is the node with the fewest connections.
+Two competing new ideas have been presented, the decreasing sequence property (DSP) and the minimal degree node.
+A crucial starting point for our approach and degree values is a minimal degree node.
 It acts as the anchor by being chosen as the beginning point.
 The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node.
-It is serving as a stimulant.
+It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
 In contrast, the DSP introduces constraints on connectivity.
 This concept is akin to a bottleneck.
+A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP.
 Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
@@
 Total Order Induced by Containers Relative to the Minimum Degree Node
-Let $G$ be a graph with a minimum-degree-node $v_{0}$.
+Let $G$ be a graph with a minimum-degree node $v_{0}$.
 Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$.
@@
 While imposing order on our graph is a solid first step, the containers do more than just enforce overall order.
-In essence, they enable group comparisons between nodes that are equally spaced from a reference node, $v_{0}$.
-A typical total order, such as lexicographical order, does not take this into consideration.
+In essence, they allow for the grouping of comparisons between nodes that are equally spaced from a reference node, $v_{0}$.
+A typical total order, like lexicographical order, does not take this into consideration.
 This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
@@
     • $u$ is equal to $v$ (denoted as $u=v$).
-  • Leveled Structure: The elements of $S$ are partitioned into a leveling of containers.
-$C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$ Each container itself forms a subset of $S$.
-This maintains a total order.
+  • Leveled Structure:
+    • The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$
+    • Each container itself forms a subset of $S$.
     • The containers are sequentially ordered relative to each other.
-This establishes a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
-  • Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$.
-This avoids trivial total orderings where all elements are isolated in singleton containers.
   • Comparability within Containers: For any two elements $u,v\in C_{i}$, their order can be determined based on a specific criterion.
   • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
-This enforces that the order of elements between containers aligns with the order of the containers themselves.
 This order means that we have order in two ways.
@@
 This ordering must also be consistent within the containers.
-The notion of a non-triviality requirement is to distinguish this ordering from other traditional total orderings on graphs.
 Lemma 5.2.
@@
 If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$.
-In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed.
+In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed.
 We will now give terms to these concepts.
@@
 Interior neighbors will be important to our study on SSNC and our search for a counterexample.
-In particular, we saw things about interior neighbors in 4 (Minimum Out-Degree 3 with Neighbors 1).
+In particular, we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1).
 The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
@@
 The number of exterior neighbors is the exterior degree of $y$ with respect to $x$.
-We contrast interior neighbors with exterior neighbors, which reveal how a container interacts with adjacent containers.
-This interaction is again critical for the Seymour Second Neighborhood Conjecture, as the nodes within the current container depend on the connections to the next container.
+We contrast interior neighbors with exterior neighbors, which reveals how a container interacts with adjacent containers.
+This interaction is again critical for the SSNC, as the nodes within the current container depend on the connections to the next container.
 As an exercise to visualize interior and exterior neighbors, consider example 5.2.
@@
 We would like to continue this.
-In order to do so, though, we need to consider the option of back-arcs and see how they might hurt our algorithm.
+To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm.
 These have the possibility of altering a container’s distance to the minimum degree node.
@@
 Leveled Structure: The elements of $S$ are partitioned into levels of containers $C_{1},C_{2},\ldots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
-Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$.
 Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion.
@@
   • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
-To prove that any oriented graph can be represented in a GLOVER, we’ll need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
+Next we will prove that any oriented graph can be represented in a GLOVER.
+We will need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
+This includes maintaining a total order of the containers, allowing for the comparison of any two nodes, establishing a total order of those nodes within the containers, and establishing interior and exterior neighbors.
 Theorem 5.1.
-Assume we have an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER.
+Given an oriented graph $G$. Assuming that there are no back arcs in $G$, this oriented graph can be represented by a GLOVER data structure.
 Proof.
-To establish a Graph Level Order, we first need a minimum degree node. We can obtain this by taking the minimum value on the degree vector of nodes. We will call that in the node $v_{0}$.
-Now that we know the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
-Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes whose distance from the node $v_{0}$ is equal.
+To establish a graph level order, we first need a minimum degree node.
+So we take the minimum value on the degree vector of nodes.
+We will call that node $v_{0}$.
+Knowing the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
+Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$.
 These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
-  • Total Order on Elements within $G$ Because nodes are partitioned based on their distance from $v_{0}$, we have a total order on the elements of $G$.
+  • Total Order on Elements within $G$ Nodes are partitioned based on their distance from $v_{0}$.
+This is a total order on the nodes of $G$.
 For any two nodes $u$ and $v$, this total order is determined by their containers.
-  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ form ranked levels, each corresponding to nodes at a specific distance from $v_{0}$.
-  • Non-triviality Requirement - Given that $v_{0}$ is a minimum degree node, there is at least one container (specifically, $C_{1}$) that has more than one node.
-  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, the order can be determined by their relationship to $v_{0}$.
+  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
+  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
   • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
   • Interior and Exterior Neighbors:
-    • Nodes $u$ and $v$ of the same distance from $v_{0}$ would be in the same container and thus be interior neighbors.
+    • Nodes $u$ and $v$ of the same distance from $v_{0}$ are in the same container and thus are interior neighbors.
     • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1.
@@
 What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers.
-This ordering is also ordering the nodes within the containers, so the GLOVER data structure is a multi-level total order.
+This ordering is more than just representative, though.
+The containers help to organize the nodes.
+What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from the chosen minimum degree node.
+This is a clear distinction from standard graph theory.
+Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space.
+However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from the chosen minimum degree node.
 The one thing that could disrupt this total ordering is a back arc.
@@
 By assuming no back arcs, we ensure that each container maintains its place in the total ordering.
+However, it is a bold assumption to claim that oriented graphs have no back arcs.
+Oriented graphs come in various types; surely some exist with back arcs.
+This is an important claim to investigate, and the next section will look into it in depth.
+We will prove that these back arcs do not hurt the algorithm.
+Instead, they only lead to the faster path to a degree-doubling node.
+In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node.
 Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering.
+Mathematical induction is a proof technique that uses inference to prove statements.
+Remember that a well ordering is a concept in discrete mathematics where a total ordering on a set $S$ has the property that every non-empty subset of $S$ has a least element.
 Lemma 5.3.
-GLOVER Supports Induction
-In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
+GLOVER Supports Induction In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
 Proof.
-We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
+We have already shown that the containers in GLOVER form a total order.
+Further, they are a discrete total order.
+Because of this, they are a well ordering.
 A set is considered well-ordered if every non-empty subset of that set has a least element.
@@
 Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set.
+∎
+This becomes an important concept because now, in addition to the nodes, we also have the containers as subsets of $S$.
+These containers are totally ordered and thus can also be used as a means of mathematical induction.
+We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from the chosen minimum degree node $v_{0}$.
 Example 5.3.
@@
 Lemma 5.4.
-Interior Degree at least 1
-Suppose we have an oriented graph $G$, represented by a GLOVER.
+Interior Degree at least 1 Suppose we have an oriented graph $G$, represented by a GLOVER.
 Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property.
-In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $N^{+}(v_{0})$ must have an interior degree of at least 1.
-That is, for $x\in N^{+}(v_{0})$,
+In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1.
+That is, for $x\in C_{1}$,
 $|\text{int}(v_{0},x)|\geq 1.$
 Proof.
-By definition, a node satisfies the decreasing sequence property if
-$|N^{+}(x)|>N^{++}(x)|$
-For $v_{0}$, the second neighbors are the nodes reachable by following an arc from $v_{0}$ to a neighbor $x$, and then from $x$ to another node $z$, i.e., $z\in N^{++}(v_{0})$.
-Since $v_{0}$ is the minimum out-degree node, we know that $d^{+}(v_{0})$ is smaller than or equal to the out-degree of any other node. Let $x\in N^{+}(v_{0})$.
-If $x$ is connected to $d^{+}(v_{0})$ or more nodes via exterior arcs, this would cause $v_{0}$’s degree to double through $x$’s neighbors, which are $v_{0}$’s second neighbors.
-If $x$ connects to $v_{0}$ through a back arc, again this would cause $v_{0}$’s degree to double through $x$’s first neighbors.
-To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
+By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$ If $x$ has $d^{+}(v_{0})$ exterior arcs, this would make $v_{0}$’s degree double because $x$’s neighbors are $v_{0}$’s second neighbors.
+To avoid this doubling, $x$ must connect to at least one interior arc.
+This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
 $|int(v_{0},x)|\geq 1$
-This argument holds for any neighbor $x\in N^{+}(v_{0})$. This ensures that every node in $N^{+}(v_{0})$ has an interior degree of at least 1.
-Thus, for all $x\in N^{+}(v_{0})$, $|int(v_{0},x)|\geq 1$.
-Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in general oriented graphs that may coincide with degree doubling.
-In particular, we see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree to double.
-Example 5.4.
-0
-$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
+This is true for any neighbor $x\in C_{1}$. This ensures that every node in $C_{1}$ has an interior degree of at least 1. Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
+Lemma 5.4 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree doubling.
+We see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree doubled.
 Lemma 5.5.
-Exterior Less than Degree
-Suppose we have an oriented graph $G$, represented by a GLOVER.
-In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i}),|\text{ext}(x,y)|<d^{+}(x)$.
+Exterior Less than Degree Suppose we have an oriented graph $G$, represented by a GLOVER.
+In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
 Proof.
-Assume, for the sake of contradiction, that there exists a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$.
-By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$
-Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$, then:
+Assume, for the sake of contradiction, that there is a node $y\in N^{+}(x)$ such that $|\text{ext}(x,y)|\geq d^{+}(x)$.
+By definition, $\text{ext}(x,y)=N^{++}(x)\cap N^{+}(y)$.
+Thus, $|\text{ext}(x,y)|\subseteq N^{++}(x)$.
+Then:
 $|N^{++}(x)|\geq|N^{+}(x)|\geq d^{+}(x),$
-suggesting that the degree of outwardness of $x$ would at least double on graph $G^{2}$. This is a contradiction to the assumption that $x$ satisfies the decreasing sequence property.
-Therefore, we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$.
+This means that the out-degree of $x$ at least doubles on the graph $G^{2}$. This contradicts the assumption that $x$ has the decreasing sequence property.
+This must be true for all $x$ in $G$, so we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
 Lemma 5.5 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 5.4 (Interior Degree at least 1), with the notable exception that Lemma 5.5 refers to any node with the DSP, whereas we only made Lemma 5.4 true for the minimum degree node.
@@
 There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
-Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an exhaustive search solution with the algorithm we are developing.
+Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an global search solution with the algorithm we are developing.
 Lemma 5.6.
-Weak Decreasing Exteriors Lemma
-Suppose that $G$ is an oriented graph represented by a GLOVER, and that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
+Weak Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a GLOVER.
+Suppose further that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
 Suppose that:
   1. 1.
-    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the subgraph $G[N^{+}(x)]^{2}$.
+    The interior degree of $y$ with respect to $x$, $|int(x,y)|$, doubles in the induced subgraph $G[N^{+}(x)]^{2}$.
   2. 2.
@@
 Proof.
-We begin by recalling the conditions for a node to have the decreasing sequence property (DSP).
+We begin by recalling the conditions for a node to have the decreasing sequence property (DSP) because it is very important for this proof.
 A node has the DSP if the size of its out-neighborhood is strictly greater than the size of its second out-neighborhood, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
@@
   1. 1.
-    The interior degree $|int(x,y)|$, doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
+    The interior degree $int(x,y)$ doubles in $G[N^{+}(x)]^{2}$, indicating that shared neighbors of $x$ and $y$ increase significantly in $G[N^{+}(x)]$.
   2. 2.
-    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$, ensuring that no back arcs exist.
-Now, suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$.
+    For all $z\in\text{ext}(x,y)$, $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$; no back arcs exist.
+Suppose for contradiction that there exists $z\in\text{ext}(x,y)$ such that $|\text{ext}(y,z)|\geq|\text{ext}(x,y)|$.
 The node $y$ has a degree that can be broken down into two components, its interior degrees and its exterior degrees.
@@
 Thus, the assumption $|\text{ext}(y,z)|\allowbreak\geq\allowbreak|\text{ext}(x,y)|$ leads to a contradiction, confirming that $|\text{ext}(y,z)|\allowbreak<\allowbreak|\text{ext}(x,y)|$ for all $z\allowbreak\in\allowbreak\text{ext}(x,y)$.
-Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$.
-The first thing to notice in this proof is the assumptions. We assume that there are no back arcs and that a node’s interior degree doubles. Both of these assumptions are important.
+Therefore, the lemma holds: for all $z\in ext(x,y)$, the exterior degree of $y$ is strictly smaller than that of $x$. ∎
+The first thing to notice in this proof is the assumptions.
+We assume that there are no back arcs from $x$’s second neighbors to $x$’s first neighbors and that a $y$’s interior degree doubles.
+Both of these assumptions are important.
 First, we will talk about the lack of back arcs.
@@
 This is one of the main points where they can be a hindrance.
-In particular, there exists a way to double count nodes as both neighbors and second neighbors if back arcs exist.
+There exists a way to double count nodes as both first neighbors and second neighbors if back arcs exist.
 We will speak more about this in the next section, but for now we have simply assumed that back arcs do not exist.
 Secondly, we assumed that the interior degree of a node doubled.
-This is similar to the notion of solving the containers separately, then joining them together again to solve the main problem.
+This is like the notion of solving the containers separately, then joining them together again to solve the main problem.
 We are assuming that this node $y$ was in another container and its interior degree doubled in that container.
 In doing so, $y$ is now in a better position to have its overall degree double, only needing to have its exterior degree double.
-The more popular representations of graphs are adjacency matrices, arraylists and edge lists.
-These suffer from the same limitations of graphs as mentioned above, They scatter the nodes along the two-dimensional plane without attempts to group into an ordering.
+Notice the wording of interior degree doubling.
+That is a separate version of the SSNC, but localized to the containers.
+The conjecture in this case is can we find a node in each container whose degree doubles.
+If so, then it gives light to this divide-and-conquer approach, where we partition the containers, solve them individually, and then combine them again for this lemma 5.6 (Weak Decreasing Exteriors Lemma).
+The more popular representations of graphs are adjacency matrices, arraylists, and edge lists.
+These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering.
 The Graph Level Order (GLOVER) improves this by partitioning the nodes in a reasonable way.
 For the SSNC, that reasonable way was the distance metric.
-In order to represent graphs in a GLOVER we would use any markup language that can encode graphs.
-There needs to be three types of data- containers, nodes, and arcs.
-The containers will be given ids, the nodes will then be assigned to their unique containers, and arcs will be determined to be interior, exterior, or back, depending on the containers of the endpoints of the arcs.
-This determination of the number of containers and the assignment of nodes to containers does not need to be declared beforehand.
-Instead, just as we have been talking about the distance of a node from a minimum degree node $v_{0}$ as the determination of container assignment, we would have the same or a similar function to assign nodes to containers.
+To represent graphs in a GLOVER, we can use any data serialization that can encode graphs, including JSON, YAML, and XML.
+There needs to be three types of data: containers, nodes, and arcs.
+The containers will be given IDs.
+The nodes will then be assigned to their unique containers.
+Arcs will also be determined to be interior, exterior, or back, depending on the endpoints of the nodes of the arcs.
+The determination of the number of containers and the assignment of nodes to containers also do not need to be declared beforehand.
+Instead, just as the data structure uses the minimum function to find a node representing the minimum degree node $v_{0}$.
+This same function will return a value $\delta=d^{+}(v_{0})$, and $\delta$ will represent the maximum number of containers.
+We would have a similar function to assign nodes to containers.
 With this container representation of the SSNC, by way of the GLOVER data structure, we can begin to visualize elements of these lemmas.
@@
 […]
-To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another but no-one follows back.
+Figure 7: Illustration of the Weak Decreasing Exteriors Lemma.
+We have a node $x$ in container $C_{i}$ connecting to three nodes in container $C_{i+1}$.
+In order to prevent $x$’s degree from doubling, we need to have $|C_{i+2}|<|C_{i+1}|$.
+To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another, but no one follows back.
 This network is then an oriented graph.
-Running the Seymour Second Neighborhood Conjecture in this context means looking for an account whose indirect sphere of influence is greater than their direct sphere of influence.
-Here, we are defining the indirect sphere of influence as those accounts they can reach through others, and we are defining the direct sphere of influence as those accounts they can directly connect with.
-Essentially, these are the same definitions as the second neighbor and first neighbor respectively that we have been working with.
-With this in mind, after finding the person who connects to the fewest accounts in this social media graph, we could set up containers just as before.
+Running the SSNC in this context means looking for an account whose indirect sphere of influence is at least as large as their direct sphere of influence.
+Here, we are defining the indirect sphere of influence as those accounts they can reach through others.
+Similarly, we are defining the direct sphere of influence as those accounts they can directly connect with.
+Essentially, these are the same definitions as the second neighbor and first neighbor, respectively, that we have been working with.
+After finding the person who connects to the fewest accounts in this social media graph, we could set up containers just as before.
 Now, lemma 5.4 (Interior Degree at least 1) says that each person in container 1 must follow another person in container 1.
-What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about, or a number of disjoint cycles where messages are spread among themselves.
+What this is saying is that, inside this first container, we will have one cycle of users where the messages are traversed about or a number of disjoint cycles where messages are spread among themselves.
 Considering example 5.5, which takes the former of these approaches.
@@
 This brings us to the next container.
-Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people - (Hugo, Isla, Jade, Kai, and Lei).
+Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people—Hugo, Isla, Jade, Kai, and Lei.
 If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3).
-From this we would conclude that a person in container 1, whose indirect sphere of influence is this persons sphere of influence, would have its first degree double and its decond degree double, implying that its overall degree doubles.
+From this we would conclude that a person in container 1, whose indirect sphere of influence is this person’s sphere of influence, would have its first degree double and its second degree double, implying that its overall degree doubles.
 Example 5.5.
 Andi
-$\Rightarrow$ BellChadDawnElseFayeGus $\Rightarrow$ HugoIslaJadeKaiLei $\Rightarrow$ H
-Figure 7: This is an example of a social media setup of the Seymour Second Neighborhood Conjecture.
-The Second Container has ten arcs and 5 nodes, which represents a perfect graph.
-This implies that in the next container $C_{3}$ (not shown), with 4 people can only have 6 arcs.
-Due to this fact, at least one person in $C_{3}$ will relate to at least five people in $C_{4}$.
-We can conclude then that at least one person in $C_{1}$ will be an influencer and have their degree double.
-Table 1: Intersection Results for Triplets of Nodes in 5.6
-Triplet(A, B, C) int(A, B) ext(B, C) Intersection Result
-[…]
-[…]
-[…]
-[…]
-[…]
+$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$ H
+Figure 8: This is an example of a social media setup of the SSNC. The second container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
 We have now defined the data structure that we will use throughout the rest of this paper.
@@
 6 Back Arcs
-With a theoretical foundation underway, we are ready to launch an investigation into back arcs.
-We will look into whether these arcs can exist in a counterexample and what other properties of oriented graphs must be true.
-It is important to restate that back arcs and exterior neighbors are distinct concepts.
+With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate whether these arcs can bring about a counterexample.
+It is important to note that the back arcs and the exterior neighbors are distinct concepts.
 Back arcs are not a subset of exterior neighbors.
-While back arcs connect to any previous container, exterior neighbors are limited to connecting with the next container.
-Additionally, exterior neighbors do not represent all possible external neighbors.
-Rather, an exterior neighbor ext(x, y) specifically refers to nodes at a distance of two from $x$ that are also neighbors of $y$.
+While back arcs may connect to any previous container, exterior neighbors are limited to connecting with the next container.
+In addition, exterior neighbors do not represent all possible future containers.
+Rather, an exterior neighbor $ext(x,y)$ specifically refers to nodes at a distance of two from $x$ that are also distance one from $y$.
 If we look at Example 4.2 again, we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$.
 Similarly, we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
-One of the things we notice about Lemma 4 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4 allowed for a cycle to exist in the out-neighbors of $v_{0}$.
+One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$.
 In such a situation, all $x\in N^{+}(v_{0})$ can still have $d^{+}(x)\geq d^{+}(v_{0})$ without the out-neighbors of $x$ causing $v_{0}$’s out-degree to double.
-This next lemma formalizes that concept.
+The next lemma formalizes that concept.
 Example 6.1.
@@
 Proof.
-Assume that $u$ and $v$ are two nodes in container $C_{i}$, and that they share a parent $v_{p}$ in container $C_{i-1}$. Suppose that $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$
-In order to preserve the necessary degree limits in the graph, $v$ will need fewer external connections if its inner degree is greater than $u$.
-$u$, on the other hand, will need more external connections to make up for its lack of inner connections.
-As a result of the imbalance in internal degrees, the external connections of $u$ are essentially raising the second neighbors of degree of $v_{p}$.
-As a result, the exterior degree of $v_{p}$ will double to make room for the additional connections that $u$ in the graph requires.
-Thus, these different interior degrees of $u$ and $v$ in container $C_{i}$, will cause the exterior degree of $v_{p}$ in $C_{i-1}$ to double. This leads to a contradiction, so the claim holds. ∎
-Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within containers of these counterexamples.
-In particular, it shows that all the out-neighbors of interior arcs within a container must have the same degree.
-In Example 5.6 and Table 1, we examined a minimum degree node $v_{0}$ with degree 6 in a regular graph.
-The interior degrees of nodes increase outwardly: $v_{0}$ has an interior degree of 0, nodes in $C_{1}$ have 1, nodes in $C_{2}$ have 2, and so forth.
+Assume that $u$ and $v$ are two nodes in container $C_{i}$. Suppose that they share a parent $v_{p}$ in container $C_{i-1}$. Suppose also that $|\text{int}(v_{p},v)|\geq|\text{int}(v_{p},u)|+1$.
+If $|int(v_{p},v)|>|int(v_{p},u)|+1$, then $v$ will have fewer exterior neighbors than $u$. Similarly, $u$ will have more external neighbors because of its fewer interior neighbors.
+These exterior neighbors of $u$ are raising the second neighbors of $v_{p}$.
+As a result, the exterior degree of $v_{p}$ will double. Because $v_{p}$ has already had its interior degree doubled, it will have its overall degree double
+Thus, the different interior degrees of $u$ and $v$ in container $C_{i}$ leads to a contradiction, so the claim holds. ∎
+Lemma 6.1 (Regular Interior Degrees) provides another insight into the structure within the containers of these counterexamples.
+It shows that all the out-neighbors of interior arcs within a container must have the same degree.
+This may not be surprising, as the examples we have been showing have all had this property.
+For example, both containers 1 and 2 in example 5.5 are orientations of regular graphs.
+The second container is actually an orientation of $K_{5}$, the complete graph on 5 nodes.
+This result begins to lend itself to some of the mathematical symmetry and structure that must be present in these containers and the SSNC.
 The definition of a back arc is an edge $(y,z)$, where $y\in C_{i}$, $z\in C_{j}$, and $j<i$ respectively.
 It is implied that the source node $y$ acquires new neighbors at a lower level, maybe doubling its degree, if such a back arc were included.
-The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which is in contrast to this.
+The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which contrasts with this.
+Next, we begin to look at how back arcs interact with the data structure and the development of the algorithm.
+Up to this point, we have gotten around back arcs by assuming that they do not exist.
+However, that will not last forever.
+These next few lemmas address the possibility of what happens when we try to construct a GLOVER and encounter a back arc.
+We will see that although we are not able to construct a GLOVER, we are still able to execute our algorithm and find a degree-doubling node.
+Consider the following table based on example 5.5.
+Table 1: Consequences of Back Arcs in 5.5
+Back Arc From(A) Back Arc To (B) Back Arc New Neighbors via $N^{+}(B)$
+Sage Faye Gus, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
+Tate Gus Hugo, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
+Ugo Hugo Isla, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
+Vera Bell Chad, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
+Wade Chad Dawn, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
+Xia Dawn Else, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
+Yoel Else Faye, Kai, Lei, Mia, Nova, Owen, Pat, Qin, Ryan
+This table looks at the possible back arcs from container 3 in example 5.5 and how they would impact the degree doubling of the node in question.
+The back arcs constructed in this table are attempts to go from container 3 $C_{3}$ to container $C_{1}$.
+In each case, we can see that $N^{+}(B)$ causes that degree of the source node to double.
+What follows is a lemma that generalizes this table.
 Lemma 6.2.
@@
 $|ext(u_{i},v_{k})|=|N^{+}(v_{k})\cap N^{++}(u_{i})|$
-Since $N^{+}(v_{k})$ includes all direct neighbors of $v_{k}$, it follows that:
-$|N^{+}(v_{k})|=deg(v_{k}).$
-By the structure of GLOVER and the absence of back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, the size of $ext(u_{i},v_{k})$ satisfies:
+We know there are no back arcs from $u_{i}$ to $v_{0},\dots,v_{k-1}$, so the size of $ext(u_{i},v_{k})$ satisfies:
 $|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
-Each node in $u_{i}$’s exterior is matched up with a node in $v_{k}$’s neighborhood.
 This means that the back arc will double $u_{i}$’s exterior degree.
@@
 This is true because its interior degree and its exterior degree will double.
+∎
 Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container-structured graph.
@@
 Lemma 6.3.
-Back Arcs Special Case
-Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
+Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
   1. 1.
@@
 $|\text{ext}(y,z)|<|\text{ext}(x,y)|$
-First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e.
-from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e.
-not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
+First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
 This gives the only possibility for $z$ as being an interior neighbor of $x$.
@@
 Lemma 6.4.
-First Back Arc Between Consecutive Containers in an Oriented Graph:
-Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$.
+First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$.
 Suppose that we represent this graph by a GLOVER.
@@
 Then since the node $v_{i-1}$ is in $C_{i-i}$, $v_{i-1}$ also maps to these exterior nodes.
-If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist{v_{0},u_{i}}-i$ so $u_{i}$ cannot belong to container $C_{i}$.
-Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$.
+If $u_{i}$ is not mapped to by any node in $C_{i-1}$, then $dist(v_{0},u_{i})-i$ so $u_{i}$ cannot belong to container $C_{i}$.
+Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$. ∎
 Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees).
-From this, we can see that a back arc to a parent container violates the orientated nature of the graph.
+From this, we can see that a back arc to a parent container violates the oriented nature of the graph.
 Lemma 6.5.
-No Back Arcs Non-Existence of Back Arc
-In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
+No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
 Proof.
-This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers).
+This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers). ∎
 […]
@@
 Each node inside a container represents a node whose degree is doubled by the inductive hypothesis.
-This process cannot continue forever so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
+This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
 […]
@@
 Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
-[…]
-[…]
-We stated that back arcs could be detrimental to our algorithm, so the removal of them in this section was important. Now we can better navigate the other properties of oriented graphs.
+We stated that back arcs could be detrimental to our data structure.
+Now that we have shown that even if they are present, they will not hinder the development of our pathfinding algorithm, as a back arc will simply lead to a degree-doubling node faster.
+Now we can better navigate the other properties of oriented graphs.
 7 Interior Degrees Doubling
 This section introduces the concept of interior degree-doubling.
-This is similar to solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container.
+This is like solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container.
 This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
 Lemma 7.1.
-Indistinguishable Regular Interior Nodes
-Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER.
+Indistinguishable Regular Interior Nodes Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER.
 We are attempting to prevent a degree-doubling node.
-If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree doubled.
+If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
 Proof.
-Let $C_{i}$ be a regular container of size $n$.
+Let $C_{i}$ be a regular container.
 Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$.
-The interior degree of a node $v_{i}$ in a container is defined as the number of arcs shared by $v_{i}$ between $v_{i}$ and a neighbor or parent node.
 In a regular container, all nodes are indistinguishable in terms of their degree.
-They also have the same parents because of the DSP.
+They also have the same parents because of containers.
 Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures.
-This means any node will have its interior degree doubled.
+This means any node will have its interior degree doubled under the automorphism $\phi$.
+∎
 The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others.
@@
 That would be the key.
-What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the node’s interior degrees double.
+What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the nodes within that container will have their interior degrees double.
 That means that we can select any of those nodes.
+We still need to prove that for each container there exists a node whose interior degree doubles.
+We have seen this in the example 5.5, but those could be too simple to observe any real patterns.
+In similar structure, what we saw in 4.1 (Minimum Out-Degree $<$ 3) was that it was easy to find degree-doubling nodes in graphs with a minimum degree of 0, 1, or 2.
+However, when that minimum degree was increased to 3, there were more cases that needed to be considered.
+[…]
 Lemma 7.2.
-Interior Degrees Double
-Let $G$ be an oriented graph represented by a GLOVER.
+Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER.
 We are attempting to prevent a degree-doubling node.
@@
 We prove this by strong induction on the size of the graph.
-Base Case: In Lemmas 4.1 (Minimum Out-Degree ¡ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree.
+Base Case: In Lemmas 4.1 (Minimum Out-Degree $<$ 3), 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0), and 4.3 (Minimum Out-Degree 3 with Neighbors 1), we establish that for graphs with minimum degrees of 0, 1, 2, and 3, there exists at least one node in each container with a doubled interior degree.
 These cases form our base, supporting the induction.
@@
 $H_{0}=\{u_{i}\}$
-, where $u_{i}$ is the minimum degree node in $C_{i}$.
+where $u_{i}$ is the minimum degree node in $C_{i}$.
 $H_{1}=N^{+}(u_{i})$
@@
 By Lemma 5.6 (Weak Decreasing Exteriors Lemma), we know that $|H_{i}|>|H_{i+1}|$ for all $i>0$ resulting in a strictly decreasing sequence of positive integers.
-An increase in $int(x,y)$ requires a proportional reduction in $ext(y,z)$.
 This construction yields a strictly decreasing sequence of subgraph sizes: $|H_{1}|>...>|H_{k}|$.
 The strictly decreasing nature of this sequence implies that it must terminate when $H_{i}\leq 2$.
-At this point, there are not enough arcs (1) in $H_{i}$ for both nodes to have an interior degree of 2.
-So, one node must have all its degree be exterior.
+At this point, there are not enough arcs in $H_{i}$ for both nodes to have an interior degree of 2.
+So one node must have all of its degree be exterior.
 Consequently, for some node $y^{*}\in H_{i}$, we find that $int(x,y^{*})=0$ for some $x\in H_{i-1}$, leading to $ext(x,y^{*})=d^{+}(y^{*})$.
-Thus, the parent of $y^{*}$, denoted $y^{*}-1$, in $G$ achieves a doubled degree within $C_{i}$.
-Therefore, for a graph of size $n$, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step.
+Thus, the parent of $y^{*}$, denoted $y^{*-1}$, in $G$ achieves a doubled degree in $C_{i}$.
+Therefore, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
 Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture.
 We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles.
-By finding a positive answer to this, and not just a positive answer, but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
+By finding a positive answer to this, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
 The way the proof was iterated was through both strong induction and a monotonically decreasing sequence of subsets.
 This cannot continue indefinitely.
-Once again, we see why the Decreasing Sequence Property (DSP) received its name.
-This double of interior degrees will help us to build our algorithm in the search for a counterexample. It does, however, give us a bit of light into the side of the Conjecture being true.
+This gives rise to why the Decreasing Sequence Property (DSP) received its name.
+Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples.
+The questions of whether this was always true or whether back arcs like lemma 6.3 could force this into being a false positive.
+This proof also helps with explaining the concept of the SSNC.
+Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER, as long as there are no back arcs.
+Once we have done this, there are containers spread out from a chosen minimum degree node.
+Now for each container we can simply run the conjecture on that container.
+The difference is that with lemma 7.2 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container.
+Then with lemma 7.1 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have their interior degree double.
+This notion of interior degrees will help us to build our algorithm in the search for a counterexample. It also gives us a bit of light into the possibility of the conjecture being true.
 8 Container Density
@@
 The Weak Decreasing Exteriors Lemma implies that containers will decrease in size.
-We need to see how that decrease in size will impact the degrees of the nodes.
+We need to see how that decrease will impact the degrees of the nodes, the other properties of containers, and the development of our algorithm.
 Lemma 8.1.
-Regular Exterior Degrees
-Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node.
-Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in $G$ has the Decreasing Sequence Property (DSP).
-Then the nodes in a container in $G$ will result in regular exterior degrees.
+Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node.
+Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in a container $C_{i}$ has the Decreasing Sequence Property (DSP).
+Then the nodes in $C_{i}$ will result in regular exterior degrees.
 Proof.
@@
 Assume, for contradiction, that the sizes of their exterior sets differ by at least one.
-Without loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|=|ext(x_{j},y)|+1$.
+Without the loss of generality, let there exist a node $y$ such that $|ext(x_{i},y)|\geq|ext(x_{j},y)|+1$.
 This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
-By Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs), we can imply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here.
-This says that with each distance from the minimum degree node, our exterior size is decreased by 1.
-This stopped at the container $C_{i}$.
+Because Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs) both hold, we can apply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here.
+This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
+This decreasing of the exteriors (and thus the containers) is stopped at the container $C_{i}$.
+This is where our assumption of the different exterior sizes comes into play.
 That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$.
@@
 By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
-So, this assumption led to a contradiction.
-This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the Weak Decreasing Exteriors Lemma in this proof because the two go hand in hand.
-Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double. The DSP is a bottleneck in this way because it is really limiting what can take place in this graph.
+So, this assumption led to a contradiction. ∎
+This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 5.6 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
+Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double.
+The DSP is a bottleneck in this way because it is really limiting what can happen in this graph in the instance of not having a node’s degree double.
 Lemma 8.2.
-Strong Decreasing Exteriors Lemma
-Let $G$ be an oriented graph represented by a GLOVER.
+Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER.
 Let $v_{0}$ be a minimum degree node in $G$.
-We are attempting to prevent a degree-doubling node, so for any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors must be bounded by
+We are attempting to prevent a degree-doubling node.
+For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors, and thus $C_{i}$ must be bounded by
 |$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
@@
 We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
-Base Case ($i=1$):
-Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$). The size of the exterior set
+Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$).
+The size of the exterior set
 $\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
@@
 Lemma 8.2 (Strong Decreasing Exteriors Lemma) uses Lemma 5.6 (Weak Decreasing Exteriors Lemma) to give us a way of calculating the maximum size of a container based only on its distance from the minimum degree node and the minimum degree node’s degree.
-The immediate takeaway from this lemma is the decreasing sequence of exteriors that we were attempting to prove.
-It is for this sequence that the term ”decreasing sequence property” was originally named.
+In Section 5 (Theoretical Foundation), we demonstrated that mathematical induction could be applied to our framework.
+This proof highlights the utility of the lemma.
+It allowed us to construct a proof based on the container distance from the chosen minimum-degree node.
+It is possible that this result could have been proven without relying on induction.
+We could have used a different approach or simply avoided induction based on container distance from the minimum degree node entirely.
+Still, the use of induction provides an important method, particularly for graphs where the usage of induction is more limited.
+Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
 Lemma 8.3.
-Interior Degrees Lower Bound
-Let $G$ be an oriented graph represented by a GLOVER.
+Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER.
 Let $v_{0}$ be a minimum degree node in $G$.
 Suppose every node in $G$ has the Decreasing Sequence Property (DSP).
-Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a node at distance $i-1$ from $v_{0}$.
-Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$, provided $(v_{0},x)\in G$.
+Let $x$ be a node at distance $i$ from $v_{0}$, and $v_{i-1}$ be a parent of $x$.
+Then the interior degree of $x$ with respect to $v_{i-1}$, denoted as $|int(v_{i-1},x)|$, is at least $i$.
 Proof.
@@
 This places a limit on the number of nodes $x$ can map to at distance $i+1$ from $v_{0}$.
-Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not asked, $x$ cannot map to nodes at earlier distances.
-Therefore, the only remaining possibility is that $x$ maps to interior nodes.
-Since $x$’s degree must account for at least $i$ mappings.
-These must be to interior nodes.
-So, we conclude that $x$ must map to at least $i$ interior nodes.
-Thus, $|int(v_{i-1},x)|\geq i$, as required.
-Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is an opposite bound for the interiors.
-Once again, the minimum degree node is the driving force behind this.
+Now, since $x$ has at least the same out-degree as $v_{0}$, $x$ must map to at least $i$ other nodes. Given that back arcs are not allowed, $x$ cannot map to nodes at earlier distances.
+Therefore, the only remaining possibility is that $x$ maps to interior nodes. These must be to interior nodes. So, we conclude that $x$ must map to at least $i$ interior nodes.
+Thus, $|int(v_{i-1},x)|\geq i$, as required. ∎
+Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is a contrasting bound for the interiors.
+Once again, the minimum degree node is the driving force, the catalyst, behind this.
 The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
 Lemma 8.4.
-Regular Graph
-All Nodes Have Equal Out-Degree
-Let $G$ be an oriented graph represented by a GLOVER.
+Regular Graph All Nodes Have Equal Out-Degree Let $G$ be an oriented graph represented by a GLOVER.
 Suppose $v_{0}$ is a minimum degree node of $G$.
@@
 Proof.
-With the proofs on the lack of back arcs and the regularity of interior arcs and the regularity of exterior arcs, we arrive at the following equation for each container:
-$\delta=|int(v_{p},u_{i})|+|ext(v_{p},u_{i})|$
-where $u_{i}$ is in container $C_{i}$, and $v_{p}$ is a parent of $u_{i}$ in container $C_{i-1}$.
 For a node $u_{i}$ in container $C_{i}$, with degree $\delta+1$, there would be the possible places for this degree to be located—back arcs, interior arcs, or exterior arcs.
@@
 While this is not a problem, it will cause the interior degrees to increase at a faster rate than we just showed in Lemma 8.3 (Interior Degree Lower Bound), as this would increase the number of interior arcs.
-This more rapid reduction in the interior degree leads to container density faster. To avoid this, regularity is key.
-Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph.
+This more rapid reduction in the interior degree leads to container density faster, which would simply lead to a degree doubling node faster. To avoid this, overall regularity is key.
+Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph. ∎
 We are incrementally refining the properties that must hold for our counterexamples.
@@
 This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
+Example 8.1.
+0
+$\Rightarrow$ 111122223333444455556666777788889999101010101111111112121212131313131414141415151515161616161717171718181818191919192020202021212121222222222323232324242424H
+[…]
+Example 8.1 shows some of the patterns that emerge with this regularity in the SSNC.
+Starting from the outside in, going inward represents the order of containers as they increase in distance from $v_{0}$.
+This example has visualized only the interior arcs, as the exterior arcs would complicate this visualization.
+However, just as in previous examples, all nodes in a previous container are mapping to all nodes in the next container.
+Some of those patterns that emerge are first the equal out-degree of nodes.
+For even-sized containers, we are able to see symmetry about balance, uniformity, and a sense of order within each container.
+Also, while it is not a point of the conjecture, we do see the presence of cycles in each container.
 9 Transitive Triangles and Seymour Diamonds
-Before we continue with the development of our DSP algorithm and show how it solves the Seymour Second Neighborhood Conjecture, we need to address some potential problems and show that they will not hinder our algorithm.
+Before we continue with the development of our algorithm and show how it solves the SSNC, we need to address some potential problems and show that they will not hinder our algorithm.
 Definition 9.1.
@@
 The concept of a transitive triangle is an important one because it involves the risk of double counting.
-The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$, so if we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
+The node $u$ is both a first neighbor of $x$ and a second neighbor of $x$.
+If we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
 Definition 9.2.
-Let $x\in G$ and $u_{1},u_{2}\in N^{+}(x)$. Then $x$, $u_{1}$, $u_{2}$, and $y$ form a Seymour diamond if $(u_{1},y)$ and $(u_{2},y)\in G$.
+Let $x\in G$ and $y,w\in N^{+}(x)$. Then $x$, $y$, $w$, and $u$ form a Seymour diamond if $(y,u)$ and $(w,u)\in G$.
 The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$.
@@
 What this situation presents is redundant second neighbors.
-There is a path $x\rightarrow u_{1}\rightarrow u_{2}$ in this diamond, and there is a path $x\rightarrow u_{2}\rightarrow u_{1}$ in this diamond.
+There is a path $x\rightarrow y\rightarrow u$ in this diamond, and there is a path $x\rightarrow w\rightarrow u$ in this diamond.
 As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
@@
 xyu
+Figure 12: Illustration of a transitive triangle with parent node $x$, child nodes $y$ and $u$, where $u$ is also a child of $y$
 Lemma 9.1.
-Transitive Triangles from Containers
-Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP).
+Transitive Triangles from Containers Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP).
 For any two consecutive containers $C_{i}$ and $C_{i+1}$, there exist two parent nodes $u\in C_{i}$ and $v\in C_{i}$, and a shared child node $w\in C_{i+1}$, such that the nodes $u$, $v$, and $w$ form a transitive triangle.
-Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers. As such, they do not hinder the algorithm.
+Lemma 9.1 (Transitive Triangles from Containers) informs us that transitive triangles are intricate parts of neighboring containers.
+As such, they do not hinder the algorithm.
+Transitive triangles represent a common situation with our containers.
+A parent node $x$ has two children $y$ and $u$ that are also connected by an arc.
+The purpose of the interior arcs is to ensure that the parent node $x$’s degree does not double.
+We have already shown through 8.3 (Interior Degrees Lower Bound) that every node in Container $i$ has an interior degree at least $i$.
+The two arcs from a parent node and the interior nodes are what create these transitive triangles.
+What is not acknowledged in lemma 9.1 (Transitive Triangles from Containers) is that every node in the container takes both non-parent roles in different transitive triangles.
+Every node is both a child of a parent node in the previous container and has a degree of at least one.
+This is what creates the transitive triangles in every container.
 Lemma 9.2.
@@
 $u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
-The following arcs exist:
-1. There are forward arcs $x\to y$ and $y\to u$,
-2. There are forward arcs $x\to w$ and $w\to u$,
-3. No back arcs exist between containers.
+The following arcs exist: 1. There are arcs $x\to y$ and $y\to u$, 2. There are arcs $x\to w$ and $w\to u$, 3. No back arcs exist between containers.
 Similar to Lemma 9.1, Lemma 9.2 (Seymour Diamonds from Containers) lets us know that we have been interacting with Seymour Diamonds all along without knowing it.
 We didn’t know it because they did not disrupt our algorithm.
+This is exactly what we are seeing happen every time there is a Seymour diamond present between three containers.
+There is a parent node, a choice of at least 2 interior nodes, and an exterior node.
+Not only is this not a problem; it’s one of the benefits of the GLOVER data structure.
+Those nodes $y$ and $w$ represent equidistant nodes from $v_{0}$ whose job is to ensure that $x$’s degree does not double.
+The arcs from $y$ to $u$ and $w$ to $u$ are exterior arcs, connecting these nodes to the next container.
+The purpose of these arcs is first to meet the degree requirements of each node and second to ensure connectivity.
+Once each node $y$ and $w$ has connected to the third container, it can take a role as a parent and form its own Seymour diamond.
 xywu
+Figure 13: Illustration of a Seymour diamond with parent node $x$, child nodes $y$ and $w$, and grandchild node $u$
 Lemma 9.3.
-Seymour Diamonds and Decreasing Path Algorithm
-Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP).
+Seymour Diamonds and Decreasing Path Algorithm Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP).
 We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm.
@@
 None of these cases are a deterrent to the execution of the DSP algorithm.
+Lemma 9.3 (Seymour Diamonds and Decreasing Path Algorithm) shows how we have been proceeding through the GLOVER data structure through our Seymour diamonds.
+These diamonds are constructed to be presented in our containers.
+As a result, we simply select an interior node within the container and continue.
+This is what we saw in our examples in section 4, where Seymour diamonds were also present, for example 4.3.
+This is also what we saw in our social media example 5.5 in section 5.
+This is also what we saw in our regular graph example in 8.1 in section 8.
 Lemma 9.4.
@@
 Suppose $G$ is an oriented graph represented by a GLOVER.
-Let $v_{0}$ be a minimum degree node with degree $d$.
-Consider a transitive triangle formed by nodes $z$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(z)=0$.
+Let $v_{0}$ be a minimum degree node with degree $\delta$.
+Consider a transitive triangle formed by nodes $u$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(u)=0$.
 Then the following cases are possible:
-  • $d^{+}(y)=d$ and $d^{+}g(z)=d$
-  • $d^{+}(y)=d$ and $d^{+}(z)>d$
-  • $d^{+}(y)>d$
-  • $d^{+}(y)<d$
+  • $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$
+  • $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$
+  • $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$
+  • $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$:
 Proof.
-  • Case 1: $d^{+}(y)=d$ and $d^{+}(z)=d$:
-    Both $y$ and $z$ have exactly $d$ outgoing connections.
-For the triangle to be transitive, they must not share all neighbors.
-This means that there exists at least one neighbor of $y$ that is not a neighbor of $z$.
-Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
-  • Case 2: $d^{+}(y)=d$ and $d^{+}(z)>d$:
-    Then $z$ may cover some neighbors of $N^{+}(y)$. Since $y$ must map to at least one neighbor unique from those of $z$, it follows that $y$ has a neighbor that $z$ does not connect to.
-  • Case 3: If $d^{+}(y)>d$:
-    $y$ has more outgoing connections than required.
-It can utilize these additional connections to map to unique neighbors without impacting the minimum degree condition, ensuring $y$ retains neighbors not covered by $z$.
-  • Case 4: If $d^{+}(y)<d$:
-    This scenario contradicts the minimum degree requirement since each node must have at least $d$ outgoing connections. Therefore, $y$ cannot be part of the triangle involving $v_{0}$.
-Now that we see that transitive triangles and Seymour diamonds are not a problem for our algorithm, we can continue our investigation of oriented graphs.
+  • Case 1: $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$: Both nodes $y$ and $u$ out-degree $\delta$.
+The triangle has the arc $(y,u)$.
+Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$.
+This means that $u$ has $\delta$ out-neighbors that are not in the triangle.
+Since we assumed that $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$, we know that $y$ cannot connect to all not $u$’s neighbors since $y$ has a neighbor in the triangle ($u$).
+This node that is $u$’s neighbor, but not $y$’s neighbor, is a candidate to extend the path.
+  • Case 2: $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree $\delta$.
+The node $u$ has out-degree greater than $\delta$.
+Again, the triangle has the arc $(y,u)$.
+Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$.
+This means that $u$ has more than $\delta$ out-neighbors that are not in the triangle.
+We assumed that $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$.
+We know that $y$ cannot connect to all not $u$’s neighbors since $d^{+}(u)>d^{+}(y)$.
+We can then select one of these nodes that is a neighbor of $u$, but not a neighbor of $y$ to be a candidate to extend the path.
+  • Case 3: If $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$: The node $y$ has out-degree greater than $\delta$.
+The node $u$ has out-degree $\delta$.
+The triangle has the arc $(y,u)$.
+Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$.
+The node $y$ has more than $\delta$ out-neighbors that are not in the triangle.
+The node $u$ has exactly $\delta$ out-neighbors that are not in the triangle.
+The maximum number of neighbors that $y$ and $u$ can share is $\delta$, that being $u$’s out-degree.
+Then we can select one of the nodes that is a neighbor of $y$, but not a neighbor of $u$ to be a candidate to extend the path.
+  • Case 4: If $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree greater than $\delta$.
+The node $u$ has out-degree greater than $\delta$.
+We have a few cases because we have to remember that in this data structure the purpose of our containers and the interior nodes was to prevent parent nodes from having their degrees double.
+The case where both nodes in a transitive triangle have more than $\delta$ out-neighbors means that $x$ has up to $2*\delta$ out-neighbors, which would cause its degree to double.
+Thus, to prevent $x$’s degree from doubling, there must be some overlap between the out-neighbors of $y$ and the out-neighbors of $u$.
+These two sets cannot be equal to one another, though.
+The most overlap we can have is $N^{+}(u)\subsetneq N^{+}(y)$ or $N^{+}(y)\subsetneq N^{+}(u)$.
+Both of these situations leave room for a node in the larger set that is not in the smaller set as a candidate to extend the path.
+As we allow for less overlap, the number of candidate nodes to extend the path grows.
+We can see that there will always be a node in the larger set to extend the path.
+Now that we have a better understanding of transitive triangles and Seymour diamonds and how they interact with our data structure.
+We can see that they are not a problem for the development of our algorithm.
+We will continue our investigation of oriented graphs.
 Next, we will blend everything together into our main results.
@@
 10.1 Introduction
-Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some final results.
-So far, this battle has been a battle between a catalyst, the chosen minimum degree node and the bottleneck, the Decreasing Sequence Property (DSP).
+Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some results.
+So far, this battle has been a battle between a catalyst, the chosen minimum degree node, and the bottleneck, the Decreasing Sequence Property (DSP).
 We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
@@
 Proof.
-We have set up the containers by a GLOVER data structure.
-There are no back arcs, else we have a degree-doubling node.
+We have set up the containers using a GLOVER data structure.
+There are no back arcs; else we will have a degree-doubling node.
 Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
@@
 10:     else
-11:         Since there is no back arc, the container $C_{i+1}$ will be bound in size to be 1 less than the size of container $C_{i}$.
+11:         Since there is no back arc, the container $C_{i+1}$ will be limited in size to be 1 less than the size of container $C_{i}$.
 12:         Repeat from Step 4.
@@
 14:end for
-Start with graph size $n$Select minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots$Identify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to prior containerLocate child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
+Start with graph size $n$Select a minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots C_{k}$dentify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to a prior containerLocate the child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
 Figure 14: Flowchart depicting the degree-doubling algorithm applied within containers, demonstrating the process of identifying nodes with the Decreasing Sequence Property.
 Corollary 10.1.
-Degree Doubling by Container Density
-In an oriented graph $G$ with minimum degree node $v_{0}$ and containers $C_{1},\allowbreak C_{2},\allowbreak\ldots\allowbreak C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
+Degree Doubling by Container Density In an oriented graph $G$ with a minimum degree node $v_{0}$ and containers $C_{1},\allowbreak C_{2},\allowbreak\ldots\allowbreak C_{n}$, as the containers shrink in size as their distance from the minimum degree node increases, and as the required interior degrees rise, a collision is inevitable.
 There are two possibilities:
   • Case 1: The containers continue to shrink in size.
-Simultaneously, the interior degrees continue to grow in size.
+Simultaneously, the interior degrees continue to grow.
 We reach a point where no feasible oriented graph is possible to accommodate the required set of nodes with interior degrees.
   • Case 2: For small minimum degree nodes, the container’s shrinkage is not fast enough to cause this collision.
-Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and increased number of nodes.
+Corollary 10.1 (Degree Doubling by Container Density) captures the essence of shrinking container sizes and an increased number of nodes.
 This was represented by the yin and yang effects early on.
@@
 Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbors represented the size of the containers.
-In small graphs like 4 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur.
-To understand the reasoning behind this, understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1.
+In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur.
+To understand the reasoning behind this, one must understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1.
 The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$.
@@
 Corollary 10.2.
-Occurrence of Last Dense Container
-Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$.
+Occurrence of Last Dense Container Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$.
 Then there will be a container that reaches maximal density.
@@
 While these are called interior degrees, these are the out-degrees of the nodes in a container.
-This lemma says that every container must have regular out-degrees.
+This lemma indicates that every container must have regular out-degrees.
 The second part of this proof is lemma 8.2 (Strong Exterior Decreasing), which says that in container $i$, the size of the exterior is at most $d^{+}(v_{0})-i$.
@@
 This is progressively getting to a denser container.
-As Corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever.
+As described in corollary 10.1 (Degree Doubling by Container Density), this cannot continue forever.
 There must be a last, most dense container.
-The Seymour Second Neighborhood Conjecture is made more beautiful by the appearance of these orientations of out-regular graphs as containers get smaller.
+∎
+The SSNC is made more aesthetic by the appearance of these orientations of out-regular graphs as containers get smaller.
 Not only has the conjecture been proven true, but with an inherent symmetry and order.
-There is a visual pattern emerging inside these containers that is not just beautiful to graph theorists, but there is a structure that is easily appealing to non-mathematicians.
+There is a visual pattern emerging inside these containers that is not just important to graph theorists but also possesses a structure that is easily appealing to non-mathematicians.
 Corollary 10.3.
-Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $k$, and let the most dense graph in container $C_{i}$.
+Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $\delta$, and let the most dense graph be in container $C_{i}$.
 Then every node in container $C_{i}$ has its degree doubled in the representation.
-The discovery of multiple degree-doubling nodes lifts the Seymour Second Neighborhood Conjecture from a mere question of existence to one in abundance.
-There were many who doubted that one even existed.
-Now there are many.
-Second is the way we find it. It’s the presence of an oriented regular graph as these containers decrease in size. There’s an element of mathematical beauty there.
+The discovery of multiple degree-doubling nodes elevates the SSNC from a mere question of existence to one in abundance. There were many who doubted that one even existed. Now there are many.
+Second is the way we find it.
+It’s the presence of a last dense container with an oriented regular graph as these containers decrease in size.
+We see this occur as the interior degrees are forced to increase as the distance from $v_{0}$ is increased.
+These interior degrees are set up to prevent a parent node’s degree from doubling.
+Ultimately though, they are unable to do that, as we see the collision of the container shrinkage and interior degree requirements.
 Then third we can look at the applications.
 We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group.
+Dense regular graphs have many applications in themselves, from areas like expander graphs to network topology design to coding theory.
 Theorem 10.2.
@@
 This makes the search for the minimum degree node $O(|V|)$.
-Once we have this value, we will place every node and every edge of $G$ into a preprocessing array ready to be partitioned.
+Once we have this value, we will place every node and every edge of $G$ into a pre-processing array ready to be partitioned.
 Each edge is examined, contributing $O(|E|)$ complexity, while nodes contribute $O(|V|)$.
@@
 Third, lemma 7.1 (Indistinguishable Regular Interior Nodes) lets us know that nodes within a container are indistinguishable from one another.
-The combination of these three lemmas let us know that every node inside a container has their interior degree double.
+The combination of these three lemmas lets us know that every node inside a container has their interior degree doubled.
 This gives the complexity of finding a node whose interior degree has doubled $O(1)$.
@@
 This gives the run time a complexity of $O(|V|+|E|)$.
+∎
 This complexity analysis shows that the Decreasing Sequence Property Algorithm can effectively find a degree-doubling node in polynomial time.
@@
 It was not spoken about in the paper because the focus was entirely set on proving the SSNC, but this is another example of multiple degree-doubling nodes, just a localized instance.
+We opened by showing several applications of the SSNC, including social media, social network modeling and network analysis.
+This highlights the necessity for an algorithm that can work efficiently in practice.
+This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph is important.
+This algorithm’s linear time complexity shows that even enormous graphs may be processed effectively.
+That makes it potentially applicable to real-world networks with millions or billions of nodes and edges.
+This suggests that the algorithm is not only theoretically sound but also practically feasible for implementation and use in real-world scenarios, such as analyzing social networks, identifying influential nodes in communication networks, or understanding the spread of information.
 11 Conclusion and Future Work
-In this work, we explored the Seymour Second Neighborhood Conjecture in oriented graphs.
+In this work, we explored the Seymour Second Neighborhood Conjecture (SSNC) in oriented graphs.
 We did so through the lenses of graph theory, combinatorics, algorithms, and data structures.
@@
 One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph.
-This effectively solves the Seymour Second Neighborhood Conjecture with an answer in the affirmative.
+This effectively solves the SSNC with an answer in the affirmative.
 The interesting thing about this algorithm was that it did not require any assumptions about the oriented graph or where the degree-doubling node was in the graph.
-Instead, the algorithm selects a minimum degree node and is able to partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
-This approach introduced a new perspective on the Seymour Second Neighborhood Conjecture.
+Instead, the algorithm selects a minimum degree node and can partition the graph into containers and form a decreasing sequence of exteriors that is eventually bound to lead us to the solution.
+This approach introduced a new perspective on the SSNC.
 It was no longer framed as merely a graph theory problem.
-By utilizing the divide-and-conquer paradigm and solving different parts independently, this algorithm shows that the Seymour Second Neighborhood Conjecture could effectively be treated with computer science techniques.
+By utilizing the divide-and-conquer paradigm and solving different parts independently, this algorithm shows that the SSNC could effectively be treated with computer science techniques.
 This suggests that there may be other open problems in mathematics that are not being given enough attention by computer scientists.
 Alternatively, it may be that there is not enough interdisciplinary communication about these problems and the potential techniques to solve them.
-The Seymour Second Neighborhood Conjecture started as a question in 1990.
+The SSNC started as a question in 1990.
 Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers.
@@
 With a solution now in hand, it’s time to begin looking forward.
-The Seymour Second Neighborhood Conjecture can be represented as a social media problem instead of an oriented graph.
+The SSNC can be represented as a social media problem instead of an oriented graph.
 Here, the nodes are people, where no one follows back.
@@
 There are too many to name in this short paper.
+We have presented a linear time algorithm that can work on large graphs in these problems.
+By constructing the algorithm for the SSNC and proving that the algorithm will find a degree-doubling node, we have proved that the algorithm is theoretically sound.
 Secondly, we introduced a data structure, Graph Level Order-GLOVER.
@@
 Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
+Finally, this research offered an alternative method for the SSNC.
+We focused on identifying a node whose out-degree doubles in the graph’s square rather than just showing that it exists.
+To do this, we developed an effective algorithm and proved its complexity.
+Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions.
+This may encourages researchers to explore constructive approaches with a focus on practical applications.
+Hopefully, this work will inspire further research on constructive approaches to other graph theory problems.
+Also, we encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
 12 Acknowledgments
@@
 A back arc is defined as an arc $(y,z)$ such that $z\in N^{+}(y)$ and $x\in C_{j}$, where $j<i$.
-Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$. Then $x$, $y$, $u$ form a transitive triangle.
+Definition: Let $x\in G$ and $y,u\in N^{+}(x)$ and $(y,u)\in G$.
+Then $x$, $y$, $u$ form a transitive triangle.
 Definition 13.7.
@@
 14 Appendix 2: Applicability of the Graph Level Order
-The deciding factor that held the algorithm together and helped solve the Seymour Second Neighborhood Conjecture was the Graph Level Order (GLOVER) data structure.
+The deciding factor that held the algorithm together and helped solve the SSNC was the Graph Level Order (GLOVER) data structure.
 This research did not begin as an investigation into data structures.
-Instead, a series of operations were conducted on the oriented graphs - partitioning, ordering those partitions, adding interior and exterior arcs, until what was left was an ordering of both the partition and nodes within the partition.
+Instead, a series of operations were conducted on the oriented graphs—partitioning, ordering those partitions, and adding interior and exterior arcs—until what was left was an ordering of both the partition and nodes within the partition.
 Further, these interior and exterior arcs allow us to define relationships based on two different metrics, an outer metric and an inner metric.
-Unlike the Seymour Second Neighborhood Conjecture, exterior arcs do not always have to only be defined to the next container.
-That was a problem specific definition.
+Unlike the SSNC, exterior arcs do not always have to only be defined to the next container.
+That was a problem-specific definition.
 Moving forward, the GLOVER data structure should be well suited to tackle problems in graph theory and beyond.
@@
 The concept of divide-and-conquer, which has dominated much of computer science, shows up in this paper as well.
-We are able to treat these containers independent of one another and solve their problems locally before bringing them back to the global problem.
+We can treat these containers independently of one another and solve their problems locally before bringing them back to the global problem.
 This shows how versatile the GLOVER data structure is moving forward.
@@
 This is hard to measure on a single scale.
-Next these nodes are not just treated like clusters inside these containers, but like nodes in graph theory.
-The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth first search, or centrality to be called.
+Next, these nodes are not just treated like clusters inside these containers but like nodes in graph theory.
+The interior and exterior arcs allow for common graph theory algorithms and techniques like Prim’s algorithm, breadth-first search, or centrality to be called.
 What is even more interesting is that because we have identified the interior and exterior arcs already, we can choose to run these algorithms only on interior arcs or only on exterior arcs.
@@
 When we take a course on graph theory, oriented graphs are probably not on the course load.
-Second neighbors or squares of graphs are probably not, either.
+Second neighbors or squares of graphs are probably not either.
 To fully grasp this problem, though, we need to be able to delve into these definitions through visualizations.
-For example, research on the Seymour Second Neighborhood Conjecture led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs.
+For example, research on the SSNC led this paper into the terms decreasing sequence property, induced subgraphs, containers, interior neighbors, exterior neighbors, and back arcs.
 The paper addressed how these terms were intertwined with each other.
 An example is where the paper says that there is a monotonically decreasing sequence.
-James Robert Brown’s book Philosophy of Mathematics speaks of the importance of visualizations in mathematics.
+James Robert Brown’s book ”Philosophy of Mathematics” speaks of the importance of visualizations in mathematics.
 Proofs are still the foundation, but sometimes visualizations can serve as inspiration or a spark for those proofs.
@@
 Where possible, things like HTML canvas and JavaScript D3.js are used to visualize these concepts and help improve understanding.
-Readers are encouraged to explore the website for additional information and examples at: https://glovermethod.com This is an evolving project.
+Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com.
+This is an evolving project.
 Additional examples and applications will continue to be added.
@@
 Please provide feedback on any aspects of the site or content for improvement.
-Our goal here was to help users connect with the Seymour Second Neighborhood Conjecture.
-Hopefully, this tool not only serves as support for this paper, but as an educational resource in the fields of graph theory and computer science.
-Thank you for your engagement
+Our goal here was to help users connect with the SSNC.
+Hopefully, this tool not only serves as support for this paper but also as an educational resource in the fields of graph theory and computer science.
+Thank you for your engagement.

### v8

@@
 Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six.
-Brantner et al.
-[1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist.
+Brantner et al. [1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist.
 Chen et al.
@@
 In recent work, Diaz [5] showed that almost all orientations of random graphs satisfy the conjecture.
-This was an extension of the work done by Botler et al.
-[2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
+This was an extension of the work done by Botler et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
 The concept of $m$-free graphs is particularly relevant to this paper, as it closely aligns with the work of Daamouch [7].
@@
 Section 3 outlines our approach to the problem, while Section 4 provides illustrative examples.
-Section 5 focuses on conditions under which a node’s out-degree may fail to double, and Section 6 examines the role of back arcs in oriented graphs.
+Section 5 introduces a data structure that will be instrumental for the proof of the conjecture.
+Section 6 begins to use that data structure to establish some foundational lemmas and properties.
+Section 7 examines the role of back arcs in oriented graphs.
 Section 8 explores the conditions that lead to doubling of interior degrees, examining their significance in relation to the SSNC.
@@
 Conjecture 2.1.
-(Square version of Seymour’s Second Neighborhood Conjecture). Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
+(Square version of Seymour’s Second Neighborhood Conjecture).
+Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
 Now that we have two different conjectures, we can investigate how they relate to one another.
@@
 Similarly, conjecture 2.1 relates this same concept to the square graph.
-The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling, and its second neighborhood being greater than or equal to the first.
+The defining characteristics of these conjectures are the relationship between a node’s out-degree at least doubling and its second neighborhood being greater than or equal to the first.
 By the definition of an arc in the square graph, we can see that an arc is defined in this square graph only for paths of length two in the original graph.
@@
 Then identify a node whose second neighborhood is at least as large as the first, as required by the conjecture.
-A similar definition, a Seymour vertex, fulfills conjecture 2.1 - it’s out-degree in the squared graph is at least double its original out-degree.
+A similar definition, a Seymour vertex, fulfills conjecture 2.1—its out-degree in the squared graph is at least double its original out-degree.
 We could equivalently search for a Seymour vertex or a degree-doubling node.
 3 Approach and Motivation
-We begin our analysis by considering general oriented graphs.
+We begin our analysis by considering generally oriented graphs.
 We make no initial assumptions about node degrees or graph properties.
@@
 When the graph is partitioned, even larger graphs, often with higher value minimum degree nodes and thus more partitions, are able to be broken into subproblems.
-These subproblems are then able to be solved independent of one another.
+These subproblems are then able to be solved independently of one another.
 These subproblems are each their own instances of the SSNC, just a localized version.
@@
 Example 4.1.
-$v_{0}$$v_{2}$$v_{3}$$v_{1}$$v_{6}$$v_{5}$$v_{4}$
-$\Rightarrow H$
+[…]
 Figure 2: Example illustrating how a minimum out-degree node in $G$ has its degree doubled in $G^{2}$ under certain conditions.
 Lemma 4.1.
-Minimum Out-Degree $<$ 3 Suppose that a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
+Minimum Out-Degree $<$ 3
+Suppose that a minimum out-degree node $v_{0}$ in an oriented graph has an out-degree less than 3. Then $v_{0}$’s out-degree will at least double.
 Proof.
@@
 The node $v_{0}$ cannot be one of these nodes.
-Hence $v_{0}$ will have at least two second neighbors - $v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first.
+Hence $v_{0}$ will have at least two second neighbors—$v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first.
 Thus, its out-degree will double in $G^{2}$.
-Therefore, in all cases, $v_{0}$’s out-degree at least doubles.
+Therefore, in all cases, $v_{0}$’s out-degree at least doubles. ∎
 Lemma 4.1 Minimum Out-Degree $<$ 3 begins to tell us about the structure of oriented graphs.
@@
 0
-$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$ H
+$\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$
 Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$.
@@
 These lemmas offer a theoretical framework and insight into our rationale. These findings will be expanded upon and referenced in the upcoming sections.
-5 Theoretical Foundation
-We will introduce some new terms to help with the rest of this paper.
-The partitions we discussed earlier will now be formally referred to as containers.
-Additionally, we shall discuss the decreasing sequence property and its significance.
-After that, we will discuss interior, exterior, and back arcs.
-We will show the importance of these arcs with some foundational lemmas.
-The GLOVER data structure will then be used to wrap up.
+5 Building a Data Structure
+In this section, we develop a new data structure for oriented graphs (Graph Level Order).
+This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond.
+In order to define this data structure, we will need to formally refer to the partitions we discussed earlier as containers.
+Additionally, we shall discuss interior, exterior, and back arcs and show the influence they have on our data structure.
+Finally, we will speak about how this data structure can use the divide and conquer paradigm to help solve the SSNC.
 Example 5.1.
 […]
-Figure 4: Illustration of containers, each enclosed in a checkered box.
-Containers group nodes based on their distance from the minimum degree node, with no assumptions on container relationships or sizes.
+Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from $v_{0}$, with no assumptions on container relationships or sizes.
 Definition 5.1.
@@
 $C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where
-$V(G)$ denotes the graph $G$’s vertex set and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
+$V(G)$ denotes the graph $G$’s vertex set, and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
 Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [7].
 The neighboring containers (see example 5.1), which will be crucial in the proofs, are the main distinction here.
-We will investigate the structure of oriented graphs from two angles by using the idea of containers.
-Alternatively, we will investigate it from the viewpoint of containers or the minimum degree node.
-Containers are seen based on their distance from the minimum degree node.
-Nodes are seen based on their position relative to containers.
+We will investigate the structure of oriented graphs from two angles—from the viewpoint of containers or $v_{0}$.
+Containers are analyzed based on their distance from $v_{0}$.
+Nodes are seen based on their position within these containers.
 They are either before a container, in a container, or after a container.
-Specifically, the case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve.
-A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in $G^{2}$.
-A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph.
-Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the nodes adjacent to $v_{0}$.
-If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double.
-This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
-However, we will need to define new concepts to proceed with this line of reasoning.
-To prevent a node’s out-degree from doubling, we must first define a term.
-The decreasing sequence property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$.
-There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen.
-We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
-Example 5.2.
-0
-$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$ H
-Figure 5: This is an illustration where node $v_{0}$ has the Decreasing Sequence Property (DSP). There are more first neighbors $d$ of $v_{0}$ than second neighbors.
 Definition 5.2.
-If the size of a node $x\in V(G)$ in an oriented graph $G$ is strictly bigger than the size of its second out-neighborhood, that is, $|N^{+}(x)|>|N^{++}(x)|$, then $x$ is said to have the decreasing sequence property.
-Two competing new ideas have been presented, the decreasing sequence property (DSP) and the minimal degree node.
-A crucial starting point for our approach and degree values is a minimal degree node.
-It acts as the anchor by being chosen as the beginning point.
-The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node.
-It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
-In contrast, the DSP introduces constraints on connectivity.
-This concept is akin to a bottleneck.
-A bottleneck in graph theory represents a point that limits flow or connectivity within the network, a role precisely filled by the DSP.
-Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
-Lemma 5.1.
-Total Order Induced by Containers Relative to the Minimum Degree Node
-Let $G$ be a graph with a minimum-degree node $v_{0}$.
-Construct a series of containers $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{k}$ ordered by distance relative to $v_{0}$.
-Each container $C_{i}$ is a set of nodes at distance $i$ from $v_{0}$.
-This imposes a total order on the nodes in $G$.
-This means that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
-Proof.
-To prove a total order, we need to show that any two nodes in $G$ can be compared.
-Let $u,v$ be two nodes in $G$ such that $u\neq v$.
-We assume that both $u$ and $v$ are reachable from $v_{0}$.
-Distance is a total ordering, so one of three things is true: either $dist(v_{0},u)<dist(v_{0},v)$, $dist(v_{0},u)=dist(v_{0},v)$, or $dist(v_{0},u)>dist(v_{0},v)$.
-The situation where $dist(v_{0},u)<dist(v_{0},v)$ implies that $u$ is in a closer container than $v$.
-The situation where $dist(v_{0},u)=dist(v_{0},v)$ implies that the nodes $u\text{ and }v$ are in the same container.
-And the situation $dist(v_{0},u)>dist(v_{0},v)$ implies that $u$ is in a further container.
-This shows that it is a total order.
-∎
-While imposing order on our graph is a solid first step, the containers do more than just enforce overall order.
-In essence, they allow for the grouping of comparisons between nodes that are equally spaced from a reference node, $v_{0}$.
-A typical total order, like lexicographical order, does not take this into consideration.
-This ”containerized ordering” allows us to differentiate nodes based on particular adjacency qualities inside each distance group in addition to their distance from $v_{0}$.
-Definition 5.3.
-A Level Total Order (LTO) on a set $S$ can be defined as follows:
-  • Total Order: For any two elements $u$ and $v$ within the set $S$, exactly one of the following must hold:
-    • $u$ is less than $v$ (denoted as $u<v$),
-    • $u$ is greater than $v$ (denoted as $u>v$), or
-    • $u$ is equal to $v$ (denoted as $u=v$).
-  • Leveled Structure:
-    • The elements of $S$ are partitioned into a leveling of containers. $C_{1},\allowbreak C_{2},\allowbreak\ldots,\allowbreak C_{n}$
-    • Each container itself forms a subset of $S$.
-    • The containers are sequentially ordered relative to each other.
-  • Comparability within Containers: For any two elements $u,v\in C_{i}$, their order can be determined based on a specific criterion.
-  • Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
-This order means that we have order in two ways.
-First, the containers themselves are ordered.
-Second, the nodes within the containers are ordered.
-This ordering must also be consistent within the containers.
-Lemma 5.2.
-Decreasing Sequence Property Lemma Suppose a node in an oriented graph $G$ does not have the decreasing sequence property. Then this node will have its out-degree at least double in $G^{2}$.
-If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$.
-In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed.
-We will now give terms to these concepts.
-Definition 5.4.
-Let $x$ be a node with the decreasing sequence property and let $y\in N^{+}(x)$.
-The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
+Let $x$ be a node in $G$ and let $y\in N^{+}(x)$. The interior neighbors of $y$ with respect to $x$, denoted $int(x,y)$, are the common out-neighbors of both $x$ and $y$:
 $int(x,y)=N^{+}(x)\cap N^{+}(y)$
-The number of interior neighbors is the interior degree of $y$ with respect to $x$.
+The number of interior neighbors is the interior degree of
+$y$ with respect to $x$.
 Now that we have introduced containers, we can talk about arcs where both nodes are inside the containers; these are interior neighbors.
@@
 Definition 5.3.
-Let $x$ be a node with the decreasing sequence property, and let $y\in N^{+}(x)$.
-The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
+Let $x$ be a node in $G$, and let $y\in N^{+}(x)$. The exterior neighbors of $y$ with respect to $x$, denoted $ext(x,y)$, are the out-neighbors of $y$ that are second neighbors of $x$:
 $ext(x,y)=N^{+}(y)\cap N^{++}(x)$
@@
 To do so, though, we need to consider the option of back arcs and see how they might hurt our algorithm.
-These have the possibility of altering a container’s distance to the minimum degree node.
+These have the possibility of altering a container’s distance to $v_{0}$.
 Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
-Having defined our arcs, we can talk about the decreasing sequence property.
-This property is explaining the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from the minimum degree node.
-Exterior degrees, though, are a part of a node’s degree equation.
-There’s another part of it though.
-As we move further from the minimum degree node, the size of the interior degrees increases.
-This phenomenon reflects a yin and yang relationship between the interior and exterior degrees.
-In all of this, the minimum degree node is serving as a catalyst and driving force for the process.
-At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees.
-With this definition of interior and exterior nodes now introduced, we are now ready to introduce the Graph Level Order (GLOVER) data structure.
+We will speak more about back arcs in the next section.
 Definition 5.5.
@@
 Proof.
-To establish a graph level order, we first need a minimum degree node.
-So we take the minimum value on the degree vector of nodes.
-We will call that node $v_{0}$.
+To establish a graph level order, we first need a minimum degree node. So we take the minimum value on the degree vector of nodes. We will call that node $v_{0}$.
 Knowing the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
@@
 For any two nodes $u$ and $v$, this total order is determined by their containers.
-  • Leveled Structure with Containers - By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
-  • Comparability within Containers - For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
+  • Leveled Structure with Containers—By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
+  • Comparability within Containers—For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
   • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
@@
 Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a GLOVER under the distance total order, completing the proof.
-What the GLOVER data structure does is transform the partitions and distance metric into an ordering for containers.
+What the GLOVER data structure does is transform the containers and distance metric into an ordering for containers.
 This ordering is more than just representative, though.
 The containers help to organize the nodes.
-What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from the chosen minimum degree node.
+What this suggests is that each partition, or set of nodes, is being totally ordered as we move away from $v_{0}$.
 This is a clear distinction from standard graph theory.
 Graph nodes are generally thought of as able to be placed anywhere in 2-dimensional space.
-However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from the chosen minimum degree node.
+However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$.
 The one thing that could disrupt this total ordering is a back arc.
@@
 Lemma 5.1.
-GLOVER Supports Induction In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
+GLOVER Supports Induction
+In any oriented graph, without back arcs, the minimum degree node $v_{0}$ within a GLOVER structure provides a well-ordering, enabling an inductive analysis of all other nodes in the graph.
 Proof.
-We have already shown that the containers in GLOVER form a total order.
-Further, they are a discrete total order.
-Because of this, they are a well ordering.
+We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
 A set is considered well-ordered if every non-empty subset of that set has a least element.
@@
 Every non-empty subset will have a least element, which is the defining characteristic of a well-ordered set.
-∎
 This becomes an important concept because now, in addition to the nodes, we also have the containers as subsets of $S$.
 These containers are totally ordered and thus can also be used as a means of mathematical induction.
-We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from the chosen minimum degree node $v_{0}$.
+We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from $v_{0}$.
+One of the additional benefits of partitioning the oriented graph into containers with interior arcs is that each container resembles its own local SSNC.
+In this instance, the interior arc has an interior degree, and if we are able to find a positive solution to this instance, we can return to the global problem with a node whose interior degree has doubled.
+Then, under the assumption of no back arcs, we would only need to prove that this node has its exterior degree double to prove that it is a solution to the SSNC.
+This is another benefit of the GLOVER data structure, as it provides us with a means to execute divide-and-conquer.
+This does not only help with the search for the degree-doubling node.
+As we shall see in the coming sections, there are multiple lemmas that will be dependent on the condition of an interior degree doubling.
+What that is stating is that the local subproblem has a solution to the SSNC.
+This overall proof strategy is through contradiction.
+We are attempting to prove the SSNC by assuming that it is not true and observing the patterns that must be present in the resulting oriented graphs.
+As we observe these patterns that must be present, we will continue to keep a log until we arrive at a contradiction.
+6 Using the Data Structure
+Having defined the GLOVER data structure, we are ready to begin using it.
+The first and most important concept that we will define that will guide our research is the decreasing sequence property.
+With that and the GLOVER data structure, and the ammunition of knowing that any oriented graph without back-arcs can be represented in a GLOVER, we can assume we are working in this environment.
+The decreasing sequence property explains the nature of how the exterior degrees, and thus the sizes of the containers, decrease as we get further from $v_{0}$.
+Exterior degrees, though, are a part of a node’s degree equation.
+There’s another part of it, though.
+As we move further from $v_{0}$, if there are no back arcs, the size of the interior degrees must increase.
+This phenomenon reflects a yin and yang relationship between the interior and exterior degrees.
+In all of this, the minimum degree node, $v_{0}$, is serving as a catalyst and driving force for the process.
+At the other end of the spectrum, the decreasing sequence property is the bottleneck limiting the expansion of the graph’s out-degrees.
+The case where $d^{+}(v_{0})=3$ and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) are important because they provide information about what a counterexample to the conjecture would involve.
+A graph $G$ was constructed in which the out-degree of $v_{0}$ does not double in the square $G^{2}$.
+A simple graph of degree three corresponds to the smallest cycle that can exist in an undirected graph.
+Therefore, if the edges are oriented in the same direction, a cycle of order three can be created among the container adjacent to $v_{0}$.
+If we look at this situation further, we see that in this example, a node that is an out-neighbor of $v_{0}$ had its out-degree (at least) double.
+This section will initiate the examination of the adequate conditions for other nodes in the graph to not have their degrees double.
+However, we will need to define new concepts to proceed with this line of reasoning.
+To prevent a node’s out-degree from doubling, we must first formally define the decreasing sequence property.
+This property is attributed to a node in $G$ whose out-degree does not double in $G^{2}$.
+There are instances where $v_{0}$ lacks its out-degree double in $G^{2}$, as we have seen.
+We’ll investigate which other nodes can possess the decreasing sequence property and which network features are necessary for this to occur.
 Example 6.1.
-[…]
-Figure 6: This example illustrates back arcs in a graph.
-The two blue arcs from $x_{3}$ to $y_{3}$ and $y_{3}$ to $z_{3}$ represent exterior neighbors indicating forward-directing arcs.
-Additionally, a red arc points backwards from $z_{5}$ to $x_{1}$.
-Although the containers are not outlined, they are implied by the difference in variables for each group.
+0
+$\Rightarrow$ 1111222233334444d $\Rightarrow$ d + $1$d + $2$d + 33332*d-1 $\Rightarrow$
+Figure 5: This is an illustration where node $v_{0}$ has the Decreasing Sequence Property (DSP). There are more first neighbors $d$ of $v_{0}$ than second neighbors.
+Definition 6.1.
+For a node $x\in V(G)$ in an oriented graph $G$, we say that $x$ has the decreasing sequence property if the size of its first neighbors is strictly larger than the size of its second neighbors, i.e., $|N^{+}(x)|>|N^{++}(x)|$.
+The SSNC asks for a node whose second neighborhood is at least as large as its first.
+Our strategy is to prove the conjecture by contradiction.
+We assume the decreasing sequence property holds for all nodes in $G$.
+The definition of the decreasing sequence property is the negation of what the conjecture asks for.
+This assumption allows us to shift our focus from directly searching for a degree-doubling node to analyzing the graph’s global composition.
+One notable change is that the decreasing sequence property calls for the second neighborhood to be strictly less than the first.
+This could set up a decreasing exterior condition, which could prove strong.
+We anticipate that this exploration will reveal constraints on the graph’s arrangement that will ultimately lead to a contradiction.
+Two competing new ideas have been presented: the decreasing sequence property (DSP) and the minimal degree node.
+A crucial starting point for our approach and degree values is a minimal degree node.
+It acts as the anchor by being chosen as the beginning point.
+The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node.
+It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
+In contrast, the DSP introduces constraints on connectivity.
+This concept is akin to a bottleneck.
+A bottleneck in graph theory represents a point that limits flow or connectivity within the network[6], a role precisely filled by the DSP.
+Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
 Lemma 6.1.
-Interior Degree at least 1 Suppose we have an oriented graph $G$, represented by a GLOVER.
+Decreasing Sequence Property Lemma Suppose a node in an oriented graph $G$ does not have the decreasing sequence property. Then this node will have its out-degree at least double in $G^{2}$.
+If we have a node $x$ with the decreasing sequence property, and $y$ is an out-neighbor of $x$, we will partition $y$’s neighbors into two sets, those that are also neighbors of $x$ and those that are not neighbors of $x$.
+In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed.
+We will now give terms to these concepts.
+In this work, we assume that the input graph is oriented and unweighted.
+We will assume that all nodes in this graph have the decreasing sequence property.
+What that means is that every node’s second neighborhood is strictly smaller than its first neighborhood.
+In the case of Seymour’s Second Neighborhood Conjecture (SSNC), this is still the setup for a proof by contradiction.
+In the previous section, we showed that any unweighted oriented graph without back arcs could be represented in a GLOVER data structure.
+Instead of allowing nodes to be placed anywhere across the plane, we will take advantage of this structure here.
+The nodes will be partitioned into containers with interior arcs to the current container and exterior arcs to the next container.
+For the remainder of this paper, when we speak of oriented graphs, the underlying assumption is that it is an oriented graph in a GLOVER data structure.
+Lemma 6.2.
+Interior Degree at least 1 Suppose we have an oriented graph $G$.
 Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property.
@@
 Proof.
-By definition, a node satisfies the decreasing sequence property if $|N^{+}(x)|>N^{++}(x)|$ If $x$ has $d^{+}(v_{0})$ exterior arcs, this would make $v_{0}$’s degree double because $x$’s neighbors are $v_{0}$’s second neighbors.
-To avoid this doubling, $x$ must connect to at least one interior arc.
-This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
+By definition, a node satisfies the decreasing sequence property if
+$|N^{+}(x)|>N^{++}(x)|$
+If $x$ has $d^{+}(v_{0})$ exterior arcs, this would make $v_{0}$’s degree double because $x$’s neighbors are $v_{0}$’s second neighbors.
+To avoid this doubling, $x$ must connect to at least one interior arc. This means that the interior degree of $x$ with respect to $v_{0}$ must satisfy
 $|int(v_{0},x)|\geq 1$
-This is true for any neighbor $x\in C_{1}$. This ensures that every node in $C_{1}$ has an interior degree of at least 1. Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
+This is true for any neighbor $x\in C_{1}$. This ensures that every node in $C_{1}$ has an interior degree of at least 1.
+Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
 Lemma 6.2 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree doubling.
@@
 Lemma 6.3.
-Exterior Less than Degree Suppose we have an oriented graph $G$, represented by a GLOVER.
-In an attempt to prevent a degree-doubling node, if $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
+Exterior Less than Degree
+Suppose we have an oriented graph $G$. If $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
 Proof.
@@
 Lemma 6.3 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 6.2 (Interior Degree at least 1), with the notable exception that Lemma 6.3 refers to any node with the DSP, whereas we only made Lemma 6.2 true for the minimum degree node.
-The results presented in 5.4 (Interior Degree At Least 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1.
+The results presented in 6.2 (Interior Degree At Least 1) and 6.3 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1.
 There, he demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles.
-Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards an global search solution with the algorithm we are developing.
+Despite the overlap, this paper will be moving in a different direction with the containers and the decreasing sequence property towards a global search solution with the algorithm we are developing.
 Lemma 6.4.
-Weak Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph represented by a GLOVER.
+Weak Decreasing Exteriors Lemma Suppose that $G$ is an oriented graph.
 Suppose further that $x$ is a node in this graph, with $y\in N^{+}(x)$ such that both $x$ and $y$ satisfy the decreasing sequence property (DSP), and suppose that we are trying to prevent a degree doubling node.
@@
   2. 2.
-    For all $z\in ext(x,y)$, we have $ext(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
+    For all $z\in\text{ext}(x,y)$, we have $\text{ext}(y,z)\cap N^{+}(x)=\emptyset$, i.e., no back arcs from the second neighbors of $x$ to the first neighbors of $x$ exist.
 Then, for all $z\in\text{ext}(x,y)$, we have $|\text{ext}(y,z)|<|ext(x,y)|$. That is, the exterior degree of $y$ with respect to $x$ decreases for every second neighbor $z$ of $x$.
@@
 That is a separate version of the SSNC, but localized to the containers.
-The conjecture in this case is can we find a node in each container whose degree doubles.
+The conjecture in this case is, can we find a node in each container whose degree doubles?
 If so, then it gives light to this divide-and-conquer approach, where we partition the containers, solve them individually, and then combine them again for this lemma 6.4 (Weak Decreasing Exteriors Lemma).
-The more popular representations of graphs are adjacency matrices, arraylists, and edge lists.
+The more popular representations of graphs are adjacency matrices, array-lists, and edge lists.
 These suffer from the same limitations of graphs as mentioned above; they scatter the nodes along the two-dimensional plane without attempts to group them into an ordering.
@@
 To represent graphs in a GLOVER, we can use any data serialization that can encode graphs, including JSON, YAML, and XML.
-There needs to be three types of data: containers, nodes, and arcs.
+There need to be three types of data: containers, nodes, and arcs.
 The containers will be given IDs.
@@
 Each container has fewer nodes.
-The fewer nodes implies that the container circle can be drawn with a smaller radius, giving light to the decrease in size.
-Example 5.4.
-[…]
-Figure 7: Illustration of the Weak Decreasing Exteriors Lemma.
-We have a node $x$ in container $C_{i}$ connecting to three nodes in container $C_{i+1}$.
-In order to prevent $x$’s degree from doubling, we need to have $|C_{i+2}|<|C_{i+1}|$.
+The fewer nodes imply that the container circle can be drawn with a smaller radius, giving light to the decrease in size.
 To further illustrate the Graph Level Order (GLOVER) data structure, imagine a social media example where users can follow one another, but no one follows back.
@@
 Considering example 6.3, which takes the former of these approaches.
-These cycles in the first container are what stops the user Andi from being an influencer.
+These cycles in the first container are what stop the user, Andi, from being an influencer.
 This brings us to the next container.
-Lemma 5.6 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 5 people—Hugo, Isla, Jade, Kai, and Lei.
-If any person in this container followed less than two other people in this same container, then they would need to follow at least five people in the next container (container 3).
+Lemma 6.4 (Weak Decreasing Exteriors Lemma) lets us know that this container will have 8 people—Kai, Lei, Mia, Nova, Owen, Pat, Qin, and Ryan.
+If any person in this container followed less than two other people in this same container, then they would need to follow more people in the next container (container 3).
 From this we would conclude that a person in container 1, whose indirect sphere of influence is this person’s sphere of influence, would have its first degree double and its second degree double, implying that its overall degree doubles.
 Example 6.2.
+[…]
+Figure 6: Illustration of the Weak Decreasing Exteriors Lemma.
+We have a node $x$ in container $C_{i}$ connecting to three nodes in container $C_{i+1}$.
+In order to prevent $x$’s degree from doubling, we need to have $|C_{i+2}|<|C_{i+1}|$.
+Example 6.3.
 Andi
-$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$ H
+$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$
 Figure 7: This is an example of a social media setup of the SSNC. The second container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
@@
 If we look at Example 4.2 again, we see that $N^{+}(v_{0})=\{v_{1},v_{2},v_{3}\}$, $|int(v_{0},v_{1})|=|\{v_{2}\}|=1$ since $(v_{1},v_{2})\in G$.
-Similarly, we see that $|ext(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
+Similarly, we see that $|\text{ext}(v_{0},v_{1})|=|\{v_{4},v_{5}\}|=2$ since $(v_{1},v_{4})\in G$ and $(v_{1},v_{5})\in G$ and $(v_{0},v_{4})\in G$ and $(v_{0},v_{5})\in G$.
 One of the things we notice about Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) when $v_{0}$’s out-degree did not double in $G^{2}$, compared to the previous examples when it did, is that $v_{0}$ in Lemma 4.3 allowed for a cycle to exist in the out-neighbors of $v_{0}$.
@@
 […]
-Figure 9: Graph representation of back arcs (shown in red) and forward-directed arcs (in blue) between containers, illustrating how these arcs impact connectivity patterns in an oriented graph.
+Figure 8: This example illustrates back arcs in a graph.
+The two blue arcs from $x_{3}$ to $y_{3}$ and $y_{3}$ to $z_{3}$ represent exterior neighbors indicating forward-directing arcs.
+Additionally, a red arc points backwards from $z_{5}$ to $x_{1}$.
+Although the containers are not outlined, they are implied by the difference in variables for each group.
 Lemma 7.1.
-(Regular Interior Degrees) Let $G$ be an oriented graph represented by a GLOVER.
+(Regular Interior Degrees) Let $G$ be an oriented graph.
 Let $v_{0}$ be a minimum degree node of $G$.
@@
 These exterior neighbors of $u$ are raising the second neighbors of $v_{p}$.
-As a result, the exterior degree of $v_{p}$ will double. Because $v_{p}$ has already had its interior degree doubled, it will have its overall degree double
-Thus, the different interior degrees of $u$ and $v$ in container $C_{i}$ leads to a contradiction, so the claim holds. ∎
+As a result, the exterior degree of $v_{p}$ will double. Because $v_{p}$ has already had its interior degree doubled, it will have its overall degree doubled.
+Thus, the different interior degrees of $u$ and $v$ in container $C_{i}$ lead to a contradiction, so the claim holds. ∎
 Lemma 7.1 (Regular Interior Degrees) provides another insight into the structure within the containers of these counterexamples.
@@
 The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which contrasts with this.
+The assumptions for this letter are that we have one parent node whose interior degree doubles, and it maps the two child nodes that have different interior degrees.
+What happens is that because these child nodes have different interior degrees, it causes this pair node to have an exterior degree that doubles by the week exterior decreasing lemma.
+We spoke previously about how the divide and conquer approach is allowed by the GLOVER data structure.
+This is now the second lemma that we’ve seen call for interior degrees doubling.
+That is going to be a staple of this algorithm as we see the divide and conqueror approach called for more often as this interior degree doubling is part of the nodes being called back into the global solution.
+Remember that interior degrees doubling are local instances of solving the Seymour second neighborhood conjecture on the container itself.
+Instead of seeking a degree-doubling node, these containers seek an interior degree-doubling node.
+What remains to be seen is if interior degree doubling is an effective solution.
+That will show in the next section.
 Next, we begin to look at how back arcs interact with the data structure and the development of the algorithm.
@@
 Consider the following table based on example 6.3.
-Table 1: Consequences of Back Arcs in 5.5
+Table 1: Consequences of Back Arcs in 6.3
 Back Arc From(A) Back Arc To (B) Back Arc New Neighbors via $N^{+}(B)$
@@
 Lemma 7.2.
-No Back Arc to Early Container Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node.
+No Back Arc to Early Container Let $G$ be an oriented graph.
 If a node $u_{i}$ in a container $C_{i}$ has its first (earliest) back arc to a node $v_{k}$ in container $C_{k}$ where $k<i-1$, then the size of the exterior neighbors $\text{ext}(u_{i},v_{k})$ is at least the degree of $v_{k}$.
@@
 ∎
+The proof above demonstrates that a node cannot have a back arc to an early container.
+To understand this proof, we must first investigate the assumptions.
+Here, we are assuming two things.
+That the node $u_{i}$ has its interior degree doubled and that the first back arc from $u_{i}$ is to the node $v_{k}$.
+As we saw in the proof, $v_{k}$ has more exterior neighbors than you survive, which is what causes $u_{i}$’s degree to double.
+The question is whether we need the condition of $u_{i}$ interior degree doubling.
+The weak decreasing exterior lemma tells us that we have a relationship between the exterior degrees of use by and $v_{k}$, not the overall degrees of these nodes.
+By assigning the interior degree to double, we can leverage this lemma.
+Secondly, we look at the notion of the first back arc.
+The containers are totally ordered.
+By mapping $u_{i}$ to its first back arc and showing that this back arc cannot exist, we have shown that $u_{i}$ cannot have a first back arc.
+Because $u_{i}$ cannot have a first back arc, this back arc of this type cannot exist.
 Lemma 7.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container-structured graph.
@@
 Lemma 7.3.
-Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
+Back Arcs Special Case
+Let $G$ be an oriented graph, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$.
   1. 1.
@@
 $|\text{ext}(y,z)|<|\text{ext}(x,y)|$
-First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
+First, we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
 This gives the only possibility for $z$ as being an interior neighbor of $x$.
-Consequently, the arc $w,z$ is an early back arc as we saw in 6.2 (No Back Arc to Early Container).
+Consequently, the arc $w,z$ is an early back arc as we saw in 7.2 (No Back Arc to Early Container).
 These arcs can lead to double counting.
@@
 First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$.
-Suppose that we represent this graph by a GLOVER.
-Suppose we are attempting to prevent a degree-doubling node and that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$.
+Suppose that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$.
 If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
@@
 Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$. ∎
-Lemma 6.4 (No Back Arc to Early Container) is a special case where we can apply Lemma 6.1 (Regular Interior Degrees).
+Lemma 7.4 (No Back Arc Between Consecutive Containers) is a special case where we can apply Lemma 7.1 (Regular Interior Degrees).
 From this, we can see that a back arc to a parent container violates the oriented nature of the graph.
 Lemma 7.5.
-No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, represented by a GLOVER in an attempt to prevent a degree-doubling node, with a minimum degree node $v_{0}$ and containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$, there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
+No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, with a minimum degree node $v_{0}$.
+Let containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$.
+Then there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
 Proof.
-This was proved in the lemmas 6.2 (No Back Arc to Early Container) and 6.4 (First Back Arc Between Consecutive Containers). ∎
-[…]
-Figure 10: Illustration of the proof of 7.2 (Interior Degrees Double).
-This is another illustration of the Interior Degree Doubling.
-Each node inside a container represents a node whose degree is doubled by the inductive hypothesis.
-This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
-[…]
-Figure 11: Illustration of the proof of 7.2 (Interior Degrees Double).
-Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
+This was proved in the lemmas 7.2 (No Back Arc to Early Container) and 7.4 (First Back Arc Between Consecutive Containers). ∎
 We stated that back arcs could be detrimental to our data structure.
@@
 This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
-Lemma 7.1.
-Indistinguishable Regular Interior Nodes Suppose we have a regular container $C$ of size $n$ of an oriented graph $G$ represented by a GLOVER.
-We are attempting to prevent a degree-doubling node.
-If a node $v_{p}$ ’s interior degree doubles, then every other node $v_{i}$ in the same container as $v_{p}$ will also have its interior degree double.
-Proof.
-Let $C_{i}$ be a regular container.
-Let $v_{p}$ be a node in this container, where $int(u,v_{p})$ doubles in $C_{i-1}$ for some $u\in C_{i-1}$.
-In a regular container, all nodes are indistinguishable in terms of their degree.
-They also have the same parents because of containers.
-Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve the degree structures.
-This means any node will have its interior degree doubled under the automorphism $\phi$.
-∎
-The crucial takeaway from Lemma 7.1 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others.
-Previously we were searching for a node in a container whose interior degree had doubled.
-That would be the key.
-What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the nodes within that container will have their interior degrees double.
-That means that we can select any of those nodes.
+Example 8.1.
+[…]
+Figure 9: Illustration of the proof of 8.1 (Interior Degrees Double).
+This is another illustration of the Interior Degree Doubling.
+Each node inside a container represents a node whose degree is doubled by the inductive hypothesis.
+This process cannot continue forever, so we must reach the last node, and the parent of this last node in the sequence is the node whose exterior degree will double.
 We still need to prove that for each container there exists a node whose interior degree doubles.
 We have seen this in the example 6.3, but those could be too simple to observe any real patterns.
-In similar structure, what we saw in 4.1 (Minimum Out-Degree $<$ 3) was that it was easy to find degree-doubling nodes in graphs with a minimum degree of 0, 1, or 2.
+In a similar structure, what we saw in 4.1 (Minimum Out-Degree $<$ 3) was that it was easy to find degree-doubling nodes in graphs with a minimum degree of 0, 1, or 2.
 However, when that minimum degree was increased to 3, there were more cases that needed to be considered.
+Example 8.2.
 […]
 Lemma 8.1.
-Interior Degrees Double Let $G$ be an oriented graph represented by a GLOVER.
-We are attempting to prevent a degree-doubling node.
+Interior Degrees Double Let $G$ be an oriented graph.
 Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP).
-Then, at least one node in each container $C_{i}$ will have its interior degree doubled.
+Then, at least one node in each container $C_{i}$ will have a node with its interior degree doubled.
 Proof.
@@
 By the Induction Hypothesis (I.H.), there exists a node $y_{1}\in H_{1}$ such that the interior degree of $y_{1}$ with respect to $u_{i}$ doubles in $H_{1}$.
-We can now define $H_{2}=ext(u_{i},y_{1})$.
-Likewise, by the I.H., there will be a node $y_{2}\in H_{2}$ whose interior degree doubles in $H_{2}$.
+We can now define $H_{2}=ext(u_{i},y_{1})$. Likewise, by the I.H., there will be a node $y_{2}\in H_{2}$ whose interior degree doubles in $H_{2}$.
 We can define $H_{3}=ext(y_{1},y_{2})$.
@@
 Therefore, at least one node in each $C_{i}$ attains a doubled interior degree, completing the inductive step. ∎
-Lemma 7.2 (Interior Degree Double) is a local version of Seymour’s Second Neighborhood Conjecture.
+Lemma 8.1 (Interior Degrees Double) is a local version of Seymour’s Second Neighborhood Conjecture.
 We were able to isolate the containers as an oriented graph and searched for a node whose (interior) degree doubles.
+We were able to set up a GLOVER data structure inside this local version starting with a minimum degree node of the container.
+We can then form local containers that are each their own respective distance from the minimum degree node.
+By strong induction, we have that each of these containers will have a node whose interior degree doubles.
+Then by the weak decreasing exteriors lemma, we have a decreasing set of exteriors inside this container.
+When this sequence gets down to size two, that node will have all exterior neighbors.
+Therefore, a parent of that node will have its degree doubled.
 By finding a positive answer to this, and not just a positive answer but a constructive algorithm that produces a positive answer, it gives us promise towards progress resolving the overall conjecture.
@@
 Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples.
-The questions of whether this was always true or whether back arcs like lemma 6.3 could force this into being a false positive.
+The question is whether this was always true or whether back arcs like lemma 7.3 could force this into being a false positive.
 This proof also helps with explaining the concept of the SSNC.
@@
 Once we have done this, there are containers spread out from a chosen minimum degree node.
-Now for each container we can simply run the conjecture on that container.
+Now for each container, we can simply run the conjecture on that container.
 The difference is that with lemma 8.1 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container.
-Then with lemma 7.1 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have their interior degree double.
+Then with lemma 9.2 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have its interior degree doubled.
+Example 8.3.
+[…]
+Figure 10: Illustration of the proof of 8.1 (Interior Degrees Double).
+Each circle represents a new container; by the I.H., there exists a $y$ whose interior degree doubles, leading to decreasing exteriors.
 This notion of interior degrees will help us to build our algorithm in the search for a counterexample. It also gives us a bit of light into the possibility of the conjecture being true.
@@
 Lemma 9.1.
-Regular Exterior Degrees Let $G$ be an oriented graph represented by a GLOVER in an attempt to prevent a degree-doubling node.
+Regular Exterior Degrees Let $G$ be an oriented graph.
 Let a minimum degree node of $G$ be $v_{0}$, and suppose every node in a container $C_{i}$ has the Decreasing Sequence Property (DSP).
@@
 This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
-Because Lemmas 7.2 (Interior Degree Doubling) and 6.5 (No Back Arcs) both hold, we can apply Lemma 5.6 (Weak Decreasing Exteriors Lemma) here.
+Because Lemmas 8.1 (Interior Degrees Doubling) and 7.5 (No Back Arcs) both hold, we can apply Lemma 6.4 (Weak Decreasing Exteriors Lemma) here.
 This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
@@
 That means there will be two neighboring containers of equal size, $C_{i-1}$ and $C_{i}$.
-So, a node in $C-{i-2}$ can have its exterior degree double.
-By 7.2 and 7.1 (Indistinguishably of Regular Nodes), we know that this node has its interior degree double as well.
+So, a node parent node in $C-{i-2}$ can have its exterior degree double.
+By 8.1, we know that this node has its interior degree double as well.
 So, this assumption led to a contradiction. ∎
-This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 5.6 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
+This proof begins with two nodes in the same container.
+By that assumption and the regularity lemma on interior degrees, we know that these nodes have the same interior degree.
+Our next assumption is that these two nodes have different exterior degrees.
+We simply want to observe what effect that has on a parent of these nodes.
+The irregular exterior degree causes a ripple in the decrease in exteriors, which should be guaranteed by the weak Decreasing Exteriors Lemma.
+The container $C_{i+1}$ is now the same size as $C_{i}$.
+This means that a node in $C_{i-1}$ will have their exterior degree double.
+This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 6.4 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
 Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double.
@@
 Lemma 9.2.
-Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph represented by a GLOVER.
-Let $v_{0}$ be a minimum degree node in $G$.
-We are attempting to prevent a degree-doubling node.
-For any node $x_{i}$ in container $C_{i}$ relative to $v_{0}$, the size of the exterior neighbors, and thus $C_{i}$ must be bounded by
-|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1)$
-for some $z\in N^{+}(x_{i})$.
+Indistinguishable Regular Interior Nodes Suppose we have a regular container $C_{i}$ of size $n$ of an oriented graph $G$.
+If a node $v_{i}$’s interior degree doubles, then every other node $v_{j}$ in the same container as $v_{i}$ will also have its interior degree double.
 Proof.
-We will prove the statement by induction on the containers $C_{i}$ relative to $v_{0}$.
-Base Case ($i=1$): Consider a first neighbor $x_{1}\in C_{1}$ (the nodes at distance 1 from $v_{0}$).
-The size of the exterior set
-$\{y\in\text{ext}(x_{1},z)\text{ s.t. }dist(v_{0},y)=2\}$
-is bounded by $d^{+}(v_{0})-1$. Each first neighbor $x_{1}$ can connect to at most $d^{+}(v_{0})-1$ nodes at distance 2. We have already proved that there are no back arcs to $v_{0}$.
-Inductive Hypothesis: Assume that for any node $x_{i}$ in container $C_{i}$, the size of its exterior neighbors is bounded by $d^{+}(v_{0})-(i-1)$. Specifically, we assume:
-|$|\{y\in\text{ext}(x_{i},z)\text{ s.t. }dist(v_{0},y)=i+1\}|\leq d^{+}(v_{0})-(i-1).$
-Inductive Step: We need to show that this bound holds for nodes at distance $i+1$ from $v_{0}$. Let $x_{i+1}$ be a node in container $C_{i+1}$ relative to $v_{0}$. Suppose, for contradiction, that
-|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|>d^{+}(v_{0})-i.$
-This would imply that the exterior neighbors of $x_{i+1}$ exceed the allowed bound, violating the decreasing sequence property (DSP) conditions of $x_{i}$.
-By Lemmas 7.2 (Interior Degrees Double), 6.1 (Regular Interior Degree), and 7.1 (Indistinguishable Regular Interior Nodes), this would cause the interior degree of $x_{i}$ to double, thereby forcing an overall doubling of the degree for $x_{i}$.
-This contradiction implies that our assumption must be false. Therefore, we conclude:
-|$|\{y\in\text{ext}(x_{i+1},z)\text{ s.t. }dist(v_{0},y)=i+2\}|\leq d^{+}(v_{0})-i.$
+Let $C_{i}$ be a regular container, and Let $v_{i}$ be a node in this container.
+Further, let $v_{p}$ be a parent of $v_{i}$.
+Then $\text{int}(v_{p},v_{i})$ doubles in $C_{i}$ by 8.1 (Interior Degrees Doubling).
+In a regular container, all nodes have the same interior by 7.1 (Regular Interior Degrees) and exterior degrees by 9.1 (Regular Exterior Degrees).
+They also have the same parents because of containers, also by 9.1.
+Any automorphism we construct, including a random automorphism $\phi$ of the container $C_{i}$, will preserve this degree structure.
+This means any node will have its interior and exterior degree doubled under the automorphism $\phi$.
+Therefore, the node $v_{j}$ which is mapped to by $\phi$ will have its interior degree doubled.
+∎
+Now that we have a collection of nodes inside a container that all have equal interior and exterior degrees, as well as the same parents.
+There becomes a question of how we can distinguish between these nodes.
+The truth is that we cannot distinguish between them.
+As a result, we are led to conclude that if one node in a container has its interior degree double, then all nodes in that container will also have this property.
+The crucial takeaway from Lemma 9.2 (Indistinguishable Regular Interior Nodes) is that regularity allows us to treat any node within a regular container as indistinguishable from others.
+Previously we were searching for a node in a container whose interior degree had doubled.
+That would be the key.
+What indistinguishable tells us is that if one node’s interior degree doubles inside a container, then all the nodes within that container will have their interior degrees double.
+That means that we can select any of those nodes.
+Lemma 9.3.
+Strong Decreasing Exteriors Lemma Let $G$ be an oriented graph, where $v_{0}$ is a minimum degree node in $G$.
+Then, for any container $C_{i}$, the size of the exterior set leading to container $C_{i+1}$ is bounded by
+|$|\{y\in\text{ext}(C_{i},z)\text{ s.t. }y\in C_{i+1}|\leq d^{+}(v_{0})-i$
+Proof.
+We will prove the statement by induction on the container index $i$.
+Base Case ($i=1$):
+Container $C_{1}$ consists of the out-neighbors of $v_{0}$, which are distance 1.
+By definition, the exterior set at distance 2 is composed of the neighbors of nodes in $C_{1}$.
+Since $v_{0}$ has a degree of at least 1, we obtain:
+$|{y\in\text{ext}(C_{1},z)\text{ s.t. }y\in C_{i}}\leq d^{+}(v_{0})-1.$
+Inductive Hypothesis:
+Assume that for some $i$, the bound holds for container $C_{i}$, i.e.,.,
+|$|\{y\in\text{ext}(C_{i},z)\}\text{ s.t. }y\in C_{i+1}\}\leq d^{+}(v_{0})-i.$
+Inductive Step:
+Consider container $C_{i+1}$.
+These are the out-neighbors of nodes in $C_{i}$.
+Each node in $C_{i}$ satisfies the Decreasing Sequence Property (DSP), its second neighborhood is smaller than its first neighborhood.
+This implies that the number of new nodes in $C_{i+1}$ must be strictly smaller than the size of $C_{i}$.
+Alternatively, we could apply the 6.4 Weak Decreasing Exteriors Lemma To $C_{i}$.
+We observe that since there are no back arcs and by 8.1 (Interior Degrees Doubling) there exists a node whose interior degree doubles, it follows that $C_{i+1}$ is strictly smaller than $C_{i}$.
+Thus, the size of the exterior set satisfies:
+|$|\{y\in ext(C_{i+1},z)\text{ s.t. }y\in(C_{i+2}\}|\leq d^{+}(v_{0})-(i+1).$
+By induction, the bound holds for all container levels $i$
 Lemma 9.3 (Strong Decreasing Exteriors Lemma) uses Lemma 6.4 (Weak Decreasing Exteriors Lemma) to give us a way of calculating the maximum size of a container based only on its distance from the minimum degree node and the minimum degree node’s degree.
-In Section 5 (Theoretical Foundation), we demonstrated that mathematical induction could be applied to our framework.
+In Section 5 (Building a Data Structure), we demonstrated that mathematical induction could be applied to our framework.
 This proof highlights the utility of the lemma.
@@
 It is possible that this result could have been proven without relying on induction.
-We could have used a different approach or simply avoided induction based on container distance from the minimum degree node entirely.
-Still, the use of induction provides an important method, particularly for graphs where the usage of induction is more limited.
+We could have used a different approach or simply avoided induction based on container distance from $v_{0}$ entirely.
+Still, the use of induction provides an important method, particularly for graphs, where the usage of induction is more limited.
 Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
+Similar to the proof of lemma 6.4 (Weak Decreasing Exteriors Lemma), we assume here that there is a node in each container whose degree doubles.
+The difference is that that assumption does not need to be explicitly stated because we have already proved that there is a node whose interior degree doubles in each container.
+Similarly, this lemma requires there to be no back arcs, like the weak decreasing exteriors lemma.
+We’ve already proved that back arcs cannot exist without creating a degree doubling node, though.
+As a result, we do not need to state that assumption.
+The proof then simply iteratively executes lemma 6.4 (Weak Decreasing Exteriors Lemma), counting the size of the exteriors as we extend our distance from $v_{0}$.
+This can be summed up by the formula $|C_{i+i}|<d^{+}(v_{0})-i$.
 Lemma 9.4.
-Interior Degrees Lower Bound Let $G$ be an oriented graph represented by a GLOVER.
+Interior Degrees Lower Bound Let $G$ be an oriented graph.
 Let $v_{0}$ be a minimum degree node in $G$.
@@
 Having already placed an upper bound on the exterior sizes and containers, it is only natural that there is a contrasting bound for the interiors.
-Once again, the minimum degree node is the driving force, the catalyst, behind this.
+Once again, the minimum degree node $v_{0}$ is the driving force, the catalyst, behind this.
 The degree of these interior nodes will also be impacted by their distance to the minimum degree node.
-Lemma 8.4.
-Regular Graph All Nodes Have Equal Out-Degree Let $G$ be an oriented graph represented by a GLOVER.
-Suppose $v_{0}$ is a minimum degree node of $G$.
-Further, let all nodes satisfy the Decreasing Sequence Property (DSP).
-We are attempting to prevent a degree-doubling node.
-Then for any node $v$ in $G$, the degree of $v$ is equal to $d$, where $\delta=d^{+}(v_{0})$.
-Proof.
-For a node $u_{i}$ in container $C_{i}$, with degree $\delta+1$, there would be the possible places for this degree to be located—back arcs, interior arcs, or exterior arcs.
-We have already shown that back arcs cannot exist.
-If this extra degree went to the exterior arcs, then it would cause the degree of a parent node to double.
-The only remaining possibility is for this extra degree to go to the interior arcs of $u_{i}$.
-While this is not a problem, it will cause the interior degrees to increase at a faster rate than we just showed in Lemma 8.3 (Interior Degree Lower Bound), as this would increase the number of interior arcs.
-This more rapid reduction in the interior degree leads to container density faster, which would simply lead to a degree doubling node faster. To avoid this, overall regularity is key.
-Thus, every node in every container’s degree is set by the minimum degree node. This leads to the regularity of the graph. ∎
-We are incrementally refining the properties that must hold for our counterexamples.
-The two most crucial aspects are the decreasing sequence property and the identification of the minimum degree node.
-However, additional constraints, such as the overall graph regularity established by Lemma 8.4 (Regular Graph), help simplify future arguments.
-This lemma restricts the graph’s structure, allowing us to consider fewer cases as we move forward with our proofs.
+Once again, we see a proof with not much that needs to be assumed.
+The only real assumption here is that we have a container $C_{i}$ in the oriented graph $G$ that is distance $i$ from $v_{0}$.
+Lemma 9.3 (Strong Decreasing Exteriors Lemma) tells us that $|C_{i}|<d^{+}(v_{0})-i$.
+By the regularity of exterior degrees, we know they all nodes in $C_{i}$ map to $C_{i+1}$.
+With the absence of back arcs, the only way for nodes in $C_{i}$ to meet the minimum degree requirement is to add more interior ones.
+Each node would need at least $i$ interior arcs to satisfy this inequality.
 Example 9.1.
@@
 The purpose of the interior arcs is to ensure that the parent node $x$’s degree does not double.
-We have already shown through 8.3 (Interior Degrees Lower Bound) that every node in Container $i$ has an interior degree at least $i$.
+We have already shown through 9.4 (Interior Degrees Lower Bound) that every node in Container $i$ has an interior degree at least $i$.
 The two arcs from a parent node and the interior nodes are what create these transitive triangles.
@@
 Lemma 10.2.
-Seymour Diamonds from Containers
-Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP). For any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
+Seymour Diamonds from Containers Suppose we have a graph $G$ represented by a GLOVER, where all nodes satisfy the Decreasing Sequence Property (DSP).
+For any three consecutive containers $C_{i}$, $C_{i+1}$, and $C_{i+2}$, let:
 $u\in C_{i},\quad v,w\in C_{i+1},\quad y\in C_{i+2}$
@@
 Lemma 10.3.
-Seymour Diamonds and Decreasing Path Algorithm Suppose $G$ is an oriented graph represented by the GLOVER, where all nodes have the Decreasing Sequence Property (DSP).
-We are attempting to prevent a degree-doubling node, where we are applying the DSP algorithm.
+Seymour Diamonds and Decreasing Path Algorithm Suppose $G$ is an oriented graph, where all nodes have the Decreasing Sequence Property (DSP).
+We are applying the DSP algorithm.
 Suppose $G$ contains a subgraph $D$ that forms a Seymour diamond.
@@
 Consequently, these additional connections offer alternative routes without disrupting the algorithm’s operation.
-None of these cases are a deterrent to the execution of the DSP algorithm.
+None of these cases are a deterrent to the execution of the DSP algorithm. ∎
 Lemma 10.3 (Seymour Diamonds and Decreasing Path Algorithm) shows how we have been proceeding through the GLOVER data structure through our Seymour diamonds.
@@
 As a result, we simply select an interior node within the container and continue.
-This is what we saw in our examples in section 4, where Seymour diamonds were also present, for example 4.3.
+This is what we saw in our examples in section 4, where Seymour diamonds were also present, for example, 4.3.
 This is also what we saw in our social media example 6.3 in section 5.
-This is also what we saw in our regular graph example in 8.1 in section 8.
+This is also what we saw in our regular graph example in 9.1 in section 8.
 Lemma 10.4.
-Interaction of Minimum Degree Nodes and Transitive Triangles
-Suppose $G$ is an oriented graph represented by a GLOVER.
-Let $v_{0}$ be a minimum degree node with degree $\delta$.
-Consider a transitive triangle formed by nodes $u$, $y$, and $x$, where $d^{+}(y)=1$ and $d^{+}(u)=0$.
-Then the following cases are possible:
-  • $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$
-  • $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$
-  • $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$
-  • $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$:
+Node in GLOVER’s Transitive Triangles Suppose $G$ is an oriented graph.
+Let $v_{0}$ be a minimum degree node and let $u$ be a node in container $C_{i}$.
+Then $u$ is in three possible locations of two different transitive triangles, which do not hurt the algorithm.
 Proof.
-  • Case 1: $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$: Both nodes $y$ and $u$ out-degree $\delta$.
-The triangle has the arc $(y,u)$.
-Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$.
-This means that $u$ has $\delta$ out-neighbors that are not in the triangle.
-Since we assumed that $d^{+}(y)=\delta$ and $d^{+}(u)=\delta$, we know that $y$ cannot connect to all not $u$’s neighbors since $y$ has a neighbor in the triangle ($u$).
-This node that is $u$’s neighbor, but not $y$’s neighbor, is a candidate to extend the path.
-  • Case 2: $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree $\delta$.
-The node $u$ has out-degree greater than $\delta$.
-Again, the triangle has the arc $(y,u)$.
-Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$.
-This means that $u$ has more than $\delta$ out-neighbors that are not in the triangle.
-We assumed that $d^{+}(y)=\delta$ and $d^{+}(u)>\delta$.
-We know that $y$ cannot connect to all not $u$’s neighbors since $d^{+}(u)>d^{+}(y)$.
-We can then select one of these nodes that is a neighbor of $u$, but not a neighbor of $y$ to be a candidate to extend the path.
-  • Case 3: If $d^{+}(y)>\delta$ and $d^{+}(u)=\delta$: The node $y$ has out-degree greater than $\delta$.
-The node $u$ has out-degree $\delta$.
-The triangle has the arc $(y,u)$.
-Since the graph is oriented, the arc $(u,y)$ cannot belong to $G$.
-The node $y$ has more than $\delta$ out-neighbors that are not in the triangle.
-The node $u$ has exactly $\delta$ out-neighbors that are not in the triangle.
-The maximum number of neighbors that $y$ and $u$ can share is $\delta$, that being $u$’s out-degree.
-Then we can select one of the nodes that is a neighbor of $y$, but not a neighbor of $u$ to be a candidate to extend the path.
-  • Case 4: If $d^{+}(y)>\delta$ and $d^{+}(u)>\delta$: The node $y$ has out-degree greater than $\delta$.
-The node $u$ has out-degree greater than $\delta$.
-We have a few cases because we have to remember that in this data structure the purpose of our containers and the interior nodes was to prevent parent nodes from having their degrees double.
-The case where both nodes in a transitive triangle have more than $\delta$ out-neighbors means that $x$ has up to $2*\delta$ out-neighbors, which would cause its degree to double.
-Thus, to prevent $x$’s degree from doubling, there must be some overlap between the out-neighbors of $y$ and the out-neighbors of $u$.
-These two sets cannot be equal to one another, though.
-The most overlap we can have is $N^{+}(u)\subsetneq N^{+}(y)$ or $N^{+}(y)\subsetneq N^{+}(u)$.
-Both of these situations leave room for a node in the larger set that is not in the smaller set as a candidate to extend the path.
-As we allow for less overlap, the number of candidate nodes to extend the path grows.
-We can see that there will always be a node in the larger set to extend the path.
+Suppose we have an oriented graph $G$ with a node $u$ in container $C_{i}$.
+Scenario 1 is that $u$ is a parent to children $v$ and $w$ who are both in $C_{i+1}$, where $v,w)\in G$. Then $(u,v)$ and $(u,w)$ are exterior arcs, and $(v,w)$ is an interior arc.
+Scenario 2 is that $u,v$ are children of $x$ and $u$ is a child of $v$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(v,u)$ is an interior arc.
+Scenario 3 is that $u,v$ are children of $x$ and $v$ is a child of $u$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(u,v)$ is an interior arc.
+In every scenario, we have shown in preceding sections how to proceed through these situations. So these do not hinder our algorithm. ∎
 Now that we have a better understanding of transitive triangles and Seymour diamonds and how they interact with our data structure.
@@
 Here, we will sum up all the lemmas we have derived from our analysis of oriented graphs. We will state the final algorithm and some results.
-So far, this battle has been a battle between a catalyst, the chosen minimum degree node, and the bottleneck, the Decreasing Sequence Property (DSP).
+So far, this battle has been a battle between a catalyst, $v_{0}$, and the bottleneck, the Decreasing Sequence Property (DSP).
 We have proved some important things: Back arcs cannot exist; interior degrees must double inside each container, and the exterior degrees decrease in size; interior degrees proportionally increase in size.
@@
 Lemma 11.1.
-Let $G$ be an oriented graph represented by a GLOVER, with all nodes satisfying the Decreasing Sequence Property (DSP).
+Let $G$ be an oriented graph, with all nodes satisfying the Decreasing Sequence Property (DSP).
 Let $v_{0}$ be a minimum degree node in $G$.
-We are attempting to prevent a degree-doubling node.
 If there are back arcs in $G$, then a node’s degree will double.
@@
 Theorem 11.1.
-Let $G$ be an oriented graph represented by a GLOVER with all nodes having the Decreasing Sequence Property (DSP).
-Let $v_{0}$ be a minimum degree node where we set up our partition.
-Then Seymour’s conjecture holds true.
+Let $G$ be an oriented graph, with all nodes having the Decreasing Sequence Property (DSP). Let $v_{0}$ be a minimum degree node where we set up our partition. Then Seymour’s conjecture holds true.
 Proof.
@@
 There are no back arcs; else we will have a degree-doubling node.
-Next is lemma 7.2 (Interior Degree Doubling), which states that in every container we will have a node whose interior degree doubles.
+Next is lemma 8.1 (Interior Degrees Doubling), which states that in every container we will have a node whose interior degree doubles.
 These lemmas are what allowed us to set up and use lemma 6.4 (Weak Decreasing Exteriors Lemma).
@@
 14:end for
+Example 11.1.
 Start with graph size $n$Select a minimum degree node $v_{0}$ivide into containers $C_{1},C_{2},\ldots C_{k}$dentify node $x_{i}$ in $C_{i}$ whose interior degree doublesMove to neighboring container $C_{i+1}$f back arc to a prior containerLocate the child of $x_{i}$ doubling $x_{i}$’s degreeHaltElse bound $|C_{i+1}|<|C_{i}|$Repeat from step 4
@@
 Ultimately, the interior neighbors represented the number of interior nodes, while the exterior neighbors represented the size of the containers.
-In small graphs like 4.3 (Minimum Out-Degree 3 with Neighbors 1), we saw that this collision did not occur.
-To understand the reasoning behind this, one must understand that the container sizes are decreasing at the rate of 1, and the interior degrees are increasing at a rate of 1.
-The problem is that the number of arcs inside an oriented graph is growing at a rate of $O(n^{2})$.
-So, for small graphs, we are just not going out far enough to catch up.
-For larger graphs, though, we do start to see this.
+In small graphs, such as the one shown in Figure 4.3 (Minimum Out-Degree 3 with Neighbors 1), the collision phenomenon was not observed.
+This is because the container sizes decrease linearly (at a rate of 1), while the interior degrees increase linearly (also at a rate of 1).
+However, the number of arcs within the oriented graph grows quadratically, at a rate of $²O(n\texttwosuperior)$.
+Consequently, while the number of nodes increases linearly, the number of edges increases quadratically.
+In small graphs, the growth hasn’t progressed far enough for the quadratic edge growth to outpace the linear node growth and cause collisions.
+In larger graphs, however, this difference in growth rates becomes significant, and collisions begin to occur.
 Corollary 11.2.
-Occurrence of Last Dense Container Let $G$ be an oriented graph represented by a GLOVER with minimum degree $k$.
+Occurrence of Last Dense Container Let $G$ be an oriented graph, with minimum degree $k$.
 Then there will be a container that reaches maximal density.
@@
 Lemma 9.4 (Interior Degree Lower Bound) tells us that each node must have degree $i$ in this container.
-Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree none.
+Thus, we have a series of regular containers that are getting smaller as we get further from the minimum degree node.
 At the same time, these containers are expected to hold fewer nodes of higher degree, i.e., a graph of more arcs.
@@
 Corollary 11.3.
-Multiple Degree Doubling Nodes Let $G$ be an oriented graph represented by the GLOVER structure, with minimum degree $\delta$, and let the most dense graph be in container $C_{i}$.
+Multiple Degree Doubling Nodes Let $G$ be an oriented graph, with minimum degree $\delta$, and let the most dense graph be in container $C_{i}$.
 Then every node in container $C_{i}$ has its degree doubled in the representation.
@@
 Ultimately though, they are unable to do that, as we see the collision of the container shrinkage and interior degree requirements.
-Then third we can look at the applications.
+Then third, we can look at the applications.
 We may be looking for the person who has the most influence, but this is saying that there may not be just one person; there may be a whole set of people, a tightly connected group.
@@
 Next, we will look into the complexity of the run-time analysis of the Decreasing Sequence Algorithm.
-First, remember that we proved in lemma 7.2 (Interior Degree Doubles) that every container has a node whose interior degree doubles.
+First, remember that we proved in lemma 8.1 (Interior Degrees Doubles) that every container has a node whose interior degree doubles.
 Secondly, we also note that the lemma 7.1 (Regular Interior Degrees) implies that all the nodes are regular inside a container.
@@
 It was not spoken about in the paper because the focus was entirely set on proving the SSNC, but this is another example of multiple degree-doubling nodes, just a localized instance.
-We opened by showing several applications of the SSNC, including social media, social network modeling and network analysis.
+We opened by showing several applications of the SSNC, including social media, social network modeling, and network analysis.
 This highlights the necessity for an algorithm that can work efficiently in practice.
-This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph is important.
+This proof that the algorithm has a complexity of $O(|V|+|E|)$, where $|V|$ is the number of vertices and $|E|$ is the number of edges in the graph, is important.
 This algorithm’s linear time complexity shows that even enormous graphs may be processed effectively.
@@
 We did so through the lenses of graph theory, combinatorics, algorithms, and data structures.
-This research helped us to use many mathematical tools like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
+This research helped us to use many mathematical tools, like both total orders, strong and weak induction, recurrence relations, and the divide-and-conquer paradigm.
 One of the key insights of this paper was the development of an algorithm that identifies nodes whose degree must double in any oriented graph.
@@
 The SSNC started as a question in 1990.
-Over the next 34 years it has inspired hundreds of mathematicians into graph theory, research, and some into computers.
+Over the next 34 years, it has inspired hundreds of mathematicians into graph theory, research, and some into computers.
 It has been talked about in books.
@@
 Many of the things inherent to SSNC made it easier for GLOVER—the well ordering and the dual metrics of distance and degrees.
+Still, the ability for the data structure to partition the nodes of an oriented graph into containers, which then allowed for powerful tools like divide-and-conquer and mathematical induction, can have further applications.
+This would then allow us to implement resources from the field of graph theory into subject matter dealing with machine learning, data science, and network flow.
 Future research and development should explore questions of complexity and scalability.
@@
 Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions.
-This may encourages researchers to explore constructive approaches with a focus on practical applications.
+This may encourage researchers to explore constructive approaches with a focus on practical applications.
 Hopefully, this work will inspire further research on constructive approaches to other graph theory problems.
@@
 This shows how versatile the GLOVER data structure is moving forward.
-What makes GLOVER special are things like total order on containers.
+What makes GLOVER special are things like a total order on containers.
 Now there is a two-way ranking of nodes, as opposed to traditional one-way rankings like lexicographical sorting.
@@
 Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com.
-This is an evolving project.
-Additional examples and applications will continue to be added.
-Readers are invited to explore the website.
-Please provide feedback on any aspects of the site or content for improvement.
+This is an evolving project. Additional examples and applications will continue to be added.
+Readers are invited to explore the website. Please provide feedback on any aspects of the site or content for improvement.
 Our goal here was to help users connect with the SSNC.

### v9

@@
 This finding extends the SSNC to the discovery of multiple nodes satisfying the degree doubling property.
-Beyond theoretical implications, the algorithm and data structure have practical applications in network optimization and algorithm design.
+Beyond theoretical implications, the algorithm and data structure have practical applications in data science, network optimization and algorithm design.
+Keywords: degree doubling, oriented graph, square, algorithm
 1 Introduction
@@
 For every oriented graph $G=(V,E)$, there exists a vertex $v\in V$ such that $|N^{++}(v)|\geq|N^{+}(v)|$, where $N^{+}$ and $N^{++}(v)$ denote the first and second neighborhoods of $v$, respectively.
-Despite decades of research, the SSNC remains one of graph theory’s challenging open problems.
+Despite decades of research, the SSNC has remained one of graph theory’s challenging open problems.
 Special cases of the conjecture have been explored extensively.
 Kaneko and Locke [11] showed that the SSNC holds for graphs with a minimum degree of at most six.
-Brantner et al. [1] demonstrated that if the conjecture were false, infinitely many counterexamples would exist.
+Brantner et al.
+[2] demonstrated that if the conjecture were false, infinitely many counterexamples would exist.
 Chen et al.
-[3] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
+[4] established a lower bound, proving that every oriented graph contains a vertex whose second neighborhood is at least $\lambda$ times larger than its first neighborhood, where $\lambda$ is the real root of the polynomial $12x^{3}+x^{2}-$ 1.
 In recent work, Diaz [6] showed that almost all orientations of random graphs satisfy the conjecture.
-This was an extension of the work done by Botler et al. [2], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
-The concept of $m$-free graphs is particularly relevant to this paper, as it closely aligns with the work of Daamouch [7].
+This was an extension of the work done by Botler et al.
+[3], where their work confirms that almost all orientations of $G(n,p)$ satisfy the SSNC.
+The concept of $m$-free graphs is particularly relevant to this paper, as our work closely aligns with the work of Daamouch [1].
 A directed graph $D$ is $m$-free if it contains no directed cycles of length at most $m$.
-In his paper, Daamouch introduces induced subgraphs and explores their role in the degree properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
+In his paper, Daamouch works with the concept of induced subgraphs and explores their role in the degree properties of vertices—an idea that will also play a significant role in the discussions and results presented here.
 This paper builds on these foundations, generalizing an approach initially developed for the minimum degree 7 case [9].
@@
 Our approach not only proves the conjecture but also identifies the specific node(s) whose out-degree doubles in the graph’s square.
-Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, and algorithm design.
-For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections surpasses their direct contacts, offering insights into the spread of infections.
+Beyond its theoretical significance, the SSNC has practical applications in domains like social networks, epidemiology, algorithm design, and the development of data-driven methods for networked systems, including A/B testing.
+For instance, in the context of disease control, the conjecture helps identify individuals whose influence in their second-degree connections matches or surpasses their direct contacts, offering insights into the spread of infections.
+In the development of algorithms for A/B testing within networked environments, understanding influence propagation is critical for designing methods that account for dependencies between individuals, ultimately leading to more accurate assessments of interventions.
 Unlike traditional proof-theoretic approaches, this work emphasizes the development of an efficient algorithm to locate the target node.
@@
 Let $G^{2}=(V,E^{2})$ where $G=(V,E)$ is the original graph, and $E^{2}$ is the set of edges defined as:
-$E^{2}={(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E}$
+$E^{2}=\{(u,v)|(u,v)\in E\text{ or }\exists w\in V\text{ such that }(u,w)\in E\text{ and }(w,v)\in E\}$
 We use the notation $N^{+}(v)$ to refer to the out-neighbors of a vertex $v$.
@@
 Similarly, $N^{++}(v)$ refers to the second neighbors of $v$, that is, nodes that are two steps away from $v$ in $G$.
+Definition 2.2.
+Let $G=(V,E)$, the first neighborhood of a vertex $v\in V$ is denoted
+$N^{+}(v)=\{w\in V(G)|(v,w)\in E)\}$
+Definition 2.3.
+Let $G=(V,E)$, the second neighborhood of a vertex $v\in V$ is denoted
+$N^{++}(v)=\{u\in V(G)|(v,w)\text{ and }(w,u)\in E)\}$
 Let $G=(V,E)$ be a graph and let $S\subseteq V$ be any subset of vertices.
@@
 (Square version of Seymour’s Second Neighborhood Conjecture).
-Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when you square the oriented graph.
+Prove that for every oriented graph, $G$, there exists a vertex whose out-degree at least doubles when the oriented graph is squared.
 Now that we have two different conjectures, we can investigate how they relate to one another.
@@
 This is why these two conjectures are seen as identical.
-We will focus our attention on Conjecture 1.1 and Conjecture 2.1.
+We will focus our attention on both conjectures 1.1 and 2.1.
 Conjecture 1.1 is easier to work with initially and understand why a node’s degree will double in the square graph.
 Concentrating on neighborhood analysis yielded numerous insights and features that aided in the development of our algorithm.
+This is because it is not necessary to work with a different graph to understand degree-doubling properties.
+However, conjecture 2.1 could have just as easily been used and in some of the later proofs we will see this.
 Next, we turn our attention to the term degree-doubling nodes.
@@
 We know that there is an out-degree distribution.
-Consequently, every oriented graph will have a minimum out-degree $d$, representing the lowest out-degree among all nodes.
+Consequently, every oriented graph will have a minimum out-degree $\delta$, representing the lowest out-degree among all nodes.
 By selecting a node $v_{0}$ with this minimum out-degree, we establish a stable starting point for our investigation.
-From this foundation, we can now partition the graph into $d_{v_{0}}+1$ subgraphs corresponding to $d_{v_{0}}+1=d^{+}(v_{0})$.
+From this foundation, we can now partition the graph into $\delta+1$ subgraphs corresponding to $\delta+1=d^{+}(v_{0})+1$.
 This partitioning enables us to leverage the divide-and-conquer paradigm.
-No nodes or arcs are being removed to form these partitions.
+No nodes or arcs are being removed from the graph to form these partitions.
 This is a conceptual idea.
@@
 That presents a problem because oriented graphs can be varying in nature.
-This is the first global search technique on the SSNC problem.
+This may be the first global search technique on the SSNC problem.
 Complexity enters the picture when discussing global search algorithms.
@@
 This means that $(v_{1},v_{2})\in G$.
-Hence $v_{2}$ is a second neighbor of $v_{0}$, thus making it at least as large as $v_{0}$’s first neighborhood.
+Hence $v_{2}$ is a second neighbor of $v_{0}$, thus making $v_{0}$’s second neighborhood at least as large as its first neighborhood.
 Case 3: Assume that $d^{+}(v_{0})=2$.
@@
 Let these nodes be $v_{3}$ and $v_{4}$.
-The node $v_{0}$ cannot be one of these nodes.
+Because this is an oriented graph, the node $v_{0}$ cannot be one of these nodes.
 Hence $v_{0}$ will have at least two second neighbors—$v_{3}$ and $v_{4}$, which makes its second neighborhood at least as large as its first.
@@
 We are learning that, to construct our counterexample, we need the minimum degree of our graph to be at least 3.
-In each of the cases of the lemma, the minimum degree node has few very out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
+In each of the cases of the lemma, the minimum degree node has very few out-neighbors, which couldn’t stop the increase that was coming from the influx of second neighbors.
 Lemma 4.2.
@@
 Now, assume that $v_{1}$ has an out-degree of 0 within $G[v_{1},v_{2},v_{3}]$. This implies that $v_{1}$ must connect to 3 nodes outside of $\{v_{1},v_{2},v_{3}\}$.
-We know that $v_{0}$ is connected to $v_{1}$, and $v_{1}$ has at least 3 out-neighbors outside of $\{v_{1},v_{2},v_{3}\}$.
-This means that $v_{0}$ has at least three second neighbors.
-As a result, $v_{0}$ has at least as many second neighbors as first neighbors, and equivalently, $v_{0}$’s out-degree doubles in $G^{2}$.
-∎
+This means that $v_{0}$ has at least three second neighbors. As a result, $v_{0}$ has at least as many second neighbors as first neighbors, and equivalently, $v_{0}$’s out-degree doubles in $G^{2}$. ∎
 Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) considers a scenario where $v_{0}$ has exactly 3 out-neighbors, and one of those out-neighbors, $v_{1}$, shares no neighbors with $v_{0}$.
-This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors, and $v_{1}$ must have at least 3 first neighbors.
+This means that all of $v_{1}$’s first neighbors must be $v_{0}$’s second neighbors.
+By assumption, $v_{1}$ must have at least as many first neighbors as $v_{0}$.
 Example 4.2.
@@
 $\Rightarrow$ 111122223333 $\Rightarrow$ 45678 $\Rightarrow$
-Figure 3: This is a representation of the node $v_{0}$ in $G$, $v_{0}$ being a minimum degree node, connecting to three nodes, $v_{1}$, $v_{2}$, and $v_{3}$.
-Each of these nodes ($v_{1}$, $v_{2}$, and $v_{3}$) connects to the nodes $v_{4}$ and $v_{5}$.
-The neighbor of $v_{4}$ is $v_{5}$, and the neighbors of $v_{5}$ are $v_{6}$, $v_{7}$, and $v_{8}$.
-While $v_{0}$ has degree 3, the neighbors of $v_{0}$ all have degree 1 in the subgraph $G[N^{+}(v_{0})]$, and they each have degree 2 outside $G[N^{+}(v_{0})]$.
-Thus, in $G^{2}$, the degree of $v_{1}$ will double.
+Figure 3: Illustration of Lemma 3: Minimum degree node v0 (degree 3) and its neighbors (degree 1 in G[N+(v0)], degree 2 elsewhere).
+In G2, v1’s degree doubles.
 See lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1)
@@
 In this case, the degree of the neighbor of $v_{0}$, $v_{1}$, doubles.
-We did not require the ”right” starting node.
+We did not require the right starting node.
 Therefore, this is a crucial distinction.
@@
 5 Building a Data Structure
-In this section, we develop a new data structure for oriented graphs (Graph Level Order).
+5.1 Definitions
+In this section, we develop a new data structure for oriented graphs Graph Level Order (GLOVER).
 This structure is specifically designed for the SSNC and oriented graphs but will have applications beyond.
@@
 […]
-Figure 4: Illustration of containers, each enclosed in a checkered box. Containers group nodes based on their distance from $v_{0}$, with no assumptions on container relationships or sizes.
+Figure 4: Illustration of containers, each enclosed in a checkered box.
+Containers group nodes based on their distance from the minimum degree node $v_{0}$, with no assumptions on container relationships or sizes.
 Definition 5.1.
@@
 The formula
-$C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and }dist(v_{0},u)=i,dist(v_{0},v)=i\}$, where
-$V(G)$ denotes the graph $G$’s vertex set, and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
+$C_{i}=\{u\in V(G)\text{.s.t. }A(C_{i})=\{(u,v)|u,v\in C_{i}\}\text{ and dist}(v_{0},u)=\text{dist}(v_{0},v)=i\}$
+where $V(G)$ denotes the graph $G$’s vertex set, and $dist(v_{0},u)$ is the shortest path between $v_{0}$ and $u$.
 Similar to this, Daamouch employed induced subgraphs in his research on $m$-free graphs [1].
 The neighboring containers (see example 5.1), which will be crucial in the proofs, are the main distinction here.
-We will investigate the structure of oriented graphs from two angles—from the viewpoint of containers or $v_{0}$.
-Containers are analyzed based on their distance from $v_{0}$.
+We will investigate the structure of oriented graphs from two angles—from the viewpoint of containers and the minimum degree node $v_{0}$.
+Containers will be analyzed based on their distance from $v_{0}$.
 Nodes are seen based on their position within these containers.
@@
 Interior neighbors will be important to our study on SSNC and our search for a counterexample.
-In particular, we saw things about interior neighbors in 4.3 (Minimum Out-Degree 3 with Neighbors 1).
+In particular, we saw things about interior neighbors in lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1).
 The notion of containers had not been introduced yet, but the induced subgraph of the neighbors of $v_{0}$ can be called container 1, and we see that when all the neighbors in that subgraph had an out-degree of 1, $v_{0}$’s degree did not double.
@@
 Definition 5.4.
-Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the component $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
+Let $v_{0}$ be a minimum degree node. Suppose that $x$ is a node in the container $C_{i}$. A back arc is defined as an arc $(x,y)$ such that $y\in N^{+}(x)$ and $y\in C_{j}$, where $j<i$.
 We have spoken of so much of our algorithm depending on the distance property of this oriented graph.
@@
 Additionally, these may force complications in Seymour’s Second Neighborhood Conjecture, where they may force us to double count nodes.
-We will speak more about back arcs in the next section.
+We will speak more about back arcs in section 8 (Back Arcs).
+5.2 Graph Level Order
 Definition 5.5.
@@
 Leveled Structure: The elements of $S$ are partitioned into levels of containers $C_{1},C_{2},\ldots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
+Total Order on Elements of G: We need that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
 Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion.
@@
 Interior and Exterior Neighbors:
-  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
-  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
-Next we will prove that any oriented graph can be represented in a GLOVER.
+  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors if $(x,y)\in G$.
+  • If $x\in C_{i}$ and $y\in C_{i+1}$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
+Next, we will prove that any oriented graph can be represented in a GLOVER.
 We will need to demonstrate how the graph can be systematically partitioned into containers while preserving the GLOVER properties.
@@
 Knowing the minimum degree node of this graph, we can partition the nodes of this graph into containers, $\{C_{1},C_{2},\dots,C_{k}\}$.
-Each of these containers $C_{i}=\{u|d(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$.
-These containers can then be sorted based on the distance $i$ into a total order for the GLOVER requirement.
-  • Total Order on Elements within $G$ Nodes are partitioned based on their distance from $v_{0}$.
-This is a total order on the nodes of $G$.
-For any two nodes $u$ and $v$, this total order is determined by their containers.
+Each of these containers $C_{i}=\{u|\text{dist}(v_{0},u)=i\}$ is a set of nodes at distance $i$ from $v_{0}$.
   • Leveled Structure with Containers—By construction, the containers $C_{1},C_{2},\dots,C_{k}$ each correspond to nodes at a specific distance from $v_{0}$.
-  • Comparability within Containers—For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric to distinguish $u$ from $v$.
-  • Consistency of Comparisons Across Containers - Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
+  • Total Order on Elements within $G$—Nodes are partitioned based on their distance from $v_{0}$. We know that distance is a total order on the nodes of $G$.
+    To prove the containers are a total order, we need to show that any two nodes in $G$ can be compared by their container.
+Let $u,v$ be two nodes in $G$ such that $u\neq v$.
+We assume that both $u$ and $v$ are reachable from $v_{0}$.
+    First, we consider the situation where $u$ is in a closer container than $v$. This implies that $\text{dist}(v_{0},u)<\text{dist}(v_{0},v)$.
+    Next, we will look at the situation where $u$ and $v$ are in the same container. This implies that $\text{dist}(v_{0},u)=\text{dist}(v_{0},v)$.
+    Finally, we consider the situation where $u$ is in a further container from $v_{0}$ than $v$. This implies that $\text{dist}(v_{0},u)>\text{dist}(v_{0},v)$
+    In each situation we were able to translate the container relationship to a distance relationship.
+We know that distance is already a total order.
+This shows that the container relationship is also a total order.
+  • Comparability within Containers—For any two nodes $u$ and $v$ in the same container $C_{i}$, we can select a second metric (degree) to distinguish $u$ from $v$ or compare $u$ to $v$.
+  • Consistency of Comparisons across Containers—Since the graph is oriented and there are no back arcs, the ordering between containers is respected.
   • Interior and Exterior Neighbors:
-    • Nodes $u$ and $v$ of the same distance from $v_{0}$ are in the same container and thus are interior neighbors.
-    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1.
+    • Nodes $u$ and $v$ of the same distance from $v_{0}$ are in the same container and thus are interior neighbors if there is an arc $(u,v)\in G$
+    • If the nodes $u$ and $v$ were of different distances from $v_{0}$, then they would be in different containers and would be exterior neighbors if the distance differs by 1 and if there is an arc $(u,v)\in G$
     Each criterion for a GLOVER is satisfied by partitioning nodes based on distance from $v_{0}$.
 Thus, the partitioned set of containers $C_{1},C_{2},\dots,C_{k}$ forms a GLOVER under the distance total order, completing the proof.
-What the GLOVER data structure does is transform the containers and distance metric into an ordering for containers.
+What the GLOVER data structure does is transform the oriented graph into an ordered of containers with metrics.
 This ordering is more than just representative, though.
@@
 However, in order for this data structure and algorithm to work, nodes need to be placed in proper alignment with their distance from $v_{0}$.
+Now, each node has a proper place inside the container, up to reordering.
 The one thing that could disrupt this total ordering is a back arc.
@@
 However, it is a bold assumption to claim that oriented graphs have no back arcs.
-Oriented graphs come in various types; surely some exist with back arcs.
-This is an important claim to investigate, and the next section will look into it in depth.
-We will prove that these back arcs do not hurt the algorithm.
+Oriented graphs come in all shapes and sizes; surely some exist with back arcs.
+This is an important claim to investigate, and section 8 (Back Arcs) will look into it in depth.
+We will prove that when these back arcs do exist, we are not able to construct a GLOVER data structure.
+Instead, we are able to immediately find a degree-doubling node.
+That is, a back arc immediately leads us to a node whose second neighborhood is at least as large as its first neighborhood.
+This means that back arcs do not hurt with the development of our algorithm.
 Instead, they only lead to the faster path to a degree-doubling node.
-In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node.
-Next, we will show one of the other benefits of this data structure in mathematical induction via well ordering.
+In addition, there will no longer be a need to represent the oriented graph in a GLOVER, as we will have already found the degree-doubling node(s).
+Next, we will show that one of the other benefits of this data structure is mathematical induction via well ordering.
 Mathematical induction is a proof technique that uses inference to prove statements.
@@
 Proof.
-We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order. Because of this, they are a well ordering.
+We have already shown that the containers in GLOVER form a total order. Further, they are a discrete total order.
 A set is considered well-ordered if every non-empty subset of that set has a least element.
-A discrete total order is a total order where each element has a distinct ”next” element in the order.
-Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a ”smallest” element according to the order.
+A discrete total order is a total order where each element has a distinct next element in the order.
+Since a discrete total order has a clear distinction between elements, any non-empty subset will always have a smallest element according to the order.
 This fulfills the condition of being well-ordered.
@@
 These containers are totally ordered and thus can also be used as a means of mathematical induction.
-We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from $v_{0}$.
+We now have several options of how to proceed with our proofs, but the foundation will be the Graph Level Order (GLOVER) data structure, which orders the nodes based on their distance from $v_{0}$-the minimum degree node.
 One of the additional benefits of partitioning the oriented graph into containers with interior arcs is that each container resembles its own local SSNC.
@@
 6 Using the Data Structure
-Having defined the GLOVER data structure, we are ready to begin using it.
+6.1 Decreasing Sequence Property
+Having laid the ground work for the GLOVER data structure, we are ready to begin using it.
 The first and most important concept that we will define that will guide our research is the decreasing sequence property.
@@
 Exterior degrees, though, are a part of a node’s degree equation.
-There’s another part of it, though.
+Still, there’s another part of it.
 As we move further from $v_{0}$, if there are no back arcs, the size of the interior degrees must increase.
@@
 Two competing new ideas have been presented: the decreasing sequence property (DSP) and the minimal degree node.
-A crucial starting point for our approach and degree values is a minimal degree node.
-It acts as the anchor by being chosen as the beginning point.
-The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by the minimal degree node.
-It acts as a catalyst, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
+A crucial starting point for our approach and degree values is the minimal degree node $v_{0}$.
+It acts as the anchor by being chosen as the starting point.
+The remaining nodes are currently being forced into the attributes that will surface in the counterexample construction by $v_{0}$.
+It acts as a driving force, influencing other nodes and setting up the conditions necessary for the subsequent analysis.
 In contrast, the DSP introduces constraints on connectivity.
@@
 A bottleneck in graph theory represents a point that limits flow or connectivity within the network [5], a role precisely filled by the DSP.
-Thus, the minimum degree node and the DSP together function as a catalyst and bottleneck, respectively.
+Thus, the minimum degree node $v_{0}$ and the DSP property together function as a catalyst and bottleneck, respectively.
 Lemma 6.1.
@@
 In the proofs of Lemma 4.2 (Minimum Out-Degree 3 with Neighbor Out-Degree 0) and Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1), we looked at the out-degree of a node inside the neighbor-induced subgraph (of $v_{0}$) and in the graph with the neighbor-induced subgraph removed.
-We will now give terms to these concepts.
+The neighbors of $y$ that are also neighbors of $x$ are called interior neighbors, while the neighbors of $y$ that are not neighbors of $x$ are called exterior neighbors.
+One thing we see in Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1) is that the interior neighbors of $v_{0}$ helped $v_{0}$’s degree not to double.
+They did this by all having a degree of 1.
 In this work, we assume that the input graph is oriented and unweighted.
@@
 Interior Degree at least 1 Suppose we have an oriented graph $G$.
-Suppose further that $v_{0}$ is a minimum out-degree node for $G$ satisfying the decreasing sequence property.
+Suppose further that $v_{0}$ is a minimum out-degree node for $G$.
 In an attempt to prevent $v_{0}$’s degree from doubling in $G$, every node $x$ in the neighboring container $C_{1}$ must have an interior degree of at least 1.
@@
 Thus, for all $x\in C_{1}$, $|int(v_{0},x)|\geq 1$. ∎
-Lemma 6.2 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree doubling.
-We see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree doubled.
+Lemma 6.2 (Interior Degree at least 1) is important because we can see some structural properties in generally oriented graphs that may coincide with degree-doubling.
+We see that when all of the neighbors of the minimum degree node have degree 1, then it will prevent that minimum degree node from having its degree-doubled.
+This is a generalization of Lemma 4.3 (Minimum Out-Degree 3 with Neighbors 1).
 Lemma 6.3.
-Exterior Less than Degree
-Suppose we have an oriented graph $G$. If $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
+Exterior Less than Degree Suppose we have an oriented graph $G$. If $x_{i}$ in container $C_{i}$ has the decreasing sequence property, then for all $y\in N^{+}(x_{i})|\text{ext}(x,y)|<d^{+}(x)$.
 Proof.
@@
 This must be true for all $x$ in $G$, so we conclude that for all $y\in N^{+}(x)$, the number of exterior connections $|\text{ext}(x,y)|$ must be strictly less than $d^{+}(x)$. ∎
-Lemma 6.3 (Exterior Less than Degree) is very similar to (Exterior Less than Degree), Lemma 6.2 (Interior Degree at least 1), with the notable exception that Lemma 6.3 refers to any node with the DSP, whereas we only made Lemma 6.2 true for the minimum degree node.
+Lemma 6.3 (Exterior Less than Degree) is very similar to lemma 6.2 (Interior Degree at least 1), with the notable exception that Lemma 6.3 refers to any node with the DSP, whereas we only made Lemma 6.2 true for the minimum degree node.
 The results presented in 6.2 (Interior Degree At Least 1) and 6.3 (Exterior Less than Degree) are closely related to Daamouch’s [1] Lemma 1.1.
@@
 If so, then it gives light to this divide-and-conquer approach, where we partition the containers, solve them individually, and then combine them again for this lemma 6.4 (Weak Decreasing Exteriors Lemma).
+6.2 Representation
 The more popular representations of graphs are adjacency matrices, array-lists, and edge lists.
@@
 Example 6.2.
-[…]
 Figure 6: Illustration of the Weak Decreasing Exteriors Lemma.
@@
 Example 6.3.
-Andi
-$\Rightarrow$ BellChadDawnElseFayeGusHugoIslaJade $\Rightarrow$ KaiLeiMiaNovaOwenPatQinRyan $\Rightarrow$ SageTateUgoVeraWadeXiaYoel $\Rightarrow$
-Figure 7: This is an example of a social media setup of the SSNC. The second container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
+Figure 7: This is an example of a social media setup of the SSNC. The third container has 21 arcs and seven nodes, which represents an orientation of the complete graph.
 We have now defined the data structure that we will use throughout the rest of this paper.
 In doing so, we have a better understanding of interior and exterior arcs, which will help guide us in our search for a counterexample.
+We also have an understanding of the Decreasing Sequence Property which will guide the nodes throughout the rest of this proof process.
 7 Back Arcs
-With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs. We will investigate whether these arcs can bring about a counterexample.
+With our discussion of the data structure concluded, we are ready to launch an investigation into back arcs.
+We will investigate the influence these back arcs have on our data structure, and our search for a counterexample.
 It is important to note that the back arcs and the exterior neighbors are distinct concepts.
@@
 For example, both containers 1 and 2 in example 6.3 are orientations of regular graphs.
-The second container is actually an orientation of $K_{5}$, the complete graph on 5 nodes.
+The second container is actually an orientation of $K_{7}$, the complete graph on 7 nodes.
 This result begins to lend itself to some of the mathematical symmetry and structure that must be present in these containers and the SSNC.
@@
 The shortest-path lengths from $v_{0}$ strictly define each level in the hierarchical container organization, which contrasts with this.
-The assumptions for this letter are that we have one parent node whose interior degree doubles, and it maps the two child nodes that have different interior degrees.
-What happens is that because these child nodes have different interior degrees, it causes this pair node to have an exterior degree that doubles by the week exterior decreasing lemma.
+The assumptions for this lemma are that we have one parent node whose interior degree doubles, and it maps the two child nodes that have different interior degrees.
+What happens is that because these child nodes have different interior degrees, it causes this parent node to have an exterior degree that doubles by lemma 6.4 (Weak Decreasing Exteriors Lemma).
 We spoke previously about how the divide and conquer approach is allowed by the GLOVER data structure.
@@
 $|ext(u_{i},v_{k})|\leq|N^{+}(v_{k})|=deg(v_{k}).$
+$v_{k}$’s neighbors are distinct from $u_{i}$’s because it is in a different container.
+We formed these containers by a partition, so these containers cannot overlap.
 This means that the back arc will double $u_{i}$’s exterior degree.
@@
 That the node $u_{i}$ has its interior degree doubled and that the first back arc from $u_{i}$ is to the node $v_{k}$.
-As we saw in the proof, $v_{k}$ has more exterior neighbors than you survive, which is what causes $u_{i}$’s degree to double.
+As we saw in the proof, $v_{k}$ has more exterior neighbors than $u_{i}$, which is what causes $u_{i}$’s degree to double.
 The question is whether we need the condition of $u_{i}$ interior degree doubling.
-The weak decreasing exterior lemma tells us that we have a relationship between the exterior degrees of use by and $v_{k}$, not the overall degrees of these nodes.
+Lemma 6.4 (Weak Decreasing Exteriors Lemma) tells us that we have a relationship between the exterior degrees of $u_{i}$ and $v_{k}$, not the overall degrees of these nodes.
 By assigning the interior degree to double, we can leverage this lemma.
@@
 Lemma 7.3.
-Back Arcs Special Case
-Let $G$ be an oriented graph, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$.
-  1. 1.
-    Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP).
-  2. 2.
-    The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$.
-  3. 3.
-    There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
-Then, for all $z\in\text{ext}(x,y)$, we have:
-$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
-First, we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e., from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e., not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$.
-This gives the only possibility for $z$ as being an interior neighbor of $x$.
-Consequently, the arc $w,z$ is an early back arc as we saw in 7.2 (No Back Arc to Early Container).
-These arcs can lead to double counting.
-An out-neighbor of $w$ might be an out-neighbor of $x$ and a second out-neighbor of $y$.
-This implies that under the assumption that $w$’s interior degree doubles, $w$’s overall degree will double through that early back arc.
-Lemma 7.4.
-First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $d$.
+First Back Arc Between Consecutive Containers in an Oriented Graph: Let $G$ be an oriented graph with minimum degree node $v_{0}$, with minimum degree $\delta$.
 Suppose that all arcs point forward (i.e., from container $C_{j}$ to container $C_{j+1}$ for $j<i$) with no back arcs between containers $C_{1},C_{2},\dots,C_{i-1}$.
-If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then it will introduce an arc violating $G$’s oriented property.
+If the first back arc occurs from container $C_{i}$ to $C_{i-1}$, then the first back arc will introduce an arc violating $G$’s oriented property.
 Proof.
@@
 Therefore, the existence of a back arc from $C_{i}$ to $C_{i-1}$ contradicts the oriented nature of $G$, and no back arcs can exist prior to $C_{i}$. ∎
+Example 7.2.
+Back Arcs Special Case
+Let $G$ be an oriented graph, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$.
+  • Suppose $x\in C_{i}$ and $y\in C_{i+1}$.
+  • $w$ is a second out-neighbor of $x$ and a first out-neighbor of $y$.
+  • $z$ is an out-neighbor of $w$, and an out-neighbor of $x$ and an out-neighbor of $y$.
+  • This implies that $z$ is in container $C_{i+1}$, the same container as $y$.
+  • By 7.1 (Regular Interior Degrees), we know that because the nodes in container $C_{i+1}$ cannot have back arcs, they can only have interior and exterior arcs.
+  • This means that by 7.3 (No Back Arc Between Consecutive Containers) that every node (including $y$ and $z$ in $C_{i+1}$ will map to the same set of exterior nodes in container $C_{i+2}$.
+  • Thus, $z$ cannot be a back arc of $w$ and an exterior arc of $y$.
+Example 7.3.
+[…]
+The concern with 6.4 (Weak Decreasing Exteriors Lemma) concern was about the potential for overlapping neighborhoods when considering how interior and exterior degrees change.
+If a node $w$ is both an interior neighbor (or contributes to the interior degree) and an exterior neighbor (or contributes to the exterior degree), then we couldn’t simply add the changes in interior and exterior degrees to get the change in the total degree.
+We’d be counting $w$’s contribution twice.
+The conditions regarding back arcs and the lemma 7.1 (Regular Interior Degrees) are crucial for the validity of Lemmas 7.2 (No Back Arc to Early Container) and 7.3 (No Back Arc Between Consecutive Containers).
+Lemma 7.3, which is most directly relevant to the cases considered here, relies on the property of oriented graphs that if every node in container $C_{i-1}$ has an arc going to every node in container $C_{i}$, then no back arc can exist from a node in $C_{i}$ to a node in $C_{i-1}$.
+This property is a direct consequence of the definition of oriented graphs and the container structure.
+The Regular Interior Degrees Lemma (Lemma 7.1) plays a crucial role in establishing this property.
+It shows that if there were an imbalance in interior degrees between nodes in $C_{i}$, it would force the exterior degree of a parent node in $C_{i-1}$ to double.
+Combined with the Weakly Decreasing Exteriors Lemma (Lemma 6.4), which states that exterior degrees cannot increase along a path under certain conditions, this implies that interior degrees within a container must be regular (essentially the same), and thus the oriented graph property holds.
+The assumptions in Lemmas 7.2 and 7.3—that we are considering the first back arc—reinforce that no such back arc exists at any point in the process of constructing the containers.
+Therefore, the scenarios considered in these lemmas (and any example based on a back arc) are consistent with the graph structure and the regularity properties established by Lemmas 7.1 and 6.4.
 Lemma 7.3 (No Back Arc Between Consecutive Containers) is a special case where we can apply Lemma 7.1 (Regular Interior Degrees).
 From this, we can see that a back arc to a parent container violates the oriented nature of the graph.
-Lemma 7.5.
-No Back Arcs Non-Existence of Back Arc In an oriented graph $G$, with a minimum degree node $v_{0}$.
-Let containers $C_{1}$, $C_{2}$, $\dots$, and $C_{k}$.
-Then there cannot exist back arcs between a node $u_{i}\in C_{i}$ and a node $v_{j}\in C_{j}$, where $j<i$.
-Proof.
-This was proved in the lemmas 7.2 (No Back Arc to Early Container) and 7.4 (First Back Arc Between Consecutive Containers). ∎
 We stated that back arcs could be detrimental to our data structure.
@@
 8 Interior Degrees Doubling
-This section introduces the concept of interior degree-doubling.
-This is like solving independent subproblems, where the subproblems we are attempting to solve here are the doubling of a node’s interior degree inside of a container.
-This can be seen as a localized version of Seymour’s Second Neighborhood Conjecture, but for individual containers instead of the overall graph.
-Example 8.1.
-[…]
+This section introduces the concept of interior degree-doubling, a key element in our proof strategy for the SSNC.
+Several preceding lemmas have relied on this concept.
+Our overall approach is based on a divide and conquer paradigm, where each container is treated as an independent subproblem.
+These subproblems focus on demonstrating the doubling of a node’s interior degree within a container.
+This can be viewed as a localized version of the SSNC, applied to individual containers rather than the entire graph.
 Figure 9: Illustration of the proof of 8.1 (Interior Degrees Double).
-This is another illustration of the Interior Degree Doubling.
+This is a first look at a sketch of the proof of the Interior Degree Doubling.
 Each node inside a container represents a node whose degree is doubled by the inductive hypothesis.
@@
 Lemma 8.1.
-Interior Degrees Double Let $G$ be an oriented graph.
-Let $v_{0}$ be a minimum-degree node in $G$ where all nodes satisfy the decreasing sequence property (DSP).
-Then, at least one node in each container $C_{i}$ will have a node with its interior degree doubled.
+Interior Degrees Double Let $G$ be an oriented graph. Let $v_{0}$ be a minimum-degree node in $G$. Then, at least one node in each container $C_{i}$ will have a node with its interior degree doubled.
 Proof.
@@
 We were able to set up a GLOVER data structure inside this local version starting with a minimum degree node of the container.
-We can then form local containers that are each their own respective distance from the minimum degree node.
+We can then form local containers that are each their own respective distance from the minimum degree node $u_{j}$.
 By strong induction, we have that each of these containers will have a node whose interior degree doubles.
-Then by the weak decreasing exteriors lemma, we have a decreasing set of exteriors inside this container.
+Then by lemma 6.4 (Weak Decreasing Exteriors Lemma), we have a decreasing set of exteriors inside this container.
 When this sequence gets down to size two, that node will have all exterior neighbors.
@@
 This gives rise to why the Decreasing Sequence Property (DSP) received its name.
-Prior to proofs, our research showed the Weak Decreasing Exteriors Lemma as an early observation through many examples.
-The question is whether this was always true or whether back arcs like lemma 7.3 could force this into being a false positive.
+Prior to proofs, our research showed lemma 6.4 (Weak Decreasing Exteriors Lemma) as an early observation through many examples.
+The question was whether this was always true or whether back arcs like lemma 7.2 could force this into being a false positive.
 This proof also helps with explaining the concept of the SSNC.
-Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER, as long as there are no back arcs.
+Any oriented graph in which we want to seek a degree-doubling node can be represented in a GLOVER.
 Once we have done this, there are containers spread out from a chosen minimum degree node.
@@
 The difference is that with lemma 8.1 (Interior Degrees Double), we already know that there is a node whose degree doubles inside that container.
-Then with lemma 9.2 (Indistinguishable Regular Interior Nodes), we know that any node we select in that container will also have its interior degree doubled.
 Example 8.2.
@@
 Now that we have established interior regularity and interior degree doubling, we can start to think about what those things mean.
-The Weak Decreasing Exteriors Lemma implies that containers will decrease in size.
+The 6.4 (Weak Decreasing Exteriors Lemma) implies that containers will decrease in size.
 We need to see how that decrease will impact the degrees of the nodes, the other properties of containers, and the development of our algorithm.
@@
 This implies that the exterior degree (the number of nodes at distance $i+1$ from $v_{0}$) is irregular between $x_{i}$ and $x_{j}$.
-Because Lemmas 8.1 (Interior Degrees Doubling) and 7.5 (No Back Arcs) both hold, we can apply Lemma 6.4 (Weak Decreasing Exteriors Lemma) here.
+Because Lemmas 8.1 (Interior Degrees Doubling), 7.2 (No Back Arc to Early Container), amd 7.3 (First Back Arc Between Consecutive Containers) all hold, we can apply Lemma 6.4 (Weak Decreasing Exteriors Lemma) here.
 This says that with each increase in distance from the minimum degree node, our exterior size is decreased by 1.
@@
 We simply want to observe what effect that has on a parent of these nodes.
-The irregular exterior degree causes a ripple in the decrease in exteriors, which should be guaranteed by the weak Decreasing Exteriors Lemma.
+The irregular exterior degree causes a ripple in the decrease in exteriors, which should be guaranteed by lemma 6.4 (Weak Decreasing Exteriors Lemma).
 The container $C_{i+1}$ is now the same size as $C_{i}$.
 This means that a node in $C_{i-1}$ will have their exterior degree double.
-This is an important property because it ties into the Decreasing Sequence Property (DSP). We were able to use the 6.4 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
+This is an important property because it ties into the Decreasing Sequence Property (DSP).
+We were able to use the 6.4 (Weak Decreasing Exteriors Lemma) in this proof because the two go hand in hand.
 Anything that leads to a hiccup in the decrease in exteriors would cause a node’s degree to double.
@@
 This proof highlights the utility of the lemma.
-It allowed us to construct a proof based on the container distance from the chosen minimum-degree node.
+It allowed us to construct a proof based on the container distance from the minimum-degree node $v_{0}$.
 It is possible that this result could have been proven without relying on induction.
 We could have used a different approach or simply avoided induction based on container distance from $v_{0}$ entirely.
-Still, the use of induction provides an important method, particularly for graphs, where the usage of induction is more limited.
+Still, this use of induction provides an important method, particularly for graphs, where the usage of induction is more limited.
 Induction may not always be necessary, but having a variety of alternative approaches to proofs strengthens our theoretical foundation and enhances our flexibility in tackling these problems.
-Similar to the proof of lemma 6.4 (Weak Decreasing Exteriors Lemma), we assume here that there is a node in each container whose degree doubles.
+Similar to the proof of lemma 6.4 (Weak Decreasing Exteriors Lemma), we assume here that there is a node in each container whose interior degree doubles.
 The difference is that that assumption does not need to be explicitly stated because we have already proved that there is a node whose interior degree doubles in each container.
-Similarly, this lemma requires there to be no back arcs, like the weak decreasing exteriors lemma.
-We’ve already proved that back arcs cannot exist without creating a degree doubling node, though.
+Similarly, this lemma requires there to be no back arcs, like lemma 6.4 (Weak Decreasing Exteriors Lemma).
+We’ve already proved that back arcs cannot exist without creating a degree-doubling node, though.
 As a result, we do not need to state that assumption.
@@
 […]
-Example 9.1 shows some of the patterns that emerge with this regularity in the SSNC.
+Example 9.1 shows some of the patterns that emerge with these interior degrees lower bound in the SSNC.
 Starting from the outside in, going inward represents the order of containers as they increase in distance from $v_{0}$.
@@
 Before we continue with the development of our algorithm and show how it solves the SSNC, we need to address some potential problems and show that they will not hinder our algorithm.
+10.1 Transitive Triangle
 Definition 10.1.
@@
 If we do not properly deal with transitive triangles, we might improperly count out second neighbors and thus not have a correct algorithm.
-Definition 10.2.
-Let $x\in G$ and $y,w\in N^{+}(x)$. Then $x$, $y$, $w$, and $u$ form a Seymour diamond if $(y,u)$ and $(w,u)\in G$.
-The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$.
-This is a four-vertex-oriented graph.
-What this situation presents is redundant second neighbors.
-There is a path $x\rightarrow y\rightarrow u$ in this diamond, and there is a path $x\rightarrow w\rightarrow u$ in this diamond.
-As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
 The notion of interior neighbors is another way of describing what [2] referred to as transitive triangles.
@@
 This is what creates the transitive triangles in every container.
+Lemma 10.2.
+Node in GLOVER’s Transitive Triangles Suppose $G$ is an oriented graph.
+Let $v_{0}$ be a minimum degree node and let $u$ be a node in container $C_{i}$.
+Then $u$ is in three possible locations of two different transitive triangles, which do not hurt the algorithm.
+Proof.
+Suppose we have an oriented graph $G$ with a node $u$ in container $C_{i}$.
+Scenario 1 is that $u$ is a parent to children $v$ and $w$ who are both in $C_{i+1}$, where $v,w)\in G$. Then $(u,v)$ and $(u,w)$ are exterior arcs, and $(v,w)$ is an interior arc.
+Scenario 2 is that $u,v$ are children of $x$ and $u$ is a child of $v$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(v,u)$ is an interior arc.
+Scenario 3 is that $u,v$ are children of $x$ and $v$ is a child of $u$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(u,v)$ is an interior arc.
+In every scenario, we have shown in preceding sections how to proceed through these situations. So these do not hinder our algorithm. ∎
+10.2 Seymour Diamond
+Definition 10.2.
+Let $x\in G$ and $y,w\in N^{+}(x)$. Then $x$, $y$, $w$, and $u$ form a Seymour diamond if $(y,u)$ and $(w,u)\in G$.
+The Seymour Diamond is a graph theory concept in graph theory where a node $x$ points to two out-neighbors $y$ and $w$, and these two out-neighbors both share the same out-neighbor, $u$.
+This is a four-vertex-oriented graph.
+What this situation presents is redundant second neighbors.
+There is a path $x\rightarrow y\rightarrow u$ in this diamond, and there is a path $x\rightarrow w\rightarrow u$ in this diamond.
+As a result, we have to investigate Seymour diamonds more to understand how they impact the graph and, more importantly, the Seymour’s Second Neighborhood Conjecture.
 Lemma 10.3.
@@
 This is also what we saw in our regular graph example in 9.1 in section 8.
-Lemma 10.4.
-Node in GLOVER’s Transitive Triangles Suppose $G$ is an oriented graph.
-Let $v_{0}$ be a minimum degree node and let $u$ be a node in container $C_{i}$.
-Then $u$ is in three possible locations of two different transitive triangles, which do not hurt the algorithm.
-Proof.
-Suppose we have an oriented graph $G$ with a node $u$ in container $C_{i}$.
-Scenario 1 is that $u$ is a parent to children $v$ and $w$ who are both in $C_{i+1}$, where $v,w)\in G$. Then $(u,v)$ and $(u,w)$ are exterior arcs, and $(v,w)$ is an interior arc.
-Scenario 2 is that $u,v$ are children of $x$ and $u$ is a child of $v$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(v,u)$ is an interior arc.
-Scenario 3 is that $u,v$ are children of $x$ and $v$ is a child of $u$ where $v\in C_{i}$ and $x\in C_{i-1}$. Then $(x,u)$ and $(x,v)$ are exterior arcs, and $(u,v)$ is an interior arc.
-In every scenario, we have shown in preceding sections how to proceed through these situations. So these do not hinder our algorithm. ∎
 Now that we have a better understanding of transitive triangles and Seymour diamonds and how they interact with our data structure.
@@
 Next, we will blend everything together into our main results.
-11 Main Theorem: Seymour’s Conjecture is True
+11 Main Theorem
 11.1 Introduction
@@
 There is a limit, though, to how large the interior degree can grow. It is bound by $\frac{n-1}{2}$ by the limit of average degree inside a graph.
-We first restate the Lemma 7.5 (Non-Existence of Back-Arcs), from section 6, but slightly rephrased.
-Lemma 11.1.
-Let $G$ be an oriented graph, with all nodes satisfying the Decreasing Sequence Property (DSP).
-Let $v_{0}$ be a minimum degree node in $G$.
-If there are back arcs in $G$, then a node’s degree will double.
 11.2 Theorem Statement
@@
 This will cause a parent of that node’s degree to double.
+∎
+Input: Graph $G$
+Output: Halt condition
+Initialization:
+Choose the minimum degree node $v_{0}$;
+Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$;
+for each container $C_{i}$ in $G$ do
+       Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles;
+       Move to a neighboring container $C_{i+1}$;
+       if there is a back arc from $C_{i+1}$ to a prior container then
+             Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double;
+             Halt;
+       end if
+      else
+             Since there is no back arc, the container $C_{i+1}$ will be limited in size to be 1 less than the size of container $C_{i}$;
+             Repeat from Step 4;
+       end if
+end for
 Algorithm 1 Algorithm for Analyzing Degree Doubling in Containers
-1:Initialization:
-2:Choose the minimum degree node $v_{0}$.
-3:Divide nodes into containers $C_{1},C_{2},\ldots$ based on their degree and proximity to $v_{0}$.
-4:for each container $C_{i}$ in $G$ do:
-5:     Locate a node $x_{i}$ in $C_{i}$ whose interior degree doubles.
-6:     Move to a neighboring container $C_{i+1}$.
-7:     if there is a back arc from $C_{i+1}$ to a prior container then
-8:         Locate child of $x_{i}$ whose exterior degree doubles, causing $x_{i}$’s overall degree to double.
-9:         Halt.
-10:     else
-11:         Since there is no back arc, the container $C_{i+1}$ will be limited in size to be 1 less than the size of container $C_{i}$.
-12:         Repeat from Step 4.
-13:     end if
-14:end for
 Example 11.1.
@@
 In small graphs, such as the one shown in Figure 4.3 (Minimum Out-Degree 3 with Neighbors 1), the collision phenomenon was not observed.
-This is because the container sizes decrease linearly (at a rate of 1), while the interior degrees increase linearly (also at a rate of 1).
-However, the number of arcs within the oriented graph grows quadratically, at a rate of $²O(n\texttwosuperior)$.
+This is because the container sizes decrease linearly (at a rate of O(n)), while the interior degrees increase linearly (also at a rate of O(n)).
+However, the number of arcs within the oriented graph grows quadratically, at a rate of $O(n^{2})$.
 Consequently, while the number of nodes increases linearly, the number of edges increases quadratically.
@@
 This suggests that the algorithm is not only theoretically sound but also practically feasible for implementation and use in real-world scenarios, such as analyzing social networks, identifying influential nodes in communication networks, or understanding the spread of information.
-12 Conclusion and Future Work
+12 Applications
+The Seymour Second Neighborhood Conjecture is not just a theoretical problem; it has practical applications.
+We have constructed an algorithm that searches for and finds nodes that satisfy this conjecture.
+Not only does our algorithm find these nodes, but it partitions the graph into a data structure, which can be beneficial for many other purposes.
+While the conjecture only asked for a single node, our approach finds a set of degree-doubling nodes, which are critical in understanding both network structure and influence dynamics.
+We have developed an algorithm that efficiently identifies degree-doubling nodes.
+This algorithm has a complexity of $O(|V|+|E|)$, making it scalable to large graphs.
+In social networks with millions of users, the efficiency of the degree-doubling node identification algorithm is crucial.
+A less efficient algorithm would make it impractical to analyze the network and identify influential users within a reasonable time-frame.
+A/B testing is often considered by companies when thinking of introducing a new product or service to customers.
+Users are randomly divided into two groups, a control group and a treatment group.
+It gives data driven decisions that focus on user feedback.
+Network A/B testing extends this methodology to social networking experiments, where users are part of an interconnected system.
+Similar to traditional A/B tests, different versions of a post or an ad are shown to different audiences to compare engagement metrics such as clicks or conversions.
+However, network A/B testing comes with significant challenges, primarily due to network interference—a phenomenon where users in the control group interact with users in the treatment group, unintentionally influencing the results.
+This limitation causes some researchers to dismiss network A/B testing entirely.[12]
+Network A/B testing often relies on assumptions about influence, which can lead to misleading results.
+The first assumption we will speak of is homogeneous influence.
+This says that all users have the same influence, which is clearly not true.
+However, Influence varies greatly across users, and failing to account for this can distort test results.
+Assuming all users have the same influence can lead to misrepresenting the results.
+A small change that affects a highly influential user can appear to have a large overall effect, even if it has little impact on the majority of users.
+[13]
+A second assumption that we will mention is random influence.
+This assumption says that influence is spread randomly throughout the network.
+If influence isn’t random, but concentrated in communities, then randomly assigning users to control and treatment groups can lead to imbalances.
+One group might inadvertently contain a larger proportion of influential users, skewing the results.
+[14]
+Finally, there is the assumption that users are independent of each other and their actions are not influenced by the actions of others.
+This is an oversimplification, as social networks are built on interconnectedness.
+This is perhaps the most critical flaw.
+Users in social networks are not independent.
+They talk to each other, share information, and influence each other’s opinions.
+Ignoring these network effects can completely invalidate the results of A/B tests.
+Some may make an assumption that only type of influential users are celebrities.
+These celebrities certainly may have a large name recognition, but they do not represent every influential user.
+Conversely, every known name is not necessarily an influential user.
+There are many potential influential users that this assumption misses like community leaders, subject matter experts, connectors between different communities and early adopters.
+The Graph Level Order (GLOVER) data structure is a way to organize an oriented graph into leveled containers.
+Nodes are placed based on their distance from a source node (normally a minimum degree node).
+Nodes are stored into ordered containers $C_{1}$, $C_{2}$, …, $C_{i}$, where $C_{i}$ contains nodes that are distance $i$ from the source node.
+Each container then connects to its next container.
+The first thing this data structure does is helps with the problem of partitioning data sets.
+By the definition of the data structure, these sets $C_{i}$ and $C_{i+2}$ have no edges in common.
+This allows us to more easily set up the control group and the treatment group.
+The partitioning ensures that there can be no spill-over or network interference between the two groups.
+The GLOVER data structure, by organizing the graph based on distance from a source node, helps to minimize network interference.
+Because nodes in different containers (e.g., $C_{i}$ and $C_{i+2}$) have no direct connections, the likelihood of users in the treatment group influencing users in the control group is significantly reduced.
+This allows for more accurate measurement of the treatment’s true effect.
+The GLOVER data structure, combined with our algorithm, allows for parallel processing of the graph.
+Because the containers in GLOVER are independent, the computation of degree-doubling nodes within each container can be performed concurrently.
+The ability to parallelize the algorithm significantly reduces the computation time, especially for very large social networks.
+This makes it feasible to analyze the network and identify influential users in real-time or near real-time, which is essential for dynamic network A/B testing.
+Secondly, we have solved the SSNC with a set of degree-doubling nodes.
+These nodes are are proven to be mathematically influential.
+Thus, we can go back into the original social network, knowing these nodes are influential with more confidence about both our control and treatment sets.
+By identifying degree-doubling nodes, we can strategically select users for our A/B test.
+These nodes are likely to be highly influential within their communities.
+Including them in the treatment group allows us to observe how the new feature spreads through the network, providing a more realistic assessment of its potential impact.
+Alternatively, we might choose to exclude them from the treatment group to see how the feature performs without the amplification effect of these key influencers, giving us a different but equally valuable perspective.
+Figure 14: Social Media Example shown with and without exterior arcs.
+The nodes in container 3 are the degree-doubling nodes, aka the influencers.
+Knowing that, we can develop a more informed treatment group, partitioned from the control group.
+While this example uses a simplified network of 30 nodes, our algorithm can efficiently handle much larger networks due to its $O(|V|+|E|)$ complexity and parallelizable nature.
+In a real-world scenario with millions of users, the computation time would be significantly reduced by parallel processing, making it possible to identify degree-doubling nodes and design network A/B tests efficiently.
+This application of the SSNC seeks to plant the seeds for future research in network A/B testing.
+There are other issues that the GLOVER algorithm does not resolve.
+Another limitation is that this algorithm will require an oriented graph, instead of a standard directed graph.
+that may limit applications as well.
+However, this research may hold a key to unlocking network A/B testing.
+13 Conclusion
 In this work, we explored the Seymour Second Neighborhood Conjecture (SSNC) in oriented graphs.
@@
 This would then allow us to implement resources from the field of graph theory into subject matter dealing with machine learning, data science, and network flow.
-Future research and development should explore questions of complexity and scalability.
+Future research and development should explore other possibilities for this data structure as well as their complexity and scalability.
 Ultimately, this paper lays the foundation for a dual-metric framework that not only reaffirms the importance of distance and degree in the Seymour conjecture but also opens pathways for innovation across graph theory, computer science, and data science.
-Finally, this research offered an alternative method for the SSNC.
-We focused on identifying a node whose out-degree doubles in the graph’s square rather than just showing that it exists.
-To do this, we developed an effective algorithm and proved its complexity.
-Other problems in graph theory and related domains may benefit from this change in emphasis from existential proofs to algorithmic solutions.
-This may encourage researchers to explore constructive approaches with a focus on practical applications.
-Hopefully, this work will inspire further research on constructive approaches to other graph theory problems.
-Also, we encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
+Finally, this research offered an alternative method for the SSNC, moving beyond existential proofs to a constructive, algorithmic solution.
+Instead of simply demonstrating the existence of a node whose out-degree doubles in the graph’s square, we focused on identifying such a node through a novel approach.
+This involved several key innovations.
+The first if these techniques is graph partitioning.
+We dissected the problem by strategically partitioning the graph.
+This revealed underlying properties that were the about these partitions and how they relate to one another that were crucial to our algorithmic solution.
+This partitioning allowed us to analyze the relationships between a vertex and its second neighbors in a new light.
+Secondly, we employed a divide-and-conquer strategy.
+After breaking down the problem into smaller, more manageable subproblems, we solved them in an independent nature.
+Returning these independent subproblems back to the overall problem gave rise to a global solution.
+This approach, combined with our partitioning technique, enabled us to develop an efficient algorithm.
+Lastly, we introduced the concept of dual metrics.
+This allowed us to simultaneously consider multiple aspects of the problem simultaneously.
+This multi-faceted perspective proved essential in connecting the local degree properties of vertices within a container to the global neighborhood property of the distance required by the conjecture.
+These techniques, particularly the graph partitioning and divide-and-conquer strategies, may be applicable to other problems in graph theory and related domains.
+While the specific degree and distance dual metrics used in the SSNC proof might not be directly transferable, the general principle of considering multiple perspectives could be valuable.
+It is possible that other problems will require the development of new data structures and metrics tailored to their specific characteristics.
+However, the core idea of strategically dissecting the problem, as demonstrated in our solution to the SSNC, offers a promising direction for future research.
+This shift in emphasis from existential proofs to algorithmic solutions, combined with the development of new data structures and analytical techniques, may encourage researchers to explore constructive approaches with a focus on practical applications.
+We hope that this work will inspire further research on such approaches to other graph theory problems.
+We also encourage greater collaboration between mathematicians and computer scientists to explore the algorithmic and practical implications of theoretical results.
 14 Acknowledgments
-I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:6 has been a guiding light throughout this journey.
+I am deeply thankful to God for the inspiration and guidance that shaped this work, as Proverbs 3:5-6 has been a guiding light throughout this journey.
 I am profoundly grateful to my family—my mother, father, and wife—for their unwavering support and encouragement.
@@
 The GLOVER data structure is protected by a provisional patent application filed with the U.S. Patent and Trademark Office.
-A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
-Ranked Structure: The elements of $S$ are partitioned into a ranking of ”containers” or ”levels,” $C_{1},C_{2},\dots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
-The containers themselves are ordered relative to each other, establishing a multi-level structure where all elements in $C_{i}$ are ordered before those in $C_{i+1}$.
-Non-triviality Requirement: At least one container $C_{i}$ must contain more than one element, i.e., $|C_{i}|>1$, to avoid trivial total orderings where all elements are isolated in singleton containers.
-Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion, such as a designated metric or property relevant to the application.
+A Graph Level Order (GLOVER) on a set $S$ can be defined as follows:
+Leveled Structure: The elements of $S$ are partitioned into levels of containers $C_{1},C_{2},\ldots,C_{n}$, where each container $C_{i}$ forms a subset of $S$ that maintains a total order.
+Total Order on Elements of G: We need that every node in $C_{i}$ is ordered before all nodes in $C_{i+1}$.
+Comparability within Containers: For any two elements $u,v\in C_{i}$, their order is determined based on a specific criterion.
 Consistency Across Containers: For any two elements $u\in C_{i}$ and $w\in C_{j}$ with $i<j$, $u$ is considered less than $w$.
-This enforces that the order of elements between containers aligns with the order of the containers themselves, establishing a coherent multi-level total order.
 Interior and Exterior Neighbors:
-  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors.
-  • If $x\in C_{i}$ and $y\in C_{j}$ where $i<j$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
+  • If $x,y\in C_{i}$ are nodes in the container $C_{i}$, then they are considered interior neighbors if $(x,y)\in G$.
+  • If $x\in C_{i}$ and $y\in C_{i+1}$, then $y$ is an exterior neighbor of $x$ if there is a connection between $x$ and $y$.
 Definition 15.4.
@@
 Where possible, things like HTML canvas and JavaScript D3.js are used to visualize these concepts and help improve understanding.
-Readers are encouraged to explore the website for additional information and examples at https://glovermethod.com.
+Readers are encouraged to explore the website for additional information and examples at
+https://glovermethod.com.
 This is an evolving project. Additional examples and applications will continue to be added.